A Bibliography of Publications in Applied Mathematics and Computation: 1977–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 December 2023 Version 1.23

Title word cross-reference

(k, n − k − 2) [AW99b]. (r, s) [PSMM93]. 1 [Ada99, Ram97a, RGL97a]. 2 [Ada93b, LS98a, Tak98, Van95, Zak89]. 2n [YK97a, RUS^{+}84]. 3 [Cha95a, Hei91, Kor97, TT86, TT88]. A [Che96a, Zho89]. A_{T,S}^{(2)} [Che96a]. A_{MN}^{+} [Wan87]. AX + XB = C [SH92]. b [KKS87, Ham94a, Rob82]. C [Dul82]. C2 [Wat79]. C^{0} [Lin98]. C^{1} [Sai88]. C^{3} [Tes80]. C^{∞} [Sas95a]. D [Pat92]. e^{nx} [Kin91]. F [MP90]. F (u(t), t, λ) = λ ∫ 10k(t, y, u(y))dy [KK83a]. f(x, y) = 0 [Mit90a]. GVψ_{0} [Kin94]. h [CW97, KKS87, DK96]. H^{+} [BMS92]. h^{2} [IST93]. hp [IBS94]. k [BR88]. K^{+} [BMS92]. L [GG91b, CG88a, GB95, Zho89]. L^{2} [Rap92].

L^{p} [GK94]. L_{1} [De 94]. L_{2} [Yür99]. l_{p} [JMS98, DEV95, Sen93]. λ [HD88]. LC^{1} [Sun98]. Lg [SW79b]. LH [KMR^{+}91]. M [ABT98, MS87b]. M(t)/G/∞ [Pat87, PM88]. M^{[x]}/C_{2}/S/N [MES94]. N [Cha93, JkC97, SMO91, AP95, CY80, Ham97, Kan97, Nar99a, SV92b, Fei87]. ∇^{2}V (x_{j}) = 0 [Yan99]. n ≥ 0 [Kin91]. O(2) [ST96]. O(n^{2}) [Wai90]. Op^{ti}A [DF96]. P [Sim92, Do96, FLG92]. P (Z < Y) [Ami99]. P 1 [YK97b]. P_{1} [KL99a]. ϕ_{0} [Sol99a]. Q [CDW91, GBS99, JkC97, LS93b, DBW88]. R^{n} [Mil87b]. S [MBS98, Sav93, ASZ94]. T [AM98b]. θ [Ram98a, Ram98b]. u_{xyz} = f [DR88]. V [XL98, Par85]. W [WH85]. W_{2}^{(1)} [Chu97]. x′ = y [RK87]. x ≤ 0 [Kin91]. y′ = 0 [RK87]. Y′′ = f(t, Y) [JMV93]. Y_{n}(z) [ZB97]. z [Tag98, VA88]. ζ(2) [CRS99]. ζ(2n + 1) [Sri98].

-Adaptive [CW97]. -analysis [DBW88, CDW91]. -approximation [DEV95, De 94]. -ary [GBS99]. -Bit [Cha93]. -city [Fei87]. -Class [Sas95a]. -Cycle [Do96, Par85, MP90, XL98]. -D [RGL97a, Cha95a, Ram97a]. -Dimensional [Kan97, BR88, CY80, Nar99a]. -Directed [PSMM93]. -equation [DK96]. -expansion [IST93]. -finite [IBS94]. -homogeneous [MS87b]. -Markov [LS93b]. -matrices [HD88, ABT98]. -methods [Ram98a, Ram98b]. -node [Pat92]. -nonconforming [KL99a]. -norm [JMS98]. -Periodic [AM98b]. -person [SMO91]. -Phase [Ham97]. -shaped [ASZ94, CG88a]. -Simplexes [JkC97]. -solvability [GK94]. -spline [Ham94a, SW79b]. -stability [Sol99a]. -stable [Zho89, Sim92]. -symmetric [ST96]. -system [Sav93, MBS98]. -systems [GB95, GG91b]. -Tori [Tak98]. -transform [VA88]. -transforms [Yür99]. -Triangulated [JkC97]. -Type [MS94c, Dul82].

1 [Jac95].

2 [CXL94, Kle82, LYX^{+}93]. 2-D [CXL94, Kle82, LYX^{+}93]. 20 [LMM86]. 200 [LMM86]. 205 [BB83b, Sca89]. 24 [LMM86, MS92c]. 29 [MS92b]. 2nd [VD97].

3 [CCSK82, Jin94, SK96]. 3-D [CCSK82, Jin94, SK96].

50 [OT93].

60 [JNF95]. 63 [Cab95]. 6400 [SM88]. 65th [Kag91a]. 66 [Sor97a].

79 [Di 98].

83 [JS98a]. 88 [Mus97b]. 89b [MS92c].

90f [MS92b]. 93d [OT93]. 95c [JNF95]. 95h [Cab95]. 97f [Di 98]. 97i [JS98a].

Abandonments [Mil95]. Abelian [Pig87]. Ablowitz [TPGV97]. Abrupt [Abu97]. ABS [ABT98]. Absolute [SBH96, Waz97c, AK80]. Absorbing [Soc88]. abstract [Du83, EL77, Opp92]. abstraction [SHR94]. Accelerating [Boy95, GW89b, LS81, SL99]. Acceleration [Bar91, Zha96b, Zha99a, BCG91, Cha90a, Joh83, Zha98b]. acceleration-pressure [Cha90a]. accelerator [CK90]. accidentally [Ten91]. Account [HZ95b, LC97]. Accumulation [WZ98]. Accuracy [Cic93b, HY99b, Sar93, HM99, Zen99]. Accurate [MEZI93, MRJ95, PR93, BTW94, BS98, Hos99, HY99a]. Across [BMS92]. Action [MS97, IA89]. actions [Mat86]. Active [IOAB95, YF91]. Actuators [Bal94]. Ada [Hus90]. Adaptive [AH94, ADS96, CW97, DSK82, HAST98, IBS94, KT91, KKH95, LL94b, MBCV95, NN87, RRB95, Sug95, SU95, SY93, TSK98, XS95, Zas93, ZZZ98, AR82, And87, And90, Bra82, Gno82, MLC87, MBS80, SB82, Shy87, Swo91, Van83, YS94, ZS98]. adaptive-grid [Gno82]. Addendum [JS98a]. Additive [MDR94, XL98, YK97b]. Adjoint [MSLB92, NR99b, NC85]. adjoint-state [NC85]. adjoints [Gol79]. admissible [AH99b]. Admission [Mil95]. Adomian [Waz98a, Waz99d]. Adsorbers [RBPK96]. advanced [JS98b]. Advances [FpDS99, Zla89, KB80]. Advection [Waz95b]. advective [Hos99, HY99a]. advective-dispersive [HY99a]. aerial [AKK99]. Aeromedical [MFS93]. Aesthetic [Mur96, Mur89a, Mur89b]. Aesthetics [Mur94]. Affects [LHK97]. affinity [Bay98a]. afterbody [Hof87, Hof88]. Age [IKP97, KP98, KR95a]. Age-Structured [IKP97, KP98]. Agents [JL97]. Aggregation [AB94, AB95b, SK86]. aggregation-disaggregation [SK86]. Aging [PLPW96]. Agriculture [SAF^{+}97]. AI [Bad93]. aided [TCS92]. AIDS [Yak92b]. Air [Sla93a]. Airflow [MKJ87]. Airfoil [Gol93, FG81, Gol88a, Sor82]. airfoils [Hal82]. Albedo [KUO95]. Alegebra [BG97]. Algebra [DS86, GSI96, Ada98, Bay98b, Bay98a, BY99, KR81a]. Algebraic [Ben92, Bra86, CWL92, Eis98, Hof88, JN95, Koč92, LP97, Man88, MBS98, NG94, Smi82, Stü83, Zas93, Bai91, DLRJ81, Eis82a, Gre87b, HT88, Hua91, Kha91, KM87, KM79, Lam99, Mär89, Sas83, Sav93, Sen90, TGK91]. Algorithm [AH94, AR93, BG93, Die96, EA94, JL97, Jin95, LAK94, LSS94, LSVS94, LL94b, Mor94, Mor95, PS96, Sch92, SS98b, SW94, WJ92, XZW96, Yak92a, Yan94, ZZZ98, dYS97, Abu99b, AK80, Bar91, CK86, Cro80, EMEH96, GS99, Ge89c, Gno82, GMGCL99, HT98, IA80, KRTW83, KMR^{+}91, KK83b, KKL99, LE85, LCE85, LW90, MLC87, MBS80, MW88a, Mor90, ONOT98, Pes91, Rav87, RG98, SK86, SW85, SW90a, SHR94, SSY99, TZ83, TS87, Wai90, Wan87, WC88, Wat79, Wat81, WI80, XZL98, ZZ98, Zho89, Zhu98, ZL98c]. Algorithmic [AS92, AS90]. Algorithms [CLC97, DHK90, Kwo93, Lia97, MZV96, MFGW97, Per96, SH92, SR97, Van83, XZL97, ADM91, CS85, CK79, Cha91, Cor89, GMS83, GT91, Gol87, HT88, HL91, HK84, JBV98, JCC98, KM86, Mel86, Par89, Sch91b, Sen86, SW79b, Szi90, TT88]. alkyl [KK91]. All-Time-Step [RTJL92]. allied [CRS99]. Allocation [BDS96, AB87, Dro90, Les89]. allowable [Jam85]. Almost [Kin97, Sim92, OM99]. ALOHA [Sch92]. Along [CFT^{+}93]. Alternate [UK95]. Alternating [Ela94, Jin94, KR95a, SL96, Szi97, CG88a, Lon88]. Alternative [EB91, GT91, Ji94, MSZ96, Ham98a]. Alternatives [MDR94, MS99, YB99].

Ambiguities [Kin97]. American [Váz98]. AMG [Stü83]. Among [ITN97, LL94a]. Amphiphilic [Kub97]. Amplitude [KSZ^{+}95]. Amplitudes [Lan92]. analyses [PM88]. Analysis [AFM97, Axe86, BDS96, CW92, CJF93, Con93c, Do96, DML98, EA94, El 92, FL94, FH95, GP94, Gra97, GL99, HB91, Hir84, JC97, JFW96, JV90, KTW81, Kon92, LS95, MZ98, MGL92, MFS93, Mei93, NW90, OL80a, Pap94, RAL97, Sas85, Sci97, SPD93, TH96, TJ95, Tuc96, Van97, Var83, WK95, ZA96, ABT98, AAES99, AHKK80, ALZ99, ALMS90, Baa88, BBD91, CXL94, CDW91, CK90, DBW88, EBR85, Eno90, HJ92, HG84, HEH81, HDP85, HF91, HC87, IBS94, Jam85, KR89, KK84, Kum92b, KL89, KR91, LS91a, LCE85, pLR87a, pLR87b, MD98, Mat86, MU90a, MU90b, Moh99b, Mor98, MH81, Pal86, PZZy99, Ram99a, SSV89, SR88, SS92, SHR94, Tes91b, VJ88, VR99a, VR99e, VR99f, VWSS94, Wan81a, Wan81c, Wan81b, Won98, WKM84, Wu94, Yag82]. analysis [YF91, YL98a, Zel87, ZA88, Zha98b]. Analytic [Ado88, ARM97, Ioa92, JV94, MN91, ZL98a, CSY98, DS99, Har87, IA85, Mil88a, SV85, SW90a]. Analytical [AP87, KR81a, Sas83, SA96, SK95, TJ95, Waz99a, van85, Got84, IA85]. Analyzing [Bay97, HKLS93, Jac95, PN97b]. Anisotropy [SR97]. anomalies [ZA88]. Answering [HKLS93]. any [Ada93b]. AOR [Bai99a, YJ99]. Aperiodic [Wan81a, Gae87]. Appearance [TyG99, Waz97a]. Appel [MS93b]. Appl [AaAaZ00b, AaAaZ00a, Cab95, Di 98, JS98a, JNF95, MS92c, Mus97b, OT93, Sor97a]. Appl. [MS92b]. applicable [Mus80]. Application [AB87, AK96, Als79, Bay98b, Bay98a, Cai93, CH96, Ioa88, Ji94, KKH95, Kni82, KA96a, Lia98, Lin96, Mer86, MZV96, Pap84, PY94, RS86, RRB95, SZ97, SKG97, SS92, SGB87, SP94, UR97, Waz95b, Yee93, Yil96, AW99a, AP87, Fym79, GT82, Har87, KKS87, KN90, Khu98b, Kle82, KK90, Kum92b, MNK99, Mur91b, RMR86, Sas83, Ser79, Smi84b, Sos88, Ste82, WW99a, WW99b, ZL98b, Zie84, van85, AM98a, HS99]. Applications [Arg96a, BH93, CS97, Cic92, DGP93, Eva90a, Eva90b, ER94b, ER94a, GS86, LS91a, MA95, MD81, Par95, SS98a, SB82, WC88, ABT98, AST86, Arg99a, Arg99b, Arg99d, CSY98, Chu97, GPAB83, LF99, McW82, Mus78, SYS94, Tha84, Wat81, Yür99, Zha98a]. Applied [Bel95, DGU96, Mus94, DK96, Faz90, Kag91b, Kam88, Mus84b, NSE81]. Applying [MCD89]. Approach [AAHM93, Arg96a, BD94, Ezz94, FHG96b, Ham97, Lei95, Mus97a, RTJL92, RR93, SEG96, SST95, Uen95, Waz95b, Arg99d, BS98, BCG91, BH80, Cha95a, CBYK99, EGH98, Fie84, FMR86, GH89, Gre87a, Ham98a, IA85, JS98b, JB89a, Khu98b, Lak81, LSP91, LB92, Mad90, Mus97b, OŞL80, ORSS83, Rap92, SK98a, Tag99, Uen89, Uen91, Váz98, Wei84, Yip92, YW85]. Approaches [MSZ96, SK98b]. approximants [Waz99a, Waz99c]. Approximate [Jód92, JPF95, KR87b, Kha81, Kin93, SS94, She90, Waz95a, YST97, AD98, CK79, KCS90, SR90, Vál88, XZL98, Zhu98]. Approximating [MZ98, SUJ98, Ada93b]. Approximation [BPZ96, Beh96b, IKP97, Jol92, MS93b, MS97, MS95b, Pap96, Pav95, Sar93, SK95, VD97, WR96, Yan94, ZW95a, vGG79, Bow78, CyGK99, CK86, DEV95, DDL93, Fun91, GZ84, GL99, Jac90, JMS98, KN90, Kum92b, MS93a, Mv98, Opp92, PVM88, De 94]. Approximations [Hal98, Jac92, Let97, RCG94, ZB97, AM81, AM98a, Gar89, Ham94a, HH99, Li98, MKJ87, NR99b, Pat91, Ser79, WS87, Waz99a]. April [Tho82d]. Aquaculture [PN97b]. Arbitrary [Wan94, ESR99, GS94b, WZ82]. arc [CK90]. Architectures [El 92, KKPL96]. arcs [Mit99]. Area [KUO95, Wan80]. areas [DG87a]. Argument [ZB97, AW87, AY98, Ino99]. Arguments [LYC93, KR87c, LG99, LF99, Pei99, WA98]. arid [Fis85]. Arising [KW96, KW97, PA94, Sas95b, Waz97b, ALZ99, DH99a, EM79, Khu98c, LP81, Mit84, NR99a, RR98, SW89]. Arithmetic [GR97, Jay96, SW90b]. arithmetics [Mus80]. ARMA [CC94, NH86]. armature [CK90]. array [CCSK82]. Arrays [DGV94]. Arrival [Mil95, MS95b]. Arrivals [Mil95, MES92]. art [SW79a]. articular [EM79]. Artificial [GKn91, MZ98, PZZy99, BTW94, JS98b, KG91, YL98b, ZSL98, ZPY99]. ary [GBS99]. Aspects [Ham92, KH96, KW86, Bog91, Die89, KR81a, MO92, NS89, YR84]. Assay [MMMK96]. Assessed [MDR94]. Assessing [KK82]. Assessment [Bha94, AS99, Got84]. asset [Ree87]. assignable [JB89b]. Assignment [MTP93, Mou96]. Assisted [Kor92]. Associated [GS97, PN97a, Ber89, CSZ98, DS99, KR88, MS95a, MR99, WLTS99]. Associative [MKT96, KKU91, KKM90]. Assumptions [Arg94, CWL92]. astrophysical [BCZ89]. Asymptotic [CC94, EM79, FW98, GS94b, GL94, HZ95a, KR89, Kar98, Kim96, Koł95, LP89, PSMM93, Ram99a, RSS89, Rua92, Sas95a, VB92, WA98, Zag92, Zha91, Ado87a, Boy89, Kur99, MN79, NR99b, VR99e, VR99f, vGG79, VR99a, VR99e, VR99f]. asymptotical [LS91b]. Asymptotically [WA96]. Asymptotics [Sch92, Doa90, Doa91]. Asynchronous [Bai98, Bai99a]. Atkinson [GC94]. atmosphere [Ten90]. atmospheric [ÖGDB98, Uen80]. atoms [RR98]. Attainability [HK96]. Attended [Mou96]. Attenuation [LR95]. attitude [HH85]. attracting [Sol99b]. attraction [Kur99]. Attractivity [PPS94, ZAES96]. Attractors [DVD95]. Attribute [MDR94, MS99, YLS93]. Attributes [YLS93]. audit [Ais88]. augmentor [Sor82]. augmentor-wing [Sor82]. autocorrelated [Hir85]. autocorrelation [BR99b]. Autocorrelations [CC94, MN79]. Automata [Ada93a, Kub94a, Ada99, Hog88, KK90, Kub94b]. Automated [MB82, Tes91b]. Automatic [Eis82a, Hus90, KS84, Rob82, SW92a, Wex87, Bai81, KRT83]. Automation [KPT87, CDW91]. Automaton [NG94, Ada97, KK91, YGH90]. Autonomic [Gun94]. Autonomous [Cha94, HAST98, IST93, KG91, Sau83]. autonomously [GKn91]. Autonomy [Gun93, Shi93]. Autoregressive [LSKH96, AAAN98, MN79]. Available [MTP93]. Average [Sch92, Kha81, MF99]. Averaged [Boy95]. Averages [HKLS93]. Averaging [HKLS93, BM89]. avoid [Mor86b]. Avoiding [JL93]. away [BD90]. axisymmetric [Ade94, HF93, Khu98a, ZL98a].

backtracking [Ye99]. backward [Ske89]. Balance [YH97, Den90, Pat89]. ball [SW90a]. balls [WW99a]. Banach [Arg90, AC94, Arg96a, Arg96b, Arg99d, Arg99e, CAQ93, Guo99, Ibr96, LS78, Liu94, PV81]. banded [Sca89]. Bandit [PY94]. Bandlimited [Kha92]. Bar [PZ94]. Bargaining [SL96, Szi98, SL98]. baroclinic [Dua99a]. barrier [AS99]. Base [RRB95, DBW88]. Base-Isolated [RRB95]. Based [Bad93, Cro95, GSK93, Kha96, Kon92, LR95, MFGW97, TSK98, XS95, AG99, Bar91, Cic99, IA89, JS98b, KM94, Mos90, MH80, PL93, RGL97a, Sla91, Sug87, VR99b, VR99c, VR99d, ZSL98]. bases [Yse86, ZL98b]. Basic [TT86, War82, CY99]. Basin [BDS96, Nar99b]. Basis [KR95b, PJ95, RK93, Boy89, FS98, HH99, IPP91, Li98, ZM91]. Bayes [ADS96, TW79]. Bayesian [AB94, AB95b, FMR86, HW80, KK84, Won98]. BDF [Wat89c]. Beam [HZ96, Lan92]. Behavior [ESS94, ES94, Ela94, Har97, Kar98, TPL95, Zag92, ES90, Gae87, GS94b, IFU98, PGHX99, RSS89, SV85, dPL84, vGG79]. behaviour [WA98]. Being [Mat93]. beliefs [Ada98]. Bell [ESR99]. Bellman [ZED91, EORA99]. Bending [KWS95, XL98]. Bergman [RMR86]. Bergman-type [RMR86]. Bernstein [CS96, Jia99]. Bessel [RK93, Sri94, ZB97]. Best [Kag91a, TM92, HQU88]. Bet [Fin95]. between [Bur96, Cha95b, Fra89, GSI96, Jam85, JY97, Van97, WZ82, Waz98a, Yam95, ZK98]. Beyond [Mur92, Ste84]. Bezout [Wam92]. Bi [MPD97]. Bi-objective [MPD97]. Bias [Jac93]. Biasing [Cic93a]. Biattribute [SPD93]. BIBO [GWL94]. Bicriterion [PL93]. bifurcate [Sey81]. bifurcating [KK83b]. Bifurcation [Cas94, DVD95, Gra97, KS92a, Le98, LLL97, PBL92, dPL84, ASZ94, GLL98, HEH81, HK84, KM79, KK83b, KR91, Lui99, Nam90a, Nam90b, Sau83, WKM84]. Bifurcations [CS88, FM96, Tak98, MR89]. biharmonics [SSB82]. bilateral [KNS81]. Bilayer [Kub97]. bin [Dro90]. Binary [DGV94, Ada98]. binding [Bay98a]. Bingham [Han92]. Binomial [Jol92, AAAN98]. Bioactive [RBPK96]. Biochemical [TW90]. bioeconomic [Gar90, Mos88b]. Biological [Shi93]. Biology [Mat93, Con89, Les89]. biomass [BD90, Ten90]. biomass-carbon [Ten90]. biomolecular [ZS98]. biomolecule [Ten91]. Biophoton [BKS98]. Biophysicalism [Con93b]. Biorthogonality [Khu98a]. Biped [IOAB95]. Biphasic [HLXZ97]. biplot [Baa88]. birthday [Kag91a]. BISTAB [WKM84]. Bit [Cha93]. Black [Den83, Den86, Den88, Zha96b, KL89]. blades [WC88]. Block [Bai97, BT97, Die96, Jin95, Bai99b, Hil80, Jin94, Zho89].

Block-Triangular [Die96]. blocks [Ada99]. Blow [CZ99]. Blow-up [CZ99]. Blurred [FG95]. Bob [Cas91, Kag91a]. bodies [GGR91, WZ82, WN91]. Body [Tuc95, AA99b, Ste82]. Boltzmann [MPS97]. Boolean [GSI96, Mit92, MO93, TTB88]. booster [NR99a]. Bootstrapping [GSI96]. Bordered [Bai97, Bai99b]. borehole [ZA88]. Born [Kin94]. bound [Wei99a]. boundaries [GKn91, Lan81, Nar99b]. Boundary [AS94a, AS94b, AR93, AC97, BKM95, BT97, Cab94, Cha96a, CW97, EL90, GG96, GM94, GC94, Gol96, Guo99, Gup98, HM94, Had95, HH96b, HH97, JNM92, KR95a, KA95, KBS96, LMS92, Mit97, Mo97, MFF92, OT92, PA94, Raj94, SZ97, SV94, ŠNG92, SS94, Ste94, TM92, VD97, VS96, Wan95a, Wan96a, Waz97b, WJ92, WCG97, YK97a, Zam88, AG85, Aga86, AW99b, AD98, AG99, Als79, AY98, Bai91, BKS84a, BKS84b, BDJR98, Bow88, Bre81, BKM79, CN90, Cab95, Cas89, Cas94, CZ99, CCSK82, Col82, CM90, Dul82, Faz90, Gup90, GK94, Har82, HSZ94, Hau82, HG99, HL88b, HL88a, IPP91, Joh82, KR84, KR87a, KR87d, KB93, KS93b, KH78, KM79, LP81, LRV84, LM94, Mär89, McW82, Mit86, ML90, MR99, NR98, NSE81, Oma86]. boundary [OT93, Pan88, Pen84, QA98, RGL98, Ram98d, Sev99b, SV83, SW89, Soc88, Tha82, UA88, Van95, VR99f, Wan99a, Wan99b, WS87, Waz99b, Wei89, WKTM88, WA99, Zam84]. boundary-conforming [Dul82]. boundary-fitted [CCSK82, Col82, Hau82, Joh82, Tha82]. Boundary-Layer [AS94a, HM94, Had95, Waz97b]. Boundary-Value [SS94, Bai91, Faz90, GK94, KB93, KH78, KM79, ML90, VR99f, WS87, WKTM88]. Bounded [Che96b, JPF95, JkC97, Liu94, Sha87, Sor99, Sos88, Wan81c]. boundedness [FL99]. bounding [SW90a, Wol89]. Bounds [CLC97, Eis98, Jód92, AK80, AG89, JTT80, QA98]. Boussinesq [DH99a, MEZI93, Siv92]. Boussinesq-Type [MEZI93]. box [Den83, Den86, Den88, FpDS99]. box-counting [FpDS99]. brain [Kje99]. branches [KK83b]. Break [Szi98]. Break-Down [Szi98]. breakthrough [Mus97c]. Brinkman [BH93, HB93, HB91]. Brittle [LK95]. Brouwer [YL96]. Brown [BK80]. Brown-Robinson [BK80]. bundle [Hof89]. Buoyancy [SK96, ZW95a, ZW95b, KS93a]. buoyancy-induced [KS93a]. Burgers [DVD95, HM98, RC91]. Burridge [Ado96a]. BVP [Die89]. BVPs [Lop92].

C [Mit92]. C-version [Mit92]. CA [BR99a]. Cable [BD95]. Cahn [Fre90]. calculating [EMEH96]. Calculation [JM93, Lan81, Lan92, YBR97, BTW94, Cau83, Jou82, LCE85, VCM91, Zen99]. Calculations [Wam92, Zas93, Abl82, Mel86]. Calculus [AK96, CS97, SS98a, TGK91, CSY98, KS84, MU90a, MU90b, TTB88]. California [Fin95]. Camassa [Boy97b]. Cancer [DGU96]. canonical [IA89]. Cantor [Bad93]. capacitated [She90]. Capacity [MZ98]. Capital [Chu92, Puu84]. Capture [Yip93]. Capture-Recapture [Yip93]. Capturing [Dav98]. Carbon [RBPK96, Ten90]. Carlo [QC98, RBR94]. Cartesian [CS99]. cartilage [EM79]. Cascades [JM93]. case [Bra86, Got85, HC85, Rin98, SL98, Wan81a]. Cases [War94, MS95a, Wan81c]. catastrophe [Ste84]. catastrophes [MR89]. categorical [AK85]. cathodic [Zam88]. Cauchy [AHKK80, Beh96b, Gol77, Gol88b, Gol89, Ioa83, Sen93, TETH98, Uen89, Uen91]. causality [NH86]. caused [zHhT99]. causes [JB89b]. Cavity [SK96, AAAaZ99b, AaAaZ00b]. Cayley [MSZ96]. CDC [BB83b, XZL97]. CDF [Ahm93]. Cell [BMS92, MMMK96, XZW96, XZL97, And90, KKL99, MH81]. Cells [CFT^{+}93]. Cellular [Ada93a, GMS96, Hog88, Kub94a, NG94, YGH90, Ada97, Ada99, KK90, KK91, Kub94b]. cellular-automaton [KK91]. Center [SST95, Lia98]. centered [KKL99]. centers [PK98]. Cerebellar [CFT^{+}93]. cerebrospinal [STD98]. Certain [AS94a, CS93, KR95b, MS95a, PGHX99, Wol98, WLTS99]. Certainty [Mal93]. Ceteris [Mat93]. chain [Ben89, JA99, KK91, LF94, Wai90, Zag91]. Chandrasekhar [DK96, Sug95, SO95, Tak85]. Change [Gho96, MBG95, WC99]. changes [Mos88a, Tur88]. changing [VB89]. Channel [Ham94b]. Channels [Ham94b, HF95, HFA98, Ram98d]. Chaos [Ebe93, WL97, LG99, TW90]. Chaotic [FG94b, MKT96, Son96, Wer93, MMC98]. Characteristic [Gu99, KP98, AST86]. Characteristics [SAL96, MW98a, QC98]. characterization [FEGL90, Smi84a, Szi99]. characterized [SW91]. charge [VJ88]. charts [JB89a, JB89b]. chasing [AG85]. Chebyshev [BPE99, Boy89, DVD95, nHyG97, HS98, yHyG99, KS93a, Lon87c, PS94a, ZB97]. Chemical [Lar97, Sel95, HEH81, MM87b, NR99a, Pat90, Shi89]. chemicals [Got84]. Chemostat [BR98, SW89]. chemotherapy [NHG83, OTA83]. Chen [Arg94]. Chen-Yamamoto-Type [Arg94]. chiral [TW90, Ten91]. Choice [EFS95, ABES84, Ge89a, GFM88, Gil89, Smi84a]. Cholesky [Zha98d]. Cholesterol [Kub94a]. Chow [Wat81]. Chow-Yorke [Wat81]. Christoffel [ADHE82, KR88]. chromatic [Kor97]. Circuit [PS94a]. circuits [MNK99]. Circulant [Jin97, JS98a, RG98]. Circular [AA94, AA95, Car88a, Car97, CCSK82, HF93, SL99]. circulation [ÖGDB98]. Circumscribed [NP92]. city [Fei87]. claim [vGG79]. claims [BS98]. Class [AK96, BS97, Che95, Eis98, Gar98, GM93, KBS96, KR95b, LLX96, Phi92, Rya96, Sas95a, TFM93, YL96, AG99, AG89, AB95b, Bai99a, Bai99b, CSY98, EC80, GS94a, Ge89b, GW89b, Ham98a, HSZ94, IP89, JN91, KS93b, Pat92, RC99, Red90b, Sen86, Wan99a, Wan99b, Wol98, WLTS99, Zha98c, dOP89]. Classes [DS99, HHT98, LS81]. Classic [HS94, Kel81]. Classical [FC93, RS92, RBR94, Bow88, CC99a, KS93b, Sai99]. classically [UK94]. classification [ÖGDB98, Wu94]. classifier [IFU98]. Climate [MBG95]. close [HKK^{+}79, HKK^{+}80]. Closed [JNM92, JNF95, Lin96, NFJ94, EL77, IP89, WH84]. Closed-Form [JNM92, IP89, WH84]. Cluster [MGL92, Baa88]. clusters [BI98]. cochlear [YF91]. code [Bar91, Jes83, SW79a]. codes [GMGCL99, GBS99, Sha89b, Wat89c]. Coefficient [Bha94, BB96, BG97, HZL97, MS89, MS92b, IPP91, RS86]. Coefficient-parameter [MS89, MS92b]. Coefficients [Ela94, FBSS96, SS94, WJO97, BF83, Boy89, Eno90, Gol88b, Gol89, KB84b, MMP90, RBA92, TT84]. Coexistence [BR98]. cognition [Pat93]. colleague [Cas91]. Collision [Sch92]. Collisions [TPS98]. Collocated [Bal94, Kum95]. Collocation [Bla96, GL94, HZ95a, Hei88, AB89, Bai91, BPE99, FS98, FG81, Ioa83, KS93a, WS87, Zam84, Zha91]. Collocation-Type [HZ95a]. color [KL89]. Combination [CH92, Hoj94, He98, Hof87, Jia99, QC98]. combinations [Hof87]. Combinatorial [GS97]. Combined [EM98, LLY96, LYF97, Lu98]. combustion [Bar91, LS98b]. Comment [Di 98]. Commonality [ITN98a]. communication [EM90, WW99a, WW99b]. communications [HKK^{+}79, HKK^{+}80]. commutation [Mur89b]. Compact [BG93, Ram98a]. Compactly [Wu97]. company [KPJ83]. Comparative [KTW81, KA96b, MH81, HK84, Waz99b]. Comparison [AS90, Beh95, Cic93a, KSZ^{+}95, MBS98, RBR94, Sca89, SV83, SV84, Zha95, AM81, Aft80, CS85, Cas89, HV90, Sch91a, VV80, VS90, Waz98a]. comparisons [HQU88, JK98, Mat86, Stü83]. compartmental [Pat89, Pat91]. compartmentalized [PM89]. Competing [MFGW97]. Competition [FW98, LG97, Zag92, ZK98, Les89, SV85, SSW99, SW89]. Competition-Diffusion [FW98]. Competitive [AM98b, LSP91, SP93]. Complementarity [Lin96, Bai98, Smi84b, Zha99b]. Complementary [Lon87a, Zag92, Moh90]. Complete [ITN98a, Koz95, RK93]. completeness [Kor97]. Complex [EHE93, SY98, Tak98, TSK98, Waz97c, ZB97, DLR85, Eva90a, Eva90b, ER94b, ER94a, HK84, Mei89, MW88b, WSL83, dPL84]. complex-valued [WSL83]. Complexity [Cha92b, Cha93, FS94, NP92, BR88, Cha92c, Cha92d, Cha92e]. Componentwise [SBH96]. composed [Kin91]. Composite [ENAAA95, AAAaZ99a, AaAaZ00a, Alb89, FH98, Tho82a]. compositions [Ada98]. compound [Ram99a, Rol98, vGG79]. compressibility [BTW94]. compressible [HRA98, VWSS94, Yag82]. compromise [GPAB83]. Comput [AaAaZ00b, AaAaZ00a, Cab95, Di 98, JS98a, JNF95, MS92c, MS92b, Mus97b, Sor97a]. Comput. [OT93]. Computation [BJTW98, Con93c, Dég96, DLR85, How94, Ioa92, Kha94, LLL97, LLX96, LUL97, MO93, Sem92, Wan96b, WSL99, Any82, Doa91, GG91b, GB95, HJ92, KS93a, Kni82, LE85, LW90, RR99, Sey81, Shy87, SJK86, Gol77]. Computational [Aga86, Bad93, BS93, CW91, CW92, Cic93a, Cos99, De 94, Ham92, HRA98, HLW94, JR93, JR95, MFGW97, Sel94, Sel95, YR84, CAP88, Col82, Dul82, Ham98b, KKS87, KR81a, KB80, Mus78, NR98, Sev99a, ZG90]. Computationally [Gro93]. Computations [Fre90, HSA97, Joh83, WI80]. Computed [PZ99]. Computer [BG97, DGU96, GMGCL99, HC87, KKPL96, Kor92, PZ94, Raj94, TW89, Zhe93, BR99a, Bay98b, Bay98a, BY99, GMS83, Gre87a, Har82, KT83, TCS92, WI80]. computer-aided [TCS92]. Computer-Assisted [Kor92]. computer-oriented [Gre87a]. Computers [CH92, DS86, Axe86, EM90, Rug93, Zhu94]. Computing [Bil96, CC94, MS87a, MS92c, Sán80, Abu99b, Che96a, Ge89b, JCC98, LB92, SY98, Wan87, Wat79, XZL98]. concave [Van95]. concavity [HS99]. concentrating [WM91]. concept [Kai91, Sch91a, SL98]. concepts [Eir89, HDP85, Mer86, Mus78, WHF89]. conceptual [RB85]. Concise [LSS94]. concordance [HDP85]. Condition [Lin96, Wu97, CM90, Gri89, Gup90, Khu98a, MM85, Shi89, Smi84b, Wei89]. conditioning [AB89, GB90, Mat89, ML90]. Conditions [AR93, Arg97a, Cab94, GG96, Ham94b, ITN98b, JPF95, KBS96, Mat89, SZ97, VD97, Waz97a, WJ92, AD98, Als79, Arg98, Arg99b, Beh89, BKS84a, BKS84b, Cab95, CZ99, FLZ91, Gyö90, Ino99, KL85, Lin94, NC85, Ram90, RGL98, Ram98d, Soc88, SP93, Van95]. Conduction [Ona97, CL98a]. Conductivity [Kub94a, CBYK99, KBÖ99, YS98b]. cones [Du83, EL77, Gyö90]. Conference [Ano89, Ano99a]. Confidence [WLB97]. Confined [CJF93]. Conflict [BDS96, HKLP97, ITN97, RPT97, HF91]. conflicts [Szi89]. Conformal [Hal82, Har82, Ive82, Hof87]. conformal-transfinite [Hof87]. conforming [Dul82, Lee98, Ste82, XL98].

Conjecture [HKLS93]. Conjugate [AH94, FT94, Mou95, GMS83, HV90, Kap88, PZ99, Yse86]. conjugate-gradient [Kap88]. conjunction [ZL98b]. connected [Har82, WN91]. Connecting [BD95, DFM94]. connectivity [CS99]. Conservation [Jim94, Li94]. Conservative [HSA97, FPGV95, VJ88]. conserving [FV91]. considerations [pLR87a]. Consistency [Beh96a, JK98, Var80]. consistency-driven [JK98]. Constant [BNN95, AW87, Gol88b, Gol89, LG99, SW91]. constants [Aft80, BL78, DT87, Gri89]. Constitutes [Gun93]. Constitutive [BNN95, LK95]. Constrained [CUK92, Fab96, KU92, KUX95, Let97, RTJL92, RL96, UK95, Yak92a, Flå92, Ge90a, KM86, Sun98, TLG90]. Constraints [Gzy97, IKU96, Shi93, Vev92, KS81]. Constructing [Arg96a, Arg98, Arg99d, Oma86]. Construction [Boy97a, EGH97, Ham94a, Jin97, JPF95, NLC95, TJ95, Eis87, EGH98, JS98a, KG91, Sas83, WC88, BK99]. Constructions [Sas95a]. Constructive [Con93c, Pen84]. Consumer [SM94, SMY95, Gae87]. Consumption [Pap92]. contact [Abd99, Dav94, Hof89]. Contaminant [LS98a, HY98, HY99b]. Contemporary [MA95]. Context [SAL96]. Context-Free [SAL96]. contingent [BS98]. Continuation [ARM97, KT91, McG98, KK83b, MMP89b, MM91, Mit86, MS87a, MS89, MS92c, MS92b, O’N90, VBH91]. Continued [MB80]. Continued-fraction [MB80]. Continuous [CT96, FS90, Jan95, Pap94, SO95, GH89, MD98, MX90a, PM89, Pat91, Pei99, RC81, VV85, ZK98, Zha99b]. Continuous-Time [SO95, Pat91]. Continuum [HK88]. contour [WC88]. contraction [Han98]. contractions [LS91b, ZA88]. contractive [dOP89]. contribution [CG88b]. Control [Avg94, Bal94, BD95, Cha95b, Che95, Che96b, Chu94, Chu96, Chu97, EM98, Fab96, Hab95, Kum95, LR95, Lei95, Lia97, Liu95, Mil95, MZV96, Pap95, Per96, Pic92, PJ95, RTJL92, RL96, RRB95, RB95, SST95, SMY95, UTL^{+}95, UHL96, VLP95, ZW95a, ZW95b, AM98a, And90, And93, BM89, BD90, BdA89, Bog91, Bra82, CK86, Enr89, Fei85, GL99, HH85, JB89a, JB89b, KS84, Lam98, NA91, ORSS83, Pat93, Sag91, Son93, Spi90, SWL98, Swo91, SP93, TLG90, Udw91, Udw92a, WSL99, YR84]. Controllability [CUK92, Chu92, Sos91, SY93, Chu98]. Controllable [Sha97]. Controlled [TH95]. Controller [vKH95, TFM93]. controllers [WH87]. Controlling [KP96]. Controls [IOAB95, HH85]. Convection [Ezz94, HM94, Had95, Yee93, YS98a]. Convection-Radiation [HM94]. Convergence [AS92, Arg94, Arg97b, Arg99a, BS97, CN97, Cao88, CL99, Cha96a, Cro93, EA94, EGH97, FHG96a, FG81, FT97, HW98, HKLS93, Hua91, JN95, KBS96, Kel81, MLB99, SSV89, UK94, YJ99, Arg99c, BCG91, Cao98, EGH98, Gol87, GB98, GW89b, HJ92, HK89, Joh83, KS78, KM87, KH87, LS81, Lon87b, pLR87a, MM87a, Man86, Par85, Par87, Szi90, Udw92b, XL99, Yan98, Yse86, ZS99]. Convergent [HSG94, HZL97, SS98b, SV92a, Wat79, Wat89a]. Converging [BG93]. conversion [Gol77]. Convex [Cro95, LYF97, PS96, Sor93, Sor94, Zha97c, Sor97a, Wol89, XZL98]. convexity [HS98, HS99, Kha81]. convexized [GQ90]. conveying [ST96]. Convolution [GS97, Jol92, WP97, BW77, KN79, RRA90]. Cooke [Fym79]. Cooling [Abd92]. Cooperation [HF91]. Cooperative [HKKT79]. Coordinate [Bra82, FHG96a, Ham97, Tho82d, CCSK82, Chr82, Col82, Eis82a, Hau82, Joh82, KK82, Kni82, Mas82, McW82, RL82, Tha82, Tha84, Tho82c, War82, Yag82]. coordinates [Bar88, KV84, WZ82]. Core [Zas93].

Corner [HHZ97]. Correction [JNF96, Bar91, JNF95, PGHX99, VR99e, Yan87]. corrections [VR99f]. corrector [Doa91]. Correlated [Kha96, Mil95, Ami99]. Correlation [Bha94]. corrosion [BDJR98]. cortex [Hof89]. Coshoidal [Boy97b]. cosine [EE89, Kel81]. Cost [Dro90, MW93, Pic92]. costs [EM90]. Couette [Uen91, Uen95]. countably [Pal78]. Counterions [Kub97]. counting [FpDS99]. countries [Zem86]. County [Tur88]. Coupled [Ado96b, FH98, Ham97, JV94, RRB95, SL92, UR97, JN91, JF91, Mei94, Pac85, dAMY89]. couples [Rin98]. coupling [Mei89]. Courant [Pap84]. Courant-Isaacson-Rees [Pap84]. Covariance [CC94, LS93b, Sug83, Uen83]. Cover [RBR94]. Coxian [MES92]. Crank [HM99, Jin94]. CRAY [LMM86]. created [Ten91]. Creation [Shi93]. CRI [Sch92]. Criteria [AFM97, FLG92, HH96b, HH96a, HH97, ES93b, FL99, HM99, Mal93, MS99, Pei99, SL83, Sol98, Wan99b]. Criterion [HKL98, KS78, Mat86]. Critical [TJ95, Yam95, HF95]. Crossing [PY94]. crowding [KR88]. Cubic [Beh92, Beh95, BG93, Ste94, Wan94, Beh89, EB99, KB93, Khu98b, Ram90]. Cubical [EGH97, EGH98]. culture [ZK98]. cumulative [Pat89]. current [CXL94, MS94a, Mus84b]. curve [MH81]. Curved [Ham92, Ham97]. Curves [KS92a, ASZ94, Eis87, Ioa88, KR81b, Knu91, MH80]. Curvilinear [Tho82d, ADHE82, HFA98, KK82, Kni82, Ste82, Tha82, Tho82c]. customer [Pat92]. cutout [HF93]. Cutting [PS96]. Cyber [BB83b, Sca89]. Cycle [Do96, MP90, Par85, TT89, XL98, XL99]. Cycles [Alw97, Fri92, Mur91b, ST96]. Cyclic [Ais88, WK95]. Cyclically [MS95b]. Cylinder [AA94, AA95, AA96, ENAAA95, AA99a]. Cylindrical [Yan99, MX90a, MX90b, MS94b, MS95a]. cylindrically [AAAaZ99a, AaAaZ00a].

D [RGL97a, Ada99, CXL94, Cha95a, CCSK82, Hei91, Jin94, Kle82, LYX^{+}93, LS98a, Ram97a, SK96, TT86, TT88, Van95, Zak89, Zha99b]. D-domain [Van95]. D-orientation [Zha99b]. dam [KK84]. Damage [LK95]. Damped [Kar98, Net89, BKS98, HG84, Jac90, JC97, UK94]. Damping [Bar95, Whi89]. Darboux [CZ99]. Darcy [BH93, HB91, HB93]. Data [BP92, GSC95, KUO95, Kha96, Koh92, Vev92, Wan81b, Ben89, BCG91, Da93c, Da93d, Da93b, EBR85, Eno90, HK81, HKKN78, HC85, IPP91, KL99b, MMC98, Mer86, MKJ87, Pat89, SW79b, SHR94, Van83]. data-fitting [MMC98]. Dawson [Let97]. Deadlock [ITN98b]. decaying [CG98]. Decell [KRTW83]. decentralized [HAST98, Lak81]. Deception [ITN97]. Decision [AK85, DGP93, HKLP97, Mal93, PN97b, RPT97, Sar93, Yak92b, YLS93, YH97, DDS85, KK84, MCD89, Rav87]. decomposable [Mal91, MS99]. Decomposing [HT98, KK82]. Decomposition [AR93, ARM97, BS97, BS93, Cai93, LB96, Waz95a, Waz95b, Waz97a, XS95, Yee93, Ado87a, DK96, DLR81, Dos90, Gol99, He96, Hei93b, Khu98b, MLB99, Waz98a, Waz99c, Waz99d, Ye98a, Ye98b]. Decomposition-Type [BS97]. decompositions [CS88]. Defect [Yan87]. Defect-correction [Yan87]. Deficient [SV94, O’N90, VBH91]. Defined [GSI96, Ibr96, Fun91, Ioa88, Kee89]. Defining [PLPW96]. Definite [LLX96, Sha94, EE89, Eva90a, Eva90b, Eva90c, ER94b, ER94a, Hua91, Man88, Udw92b]. Definition [BD94]. deformed [HW98]. Degenerate [CY98, RK87, CC99a, Uen83]. Degraded [HKK^{+}79, HKK^{+}80]. degree [EB99, Kor97, Mur91a]. Delay [FL94, HH96b, HH96a, HH97, JV94, LWLL95, LL96, She96, VB92, ZA96, ZAES96, AW87, BH80, Cha85, Chu97, Chu98, Du83, EL91, FLZ91, GS94b, HG99, Sau83, Wan99a, Wan99b]. Delay-Differential [HH96a, HH97, Chu97]. delay-independent [FLZ91]. Delayed [WZ95]. Delays [AR98b, FW98, Kum95, Lu98, Udw91, Udw92a]. delimitation [van85]. delivery [NHG83]. demography [AR90a, AR90b]. Dendrites [CFT^{+}93]. Denominator [KKH95]. density [Kap80]. dependence [Kap80, Mit86]. Dependent [MSLB92, MS94c, PS94b, Szi98, ZA96, AAES99, AKK99, Tau80, ZH86]. Deployment [BD95]. Derivation [Fis85, Jim94, IA85]. Derivative [GSC95, KPT87, Sem92, KRT83]. Derivatives [Har87, Hus90, KB84b, KT83, Wex87, dOP89]. derived [Beh89, Beh96a, OL80b]. described [Mv98]. describing [PV81, TETH98, TET99]. description [Hir84]. Design [Bay93, Che95, HZ95b, HZ96, Jes83, KS81, MZ98, SS91, BB83a, CS85, JB89b, Sug83, Sug87, WH87]. designing [YL98b]. Destabilization [OMS95]. Detecting [Fri92, JY97, Mur91b]. Detection [Gho96, KK95, SO95, WC99]. Detector [MFGW97]. Determinacy [RK87]. Determinants [SC96, CS98]. Determination [KK83b, BPE99, BJ85, Ioa88, Mor90]. Determine [SP95]. Determining [HSG94, Kru81, XZW96, SR88]. deterministic [AM81, AS81, Den90, GD87, Pat90, VV85]. Develop [KS92a]. developed [Zem86]. Developing [RK93]. development [HB93, HKWdZ83, Mei89, Sor85]. Developments [Gol96]. deviating [KR87c, LF99, Pei99, WA98]. deviation [Cha96b]. devices [Ade94]. diagonal [Hua91]. Diagonalization [LR85]. Diagram [NG94, HT98]. Diamond [Kan97, Kan99]. Dice [SM95]. diced [Lan89]. dichotomy [Bur89, ML90]. Difference [BG94b, Jim94, JNM92, JPF95, JNF96, KTZK96, KBS96, LG94, MEZI93, NFJ94, Phi92, PPS94, RBPK96, Sel94, SA94, SS97, TPL95, Wan95a, Wan96a, Zha95, AW99a, Ahm79, Asa98, BT98, BS98, BH80, Dav94, DSK82, Gar89, GP88, Gyö90, Hos99, HY99a, JF91, JNF95, KBÖ99, KKL99, Lin94, Mor98, MR99, Pan88, RGL97a, Sas88, Sim99b, Ste82, VC98, VR99a, VR99b, VR99c, VR99d, WA98, ZA88, ZL90]. Differences [JY97, Yil96]. Differencing [FHG96a]. Different [KKPL96, CL99, Mur89a]. Differentiability [Ibr96, Arg99b]. Differential [Ado97a, AW87, BKM94, BKM95, BD96, Béc98, CG96, Cha94, CH96, CL98a, Da93a, EN94, ES92, ES93a, ESI95, ESS94, ES94, Ela94, GG96, GGS97, GL92, Gra95, Gro93, GWL94, Gup98, HZL97, HH96a, HH97, Jód92, JL93, JV94, KA95, KTZK96, KW96, KW97, LMS92, LWLL95, LWX92, LLS94, LLS96, MRJ95, NS96, RBA92, RAL95, RK93, SJ96, ŠNG92, SKG97, She96, SIA96, Tho82d, VB92, Wan95a, Wan96a, WA96, Yan97, ZAES96, ZZZ98, Ada89, Ada91, Ada93b, AM81, AS81, Ado81, AR87, ASK98, AR82, AR90a, AR90b, AY98, Bai91, Bai81, Bee87, BL78, Bra78, BH80, CH99, Chu97, CM90, Cor89, Cro80, Da93b, DT87, DPTT98, Du83, ES93b, EL91, FS98, FL99, Gla78, GZ84, Guo99, HEH81, HK81, HG99, HKKN78, Heu90a, Heu90b, HK84, HL88b, HL88a, IM80]. differential [JC97, JN91, JF91, KSB89, KST80, Kha91, LS87, LL89, LS91a, LP89, Liu94, LF99, MD98, Mär89, Mik87, MB88b, Nak82, NJC99, Nie84, Nie85, OL80a, PF99, PGHX99, Ram98b, Ram98c, Ram99c, SL83, Sán80, Sau83, Ser79, Sey81, Sha90a, Sha90b, SV84, SR88, SS92, Sol98, Sor82, SK98b, TET99, TT84, TT90a, TT90b, VV80, Vau80, War82, WSL83, ZA91, Zha98c, Zla89, dAMY89]. differential-algebraic [Bai91, Kha91, Mär89]. differential-delay [BH80]. Differential-Difference [Wan95a, Wan96a]. differentiation [HL82, MM85, Ske89]. Differintegral [AK96]. Diffusion [Ado97d, AR98b, Asa92b, FW98, Gar98, Ram97a, Ram97b, SZ97, Yee93, Zas93, AG89, EM79, GS94a, GMS83, HV90, JV98, KP98, Mei89, Pac85, Ram98a, Ram99b, Sai99, mW98b, ZL98b]. Diffusion-Reaction [Gar98]. diffusive [Ram99d]. Digamma [WLTS99]. digits [IRU83, RUS^{+}84]. dilation [TyG99]. dilemma [NS89]. dilemmas [Mus84b]. Dilution [MMMK96]. Dim [KK95]. Dimension [JL93, Lap92, Sol99b, FpDS99, Tom80]. Dimensional [Asa97b, BG94c, CS92, GL94, nHyG97, HK96, JM93, Kan97, KKH95, KD95b, Mus97a, NSV92, NS97, Zha97c, AKV83, BR88, BH93, Bar91, BF83, Cau83, CY80, CH99, Dua99a, Dul82, Hol86, HV90, yHyG99, Jam83, JBV98, Jou82, KM87, KCS90, MB82, MS83, Mur91b, Mus97b, Mus97c, Nar99a, OU99, OL80b, Pat88, Shi82, SSB82, ST96, Tha82, Tha84, Tho82a, WZ82, Waz98b, WM91, ZA88, Zia89, ZMV99]. dimensions [CG87, GL99, Kap88, KW86]. dioxide [Ten90]. dipole [VJ88]. Dirac [AKV83, Doa91]. Direct [Ona97, SH92, KB85, Knu91, VR99e, bS89]. direct-correction [VR99e]. Directed [PSMM93]. Direction [KK83b]. directions [CK99]. directly [bS89]. Dirichlet [BKS84a, MM94, Van95, Won98]. disaggregation [SK86]. disaster [zHhT99]. discontinuities [Dav94]. Discontinuous [WJO97, BF83, EJNT88, RC99]. discordance [HDP85]. Discrete [CT97, CG83, EFS95, FW98, Fri97, Gol89, Gol93, Gol95, HSA97, HH96b, IA89, Jim94, JMS98, KR84, KR87a, KPP93, Kum95, Kwo93, LAK94, Lia97, Mal93, MDR94, MM91, MB95, PA94, RTJL92, SHW97, Sug95, SY93, SMY95, WP97, AT80, Aga86, AP95, DS87, DS88, GS86, GB98, Got85, He96, HK88, HL82, LS99, MS99, PM89, Pat91, SS91, Sha87, SW91, SL98, TLG90, UA88, VJ86, WH85, YR84, Zen99]. Discrete-Delay [HH96b]. Discrete-Numerical [HSA97]. Discrete-Time [Fri97, Lia97, RTJL92, Sug95, TLG90, YR84]. discrete-valued [SS91]. discretisation [Gri89]. Discretization [LY98, Ona97, Da93c, Da93d, WH87]. Discretizations [Arg97a, Gar89, Ye99]. discretized [BKS83, XL98]. discriminant [TW79]. discussion [Har82]. Disease [Alw97, MAS95, MH81, Wat89b, Yip91]. disequilibrium [Zha89]. disk [Bar88]. disparity [TETH98]. dispersive [Hos99, HY99a, OŞL80]. Displacement [YBR97, KKS87, Kai91]. displacement-rank [Kai91]. Displacements [dL95]. Dissipation [OMS95, Dua99b]. Dissipation-Induced [OMS95]. dissipative [OŞL80, WW98]. distance [Fra89, RG98]. distinguished [Pat92]. distortion [Abu99a]. Distributed [Bil96, Che96b, El 92, HSG94, RAL97, dAMY89, Ada98, BW88, Den90, LF99, Udw91, Udw92a, WA98]. Distributed-multirate [dAMY89]. distributed-parameter [Udw91, Udw92a]. Distribution [BB96, Fin95, Sla93b, BU77, Cra79, Moh99a, Won98]. Distributions [Bur96, JY97, Pat92, Eis87, GBS99, MS93a, MMT99, PM89, Rol98, Sch91b, vGG79]. Divergences [PSMM93]. Diversion [RPT97]. diversity [BU77, GG91a]. Domain [BS93, Cai93, GC94, He96, Jac95, KR95b, LB96, Ye98a, Ye98b, Dos90, Hei93b, MX90a, MX90b, MLB99, SS91, Van95, ZA88, ZL98a]. Domains [Ham92, Ham97, HHZ97, JPF95, Kur99, Mo97, Guo99]. Dominance [YLS93, Hua91, Kee89]. Double [RAL97, RRA90]. Doubling [LAK94, AB87]. Doubly [MS97]. Doubly-Stochastic [MS97]. Down [Szi98]. Drazin [Ji94, Wei98]. Drilling [DG87a]. Driven [Mil95, SK96, JK98]. droplet [SKCW94]. droplet-turbulence [SKCW94]. Drug [LL94a, NHG83]. drug-delivery [NHG83]. dual [Kel81]. Duality [Cha95b, Vál88]. Due [AA95, AA99b, DDS85]. Duffing [WL97]. during [Bar95, Mei89, Tuc96, YBR97]. Dusty [HS96, AH99b, BH90, BH93, HB93]. Dynamic [AA96, Ben92, Bil96, BGY97, CT96, CT97, Hab95, Kum95, LSY98, MPD97, SO90, SY93, SM94, SMY95, Whi93, Abl82, AM98a, Any82, DF96, EB91, Hen86, HKKT79, KL85, LS91b, SGB87, Udw87, UK94, VL83]. Dynamical [Fri97, LR95, Lei95, MKT96, OMS95, RW93, SW79b, SAF^{+}97, Son96, Tuc95, Unb96, BL87, KPJ83, LS99, MR89, NS89, RR98, Sol99b]. Dynamics [Ado97d, BG94a, Chu96, Hal98, HLW94, HSA97, IKP97, LF94, UTL^{+}95, UHL96, UR97, Uen95, ZA96, ZAES96, BD90, Bec85, BK99, CJ88, Cha88b, DS87, DS88, Dua99a, HKK^{+}79, HL91, Kaz99, Lam98, Mos90, Rin98, van85].

Early [JY97]. earning [Ree87]. Eckalbar [PBL92, Zha89]. Ecological [UR97, Den90, Hog88]. ecology [JTT80, JT80]. econometric [MP86, Pal86]. econometrics [Hen86]. Economic [Avg94, Chu96, KTW81, Lin96, SAF^{+}97, Van97, Zhe93, Chu97, Chu98, HKKN78, KP91, TETH98]. Economical [NP92]. Economics [KS92b]. Economy [Pap92, Sos88, Sos91]. Ecosphere [SAF^{+}97]. Ecotoxicology [AB95a]. eddy [CXL94]. Effect [GWL94, Ham92, Kub97, Raj94, Raj97, SEG96, AA99a, BR99b, KK91, MB88a, Tak85, TW89]. Effective [EJNT88, JB89a]. Effectiveness [Ahm93]. Effectivity [HD96]. Effects [CY98, FHG96a, Kub94a, Kum95]. Efficient [AH94, Gro93, How94, JBV98, KM86, Lia97, Ada89, ES90, HM98, RC99, SK98a, WI80]. efficiently [NH86]. EHB [SW79b]. Eigenform [HLW94]. eigensystems [AAES99]. Eigenvalue [CH96, Sha94, VD97, WJ92, AAES99, ABW99, Bai81, Gla78, Mit84, Van95, Var80]. Eigenvalues [Ben92, BG94c, GG96, How94, WA99, Lui99]. Eigenvectors [LLX96]. ejection [YDB79]. Elastic [Che96b, KD95a, KD95b, Le98, LK95, AAAaZ99b, AA94, AAA99, AA99c, AaAaZ00b, Soc88, Whi91]. elasticity [KM87, Lee98]. Elastodynamics [LM95]. electric [JS98b, NA91]. electrical [KBÖ99, Sev99b]. Electrocardiography [IM92]. Electrochemistry [Mac93, LE85]. Electromagnetic [Ado96b, LYX^{+}93]. electronic [MNK99]. Electronics [SP95]. electropotential [MX90a, MX90b]. Element [Cha96a, CW97, Da93a, EN92, El 92, GC94, Gol96, nHyG97, Hei93a, KA95, KWS95, LYX^{+}93, LC97, Mei93, Siv92, VD97, CG87, Cha88a, Cha90a, CXL94, CL98b, Da93b, DZ99, FS90, Gu99, HY98, yHyG99, IBS94, Jac90, KL99a, KP98, Lee98, Lin98, Moh99a, NC85, Ohm98, Sor82, Van95, Yan98, Ye98a, Ye98b, Zam84, Zam88]. Elementary [HZ95b, Sas95b, IP89, LS99, Ste84]. Elements [Alb89, BR99a, HZ95b, HZ96, Gri98, Pig87, XL98, YK97b]. eligible [Pig87]. Elimination [Moh94]. Ellipsoid [NP92]. Elliptic [AZK96, CG96, Cen97, EHE93, HHZ97, KA95, LWX92, Mo97, SKG97, Son93, Sri95, Tho82b, TH94, VD97, WJO97, Ada89, Ada91, Ada93b, AD98, Bar89, Cas94, CG87, GP88, GL99, KB84a, KB84c, KB84b, KB85, KM94, Kle82, LM94, Roa82, RS86, Sor82, Tha82, Tha84, TT86, Tho82a, Van83, Vas91, War87, Whi91, Yan98, Ye98a]. Elliptic-Type [Cen97, Sri95, TH94, AZK96].

Elliptical [HSG94]. ELXSI [SM88]. Embedded [HM94, MS83]. Embedding [LFS95, Lor80]. emerging [GKn91]. emission [BKS98]. Empirical [Hen86]. enclosure [Vol92]. enclosures [SW90b, Wol98]. End [Beh89, Ram90]. Endo [RR93]. endpoint [Boy89]. energy [FV91, FS90, Got85, HJ92, KS81]. Enforcement [LHK97]. engineered [AS99]. Engineering [DGP93, Wat81]. Engquist [Vul90]. Enhancements [CS92, Ada91]. Entire [KR95b]. entrainment [OS80]. Entropy [FT97, SHW97, Din98, Fie84, FT98, Tag99]. Entropy-Convergence [FT97]. Entry [Ham94b, Tuc96]. Enumerating [Pig87]. environment [Pat90, VB89]. Environmental [HKLP97, LHK97, Van97, YH97, BB83a, Got84, HC85, Jam86]. Environments [LG97]. Enzyme [Bay97, Bay98b, BY99]. Epidemic [BGY97, Kim96, MAS95, Yak92b, ALMS90, DML98, KP98, Yip92]. epidemics [Gle88, MO92, YGH90]. Epidemiological [PR93, TW89]. Epidemiology [IKP97]. Epstein [AZK96, Cen97, EHE93, Sri95, TH94]. Equation [AS94a, AS94b, Ado96c, Ado97b, Asa97a, BG97, Boy97b, BG94b, DVD95, EN94, FC93, Gol93, Gro93, HSG94, nHyG97, Jód92, KW96, KW97, LAK94, LM95, MEZI93, NSV92, Ona97, PPS94, RK93, Sen93, SP94, Tak98, UK95, Wan95a, Wan96a, Waz97b, WL97, WCG97, Yee93, Zha96b, Abd99, AR98a, AKV83, AP87, Asa98, AY98, BdA89, Bar88, Ben87, BK99, BKS83, BKS84a, BKS84b, BT98, BH80, Da93c, Da93d, DH99a, DK96, DR88, DZ99, FV91, FG81, Gol88a, GPRV86, GL94, HKKN78, Hol86, HV90, HM98, yHyG99, HL88a, Jac90, JV90, JBV98, Jin94, JPS86, KK83a, KST80, KN90, Kel81, Khu98b, Khu98c, Kin93, KV84, MS94b, OL80a, OL80b, Ram98b, Ram98c, RC91, RC99, RS86, Sai96, Sai99, Sco80, Sco81, Sim99a, Sim99b, Sug87, TET99, VR99b, VR99c, VR99d]. equation [Wan99a, Waz98b, Waz99c, Yan99, Zie84]. Equations [Abd93, Ado96b, Ado97a, Ado97c, AR98b, AK96, AC97, Ayo93, BKM95, BD96, BNN95, Béc98, Bla96, CY98, Cha94, Cha95b, Che97, CZ92, Da93a, De 94, ES92, ES93a, ESI95, ESS94, ES94, Ela94, FHG96b, GG96, GGS97, GM94, GL92, Gra95, Gup98, HZ95a, HZL97, Jac92, JL93, JV94, JN95, KA95, KT91, KTZK96, Kar98, LYC93, LG94, LL95, LY98, LWLL95, LWX92, LLS94, LLS96, LP97, MRJ95, Mor95, MBS98, Mur96, NS96, NS97, NN87, PS94a, Pap95, Phi92, Ram97a, Ram97b, RAL95, SJ96, ŠNG92, SKG97, She96, SS98b, SA94, Siv92, SIA96, SL92, SS97, TPL95, Tho82d, VB92, WK96a, WK96b, WA96, XZW96, Yan97, ZAES96, Zha95, ZZZ98, ZW95a, ZW95b, Ada89, Ada91, Ada93b, AM81, AS81, Ado81, AR87, Ado87a, Ado88, AW87, AW99a, Ahm79]. equations [AHKK80, ASK98, AP87, AR82, AR90a, AR90b, Bai91, Bai99a, Bai81, Bay98a, Bee87, BL78, BW77, Bow78, Bra78, BH80, CZ99, Cha88a, Cha90b, Cha95a, Cha96b, CH99, CC99b, Chu97, CM90, Cor89, Cro80, Da93b, DEV95, DT87, Din98, DPTT98, Du83, ES99, ES93b, EL90, FS98, FL99, Ge90b, Gol77, Gol79, Gol87, Gol88b, Gol89, GB90, Gol99, Got85, GS94b, GZ84, Gu99, Guo99, Gyö90, HG99, Heu90a, Heu90b, HK84, HL88b, Ioa83, IP89, IA89, Jam83, Jes83, JV98, JN91, JF91, KB84c, KB84b, KB85, Kad86, KKS87, KSB89, KN79, Kha91, KM79, LV87, LS87, LS91a, Lam86, Lin94, Liu94, Lor80, MMP89a, Mär89, Mer86, Mik87, Mil87a, MB88a, MB88b, Mit90b, Nak82, NJC99, NLC95, Nie84, Nie85, OM99, PV81, Pac85, Pap84, PF99]. equations [Pei99, PGHX99, RBA92, RGL97a, Ram98a, Ram99b, Ram99c, RRA90, RS86, SL83, Sán80, Sas83, Sas85, Sas88, Sau83, Sav93, Sca89, SK86, SR90, Ser79, SU90a, SU90b, Sey81, Sha89a, Sha90a, Sha90b, SV84, SR88, Shi82, SS92, Soc88, Sor82, SK98b, TETH98, TT84, TT90a, TT90b, VV80, Vau80, VS90, VR99a, VR99b, VR99c, VR99d, VR99e, VR99f, mW98b, Wan99b, WH84, WA98, Ye98b, Ye99, ZA91, Zam84, ZH86, Zha98c, ZL98b, Zia89, Zla89, ZMV99, ZL90]. equicorrelated [KB98]. Equicorrelation [BB96]. Equidistant [Abl82]. Equidistribution [And87]. Equilibria [RK87, Har86, VCM91]. Equilibrium [AD98, NP92, MM87b, Smi84b, Ten90, Wei84]. equipotential [Ioa88]. equivalence [AHKK80, BK80, GP88, Zha97b]. Equivalent [LS93b, Gol77, VR99b, VR99c]. equivalents [GD87]. ER-Materials [LR95]. Erdélyi [ASK98]. Erdélyi-Kober [ASK98]. Errata [MS92c, MS92b, Mus83a, Mus84a]. Erratum [AaAaZ00b, AaAaZ00a, Cab95, JNF95, Mus97b, OT93, Sor97a]. Error [CG88a, CW92, CLC97, CB94b, GM94, GL94, HZ95a, HG84, HLW94, HS94, Jac90, Jac92, Jód92, Let80, Mas82, Mor94, SH92, Tuc96, AC94, Ber89, Da93b, Enr89, Han92, Hua99, LCE85, MS92a, Mor90, QA98, Ske89]. error-control [Enr89]. Error-Free [Mor94, SH92, Mor90]. Errors [BBB97, Kha96]. Estimate [CLC97, Sas95b, Han92, Pet88]. Estimates [Cai93, GM94, HLW94, Jac92, MKT96, AC94, CG88a, Da93b, Hua99, Jac90, KS80, KM87, Let80, MM87a, Pat87, Pat89]. Estimating [AAAN98, Cra79, Fin95, HS94, RBR94, AT85, GFM88, Gil89, GG91a, MH80, Sch91b]. Estimation [Ahm93, Ami97, Ami99, AB94, Bay93, Cha95b, Cic92, Cic93a, CH93, Cic93b, Eis98, FBSS96, GSC95, HD96, KA96b, MFGW97, Sla93a, SU95, TA93, Whi88, Whi89, Whi91, Whi93, WJO97, Whi97, Zia89, BB83a, BY99, BH80, CSR88, CyGK99, CG98, Faz91, FpDS99, HW85, IPP91, MW88b, NC85, RB85, Sag91, Ske89, Tes91c, TT90a, TT90b, Uen83, YW85]. Estimator [Jac93]. Estimators [ADS96, FS94, HW80, Sug83]. estuarine [Joh82]. etch [QC98]. Euclidean [Pes91, RG98]. Euler [Boy95, CZ99, CSZ98, Jam83, Jes83, Mer86, SJK86, VR99a, VR99b, VR99c, VR99d]. European [CY99, Váz98]. eutrophic [GGR91]. eutrophication [Bec85, Gar90, Mos88b]. evaluate [BR99b]. evaluated [KT83]. Evaluating [FHG96b, GC94, ER94a, ZS98]. Evaluation [BG94a, CS96, EE89, Eva90c, GM93, Gol96, Hus90, Kha92, KM79, LMM86, Sar93, SC96, Zhe93, CS98, Doa90, Eva90a, Eva90b, ER94b, GMGCL99, HK84, KRT83, Wex87]. evaluations [CRS99]. Evidential [KK95]. Evolution [Cha95b, Ebe93, KS92a, Pap95, BKS98, Den90, HAST98, Kje99, SU90a, SU90b, YL98b]. evolutionarily [CV89]. evolutionary [ES99, IFU98, VB89]. Evolving [JL97, ABES84]. Exact [EHE93, Jan95, JN91, JPF95, Lan92, LL94b, RR99, RTJL92, WP97, Yak92a, Fei85, Gar89, Ham98a, IA85, KT83]. Exactly [LMSS97]. example [Ber89, Chu98]. Exception [CJZ94]. Exceptional [Zha97c]. excitation [EB99]. excited [Gre87b]. Exclusion [XZW96, XZL97]. exhibiting [NR98]. Existence [AW99b, CY98, CC99a, GS94a, Gar98, IST93, JPF95, JS96, Kim96, Koz91, KPP93, Lin96, Moo81, Nie84, Pap92, Sas88, Sor96, Sor99, VV80, Di 98, Gyö90, Har86, KP91, MR99, Pen84]. Exoatmospheric [CO96]. Expansion [GL94, HZ95a, MFS93, PS94a, IST93, RMR86, Zha91]. expansions [MB80]. Expectations [SY93]. Expected [pLR87a, pLR87b]. experience [Jou82]. experiences [Stü83]. experimental [Ben89, EBR85, Eno90, HGS88, pLR87a]. experimentally [Abu99b]. Experiments [Jac95, Yip93, HC87, OL80b, SS91, SM88]. Explanatory [Ami97]. Explicit [AR87, Ado97a, ASK98, KR95a, Sim92, Sim99a, TS98, EM90, HH85, Hos99, PF99]. Explosion [CK95]. Exponent [WL97]. Exponential [Ami97, ADM91, CMM96, Gho96, Kum92a, JMS98, MMP89a, MS93a]. exponential-polynomial [MS93a]. Exponentially [Sel94, Ram99d, Sim99a]. exponentials [CG98, MMP90]. exponents [Kaz99]. Expression [Ji94]. Extended [CB94b, CB94a, Knu91]. extension [BK80, Ge89a, GW89b, Pat90, Smi86]. Extensions [GB95]. External [Raj97, Ste82, Van95]. Extinction [AM98b]. Extraction [Bha94]. Extrapolation [BG94c, HHZ97, KKPL96, IST93, KT89]. Extremal [CN90, Oma86]. Extreme [Hir85, Hir84, HC87, Smi84a]. extreme-value [HC87]. Extremum [Sor94, Sor97a]. Eye [EM98].

facility [She90]. factor [GG91a]. factorable [Ge89c]. Factored [Vos89]. factorization [Ram98a, Ram99b, Zha98d]. factorizations [Axe86, Kap88]. Factors [Hoh97, SS91, TCS92]. Factory [CDW91]. facultative [Zag91]. Fail [SH92]. Fail-Proof [SH92]. Falkner [Asa97a, Asa98]. Families [Gho96, Zha97c, MS93a, Mus97c]. Family [Jim94, MW96, Mus97a, BT98, HZ98, KR81b, Mus97b]. Farms [HYL97]. FAS [RS86]. Fast [Dul82, Jay96, Jin95, KB84a, Mor90, PJG^{+}97, Tuc95, ZL98c, BPE99, Bar89, HKWdZ83, KB85, Mur91b]. faster [Arg98]. fatal [Wat89b, Yip91]. fate [Pat90]. Feasibility [Cro95]. feasible [CK99]. Feedback [TFM93, BL87, HH85, WSL99]. Feedforward [MZ98, WSL99]. Fériet [MS95a]. Fermi [Doa91, Waz99c]. Field [HM94, SEG96, Mur89a, Mur89b, Mus80, VJ88, VWSS94]. Fields [LYX^{+}93, SK95, Bar91, SW91]. Filariasis [TT89]. filled [Ge90a, GQ90]. Film [KH96]. films [CG88b]. Filter [CB94b, CO96, Cic99, KK95, Sug95, CB94a]. Filtering [SU95, Cha89a, Heu90a, Heu90b, DH99a]. Filters [KKH95]. Filtration [SK95]. fin [Hof87]. fin-afterbody [Hof87]. Final [PY94]. finance [Sha90a, Sha90b]. find [Bar91]. Finding [Cha94, Die96, Ge90b, LLS94, LLS96, Mil87a, Mil88a, Sav93, XZL97, HZ98, KST80, PP93, ZS99, ZG90]. Finite [Car88b, Cha88a, Cha96a, Da93a, EN92, El 92, FG94b, FG95, nHyG97, Hei93a, HK96, HZ95b, HZ96, Jim94, KBS96, LYX^{+}93, LC97, Lin94, LS98a, Mei93, MS97, Moh99a, MEZI93, Pic92, RBPK96, Sel94, Siv92, VD97, Yil96, ZA88, Zha97c, Asa98, BT98, BS98, CG87, Cha90a, CXL94, Da93b, DZ99, DSK82, EL91, FS90, Gar89, Gno82, Gol87, Gri98, Gu99, Hof88, HY98, Hos99, HY99a, yHyG99, IBS94, Jac90, KRTW83, KL99a, KP98, KBÖ99, KKL99, Lee98, Lin98, NC85, Ohm98, RGL97a, RC91, RR99, Sim99b, Ste82, Van95, Wan87, XL98, Yan98, Ye98a, Ye98b, YK97b, Zam84]. Finite-Difference [MEZI93, ZA88, Asa98, BT98, Gar89]. Finite-Dimensional [HK96, Zha97c]. Finite-Element [Siv92, Jac90, NC85]. Finite-Horizon [Pic92]. finite-span [Hof88]. finite-volume [Gno82]. fins [Hof88]. First [Abd93, Ayo93, BKM94, BKM95, Cab94, CN97, VS96, YBR97, And93, Cab95, CG87, Gol79, GL99, HG99, LS87, MR99, NSE81, Nie84, Nie85, SV83, Yan98, YL98a]. First-Order [Ayo93, Cab94, VS96, And93, Cab95, GL99, Yan98, YL98a]. Fitted [Sel94, CCSK82, Col82, Hau82, Joh82, McW82, Ram99d, Sim99a, Tha82]. Fitting [BG94c, HK81, HKKN78, BCG91, MMC98]. FitzHugh [Jac92]. Five [Zha96b]. Five-Point [Zha96b]. Fix [Wu97]. Fixed [EL77, Gun94, Wan96b, YL96, BR99b, HF93, Pes91, Wat79]. fixed-effect [BR99b]. Fixed-Point [YL96]. flares [zHhT99]. Flat [HM94, Mit97]. Flatness [JkC97]. Fleet [MFS93]. Flexible [LFS95]. flexural [Whi88]. flip [KR91]. flood [zHhT99, HW85]. Flow [Ezz94, Had95, Ham92, Ham94b, HS96, Ham97, KW96, KW97, Mei93, NG94, Raj94, RCG94, Sel95, Uen95, WZ98, AH99a, AH99b, BH93, CCSK82, Faz91, FH98, HB93, HF95, HFA98, Han92, Jam83, Joh83, Jou82, Khu98a, KL99b, Mei94, MS94a, Mel86, Mer86, Pat88, Pat89, Shy87, Ste82, Uen91, VWSS94, WN91, Wei84, Yag82, Zhu94]. Flows [BJTW98, Gre97, PN97a, Raj92, Raj97, TSK98, ADHE82, BTW94, Cas81, Cau83, Ham98a, HRA98, KS93a, Kle82, Kni82, MS83, Ohm98, Puu84, Ram98d, SJK86, ZA88]. Fluctuating [Raj97]. Fluctuon [Con93c]. Fluid [Ham97, KH96, KD95a, KD95b, KW96, Kub97, PN97a, RCG94, Waz97b, WZ98, ZW95a, ZW95b, AH99b, BH93, HB93, Han92, HRA98, Ohm98, STD98, ST96]. Fluids [Had95, BS86, Cas81]. fluorescence [HRA98]. Flux [Dav94, Cos99]. FM [Abu99a]. Focal [NS96]. Focused [Lan92]. Fog [BD94]. Fokker [Zie84, ZMV99]. Follower [OMS95]. Food [SPD93, JA99, LF94, Zag91]. food-chain [LF94]. Force [Con93c]. Forced [Ela94, Wan99a]. Forces [Che96b, OMS95, ZS98]. forcing [Dua99b]. forecast [zHhT99, ZPY99]. Forecasting [CT96, CT97, AT85, EB86, HW85, YW85]. forecasts [AK85]. Forest [Cas83, KUO95, TDK94]. Foreword [Cas91]. Form [FH95, JNM92, LY98, NFJ94, Gun92, IP89, JNF95, SW85, WH84, Ye99]. Formal [FHG96b, Bow88, Pat93]. Formalism [Kon92]. Formation [Gre97, BKM79, Mei89, Mit84]. Forms [Die96, Gun93, BR88, Mit92]. Formula [GWL94, HS94, CSZ98, DT87, Doa91, EJNT88, Fym79, Gre87b, Lan89, Zen99, Zla89]. formulae [IA85, IP89, vGG79].

Formulas [CWL92, CS93, GW89a, Ske89]. Formulation [LS93b, RW93, XZL97, BTW94, Cha90a, Cha90b, MS94a]. Formulations [KSZ^{+}95, Lam98]. FORTRAN [Ada89]. Forward [Lan92, Rav87, VR99a, VR99b, VR99c, VR99d]. Forward-Scattering [Lan92]. Foundations [Wer93]. Founding [Mat93]. four [EE89, Eva90c, JV90]. Fourier [Abu99b, Car88b, Gzy95, Gzy97, Kan99, KWS95, KL89, Lon87c, RCG94, RRA90, RAL95, RAL97, Zen99, dFG94]. Fourth [Wan95a, Gup90, GK94]. Fourth-Order [Wan95a, Gup90, GK94]. Fractal [Yan94, FpDS99, Mil88b]. fraction [MB80]. Fractional [AK96, CS97, ES92, ES93a, ESI95, ES99, Nar99a, SS98a, ASK98, CSY98, ESR99, KVT98]. Fractional-order [ES99]. Framework [PMMA97]. Fredholm [GGS97, Gol77, GB90, KN90, Lor80, NN87, RRA90, Sen93]. Free [BJTW98, CS96, Ezz94, Mat93, Mor94, Raj94, SH92, SAL96, BTW94, Cas81, Kor97, Mor90, Sai96]. free-surface [Cas81]. freedom [EB99, Mur91a]. freezing [QA98]. Frequency [Fin95, Jac95, Reg96, SEG96, SK95, BKS98, SS91, Wan81a, Wan81c, Wan81b]. frequency-domain [SS91]. frictions [Wei84]. Friedrichs [Tom80]. friend [Cas91]. Frobenius [DDL93, Din98, DZ99, MS97]. front [Mey78]. front-tracking [Mey78]. frontier [DG87a, LSY98]. fueling [CK90]. Fujitsu [LMM86]. full [Hei88]. fully [Pat88]. Function [Ayo93, BG93, BG94c, JS96, LY98, LP97, Mac93, Pap96, PJ95, RTJL92, Sor93, Sor94, ZB97, ZSL98, Car90, CSZ98, DS99, Ge90a, GT91, HH99, Jer88, JMS98, KT83, KCS90, LE85, LW90, Li98, Mal91, Mil88a, RC91, Sor97a, Uen89, Wan81a, Wan81c, Wan81b, Ye99, ZG90]. Functional [Cha95b, ES94, GL92, Gra95, KTZK96, LYC93, LWLL95, Pic92, TPL95, Yan97, DT87, GT91, HKL98, OL80a, VCM91]. functional-differential [DT87]. Functions [AAHM93, Boy97a, BP92, CS95, CS97, CMM96, Ibr96, Ioa92, KR95b, Kha92, Koł94, Koł95, MDR94, MZ98, MW96, MW93, Mus97a, Ona97, Sas95a, Sri94, Wu97, XZL97, AAES99, Aft80, Beh96a, Boy89, CSY98, DH99b, DS99, FS98, Ge89b, Ge89c, GQ90, Hol89, HL82, IM80, IA85, Ioa88, KKS87, Kin91, Kur99, LMV91, MS99, MS95a, Mit90a, Mv98, Mus97b, PVM88, RS88, RC99, Sco80, Sco81, SW90a, Son93, WLTS99, ZM91, Zha99b]. Further [Car97, MS95a, Mus78, Sri98]. fusion [CK90]. Fuzzy [BBD91, BD94, zHhT99, KK95, IFU98, ÖGDB98, WC99].

Galerkin [Asa92a, Asa97b, Bow78, Gol88b, GZ84, HY99b, HM99, HY99c, Jac92, MSLB92, PJG^{+}97, mW98b, ZM91]. gamble [DH99b]. Game [NS89]. Game-dynamical [NS89]. Gas [HS96, Uen95, Abl82, AH99a, BH90, HFA98]. gas-particulate [AH99a, HFA98]. gasdynamic [LS98b]. Gases [MPS97]. Gauss [GMGCL99, KS78, KL99b, KL89, Zha96b]. Gaussian [CMM96, Lan92, Let80, MS92a, MMT99, McG79, Mil98]. Gel [Kub97]. Gel-Fluid [Kub97]. Gel′fand [AR98a, Kin93]. Gelfand [McG98]. General [AK96, EC80, ES90, LLY96, MW96, NFJ94, RK93, SL83, SA96, TA93, Tho82c, XZL97, Alb89, Als79, AB95b, Fah88, JNF95, Kap80, Mae89, MM87a, MS87b, Sun98, Szi89, Wei84, WKTM88]. generalised [Rob82]. Generalization [Boy95, CK99, Fym79]. Generalizations [AZK96, ZED91, SB82]. Generalized [Arg94, Arg96b, Arg97a, BP92, Cen97, DT87, EHE93, Gol93, Gup98, HW80, HZL97, JS96, MS99, PP93, Sri94, Sri95, TH94, VS96, Won98, XS95, AA99b, Arg90, Arg99e, Che96a, DLR81, DS99, FG81, Gol88a, He96, HZ98, KRTW83, Mal91, MB80, MD81, Sen90, Smi84a, Sta99, TyG99, YJ99, Zha98d]. Generated [Arg97b, How94, KTZK96, ADHE82, Arg99c, KVT98, RUS^{+}84, WH84]. Generating [BI98, CS95, Sri94, SB82]. Generation [CS92, Col82, GSK93, Knu93, SW92a, SW91, SR97, Tha82, CSR88, Dul82, Eis82a, Eis82b, GT82, Hal82, Ham94a, Hof88, Ive82, Jou82, KM94, Kle82, Kni82, Knu91, Nak82, Roa82, RS86, Rob82, Shi82, SYS94, SSB82, Smi82, Son93, Sor82, Spi90, Tha84, Tho82a, Tho82b, Tho82d, War82, WZ82, War87, YS94]. generator [And90]. generators [Ada99, And87]. Genetic [ARM94, CLC97, JL97, ONOT98, RG98]. Geometric [AB95a, CWL92, MES92, Sas95b, Zak89, AB87, HK81, Zha91]. geometric-mean [HK81]. geometrical [Eis87]. geometries [AH99b, Mus80, Tha82]. Geometry [Mit97, Mil88b]. Georgia [Tur88]. geostrophic [Dua99a, Kaz99]. Gini [BU77]. Ginzberg [Tak98]. Ginzburg [Kwo93]. give [Yse86]. given [DG87b, Mil87a, Sug83]. Global [CY98, CUK92, Chu96, GLL98, KT89, KB89, Le98, LS91b, LG99, MBG95, PPS94, Sau83, SW94, Ske89, Sla93a, Sor93, Sos88, Wol96, Yam95, ZAES96, Cas89, Ge89c, Ge90a, GQ90, MBS80, Moh90, Mv98, Par89]. Globally [SS98b, Wat89a, GQ90, HK88, Wat79]. GMCR [HKLP97]. GMRES [Kwa97]. Gompertz [ADS96]. Gordon [Ado97b, BT98, FV91, GPRV86, JV90, WCG97]. Goursat [Waz93, Waz95a]. governed [PM89]. Gradient [AH94, BNN95, Jac93, Mou95, HV90, Kap88, MKJ87, Pen84, PZ99, WSL99, Yse86]. gradients [GMS83]. Gradostat [Zag92]. Graduating [BP92]. Grain [Gre97]. Gram [TM92]. Gramicidin [Kub94a]. Grande [BDS96]. granular [Ahm99]. Granule [CFT^{+}93]. Graphing [Mit90a]. graphs [BR99a, CS99, Kor97]. gravitational [BCZ88]. Gravity [TJ95, Ahm99]. Gray [GBS99]. grazing [Fis85]. Greece [MBG95]. Green [Ayo93, MS94b, Sco80, Sco81]. Greenberg [Mad90]. Grid [And90, CS92, Knu93, LS98a, Sor82, SR97, TSK98, Any82, CSR88, Eis82b, Gno82, GT82, Hal82, Ham94a, Hof88, Ive82, Joh83, KM94, Kle82, Knu91, KCS90, MB82, Nak82, Roa82, RS86, Shi82, SYS94, SSB82, Shy87, Smi82, Son93, Spi90, Tho82b, Tho82d, War87, YS94, Ye99, Zha98b]. gridding [DSK82]. Grids [SW92a, AST86, Ada93b, ADHE82, AR82, And87, Dul82, Hof87, Ste82, Tho82a]. Gronwall [ZED91]. Gronwall-Bellman-Reid [ZED91]. ground [AKK99]. ground-level [AKK99]. Groundwater [LS98a, YCZ98, Bow88, HY98, HY99b, HM99, HY99c, HY99a, NC85]. Group [KR95a, MES92, JK98, SR88, Wei99b, YB99]. Groups [Bil96, Pig87, WW99a]. Growall [EORA99]. Growall-Bellman [EORA99]. Growth [Avg94, Chu92, Chu94, Chu96, KPP93, Pap94, PN97a, ZA96, JA99, Kap80, KPJ83, KP91, Puu84, TET99, Ten91, ZSL98]. Grünwald [He98].

Haar [RJY80]. habitats [Fah88]. Hadamard [RJY80]. Hadamard-Haar [RJY80]. Half [Reg94, Reg96, dL95, AA99c, Abu99a]. Half-Space [dL95, AA99c]. Half-Wave [Reg94, Reg96, Abu99a]. Halley [CAQ93, HS99]. Hamburger [FT97]. Hamilton [YST97]. Hamiltonian [BT97, Fri92]. Hammerstein [HZ95a, HL88b, HL88a]. Hankel [Kha92, KA96a]. Hansen [Mad90]. Hansen-Greenberg [Mad90]. hard [GS86]. Harmonic [Abu97, Jac93, Wan81a, Wan81c, Wan81b, EB99, Ioa88]. Hartley [Jay96]. harvesting [Cas83]. Hausdorff [Tag99]. Having [Sas95a, vGG79]. Hawaii [RPT97]. hazard [Got84]. Head [EM98, HKL98]. health [BBD91]. health-risk [BBD91]. Heat [AA95, HM94, Ona97, SKB95, AA99b, CL98a, DSK82, Jin94, Lan81]. Heat-Conduction [Ona97]. Heavy [TPS98, AKK99]. Heavy-Ion [TPS98]. helium [Gre87b]. Helmholtz [JBV98, JPS86, OL80b]. hemisphere [MS94b]. Hemopoietic [TH96]. Hepatic [Bha94]. Hereditary [Chu96, HHT98, LS95, CY99]. Heredity [FC93]. Hermite [Doa91, LMV91]. Hermite-type [LMV91]. Heteroclinic [ST96]. Heterogeneous [LZ96]. heuristic [AT85]. HIC [HKL98]. Hierarchical [Mal93, PMMA97, Yse86, DHK90, TyG99]. hierarchy [Har87, SV85]. High [DDL93, DS86, LF99, SP95, Zen99, Hol89, TS98, VWSS94, ZS99]. High-accuracy [Zen99]. Higher [Abd93, GG96, JNM92, JNF96, LW94, NFJ94, ŠNG92, AW99a, IST93, JNF95, KT83, Mei89, Whi89, WA98, bS89]. Higher-Order [GG96, JNM92, KT83, Whi89]. Highly [MEZI93]. Hilbert [Cha95b, Cro93, IPP91, Mor95]. hill [Kru81]. Hilliard [Fre90]. History [FHG96a]. Hitachi [LMM86]. HIV [HD96, LL94a]. Hodgkin [Lam86]. Holding [PBL92]. Hollow [AA94, AA99a, ZL98a]. Holm [Boy97b]. holon [HAST98]. Homfly [GMGCL99]. Homogeneous [ENAAA95, VB92, Wam92, AAAaZ99a, AAA99, AaAaZ00a, IPP91, Kur99, MS87b, NSE81]. Homotopy [KT91, LLY96, LYF97, LLS94, LLS96, YL96, Mor86a, MS87a, MS87b, MS92c, WS87, Wat89a, ZH93]. Hooke [HLW94]. Hopf [DVD95, GLL98, HEH81, LLL97, MR89, NLC95]. Hopf-bifurcation [GLL98]. Horizon [Avg94, Pap94, Pic92, KP91, Rav87].

Horizontal [Had95]. Householder [Cic92]. Hubbell [AZK96, Cen97, EHE93, Sri95, TH94]. Hull [BJTW98]. hulls [XZL98]. Human [EM98, TET99]. Huxley [Lam86]. Hybrid [Bai97, RBPK96, Swo91, Bai99b, ONOT98, Uen89]. hydraulics [Bow88]. hydrocode [KH87]. hydrodynamic [Pap84]. hydrodynamical [HK88, HK89]. hydrodynamics [BCZ88, Joh82]. hydrogen [CK90]. hydrologic [BW88, RB85, Sor85]. hydrological [YW85]. hydrology [SU90a, SU90b]. Hydropower [MPD97]. Hyper [Mus97a, Mus97b, TyG99]. hyper-dilation [TyG99]. Hyperbolic [Ayo93, BKM95, Car97, CS92, FBSS96, KTZK96, LYC93, Car88a, Li94, Pei99, Sev99a, Tom80, Wan99a, Wan99b]. Hypercube [KN89]. hypergames [WHF89]. Hypergeometric [CS95, MSZ96, Yip93, CRS99, DS99]. hypernumber [Mus80]. Hypernumbers [Mus78, Mus80, Mus83a, Mus83b, Mus84a, Mus94]. Hyperquadrilaterals [Yeh97]. hyperspheres [Lan89].

ICAP [HGS88]. Idealized [Any82]. ideas [KP85, WK98]. Identification [Eis98, GSK93, HHT98, RAL95, SJ96, Sci97, Unb96, WW98, XS95, Bec85, CBYK99, KKU91, KS80, MMP89b, Sag91, YW85]. identifier [YB99]. Identify [PMMA97]. Identities [GS97, MSZ96, Yür99]. II [MS92b, Wan99b, Arg96b, Arg99e, BH93, BKS84b, Cha92c, CB94b, Cos99, Da93d, ES93b, Eno90, Eva90b, EGH98, GB95, Gol77, Gol87, HF95, HKK^{+}79, KD95b, MX90b, MK93, Mus78, Sco81, SW79a, VR99b, Wan81c, Zha98a]. III [Cha92d, ER94b, HKK^{+}80, VR99c, Wan81b]. ill [DH99a, Gol79, Kee89]. ill-defined [Kee89]. ill-posed [DH99a, Gol79]. image [Cha89a]. Images [KKH95]. imaging [HRA98]. Imbedding [Ayo93, NS97, Als79, Bre81, Hil80, KR84, KR87a, KH78, NSE81, QHN78, Sag91, Sug83, WKTM88]. Imbricate [Boy97a]. Impact [MBG95, Fra89]. Implantation [Ada93a]. Implementation [BS93, FpDS99, GXQ99, Har82, HGS88, Jes83, Mae89]. Implementations [SH92, KN89]. Implementing [Bre91, KMR^{+}91]. Implicit [Abd92, JNF96, NFJ94, Ram98a, AT82, Bar89, DPTT98, HL91, JF91, JNF95, Mit90a, Mit90b, Pat88, Ram98b, VR99e, Zho89]. Importance [YLS93]. Impossibility [ITN97]. Imprecisely [MDR94]. impressing [Abd99]. Improve [HYL97]. Improved [MF99, QA98, SW85]. Improvement [NR99b]. improvements [Mad90, Sha89b, Vul90]. Impulse [GWL94, MB88a]. Impulses [KTZK96, She96, KB89]. Impulsive [BKM94, BKM95, CY98, LMS92, Liu95, EL90, FL99, Guo99, HG99, LL89, LP89, RSS89, SS92]. Impulsively [Kar98]. incidence [TT89]. Inclusion [KD95a, KD95b, MW88a]. Inclusions [CG96]. Income [Chu94]. Incomplete [GSK93, Axe86, Kap88, Lak81, MS95a, SO90]. Incompleteness [Cha92a, Gun94]. Incompressible [Waz97b, BH90, BS86]. Inconsistent [GSI96]. incorporating [ONOT98]. incorporation [NR99b]. Increase [JL93]. Increasing [Lon87b, Yil96]. Incremental [GP97]. indefinite [Cao88, Man86, YK97b]. Independence [Arg97a, CJZ94, MN79]. Independent [Jol92, SA94, AP95, FLZ91]. Indeterminate [Har95]. Index [Ano95, Ano96, Ano97, Ano99b, Wei98, BU77, Kor97, Var80]. Indexed [MS95b]. Indices [HS99]. Indirect [BPZ96]. Induced [MM94, OMS95, KS93a, Mas82]. Industry [SPD93]. Inelastic [BNN95]. Inequalities [BKM94, KTZK96, PZ97, Zha97c, AW87, AT80, AP95, EORA99, GT91, LF99, PK98, ZED91, ZZ98]. Inequality [IKU96, YST97]. inertias [Udw92a]. Inexact [Arg97a, Arg99a]. Infection [HD96]. Inference [HT90a, HT90b, Yip91, Yip93, KB98, MP86, Yip92]. Infinite [Avg94, BG94c, ENAAA95, HM94, LWLL95, Pap94, SS98a, Sas95b, AAA99, Abd92, CCSK82, Doa90, Gla78, GW89b, KR84, KP91, LS81, RC91, Waz98b, WLTS99]. Infinite-Dimensional [BG94c]. infinitely [Pal78]. Infinity [Sim92, Mor86b]. Influence [Ahm99, BBB97, BR99b, DML98]. influenced [Cas81]. Information [Cha92a, Ebe93, GSK93, IL95, MGL92, Par95, Van98, GFM88, Gil89, HW80, Lak81, SO90]. Information-Theoretic [Cha92a]. inhibiting [JA99]. Inhomogeneous [AA96, Asa92b, Sai99]. Initial [BBB97, CN97, ENS96, JNM92, MKT96, RGL97b, Sim92, Sim94, WCG97, AA99a, AAA99, Bow88, Da93c, Da93d, Da93b, Enr89, GZ84, Gyö90, Kad86, KK83a, KT89, Lin94, NJC99, SV92a, SV92b, SP93, TS98, Uen83]. Initial-Boundary [WCG97, Bow88]. Initial-Value [RGL97b, Sim92, Sim94, TS98]. Injury [HKL98]. Input [AA96, Bay93, ABT98, Moh99b]. input-output [ABT98]. inputs [GW89a, KS80, Pat87, VV85]. Inscribed [NP92]. Instabilities [HSA97, TW90]. Instantaneous [AA95]. Insulin [Bha94]. Integer [Mor94, GH89, Zhu98]. integrability [Mur89b]. integrable [BK99]. Integral [Abd99, AG99, AZK96, Bla96, Cen97, De 94, EHE93, GM93, Gol95, GL94, LM95, Let97, NN87, Sen93, Sri95, TH94, WR96, WK96a, WK96b, AHKK80, ASK98, Bee87, Ben87, BL78, BW77, DEV95, EORA99, ES99, Fym79, Gol77, Gol87, Gol88b, Gol89, GB90, Gup90, Ioa83, IP89, Kad86, KK83a, KKS87, KVT98, KN79, Khu98c, Lor80, OM99, RRA90, SR90, Sug87, TTB88, ZL98c].

Integral-based [AG99]. integral-equation [Sug87]. integral-type [Gup90]. Integrals [Beh96b, BE85, CSZ98, GC94, Koz95, Waz97c, Yeh97, Doa90, Doa91, EE89, Eva90a, Eva90b, Eva90c, ER94b, ER94a, Gol87, GW89b, LCE85, LS81, MS95a, Wol98]. Integrated [AAHM93]. Integration [ARM97, BNN95, Koh92, LC97, MRJ95, RAL95, Sim92, Sim94, ZZZ98, Alb89, Beh89, GW89a, LM98, MF99, Red90a, Red91, Söd89, TS98, Wan80, Zam84, bS89]. Integro [Da93a, GGS97, Gro93, Bee87, BL78, FL99, Guo99, HL88b, HL88a, TET99, TT84]. Integro-Differential [Da93a, GGS97, Gro93, Bee87, FL99, Guo99, HL88b, HL88a, TT84]. integro-partial [TET99]. Integrodifferential [Abd93, NSV92, NS97, Rua92, Da93c, Da93d, ER88, EL90, MB88a, RSS89, Sol99a]. intelligence [HKK^{+}80]. Intelligent [MFGW97, SWL98]. intensity [Jam86]. interacting [Pat91, Zag91]. Interaction [HM94, Yam95, AA99b]. Interactions [JFW96, WZ95, SKCW94, SP93]. Interactive [AAHM93, DG87b, SYS94, TDK94, KSB89]. interconnected [Cha85]. Interconnections [SP95]. Interdependent [ITN97]. Interdimensional [Mus97a, Mus97b]. Interface [Mus94]. interference [Kle82]. intergrid [KL99a]. Interior [LLY96, LYF97]. Intermediate [RGL98]. Intermittently [Kar98]. Intermodulation [Abu97]. Internal [PJ95, ADHE82, Kni82, Mos90, Sla91]. International [Sla93a]. interpolants [EJNT88]. Interpolating [Bur96, HD88]. Interpolation [CWL92, Dég96, GSC95, ADM91, Beh89, CW89, HL82, Jia99, LMV91, Mv98, Ram90, Sai88, SW79b, Wat89c]. Interpolations [Wan92]. Interpolatory [Beh92]. interpretation [Mos90]. Interspecific [Les89]. Interval [DTW88, Eis98, HG83, LL96, Mad90, Moh94, MW88b, Sha97, SHW97, SW94, TH96, Wol96, Wol98, Cor89, Ge89a, Jam85, Kop90, LS87, LS91a, SW90b, SS92, Wol89]. Intervals [WLB97, KR84, RC91]. Intervention [BGY97]. Intravenous [LL94a]. Intrinsically [Béc98]. Introduction [SU95, Tes91a, UTL^{+}95, UHL96]. Invariant [Ayo93, Bre81, DGV94, Hil80, Kop90, NSE81, NS97, Sag91, Tak98, Als79, Gyö90, KR84, KR87a, LS99, Lor80, QHN78, Sen90, WKTM88]. invariant-imbedding [WKTM88]. invasions [CV89]. inventory [Zha89]. Inverse [HZL97, IM92, JM93, Ji94, Lam99, MB95, Pen93, RBR94, SP95, Tag98, BL91, CAP88, Kag91b, Lin98, Sen90, Sev99a, Sev99b, SY98, Wan87, Wei98, Wei99b, YS98b]. inverses [Che96a, KRTW83, Sta99]. Inversion [Hai97, Uen80, Wan93, Fym79, LB92]. Invert [LSS94, LSVS94]. invertibility [MR91]. Invertible [DGV94]. Inverting [Kwo93]. investigation [Ado87a]. Investment [CH92, Hoj94, OU99]. Inviscid [PN97a, Joh83]. Involving [Avg94, CS97, MFF92, SS98a, Waz97c, ASK98, CSY98, CSZ98, KB84b, RMR86, Rol98, Yür99, ZA88]. iodine [HRA98]. Ion [Kub94a, TPS98]. IRK [Abd92]. irrational [EE89, Eva90a, Eva90b, Eva90c, ER94b, ER94a]. Irregular [Mur94]. Irrigation [MPD97, MN91]. Isaacson [Pap84]. Isolated [RRB95]. Isoparametric [LC97]. Isotropic [AA95, AA96]. Issue [DGP93, SU95, UTL^{+}95, UHL96]. Italian [Chu98]. Iterate [BG94c, Gol93]. Iterated [GL94, ZSL98, Ber89]. iterates [Nar99a]. Iteration [Bai97, NAM93, SB92, Bai99b, SL99]. Iterations [CG96, LWX92, ZS99]. Iterative [CN97, Che96a, Che97, CZ92, EGH97, FBSS96, JN95, LS95, Lu98, MES94, QHN78, SL92, Wan95a, Zha97a, AM81, Du83, EL91, EGH98, Faz91, MB88a, Moo81, Pan88, Sca89, SK86, Udw92b, UK94, Zha98c]. IV [Cha92e, Die89, ER94a, VR99d]. IVPs [KKPL96, EJNT88].

J. [Di 98]. jackknife [Bre91]. Jacobi [CS93, CS95, EA94, YST97]. Jacobian [Chr82, KK82]. Jet [CJF93, JFW96]. jets [Ram99a]. joined [MS94b]. Jones [EMEH96]. JOR [Udw92b]. Jordan [Sen90]. Jumping [IKU96]. Junction [RB95, SEG96]. Jupiter [Tuc97].

Kadomtsev [Ado96c, BG97]. Kakutani [Wan96b]. Kalaba [Cas91, KKU91, Kag91b, Tes91b, Tes91c]. Kalman [CB94a, CB94b]. Kampé [MS95a]. KdV [Ado97c]. Kernel [NSV92, Da93c, Da93d, Kan99]. Kernels [Jin97, NS97, NN87, BW77, Bow78, JS98a, KKS87, Mor90]. kidney [BL91]. Kind [Abd93, NN87, WK96a, Gol79, NSE81, vGG79]. kinds [ESL98]. Kinetic [Bay97, BY99, OTA83, RR98]. kinetics [Bay98b, Ten91]. kinship [Sch91a]. Kirchhoff [Whi88]. Klein [Ado97b, JV90, WCG97]. knapsack [Dro90]. Knopoff [Ado96a]. Knot [Beh95, Beh92]. knowledge [DBW88]. Kober [ASK98]. Kobori [LS97]. Kornai [Sos91]. Kornai-Weibull [Sos91]. Korteweg [RC99]. Kramer [Sai96]. Krawczyk [SW85]. Krawczyk-Moore [SW85]. Kronecker [MFF92]. KSR [Zhu94]. KSR-1 [Zhu94]. Kutta [PF99, Abd92, And93, EJNT88, HS94, SW79a, Sim97, SIA96, Vos89].

labor [Puu84]. lacunary [FT98]. Ladik [TPGV97]. Lag [Boy95, Sim92, Sim94, Sim99b]. Lag-Averaged [Boy95]. Lagrange [Got85, HD88, IA89, RS88]. Lagrangian [Lam98, MM85]. Laguerre [KN79, KNS81, RS92, Sum84]. Lake [Bec85, van85]. Lakshmikantham [Zan87]. lambda [DLRJ81]. laminar [Han92]. Land [AFM97, PMMA97, RPT97, Tur88]. Land-Water [RPT97]. Landau [Kwo93, Tak98]. landscape [CJ88]. landscapes [Mil88b]. Laplace [Gzy95, Gzy97, Hai97, MNK99, Sco80, Sco81, Yan99, Yür99]. Laplacian [HB91]. Lapwood [BH93, HB91, HB93, HB91]. Large [BNN95, Cha96b, KWS95, KSZ^{+}95, Bai98, BW88, Ge89b, ÖGDB98, RB85, Sen86, She90, Zel87, dAMY89]. large-scale [BW88]. Largest [BG94c, AAES99]. latent [DLRJ81, Lan81]. lattice [Mur89a, Mur91a]. lattices [DH99a]. Law [HLW94]. Laws [Chu96, Con93a, Jim94, Li94]. Lawvere [Gun94]. Layer [AS94a, AS94b, HM94, Had95, Raj94, Waz97b, AAA99, Zam84]. Layered [ENS96, dL95]. layers [FH98, NR98]. Lead [LWLL95]. leader [SMO91]. leader-leader [SMO91]. leading [IRU83, RUS^{+}84]. Learning [AH94, GSK93, Per96, SM94, ESL98, IFU98, Lak81]. Least [AB95a, Cha96a, LYX^{+}93, TM92, Whi97, AK80, Ber89, CXL94, HW80, IA89, KRT83, SUJ98, Sha89a, Sug83, SSY99, Uen83, Yan98, YL98a, Ye98a, Ye98b]. Least-Square [LYX^{+}93, Ye98a]. Least-Squares [Cha96a, CXL94, Sha89a, Sug83, Uen83, Ye98b]. Leela [Zan87]. left [YDB79]. Legendre [Lon87c]. Leibniz [Mur92]. Length [Sch92, Faz91, HF95, KK91]. Leslie [MR91]. less [Zem86]. less-developed [Zem86]. Level [Flå92, LL95, LY98, RCG94, Zha97a, AKK99, YK97b, Zha98c]. Level-constrained [Flå92]. Levels [CB94b]. Levin [GW89b]. Levin-Sidi [GW89b]. Levison [Mor94]. Levitan [AR98a, Kin93]. Liapunov [RS88, Sol98]. Library [Hus90]. libration [Tuc97]. Lidstone [WA99]. Life [Con93c, Gun93, Gun94, GKn91, Gun92, JS98b, KG91]. Lifelike [ZSL98]. light [KL85]. Lighthill [Boy97a]. Like [Arg96b, Arg97a, Arg99a, Arg99d, Arg99e, GP97, Hua99, Yan99, Arg94, Arg96a, Beh96a]. likelihood [NC85, RB85]. Limit [Alw97, Ji94, Koł94, Sta99, AS81, CM90, KM79, Mur91b]. limitations [Pat93]. Limited [Lap92]. limiters [Dav94]. Limiting [MMMK96, Vis80, Zag92, ZK98]. Line [Hai97, dL95, AA99b, HT98, Spi90, Sug83]. line-spacing [Spi90]. Linear [Bai97, Béc98, BT97, Cab94, Cha95b, Che97, CH93, DS86, Eis98, ES93a, Hoh97, Hoj94, JY97, KKM90, LMSS97, Lop92, MS94c, Mor95, MKT96, NS96, Pap96, PN97a, RK87, RAL97, Reg94, Reg96, SS94, Sug95, Yak92a, YK97a, Abu99a, Ada89, Ahm79, Als79, Bai99b, Bre81, Cab95, Cha95a, Da93b, DPTT98, FLZ91, GS99, Ge89a, Ge91, GZ84, He98, HKK^{+}79, Jia99, KR81a, KS81, Kha81, KB98, Kop90, LR85, LV87, Lee98, LM98, MR99, PVM88, Pen84, PK98, QA98, Rin98, Sai99, SK86, SR90, Sha89a, Smi86, Sug83, Sug87, Vau80, Wei98, Wei99a, WKTM88, Whi89, ZA91]. linearities [EB99]. Linearization [Ram97a, Ram97b, RGL97a, RGL98]. Linearized [RGL97b, Ram99b, Ram99c, Uen95, Ram98a, Ram98b, Ram98c, Uen91]. lines [AST86, Ioa88, Mik87, SSV89]. Linkage [MS94c]. Liouville [ABW99, Bai81, WJ92]. Lipid [Kub97, KK90, Kub94b]. Lipschitz [Arg97a, ER88]. liquid [Ram99a]. LISP [Cha92b, Cha92c, Cha92d, Cha92e]. Living [Kon92, MS93c, MK93]. Load [dL95]. loading [Bec85]. Local [CW97, GSC95, KTW81, Kaz99, Koč92, Yam95, Sol99b]. localized [ALMS90]. locally [KM86, Mv98]. locating [DFM94]. location [Pes91, She90, Van98]. log [Smi86]. log-linear [Smi86]. logarithmic [Boy89]. Logarithmically [GM93]. Logic [MO93, TGK91]. Logistic [FC93, JS96, Ram98b, Ram98c, TET99].

Lokahi [YH97]. long [ZS98]. long-range [ZS98]. loop [Mur91a]. loss [MN91, TW79]. losses [DDS85]. Lotka [AM98b, LL96, MW98a, SA96]. Lottery [Fin95]. Love [Rin98, AAA99]. Low [BR88, RCG94, SEG96, KCS90, Ram99a]. low-dimensional [KCS90]. Lower [OT92, LS87, OT93]. lowest [CL98b]. lowest-order [CL98b]. LTDRM [SZ97, ZL98b]. Lubrication [KH96]. Lunar [TJ95]. Lyapunov [Aft80, Kaz99, WL97].

M [Di 98, Jac95, Jac95]. M/M/1 [Jac95]. Mach [RCG94]. machining [MCD89]. Maclaurin [CSZ98]. macro [Zha89]. macroeconomic [Moh99c]. macroscopic [MO92]. MACSYMA [WH84, RK87]. MACSYMA-generated [WH84]. Magnetic [HM94, VJ88]. Magnetized [AC97]. Magneto [AA94]. Magneto-elastic [AA94]. Magnetohydrodynamic [Mei93]. magnetohydrodynamics [Mei94]. magnification [Mit90b]. magnitude [LMV91]. maintenance [Ree87]. Making [DGP93, PN97b, Sar93, YLS93, YH97, MCD89]. manageable [Chr82]. Management [HKLP97, TDK94, YH97, Ass87, CW84, SGB87]. Managerial [Sar93]. Manifold [Beh90, SST95, Lia98]. manifolds [Sol99b, SK98b]. Many [Dég96, Hoh97, KT83, Pal78, RC81]. Many-Variable [Dég96]. map [RC81]. MAPPACC [Mat86]. Mapping [GF96, Sor97b, Hof87, Rob82, Wan80]. Mappings [MM94, Boy89, GT82, Har82, Sor99]. Maps [Ram98b, UR97, KR91, Mil87b, Nar99a, Szi90, Wat79]. Marcel [Kan99]. Marchenko [AR98a]. Marching [Nak82, BKS83, BKS84a, BKS84b, KB84c, KB84b, Roa82]. Market [SM94, SMY95, Mos90, Sos88, Sos91]. marketing [Zem86]. Markets [Pal78].

Markov [Ben89, Har95, LS93b, MS97, PM89, Pat90, Pat92]. Markov-renewal [Pat90, Pat92]. Markovian [MES92]. Marquardt [GPAB83]. Mars [Tuc96]. Martingale [Cha88b, Yip92]. mass [DSK82]. Massive [MO93]. Massively [Dég96, Hoh97]. Master [Ada98]. Master-slave [Ada98]. Matching [AR93]. Material [ENAAA95, HKLS93, HLW94, HSA97, SZ97, Whi93, AA99c, HL91]. Materials [HLW94, PZ94, LR95]. Math [AaAaZ00b, AaAaZ00a, Cab95, Di 98, JS98a, JNF95, MS92c, MS92b, Mus97b, OT93, Sor97a]. Mathematical [Chu92, CFT^{+}93, HD96, KG91, KSZ^{+}95, Mur94, Mur96, STD98, Zhe93, ALMS90, AR90a, AR90b, BCZ88, GD87, Ten90, ZPY99]. mathematics [Kag91b, Les89, Mus84b]. Matrices [How94, KA96a, LSVS94, LLX96, MS94b, ABT98, AB89, DLRJ81, DLR85, HD88, Hua91, JCC98, RG98, Ser79, Udw92b, Var83]. Matrix [BG94c, CC94, FHG96a, Jód92, JL93, JN95, KPT87, LSS94, LP97, MS97, MES92, RAL95, Sha94, SP94, Wan93, Axe86, BH89, Cha90b, Che96a, DLR81, Har87, JN91, JMV93, KST80, KK82, LB92, MD81, Sas85, Sha89a, SY98, Sum84, TCS92, WH84]. Matrix-Geometric [MES92]. Matrix-Sign-Function [LP97]. max [MS88, ORSS83, TZ83]. Maxentropic [Gzy95, Gzy97]. Maximal [Pap92, LS91a, Pig87]. maximal-order [Pig87]. Maximum [Cha93, FT98, NC85, SHW97, Din98, Fie84, Hir84, Kor97, Rav87, RB85, Tag99, Tom80]. maximum-entropy [Fie84]. Maximum-likelihood [NC85, RB85]. Maxwell [Ado96b, Cha88a, Cha90b]. MCDM [MTP93]. Mean [AB95a, Bha94, FLG92, GR97, Ibr96, Sha87, Ahm79, HK81]. Means [SV94, Mat86, WH87]. measles [ALMS90, DML98]. Measure [GWL94, IL95, Lap92, Van98]. Measured [Vev92]. measurement [JB89a, Pat93, Rol98, TyG99]. measurements [Jam85]. Measures [MS93b, MGL92, FL99, Kop90, LL89, RS88]. Measuring [Mil88b, PLPW96, Tuc95]. Mechanical [RRB95, UK95, ZL98a]. Mechanics [Bel95, Gre87a, McW82, YF91]. Mechanism [IOAB95, WM91]. Media [Ado97e, ENS96, HS96, LZ96, YBR97, Ado87b, AH99a, AH99b, BH90, BH93, Cha88a, HB93, OŞL80, Pat88, Zhu94]. Median [FS94]. medication [Ass87]. medicine [FEGL90]. Medium [Ezz94, HM94, KD95a, KD95b, WZ98, AAAaZ99b, AAA99, AaAaZ00b, Ahm99, Sha87, Ten91, Waz98b]. Meeting [ITN98b, Ino99]. membrane [KK90, KK91]. Membranes [BMS92, Kub94b]. Memories [MKT96, KKU91, KKM90]. Memory [El 92, ALZ99]. Mesh [Arg97a, Abl82, Jou82, KB93, MS83]. meshes [Bra82, Chr82, KK82, KKL99, SB82, WN91]. metabolism [Cas88]. metabolism-repair [Cas88]. Method [AG85, Asa92a, Asa92b, Asa97a, Asa97b, BPZ96, Bay97, BS93, Boy95, Cab94, CWL92, Cha94, CW97, CAQ93, CT96, CLC97, CT97, Cro93, Cro95, Do96, EN92, EFS95, GR97, GC94, Gol96, Gro93, HSG94, HZ95a, HG83, nHyG97, HKLS93, HLXZ97, JR93, JR95, KR95a, KA95, Koč92, Koh92, Kwa97, LV87, LYX^{+}93, LLY96, LYF97, LLL97, LS98a, LLS94, LLS96, LP97, MSLB92, Moh94, MS95b, Mou95, MBS98, NN87, OT92, PS94a, PJG^{+}97, PZ94, SJ96, Sel94, Sel95, SKG97, SHW97, Sim92, SK96, Ste94, SL92, SP95, SL96, Szi97, Tuc95, Vev92, Waz95a, Waz95b, Yee93, Yil96, YBR97, YL96, Zha96a, Zha97a, AST86, AM81, Ado87a, AD98, And93, AT85, Arg90, AC94, Arg99b, Asa98, Bai91, BM89, BL78, Bow78, BS86, BCZ88, Cab95, CG87]. method [Cha88a, Cha90a, CL98b, CK99, CH99, CC99b, Da93b, DEV95, DK96, DR88, Din98, DZ99, Faz90, Faz91, Fei85, FpDS99, FG81, GXQ99, Gol79, Gol88b, Gol99, Gre86, HW98, He96, Hei91, HGS88, HS98, HS99, Hof87, HAST98, HY98, yHyG99, Ioa83, Jam83, JV98, Jin94, JF91, KR87b, KK83a, KKS87, KS78, KST80, KL99a, KS93a, KS93b, KL99b, Kle82, Knu91, KM87, KK91, KL89, LP81, LRV84, LS87, Lee98, Li94, Lin98, LM98, Lor80, Mad90, MW98a, Mik87, MB88a, Mit90b, MLB99, MH80, NR98, NR99a, O’N90, Ohm98, ONOT98, Oma86, OT93, PV81, Pap84, PP93, PZ99, QHN78, Ram98c, RC99, Red90a, Red91, SSV89, Sch91a, SUJ98, Sev99a, SR88, Shy87, Sol98, Spi90, Sug83, Sun98, Tha82, TCS92, Udw92b, Van95, VR99a, VR99b, VR99c]. method [VR99d, VR99e, VR99f, WSL99, WS87, Waz98a, Waz99c, Waz99d, XL98, Yan98, YL98a, Ye98a, YJ99, Zam84, ZM91, Zha98d, ZL98b, ZG90, bS89]. Methodology [Fis85, Yak92b, MM87b, Sav93]. Methods [Abd92, Abd93, AC97, Arg94, Arg96a, Arg96b, Arg97b, Arg97a, BS97, Bai97, Ben92, Bla96, BT97, Che97, CZ92, Cic92, Cic93b, Da93a, Dav98, EGH97, GG96, GGS97, GM94, GP97, Hei93a, IKP97, KKPL96, KBS96, KSZ^{+}95, LL95, LC97, Lu98, LB96, Net89, NP92, Ona97, PK98, Ram97a, Ram97b, RGL97b, SW92b, SA94, Sim94, Sim97, Son96, SIA96, SBH96, TPGV97, TJ95, Wan95a, WK96b, Wol96, YK97a, ABT98, Aga86, Alb89, Arg98, Arg99a, Arg99c, Arg99d, Arg99e, Bai98, Bai99a, Bai99b, BPE99, Bee87, BT98, BH89, Cao88, CL99, Cas89, CW91, Che96a, DTW88, Dos90, EC80, EM90, EL91, EGH98, FS90, Gol89, GB90, GB98, Gri89, Gu99, Han98, HG84, Hei88, Hei93b, HKWdZ83, HV90, Hua91, HZ98, Hua99, HL82, IST93, JMV93, KB85, Kad86, KB93]. methods [Kap88, KW86, KT89, KP98, KBÖ99, MM87a, Man88, MMP89b, Mey78, Pan88, PF99, QC98, RGL98, Ram98a, Ram98b, Ram99d, Ram99c, SB82, SS86, Sim99a, Sta99, TS98, UK94, Vos89, mW98b, Wat89a, WM91, Wol89, XL99, Yan87, Ye98b, Yse86, ZH86, ZH93, Zha98c, Zho89, Zla89, dAMY89, dOP89, van85]. methotrexate [NHG83]. MG [RS86]. MG-FAS [RS86]. MHD [MS94a]. microbial [ZK98]. Microelectronics [SP94]. Microscopic [MO92]. midpoint [AC94, VR99e]. Migration [Sla93a, Sla91, TETH98, TET99]. mild [Arg99b, KN90]. min [MS88, ORSS83, TZ83]. min-max [MS88, ORSS83, TZ83]. Minesweeper [Ada97]. minicomputer [MB82]. Minimal [Sha94, Sim94, Zha97a, Ada99, Cao98, CG83, HJ92, LS91a, Mit92, Sim99b, Zha98b, Zha99a]. minimal-energy [HJ92]. Minimax [Let97, Pap95, SHW97, Ge91, Sen86]. minimization [VCM91]. Minimize [PMMA97, KT83]. minimizing [BD90, Ge89b]. Minimum [Sor93]. Minimums [BG93]. Minmax [SW96]. Mises [FH95]. Mismatched [Che95, RL96]. Mistune [WK95]. Mistuned [WK95]. Mixed [BKS84b, Gra95, HM94, Had95, KTZK96, MS88, SL92, Wan96a, Cha90a, CL98b, KB84b, LV87, Lee98, Moo81, Sco80, Sco81]. Mixture [HLXZ97, BS86]. mixtures [BE85]. mobile [SWL98]. modal [HG84]. Mode [Jan95]. Model [Ado96a, Alw97, ADS96, BG94b, CK95, CT96, CFT^{+}93, Con93c, EM98, EB86, GP94, GQS98, GMS96, HLXZ97, HD96, IM92, KD95a, KD95b, Kim96, KPP93, LL94a, LG97, MAS95, MPS97, OA97, Pap94, PN97b, PS94b, PR93, PBL92, PJ95, SKB95, SAF^{+}97, TPGV97, TH96, TJ95, Whi93, WZ98, Yak92b, Yip93, ZA96, Zhe93, AKK99, BR99b, BL91, BCZ88, BCZ89, Con89, Fah88, Fra89, Fre90, Gar90, Gar89, HK81, HY99b, HM99, Hos99, HY99a, Jam85, JTT80, JT80, Kap80, KPJ83, KP86, KK90, Kub94b, LG99, MW98a, Mit84, Mos88b, Mos90, Pat90, PZZy99, Puu84, Rol98, SSW99, Shi89, Sos88, Sos91, SHR94, SS87, SMO91, TW89, Ten90, TyG99, Tur88, VJ86, Váz98, War82, Waz99a, Whi88, Whi91, Yip92, Zha89, ZPY99]. modeled [Szi90]. Modeling [ABES84, BMS92, HY98, Hos99, LSKH96, LS98a, RBPK96, Ros93, SV85, UR97, Unb96, YF91, Bog91, BDJR98, CG83, CG88b, FMR86, Gre87b, Han92, Hen86, Hog88, Joh82, Kag91b, ÖGDB98, Pal86, Tes80, YGH90, van85]. modelling [Moh99a]. Models [Ami97, AS92, AB94, BBB97, Ben92, Gra97, Ham94b, IKP97, KA96b, Mil95, NP92, Raj92, Raj97, Ros93, RW93, TA93, AK85, Ami99, AT85, AS90, AH99a, AB95b, BH93, EC80, EB91, Fra89, HT90a, HT90b, HB91, HB93, HF95, HKKN78, HK88, HC85, HY99c, Kee89, KM94, KB98, KP98, MN91, MP86, Moh99c, NH86, Pat93, Pat91, RB85, STD98, Sla91, Smi84a, Sor85, TZ83, VS90, War87, WK98, WM91, Whi89, SO90]. Modes [WK95]. modification [GPAB83, Waz99d]. Modified [LL95, MMMK96, Sim97, TM92, Bar89, BTW94, Bow78, KMR^{+}91, KBÖ99, LS91b, MW98a, ONOT98, VR99a, VR99b, VR99c, VR99d, VR99e, VR99f, WC88, WSL99, Waz99c, Zho89]. modifier [KK91]. Modular [Car90]. Modulated [Reg94, Reg96]. molecular [Gre87a]. molecules [KK91]. MOLP [DG87b]. Moment [FT97, EC80, FT98, JTT80, Tag98, Tag99]. moments [Lon87a]. monitoring [Cas83, Jam86, Mos88a, NA91]. Monomial [MBS98]. Monotone [AS92, Cab94, CG96, Du83, EL91, Fio95, LP81, LRV84, LWX92, PV81, SL92, Vol92, WA96, Cab95, DR88, LS78, Moo81, Oma86, Pan88]. monotonic [Lon88]. monotony [LV87]. Monte [QC98, RBR94]. Moore [SW85, Wan87]. MOSI [Bar89, Bar91]. Most [Cro95, KR81b]. Motion [EM98, KU92, KUX95, RW93, UK95, VJ88, BH90, BS86, Dua99b, ST96]. motivated [TETH98]. Mountainous [KUO95]. Moving [dL95, BDJR98, Lan81, SUJ98]. moving-boundary [BDJR98]. MP/24 [LMM86]. MR [Cab95, Di 98, JS98a, JNF95, MS92c, MS92b, OT93]. MSS [KUO95]. MUDPACK [Ada89, Ada91, Ada93b]. MUDPACK- [Ada93b]. Multi [Gup98, JL97, MDR94, Wam92, YLS93, Zha97a, LR85, ÖGDB98, Söd89, Zha98c]. Multi-Agents [JL97]. Multi-Attribute [MDR94, YLS93]. Multi-Homogeneous [Wam92]. Multi-Level [Zha97a, Zha98c]. multi-objective [ÖGDB98]. Multi-Point [Gup98]. multi-purpose [Söd89]. multi-time-scale [LR85]. Multiattribute [MW93, Mal91]. Multiblock [SW92a, SR97]. multibody [SK98a]. multichannel [Ais88]. Multicriteria [MFS93, PMMA97, Van97, MCD89, Tes91c]. Multicriterion [BDS96, DGP93, Mat86, RPT97, CDW91]. multidelay [FLZ91]. Multidimensional [Kon92, Ram97b, Ram98a, Ram99b, XZL98]. multidisciplinary [SHR94]. multielement [Hal82]. multifacility [Pes91]. Multigrid [Ada93b, BF83, Cau83, CWL92, Do96, Hei93a, HKWdZ83, Kam88, Koč92, KKL99, MM87a, Man86, SS86, SR97, Tha84, YO98, Zas93, Zha96b, Ada89, Ada91, Axe86, BB83b, Bar88, Bra86, Cao88, Den83, Den86, Den88, GMS83, Gre86, Hei88, Hei91, Hei93b, HT88, HGS88, Hol86, HV90, Hua91, Jam83, Jes83, JPS86, KP85, KL99a, KW86, KM87, KM86, KL89, Man88, MP90, MS83, Mel86, Mer86, NW90, Par85, Par87, Spi90, Stü83, TT86, TT88, Van83, Vol92, XL98, XL99, Yan87, Yse86, ZH86, ZS98, Zha97b, Zha98a, Zha98b, Zha99a]. Multigrid-sinc [SS86]. Multigroup [Zas93]. Multilayered [KD95a, KD95b]. Multilevel [Mit86, Vas91, Zem86, XL98]. multinational [Baa88]. multinormal [BE85]. Multiobjective [AAHM93, EFS95, HDP85, TDK94, DDS85, SGB87, SS87]. multiparameter [ASZ94, pLR87b]. multiphase [Pat88, Zhu94]. Multiple [AFM97, JFW96, Joh83, KS92a, LM95, Lui99, Mal93, YH97, Yeh97, AS99, Bre81, Cha88a, DS87, JB89b, MS99, Moo81, Sav93, WN91, Wol98]. Multiple-criteria [Mal93]. Multiple-grid [Joh83]. Multiplication [CS96]. Multiplication-Free [CS96]. multiplicative [BM89]. multiplicity [HSZ94, Shi89]. multiply [Har82]. multipoint [AG85, Als79]. Multiprecision [Jay96]. Multiprocessors [MBCV95, Gre86]. Multiquadric [HLXZ97, SKG97, ZL98b]. multirate [dAMY89]. Multirational [Pap96]. Multisector [Pap92]. Multiserver [MES92]. Multisinusoidal [Bay93]. multisplitting [Bai98, Bai99a, CL99]. Multistep [SBH96, dOP89]. multisteps [JMV93]. multitasking [SM88]. Multivalued [ESI95, TGK91]. multivalued-logic [TGK91]. Multivariate [BB96, JB89a, AB87, JB89b]. Mutual [MGL92, Par95, GFM88, Gil89]. mutualism [Zag91].

Nagumo [Jac92]. Narrow [RW93]. Nash [Szi99]. Nashville [Tho82d]. Naslund [Sar93]. Nations [Chu92, Chu94]. Natural [Ham97, IL95, MS93a, Ada99, CW84, Hir84, SGB87]. natural-resource [CW84]. Naturality [NAM93]. Navier [Cha96b, Gno82, GP97, Ham98a, nHyG97, yHyG99, LL95, LY98, Shy87, SJK86, VWSS94, Ye99, ZH86]. Navier-Strokes-like [GP97]. Navigation [Tuc96]. Near [Let97]. Near-Minimax [Let97]. Necessary [FLZ91, Waz97a, BD90, Gun92]. negative [Sas83]. negotiable [YB99]. negotiation [LSY98, YB99]. neighborhood [Mit90b]. neighborhoods [Mil87a]. Nested [KPT87, MX90a, MX90b]. net [Pat89]. netlet [GLL98]. Network [BB83a, LSKH96, LL94b, PS94b, Wan93, CS85, Cha88b, LB92, Pat92, SY98]. Networks [AH94, MZ98, Mou95, PJ95, WZ95, YST97, HAST98, Li98, PZZy99, YL98b, Zel87, Zem86, ZPY99]. Neumann [HKLS93]. Neumann-Richtmyer [HKLS93]. Neural [AH94, LB92, LL94b, MZ98, Wan93, WZ95, YST97, GLL98, Li98, PZZy99, SY98, YL98b, ZPY99]. Neuromuscular [OS80]. Neuron [CFT^{+}93]. neuronal [HT90a, HT90b]. neurons [ESL98]. Neutral [CUK92, GL92, Gra95, HHT98, HH96a, HH97, LG94, Yan97, GS94b, MB88b]. Neutron [Zas93]. Newton [Arg94, Arg96a, Arg96b, Arg97a, Arg98, Arg99a, Arg99b, Arg99d, Arg99e, DTW88, GP97, HW98, HG83, Hua99, Mad90, Mur92, SW90a, SW92b, Sun98, SL99, SSY99]. Newton-Like [Arg94, Arg96a, Arg96b, Arg97a, Arg99a, Arg99d, Arg99e, Hua99]. Newton-Type [SW92b]. Newtonian [Had95]. Nicolson [HM99, Jin94]. Nine [Mur96, YO98]. nine-point [YO98]. No [Ham92, Cab95, Di 98, JS98a, JNF95, MS92c, MS92b, Mur89b, Mus97b, OT93, Sor97a]. No-Slip [Ham92]. node [Pat92]. Noise [KKH95, SEG96, Waz97a, WL97, TW89, TW90, VS90]. noise-in-variables [VS90]. Non [Ado97b, AH99a, Da93b, ENAAA95, MMT99, SA96, VC98, AAAaZ99a, AAA99, AaAaZ00a, AW99a, AD98, DPTT98, EB99, MR99, NC85, Sai99, SSY99, Szi99, UK94, GSI96, Had95]. Non-Boolean [GSI96]. non-classical [Sai99]. non-classically [UK94]. Non-Gaussian [MMT99]. Non-Homogeneous [ENAAA95, AAAaZ99a, AAA99, AaAaZ00a]. non-linear [DPTT98, MR99, Sai99]. non-linearities [EB99]. Non-Newtonian [Had95]. Non-Perturbative [Ado97b, SA96]. non-positive [AW99a]. Non-reactive [AH99a]. Non-smooth [Da93b, SSY99]. Non-standard [VC98, AD98]. non-steady-state [NC85]. non-symmetric [Szi99]. Nonautonomous [AM98b, IM80]. Noncollocated [Kum95, Udw91, Udw92a]. Noncolocated [LFS95]. Nonconforming [LY98, CL98b, KL99a, XL99, YK97b]. Nonconstructible [Ada99]. Nonconvergence [SB92]. Nonconvex [YL96]. Nondifferentiable [SW94]. nondispersive [Udw91, Udw92a]. nonequidistant [Vul90]. Nonhomogeneous [Koł95, NFJ94, Cas81, JNF95]. noniterative [Faz90]. Nonlinear [Abd92, Abu99a, ARM94, Ado96d, Ado97a, Asa92b, BGY97, Cab94, CK79, Cha89a, Che95, CZ92, CUK92, Chu96, Da93a, DS88, Eis98, FLG92, GP94, Gzy97, Hab95, KRT83, KT91, Kar98, Kum92a, KA96b, LLY96, Lin96, LYF97, LWX92, MW93, Mo97, MZV96, ONOT98, PS94a, Pap95, PN97b, PJG^{+}97, Pen93, PPS94, RAL95, Rya96, SZ97, ŠNG92, SA94, Ste94, SS97, TA93, TFM93, TH95, UR97, Unb96, Vev92, Waz95b, Whi93, WCG97, XZW96, Ado87a, Ado87b, Ado88, AKV83, Bai98, Bai99a, BL78, CN90, Cab95, Cas87, CL98a, Chu97, Chu98, DH99a, DT87, EM79, EC80, ES93b, ER88, EL90, FPGV95, FS90, Gar90, GH89, Ge90b, Gla78, GT91, GS94b, GW89b, Gu99, Gup90, HSZ94, HEH81, Hau82, HK81, He96, Hol89, IP89, JC97, JTT80, JT80, JV90, JBV98]. nonlinear [JPS86, KK83a, KH78, KM79, KV84, LP81, LV87, LS81, LS99, Lia98, LF99, MM91, Mit86, Mor98, OŞL80, Pac85, Pei99, PGHX99, Sav93, Sen86, SW89, Sol99a, TET99, TZ83, Vol92, mW98b, WH87, WS87, Whi91, WC99, Yan87, YF91, ZZ98, ZH93, Zha99b, Zhu98, ZL98b]. Nonlinearities [Abu97, KS92a, Ado81]. Nonlinearly [TLG90]. Nonlocal [KTW81, Mo97, VD97, CZ99]. nonmonotonic [Kru81]. nonnegative [Gyö90]. nonnested [MP86]. Nonnormal [GD87]. nonoscillation [ES93b]. Nonoscillatory [She96, TPL95]. Nonparametric [Ahm93, PY94, Sch91b, TT90a, TT90b]. nonrectangular [KB85]. Nonrecursive [Ben87]. nonseparable [KB84c]. Nonsmooth [Da93d, dYS97]. Nonsquare [LFS95]. Nonstandard [Kon92, RGL97a, SV92a, SV92b]. nonstationary [Szi90]. nonsteady [Yag82]. nonstiff [bS89]. Nonsymmetric [Jód92, Cao88, Den83, Man86, YK97b]. nontrivial [Gyö90]. nonuniform [AST86]. Norm [Cai93, Cro93, CK99, GP88, JMS98, Tom80]. norm-relaxed [CK99]. Normal [BB96, BG94c, Kha96, WR96, Cra79, Rol98]. normality [MN79]. normed [Kha81]. norms [Da93b]. Not-a-knot [Beh92, Beh95]. Note [AS94a, Cha93, CAQ93, FG94a, Gol95, Gre87a, Jin97, Mac93, Sem92, Sen93, Sri95, Szi97, Zha96a, AB89, BW77, Cao98, EM90, Gil89, Gol88a, Gol99, Ham98b, Har86, HT88, JS98a, MS92a, Par85, Red91, TT88, Udw81, Var80, WM87, XL99, Zan87]. Notes [Wan96b]. Novel [Bay97, CSY98]. NP [Kor97]. NP-completeness [Kor97]. NR [KL99b]. nuclear [FEGL90]. nuclear-medicine [FEGL90]. Null [CUK92]. Null-Controllability [CUK92]. Number [Cha93, Pav95, RCG94, Sor97b, Wam92, CB94a, KCS90, MM85]. Numbers [Fin95, Gri89, Ram99a]. Numerical [AS81, ARM97, ALZ99, AC97, Asa92b, Asa97a, CS85, Cas81, CJF93, DF96, Fab96, FPGV95, GM93, Gol96, Gre97, GPRV86, Hal98, HFA98, HEH81, HSA97, HLXZ97, HHZ97, IKP97, IPP91, JFW96, JMV93, Joh82, Kad86, KR87c, KR87d, Kha92, Koh92, KSZ^{+}95, KV84, KR91, LS93a, LC97, LLL97, LWX92, Lu98, MMP89a, MX90b, McG98, Moh99b, MPS97, MRJ95, NN87, Ohm98, OL80b, PS94b, Pav95, RBPK96, Red90b, SKB95, Sey81, SKCW94, SKG97, Sim92, Sim94, Sim97, SK96, SK98b, Tho82a, Tho82d, TPS98, TJ95, WZ82, Waz93, WM91, ZW95a, ZMV99, AR98a, AKV83, ADM91, Bra78, BS86, BCZ88, BCZ89, CBYK99, CW89, Cro80, DH99a, DEV95, DPTT98, Doa90, Eir89, FV91, Gla78, Ham94a, HK89, HK84, HM98, HL82, Ioa83, JV90, JF91, JPS86, KR84, KB84c]. numerical [KR87b, KR89, KN90, KL85, KN79, KT89, Kha91, LW90, Li94, LM98, LF94, LS98b, MLC87, MF99, Mey78, MM85, NR99b, OL80a, OŞL80, ONOT98, RC99, Red90a, Red91, Sim99a, Sim99b, Söd89, SV92a, SK98a, TT84, TS98, UA88, VJ88, Váz98, VR99b, VR99c, VR99d]. numerical-integration [Red91]. NURBS [SYS94]. nutrient [Bec85, JA99]. Nyström [PF99].

O [TW79]. O-I [TW79]. Objective [YH97, MPD97, ÖGDB98]. Objectives [Van97, DG87b]. Objects [KK95]. oblique [Wei99b]. Observability [MX90a, MX90b, MM91]. Observation [Cic93b, CB94b]. Observational [Fri97]. Observations [BBB97, CB94b, FC93, CB94a, KS80]. observed [HKKN78]. observers [CB94a, Cic99]. obstacles [LSP91]. Obtain [PJG^{+}97, Mei89]. Obtained [Mur96, Abu99b]. Obtaining [MKT96]. occupancy [WW99a, WW99b]. ocean [Kaz99]. octant [Sco80, Sco81]. Octonions [Car97, Car88a]. ODE [Ano89, BH89, Mae89, VR99f, Zha91]. ODEs [RGL97a, Sim97, SB92, SBH96, And93, Eir89, IST93, Sha89b, Söd89, bS89]. Off [MPD97, GW89a, Van97]. off-step [GW89a]. Offer [SL96, Szi97, Szi98]. Oil [YBR97, BR99b, VJ86]. Oligopolies [SY93]. oligopoly [SO90]. on-line [Sug83]. One [Asa97b, Gho96, KH78, Kum92a, Lei95, AKV83, AAAN98, DG87b, JBV98, Man86, Pat88, SSV89, Tom80, Waz98b, Zho89]. One-Dimensional [Asa97b, AKV83, JBV98, Pat88, Waz98b]. One-Parameter [Gho96, KH78]. one-step [SSV89, Zho89]. onset [DS87]. onto [Har82]. Open [Lin96]. Operated [KP96]. Operation [PL93, EB91]. Operational [CS93, RAL95]. Operations [KPT87, HKK^{+}79, HKK^{+}80, MCD89]. Operator [BD96, CZ92, Kan97, Kwo93, MS97, SL92, Bar89, DDL93, Din98, DZ99, Gol99, He98, KL99a, NLC95, RGL98, SR88]. Operator-Differential [BD96]. operator-splitting [RGL98]. Operators [Arg97b, Cai93, CS97, Gol95, MA95, ASK98, Arg99c, CSY98, GP88, LS78, Moo81, Mus83a, Mus83b, Mus84a, TT86, Wol89]. Opportunity [DDS85]. Optical [Ado97e, Tuc96, Uen80]. Optics [Pen93]. Optimal [AM98a, Avg94, BdA89, BB96, Cha96a, Chu94, CH93, Fab96, Gar90, Ge89a, GB98, HKK^{+}79, HKK^{+}80, JS96, Lia97, Mil98, Mos88b, Mos88a, Mou96, Pic92, Ree87, RTJL92, RB95, TT90a, TT90b, YDB79, ZW95a, ZW95b, Cha88b, Fei85, Fis85, GGR91, GL99, HH85, JB89b, KS80, KS84, KP91, NHG83, TLG90, TW79, WSL99, YR84]. optimal-control [KS84, TLG90]. Optimality [Cha91]. Optimally [KM95]. Optimization [AAHM93, FGLF88, HZ95b, Liu95, MBCV95, SW94, Wol96, dYS97, Cas87, DF96, Ge89c, Ge90a, GQ90, Jer88, JB89a, Kag91b, KB80, MLC87, MBS80, MS88, Moh90, ONOT98, Par89, Pet88, Sen86, Sun98, SS87, VL83]. Optimized [KKPL96, HY98]. Optimum [BBB97, CH92, Hoj94, HZ96, Kum92a, AB87, PVM88]. option [CY99, Váz98]. Options [PMMA97]. orbit [BPE99, Gri98]. Orbits [FG94b, FG95, DFM94, Sey81]. Order [Abd93, Ayo93, BKM94, BKM95, Cab94, Cha96a, Ebe93, EN94, ESS94, ES94, Ela94, FT94, GG96, GGS97, GM94, GL92, Gup98, JNM92, JL93, JV94, JPF95, JNF96, KM95, LG94, LW94, Lop92, NFJ94, NS96, OT92, SJ96, ŠNG92, Sim92, Sim94, SBH96, TPL95, VD97, VS96, Wan95a, Wan96a, WA96, YLS93, Yan97, YK97a, AW99a, AG99, AAAN98, And93, CN90, Cab95, CN97, CG87, CL98b, DDL93, ES99, ES93b, GS94b, GZ84, Gri98, GL99, Gup90, GK94, HG99, HL88a, IST93, JK98, JNF95, KB93, KT83, KVT98, KS93b, LRV84, LS87, LF99, MR99, Nie84, Nie85, Oma86, OT93, PGHX99, Pig87, Sas85, Ser79, SV83, SV92b, TS98, Van95, Whi89, Wol89, WA98, Yan98, YL98a, Ye98a, ZA91, ZS99, bS89]. order-interval [Wol89]. Ordered [JY97]. ordering [KL99b]. Orders [ES93a, Ioa92, MTP93, DG87b, ESR99]. Ordinary [Cha94, GG96, Gup98, LLS94, LLS96, SJ96, ŠNG92, SIA96, Bai81, Bra78, Cor89, Cro80, HEH81, HK84, KSB89, NJC99, Ram99c, SL83, Sey81, Zla89]. Oreste [MTP93]. Organ [FEGL90]. organisms [Mei89]. Organization [DBW88]. Organizing [JL97]. orientation [Zha99b]. oriented [Gre87a]. origin [Kai91]. Orr [MSZ96]. Orthogonal [Chr82, Eis82b, KVT98, Kin91]. Orthogonality [CS95, RS92, Sri94]. Orthonormal [RK93]. Orthotropic [AA94, ENAAA95, AAAaZ99b, AAAaZ99a, AAA99, AA99c, AaAaZ00b, AaAaZ00a]. orthotropy [AA99a]. Oscillating [BD96, Sim97]. Oscillation [ES93b, GL92, KS92b, LYC93, LG94, MB88b, SS97, WA96, Wan99a]. Oscillations [Gra95, KSZ^{+}95, Mur93, Mur94, Phi92, VLP95, WZ95, Yan97, ZAES96, ZA91, Zha89]. Oscillator [Kar98, PJG^{+}97, BKS98]. Oscillators [ARM94, MF99]. Oscillatory [ESS94, ES94, Ela94, Net89, Pei99, TPL95, Wan99b, Doa90, GS94b, PGHX99]. Osher [Vul90]. Ostrowski [PP93]. Other [Mat93, CRS99, Ioa88]. Outflow [Sem92]. Output [Whi97, ABT98, Moh99b, Wat89c]. outputs [VV85]. Overcoming [SR97]. Overhead [SW96]. overlapping [MLB99, WN91]. overlying [AAA99]. overrelaxation [GMS83]. overview [KB84a].

package [Ada91, WKM84]. Packing [Mus97a, Dro90, Mus97b, Mus97c]. Padé [Dav98, Kum92b, Waz99a, Waz99c]. Painlevé [VJ88]. pair [EJNT88]. paired [HQU88]. pairwise [JK98, Mat86]. paper [Ano99a, Di 98]. para [Gre87b]. Parabolic [CW92, CY98, Kor92, Lu98, SP93, CW91, CC99a, Da93c, Da93d, Da93b, Lin94, MB88b, Nak82, SV84]. paradigm [ESL98, Hog88]. paradox [CM90]. Parallel [BS97, Bai97, Béc98, BS93, CJ88, Cro93, Cro95, Dég96, EA94, GMS96, HGS88, HL91, Hoh97, KKPL96, Koł94, Koł95, MO93, Par89, SW96, SH92, SJK86, Axe86, Bai98, Bai99b, CL99, EM90, FH98, GS99, Ge89b, Ge89c, GS86, HZ98, KN89, KL99b, KM86, Kur99, TT88, Zhu94]. Parallel-Series [Koł94, Koł95]. Parallelizing [YCZ98]. Parameter [BY99, CSR88, CyGK99, CG98, Cic93a, Gho96, HHT98, JTT80, KA95, MMP89b, MSLB92, MB95, MKT96, Mur96, RAL97, SJ96, Sci97, SA94, Siv92, Bai81, Bee87, BH80, HW85, HC85, KS80, KH78, Mit86, MS89, MS92b, Udw91, Udw92a, van85]. Parameter-Dependent [MSLB92]. parameter-space [van85]. parameterization [DF96]. Parameters [ADS96, Fab96, GWL94, HSG94, JS96, KA96b, Lan92, PLPW96, Whi93, AAAN98, AK80, Cra79, Kee89, KR88, NC85, PM88, Pat89, Whi91, Zia89]. Parametric [VL83, BT98]. Parametrizations [HK96]. parametrized [NH86]. Parasitics [SP95]. Pareto [LSY98]. Paribus [Mat93]. Part [Arg96a, MS93c, MK93, RK87, Gol87, LV87, MS92b, Arg96b, Arg99d, Arg99e]. Partial [Ado97a, BKM94, BKM95, Béc98, BG94b, KW96, KW97, LL96, LWX92, SKG97, Tho82d, ZZZ98, Ada89, Ada91, Ada93b, AR82, Bee87, CH99, Da93b, DG87b, FS98, HEH81, JF91, KB84b, Mik87, MN79, Nak82, Sor82, TET99, WA98, Zha98c]. Partially [DPTT98, BM89]. particle [CG88b]. Particular [Gol96]. particulate [AH99a, HFA98]. Partition [SB92]. Partitioned [HW85]. Partitioning [SIA96, Cha91]. parts [Car90]. Pascal [Beh96a]. Pascal-like [Beh96a]. Passage [MFGW97]. Passage-Detector-Based [MFGW97]. patch [Mit84]. Patched [RL82]. patching [Hei93b]. patchy [Fah88]. paths [Mur91a]. Pattern [Gre97, Wu94, Mei89, YDB79]. Patterns [CV89, Mus97a, WZ95, ABES84, IRU83, MKJ87, Mus97b, ÖGDB98]. PDE [Fun91, KM94, War87]. PDEs [RGL97a, Vol92]. Peakons [Boy97b]. PEC [Cro80]. pellet [CK90]. Penalty [LHK97, RTJL92, Yak92a, Fei85]. Pendulum [KSZ^{+}95, Pet88]. Penrose [Wan87]. Perceptions [ITN98a]. percolation [BI98]. Perfect [AB94, AB95b]. Performance [Abu97, Ahm93, AS99, DS86, LMM86, Cic99, JB89a]. Period [GF96, Pet88]. Period-to-Period [GF96]. Periodic [AM98b, BKM95, CW92, Cha94, Gae87, GF96, HG99, HL88b, HL88a, KBS96, Kor92, LAK94, LWLL95, LG97, LZ96, LLS94, LLS96, OM99, OT92, PJG^{+}97, ŠNG92, Sim92, VS96, Wan81c, WJ92, CN90, CW91, Den88, Dua99b, Dul82, Eir89, LRV84, Liu94, MB88a, Mor90, Nie84, OT93, PF99, RC81, Sán80, Sau83, Sey81, TS98, Zha91]. Periodically [Fio95]. Periodicity [FW98, TT89, BJ85]. periods [WC99]. Permanence [FW98]. Perron [DDL93, Din98, DZ99, Har87, KST80, MS97]. Persistence [FM96, LL96, Zag91]. person [Lak81, SMO91]. Perturbation [BS93, JR93, JR95, SEG96, Sel94, Sem92, Wei99a, Cha96b, Gar89, KR87b, KR87c, KR89, Kam88, LO99, NR99a, NR99b, Red90a, Red90b, Red91, RR98, VC98, Waz99b, Wei99b]. Perturbations [ABT98, PN97a, Sor99]. Perturbative [Ado97b, SA96]. Perturbed [Arg96a, Arg96b, KR95a, KM95, Lop92, VB92, WL97, Arg99d, Bai91, DTW88, Gol88b, GP88, KR87d, KB93, LR85, LM98, MF99, NR98, Rol98, Arg99e]. Perturbing [RS88]. Petersburg [DH99b]. Petrov [mW98b]. Petviashvili [Ado96c, BG97]. Phase [Ham94b, Ham97, Kub97, Sim92, Sim94, Tuc97, YBR97, HF95, Lan81, Sim99b]. Phase-Lag [Sim92, Sim99b]. Phased [Bay93]. Phenomena [PA94, HC87, PV81]. photometric [Koz91]. photons [RR98]. Physical [Mus94, Sla91, Any82, BJ85, Mus84b]. Physical-based [Sla91]. physically [Mus80]. Physics [Con89, Mat93, Wer93, Ado87a]. physiology [AG89]. phytoplankton [van85]. phytoplankton-dynamics [van85]. Picard [AM81, LL95]. Piecewise [Beh92, Cha95a, RGL97b, AW87, Kop90, LG99, RGL97a, Ram98c]. Piecewise-Linearized [RGL97b]. Piedmont [Tur88]. Piezo [Bal94]. piezoelectric [Ade94]. pinning [Cos99]. Pipe [Mei93, Sca89]. pipes [BDJR98]. placement [Mil98]. plan [FGLF88]. Planar [EM98, Rya96, RC81]. Planck [Zie84, ZMV99]. Plane [KD95a, Knu93, PS96, AH99b, Ioa88]. planes [Baa88]. planetary [Tak85]. Planning [Van97, CAP88, KP91, MCD89, Rap92]. Plant [Bay93, RBR94]. Plants [OA97]. Plasma [AC97, CK90]. plasma-arc [CK90]. Plate [HM94, Had95, KWS95, RW93, Whi93, Whi88, Whi89, Whi91, XL98]. Plates [Bal94, Le98]. plays [Ada97]. PLOD [KSB89]. Plume [JFW96]. Point [Arg97b, GM94, Gho96, Gun94, Gup98, KBS96, Kum95, LMS92, LLY96, LYF97, Lop92, MFF92, SV94, Sor93, TM92, YL96, Zha96b, Arg99c, Cas89, EL77, Faz90, KR87d, KS93b, KK83b, ML90, MR99, NR98, NSE81, Pen84, Pes91, SW90a, Szi90, Tuc97, Udw91, UA88, VR99f, WS87, WKTM88, YO98, Zha98d]. Point-to-Point [Arg97b, Arg99c]. point-to-set [Szi90]. Points [FT94, JL93, LW94, Lop92, NS96, Sor94, Sor96, Wan96b, Arg99a, Di 98, HK84, KM79, KK83b, Moo81, RC81, Sor97a, Wat79, Waz99b]. Pointwise [AG89, Pic92, Eis87]. Poisson [CZ99, And87, BB83b, Bar88, BKS83, BKS84a, BKS84b, Hol86, Shi82, Zha96b, ZL98c, vGG79]. Polar [FH95, KH96, Bar88, KV84]. Polarization [Tak85]. policies [Cas83, Fis85, Gar90, Mos88b]. Policing [Mil95]. Policy [Mou96, PN97b]. pollutants [AKK99]. Pollution [GMS96, YCZ98, Tau80]. polygon [Sai88]. Polygonal [KWS95]. Polyhedral [Zha97c]. Polynomial [Beh92, CW89, Dég96, KA96a, MZ98, MS93b, Rug93, Wam92, AP87, He98, HV90, HL82, Jia99, MS99, MMP90, MS93a, MW88a, MW88b, Mor86a, Mor86b, MS87a, MS87b, MS89, MS92c, MS92b, VBH91, ZS99]. Polynomials [CS93, CS95, CS96, Hoh97, LS93a, MS93b, MM94, RS92, Sri94, dFG94, EMEH96, ESR99, GMGCL99, HZ98, KVT98, MB80, TCS92]. Population [BD90, GQS98, IKP97, Sla93b, ZA96, ZAES96, DS88, ES90, HK81, Kap80, Les89, LG99, MW98a, TET99, WK98, Waz99a]. Populations [PS94b, Fah88, KP98, LF94, PM89, Pat91, Zag91, ZK98]. poro [ZL98a]. poro-mechanical [ZL98a]. Poroelastic [Bar95]. Porous [Ezz94, HM94, Ham94b, HS96, LZ96, WZ98, YBR97, AH99a, AH99b, BH90, BH93, FH98, HB93, HF95, HFA98, Pat88, Ram98d, Zhu94]. portable [Ada89, WKM84]. posed [DH99a, Gol79, HK88]. Position [Bal94, KP96]. Position/Rate [Bal94]. Positive [LLX96, Sha94, AW99a, Har87, Hua91, KST80, MMP89a, Smi86, Udw92b]. positive-definite [Udw92b]. Positivity [LFS95]. positone [ASZ94]. possible [Gun92, Mos90]. postbifurcation [ST96]. Potential [CFT^{+}93, TJ95, Wer93, Cau83, Got84, Jou82, MS94b, WN91]. Potentials [Kin97]. Powell [ONOT98]. Power [Cha87, CS97, MZV96, SP95, BdA89, DHK90, JS98b, NA91]. Power-series [Cha87]. Practical [FLG92, FL94, Sha89b, SHR94, Zha98a]. pragmatism [Bur89]. prebiotic [Ten91]. Precipitation [MBG95, ÖGDB98]. Precisely [Mor95]. Preciseness [Yil96]. precision [Van98]. Preconditioned [GP97, Kwa97, GMS83, HV90]. preconditioner [JPS86]. Preconditioners [Jin97, JS98a, NLC95, Vas91, XL98, YK97b]. Preconditioning [CL98b, Hei93a, LC97, Dos90, Rug93]. Predator [Gra97]. Predator-Prey [Gra97]. Prediction [Kha96, Zhe93]. Predictive [Hab95, KB98]. Preface [DGP93]. Preference [ITN97]. preferences [Mat86]. preliminary [Har82]. prescribed [vGG79]. presence [JB89b, TW90]. Presentation [CS93]. preserve [FS90]. pressure [Bar91, Cha90a, MKJ87, STD98, VWSS94]. pressure-correction [Bar91]. preventive [Ree87]. Prey [Gra97]. Price [GP94, SMY95, CY99, Smi84b]. Prices [KPP93]. pricing [BS98, Váz98]. Prime [KCS90]. Principal [Baa88, Beh96b]. Principle [Arg97a, IA89, LO99, Mos90, Mur91b, Rav87]. Principles [Mur96]. prior [HW80, Sch91b]. Priori} [CH93, Jam85]. prisoner [NS89]. Probabilistic [Uen95, AK85, BBD91].

Probabilities [Szi98, Ben89, JCC98, Smi86]. Probability [Bur96, MS93b, MS93a, Smi84a]. Problem [Asa92b, Asa97b, BKM95, BPZ96, BS93, Cha96a, EN92, FT97, GP97, GWL94, Gup98, HZL97, IM92, JL93, McG98, Moh94, MB95, Pen93, Sas95b, SJ96, Sci97, Sel95, SM95, SHW97, Sim94, SK96, VS96, Wan95a, Wan96a, War94, Waz93, Waz95a, Waz95b, WCG97, YK97a, Abd99, AHKK80, AK80, AY98, Bow88, BL91, BKM79, BCZ89, CAP88, Cha90a, CXL94, DG87b, Fei87, FT98, Gar89, GMS83, Gla78, GZ84, Gup90, GK94, He96, HG99, HL88b, HL88a, IM80, IPP91, JMS98, KS84, KL85, KP91, KH87, Lak81, LM94, Lee98, Lin98, MS88, MB88a, Mit84, OU99, Pes91, RMR86, Sev99a, Sev99b, SW89, TETH98, Tag98, Tag99, Tak85, Tau80, Uen83, Uen91, Van95, Vis80, Wan99a, Wan99b, Waz99b, WW99a, WW99b, XL98, Ye98a, YS98a, YS98b]. Problems [Asa92a, Avg94, Bel95, Cab94, CN97, CW92, CH96, Cro95, Eis98, EFS95, Fab96, Gar98, GM94, HHZ97, JR93, JR95, JNM92, KR95a, KM95, KTZK96, KBS96, KR95b, KS92a, Koč92, LMS92, LMSS97, LM95, Lia97, LLY96, Lin96, Lu98, Mal93, MSLB92, Mo97, MFF92, Net89, OT92, PA94, PS96, Ram97a, Ram97b, RGL97b, RTJL92, SKB95, SZ97, SV94, ŠNG92, Sel94, Sem92, SS94, Sim92, SK95, Ste94, TFM93, TM92, VD97, WJ92, YL96, AG85, Aga86, AW99b, ABW99, AD98, AB87, AG99, Als79, ASZ94, AG89, AT82, Bai91, Bai98, Bai81, BM89, BF83, Bre81, BJ85, CN90, Cab95, Cao88, Cas89, Cas87, Cas94, CW91, CC99a, CK86, CL98a, DDS85, Den83, Den88, DTW88, DF96, DSK82, EL90, Faz90, Fei85, FS90, GS94a, Ge91, GS86, Gol77]. problems [GL99, Guo99, HSZ94, Hau82, Hei93b, HK88, HK89, KR84, KR87b, KR87a, KR87c, KR87d, KR89, KB93, Kag91b, KS84, Kam88, KS93b, KW86, KT89, KH78, KM79, LP81, LRV84, LO99, LM98, MM87a, Man86, Man88, Mär89, Mey78, Mit86, Moh99a, ML90, MR99, NR98, NR99a, NR99b, NJC99, NSE81, Oma86, OT93, ORSS83, Pan88, Pen84, PZ99, QA98, Red90a, Red90b, Red91, Sco80, Sco81, SV83, She90, Smi84b, SV92a, SV92b, SSY99, TLG90, TT88, TS98, UA88, VC98, Vas91, VV85, VR99f, WS87, Wei89, WKTM88, WA99, Yag82, Yan87, YK97b, ZL98a, Zha98d, Zha99b]. procedural [Mad90]. Procedure [EGH97, FBSS96, Sel95, Siv92, Yip93, BTW94, CG88a, EGH98, GG91a, MMC98, TW79, Yip92]. Procedures [Whi97, KL85, Moh99b, MW88b, TT86]. proceedings [Tho82d]. Process [IKU96, TM92, YBR97, AAAN98, EM79, Har87, MN79, SV85, SL98, TyG99, WH87, Wat89c]. Processes [CC94, GSK93, Hab95, LS95, Hir84, LP81, PM89, Rol98, Vis80, WH87, vGG79]. Processing [LSKH96, SW96, CJ88]. processors [GS86]. Producer [SM94, SMY95]. Producer-Consumer [SM94, SMY95]. Product [Cai93, Sas95b, Ado81, CS99, McG79, Mit92, WI80]. product-type [McG79]. production [Rol98, SS87]. Products [CMM96, MTP93, MFF92, BE85]. profile [Kru81, QC98]. profiles [Uen80]. Profitable [PMMA97]. Program [Cha92b, PZ94, VA88, Zhe93, Cha92c, Cha92d, Cha92e, KT83]. Program-Size [Cha92b, Cha92c, Cha92d, Cha92e]. Programming [BGY97, EFS95, Hoj94, LMSS97, LLY96, LYF97, MPD97, Pap96, PN97b, PS96, TDK94, AB87, EB91, Flå92, GH89, Ge91, GD87, Han98, HKKT79, KKM90, PVM88, PK98, SGB87, Zhu98]. Programs [Pap92, Yak92a]. progress [KH87]. project [CDW91]. Projection [Cro93, Cro95, Han98, PS96, Baa88, CL98b, Gol89, GB90, GB98, MB88a, Wei99b]. projection-iterative [MB88a]. Projections [Kha94]. Projective [Yan94].

Projector [Dos90]. projectors [DLRJ81]. Proof [Gun94, SH92, UK95, KRTW83, Zha97b]. Propagation [AA94, AAA99, AA99c, Ado97e, ENS96, AA99a, Ado87b, Ahm99, Ber89, OŞL80]. propeller [Jou82]. proper [Sor99]. Properties [Fio95, MM94, PSMM93, Sas95a, SZ97, EM79, Fei87, FS90, pLR87a, RUS^{+}84, Zha91, Zha99a]. Property [NG94, RS92, Jac90]. proposal [MNK99]. prospection [Sev99b]. protection [Ree87, Zam88]. proton [zHhT99]. prototype [Zam88]. Pruning [LSS94]. Pseudo [Hai97, SY98]. pseudo-inverse [SY98]. Pseudo-Inversion [Hai97]. Pseudospectral [Fun91, SK96, yHyG99]. pseudospectral-finite [yHyG99]. Pták [AC94]. Pulse [HZL97]. Pulse-Spectrum [HZL97]. pure [Lee98]. purpose [Mae89, Söd89]. pursuit [Baa88]. Putting [PZ94].

Q [HDP85]. Q-analysis [HDP85]. Quadratic [Wan95b]. Quadratically [SV92a]. Quadrature [CMM96, MW96, Ioa88, Let80, MS92a, Mil98]. quadratures [McG79]. Quadrilateral [Wan94]. Qualitative [GP94, LL94a]. Quality [HYL97, SPD93, Bog91, KK82]. Quantitative [Han92]. Quantitatively [Tuc95]. quantum [Mus80]. Quartic [SV94]. Quasi [JM93, Raj97, SL92, SSY99, Cao98, Dua99a, Kaz99, Sun98, SL99]. quasi-geostrophic [Dua99a, Kaz99]. quasi-minimal [Cao98]. Quasi-Newton [SSY99, SL99]. quasi-Newton-SQP [Sun98]. Quasi-solutions [SL92]. Quasi-steady [Raj97]. Quasi-Three-Dimensional [JM93]. Quasiconcave [AAHM93]. Quasiconcave/Quasiconvex [AAHM93]. Quasiconvex [AAHM93]. quasifixed [Moo81]. quasigeostrophic [Dua99b].

Quasilinear [JR95, WA96, Bee87, Pan88, Tho82a]. Quasilinearization [VS96, JV98, WC88, HKKN78]. Quasisteady [Raj92]. Quaternions [Car97, Car88a]. questions [CW84]. Queue [Jac95, MS95b, MFGW97, MES92]. Queueing [Mil95, Mou95, Sos88, Sos91]. Queues [Mou96, MES94]. Quick [SC96, CS98]. Quickly [BD96]. Quintic [Ste94, AP87]. quite [IP89]. Quotient [Har97].

racemic [Ten91]. Radial [HH99, PJ95, Wu97, Cas94, FS98, Li98]. radiance [Uen80]. Radiation [DGU96, HM94, Rap92, SK95, CAP88, OL80b]. radiation-therapy [CAP88]. Radiative [MA95, KKS87, Tak85]. radio [WW99a, WW99b]. radiotherapy [FGLF88]. railgun [CK90]. railway [JS98b]. rainfall [Sor85]. rainfall-runoff [Sor85]. Randles [LE85, Mac93]. Randles-Sevcik [LE85, Mac93]. Random [Ado96d, Ado97e, Ben92, Gho96, LS93a, PY94, SU90a, SU90b, WP97, BR99a, BI98, Jer88, MBS80, Mos88a, SW91, VV85]. Randomly [How94]. range [Gla78, ZS98]. Rank [JY97, Kai91]. ranking [JK98, MS99]. Rapid [CN97, VCM91]. Rarefied [Uen95]. Rate [Bal94, Cic93b, LK95, Lon87b]. Rates [Cha96a, CB94b, JN95, MS95b, Arg99a, GB98]. Rather [Mat93]. rating [MS99]. Rational [Jód92, Let97, PZ94, Ser79, CW89, Ge89c, O’N90, PVM88]. Rationale [GPAB83]. rationality [Tes80]. Rationalized [RJY80]. ratios [MB80]. Raviart [CL98b]. Rayleigh [AA99c]. Reach [ITN98b, Ino99]. reachability [CW84]. Reacting [MPS97]. Reaction [Ado97d, Gar98, Ram97a, Ram97b, Yee93, Hos99, JV98, Mei89, Pac85, Ram98a, Ram99b, mW98b]. Reaction-Convection-Diffusion [Yee93].

Reaction-Diffusion [Ado97d, Ram97a, Ram97b, Mei89, Pac85, Ram98a, mW98b]. reactive [AH99a]. Reactor [Sel95, Zas93, HEH81, NR99a, Shi89]. reactors [CK90, Faz91]. Real [GW89a, Hai97, LLX96, LUL97, Sha94, Wan93, DLR85, HW85, Mv98, MW88a, PVM88, Wan81a, Wan81c, Wan81b]. Real-Time [LLX96, LUL97, Wan93, GW89a, HW85]. real-valued [Mv98]. Realistic [AB94]. Realization [LS93b]. Reallocation [RPT97]. Reanalysis [CW97]. Recapture [Yip93]. reciprocal [Har87, Var83]. reclamation [GGR91]. recognition [Wu94]. Recombination [SW96]. Reconciliation [Vev92]. Reconstruction [FG94b, FG95, Gzy95, Gzy97, Kha94]. reconstructions [Li94]. Recourse [Yak92a]. Recourse-Constrained [Yak92a]. recovery [VJ86]. Rectification [Reg94, Reg96, Abu99a]. recurrence [TS87]. Recurrences [SW96]. Recurrent [Wan93, Ino99, SY98]. Recursive [Hai97, NAM93, YW85, SW79b, WW99b]. Red [Zha96b, KL89]. Red-Black [Zha96b, KL89]. Reduced [BH89, WSL99]. reducible [MS95a]. reducing [EM90]. Reduction [Jac93, KKH95, SKB95, Sen93, LO99, Lia98]. Reductionism [Con93a]. redundant [Udw81]. Rees [Pap84]. Refined [HM99]. Refinement [CW97, MB82]. Reflection [MA95, NSV92]. Reflectionless [Kin97]. reflector [Uen89]. regard [BD90]. regeneration [BD90]. regimen [NHG83]. Region [dYS97, Khu98a, RMR86, SSY99]. regional [SGB87, TETH98]. regionalized [CS85]. regions [CG88a, Col82, DS88, Fis85, Har82, KB85, SW91]. Regression [AB95a, BBB97, Kha96, Kum92a, KA96a, Pat87, Pat89, Pat91, TA93, AK80, Fra89, HW80, KB98, TZ83, VS90]. Regression-Based [Kha96]. Regular [JNM92, JL93, JkC97, NN87, KR87a]. Regularization [dFG94, DH99a, PZ99]. Regularized [Whi97]. regulation [YDB79]. Regulations [LHK97]. Regulator [TFM93, Sug87]. regulators [KS81]. Reid [ZED91]. Related [CS93, CS95, BCZ89, Sta99, Udw92b, Wan80]. Relation [GSI96]. Relations [CS95, SAL96, Sri94, Beh96a, Mur89b, TS87]. relative [Ahm87, DS87, DS88, Sco80, Sco81]. Relativistic [TPS98]. Relaxation [BS97, Son96, Zha96a, Bai98]. relaxed [CL99, CK99]. release [VV85]. Reliability [Ami97, AB94, Koł94, Koł95, NA91, PL93, AB95b, Kur99, Udw81]. Reliability-based [PL93]. Reliable [WK96a, Waz98b, Waz99d]. relief [VWSS94]. Remark [Shi89, Dua99a]. Remarks [Cab95, Kai91, OT92, OT93, Par87]. Remote [Cro95, Fym79]. remote-sensing [Fym79]. Remotely [KP96]. renewal [Pat90, Pat92]. repair [Cas88]. Repeated [RAL95]. replacement [Ree87]. Representation [CS96, Gol95, BJ85, ES90, Mor90, Sug87]. Representations [Sla93b, Smi86, Sri98, Sta99]. representative [KR81b]. reproducible [Mei89]. Research [DGU96]. reservoir [EB91, VJ86, Zie84]. resetting [OS80]. Residual [RBPK96, Zha97a, Zha97b, Zha98a, Cao98, Zha98b, Zha99a]. Residual-Finite [RBPK96]. Residuals [CLC97]. Residue [Cen97, EHE93, EE89, Eva90a, Eva90b, Eva90c, ER94b, ER94a]. Resolution [FG94b, FG95, GP97, Lap92, Sch92, Whi97, Ada93b, Dav94, Kha91]. Resolution-Limited [Lap92]. resolvents [Gol77]. resonance [GK94, SV83, SV84]. resource [CW84, Les89, SGB87, ZK98]. Resources [PL93, TDK94]. Respect [YLS93]. respects [MS87b]. Response [AA96, EB99, Udw87, JC97, VB89]. restoration [Cha89a]. Restricted [Che97]. restrictions [MS88]. result [Gol88a, LS99]. resultants [WM87]. resulting [WW99a, WW99b]. Results [Arg96a, Arg96b, Arg99d, Arg99e, Bla96, Car97, Gar98, Le98, dFG94, HQU88, HSZ94, Rav87, SV83, SV84, Udw92b, VV80, ZL90]. resumming [Lan89]. resuspension [Bec85]. Retarded [EN94, AY98, LS78, PF99]. Retrieval [KUO95]. revenue [Ree87]. revenue-earning [Ree87]. Reverse [NG94]. Reversible [OMS95, LE85]. reversion [Cha87]. Review [Ham94b, AR90a, AR90b, MU90a, MU90b]. Revisited [KUX95, Ahm87, Son93, VC98]. Revolution [Wer93, BK99]. Reynolds [Ram99a]. Riccati [BdA89, Die89, JN91, Jód92, JN95, LAK94, LP97, MRJ95, Sán80, Sas83, Sas85, Sas88, WH84]. Richtmyer [HKLS93]. Ridge [Cic92, Cic93a, Cic93b, HW80]. Ridge-Type [Cic92, Cic93a, Cic93b]. Riemann [Ayo93, IPP91, LS98b]. Riemann-Green [Ayo93]. Riesz [Kan99]. rigidity [Whi88]. Rio [BDS96]. Rise [PMMA97]. risk [BBD91, Got84]. river [Bog91]. RL [LCE85]. RLS [LL94b]. Robert [Tes91b, Tes91c]. Robinson [BK80]. robot [IFU98]. Robust [Che95, vKH95, LL96, RL96, TA93]. Robustness [TW79]. root [Har87, HZ98, KST80, PP93]. root-finding [HZ98, PP93]. roots [DLR85, IP89, Mil87a, Sav93, ZG90]. Ross [FG94a]. rotating [HH85, Jou82]. Rotation [BNN95, ENAAA95, Pav95, AAAaZ99a, AaAaZ00a, VCM91]. Routing [MW93, Cha88b, KL85]. Row [LSS94]. Row-Pruning [LSS94]. Rule [Yeh97]. Rules [Con93a, DH99b, Ioa88]. run [Zha89]. Runge [Abd92, And93, EJNT88, HS94, PF99, SW79a, Sim97, SIA96, Vos89]. runoff [Sor85]. Rural [Van97]. s [Ano99a]. S.N [Jia99]. S810 [LMM86]. S810/20 [LMM86]. saddle [Zha98d]. safety [KK84, VWSS94]. Saffman [BH93, HB93]. Salesman [War94, Dro90, Fei87, War91]. Salinization [PMMA97]. Sample [Bha94, BP92, CC94, GFM88, Gil89, MN79]. samples [Fie84]. Sampling [BU77, Cic93b, Jam86, MGL92, PSMM93, RBR94, AB87, CY80]. Satellite [KUO95, SST95, VLP95, HH85]. Satisfied [Mur96]. Saturated [Bar95]. Saturation [YBR97]. Scalar [KM95, JBV98, Oma86]. scale [BW88, Ge89b, LR85, ÖGDB98, Sen86, She90]. Scaled [KM95]. scales [ABW99]. scaling [Zha97b, Zha98a]. scan [HT98]. Scattered [GSC95]. scatterers [Sha87]. Scattering [Ado96b, KD95a, KD95b, Lan92, LM95, Uen89]. Scheme [CS92, CW97, HKLS93, Jol92, LR95, Lop92, MEZI93, AT82, Bar89, DG87b, FPGV95, Ge89b, HM98, JS98b, KCS90, Pan88, Pat88, Sim99b, Szi89, Tom80, VJ88, Vul90]. Schemes [FHG96a, Jim94, MW96, Sel94, And87, CL99, DPTT98, FV91, HL91, JV90, SSV89, SV92a, VC98]. Schmidt [TM92]. Schrödinger [BK99, FPGV95, He96, Khu98b, Sim99a, Sim99b]. Schrödinger-type [He96]. Schroeder [Bay93]. Schroeder-Phased [Bay93]. Schwarz [ADHE82, CG88a, KR88, YK97b, ZZ98]. sciences [Sor85]. Screening [Kim96, MDR94]. Search [Bad93, Hoh97, Cha91, Jer88, MLC87, MBS80, Par89]. Searching [Hol89, IA80]. secant [Arg90]. Second [Cha96a, ESS94, ES94, Ela94, FT94, GG96, GGS97, GM94, GL92, Gup98, JL93, JV94, JPF95, KM95, LG94, Lop92, NS96, NN87, OT92, SJ96, Sim92, Sim94, TPL95, WK96a, WA96, Yan97, AG99, CN90, ES93b, GZ84, HL88a, KS93b, LRV84, Oma86, OT93, PGHX99, Sas85, Ser79, SV83, Sol98, Van95, VR99c, Ye98a, ZA91, dOP89]. Second- [OT92, OT93]. Second-Kind [WK96a]. Second-Order [GG96, GGS97, GM94, JL93, JV94, JPF95, KM95, NS96, Sim92, Sim94, TPL95, WA96, ES93b, GZ84, Ser79, Van95]. section [Bog91]. sectional [Cha91]. Sectionally [Ioa92]. Sedenions [Car97, Car88a]. sediment [Bec85]. Seek [CH92, Hoj94]. segmented [Col82]. Seidel [KS78, KL99b, KL89, Zha96b]. Selection [HC85, CDW91, EB86, Fra89, TW90]. Self [BG94a, JL97, KS92b, MSLB92, Shi93, ZZZ98, NR99b]. Self-Adaptive [ZZZ98]. Self-Adjoint [MSLB92, NR99b]. Self-Creation [Shi93]. Self-Evaluation [BG94a]. Self-Organizing [JL97]. Self-Sustained [KS92b]. Selftuning [vKH95]. Semi [Abd92, HM94, WJ92, AAA99, He96, LCE85, RC91, PM89]. semi-discrete [He96]. Semi-Infinite [HM94, Abd92, AAA99, RC91]. semi-integrals [LCE85]. semi-Markov [PM89]. Semi-Periodic [WJ92]. Semiclassical [Gre87b]. semiconductor [ALZ99]. semidefinite [Han98, Sas83]. semiderivatives [LCE85]. Semidirect [Roa82]. Semidirect/marching [Roa82]. Semidiscrete [Jac92, Jac90]. semigroups [Opp92]. Semilinear [CW92, CY98, CW91, CZ99, CC99a]. semimartingale [TT90a, TT90b]. Semipositone [KS92a]. sensing [Fym79]. sensitivities [Tuc97]. Sensitivity [AAES99, Cic93b, CB94a, CB94b, CO96, De 94, EBR85, Eno90, KS80, Pap94, PM88, TH96, AHKK80, Tes91b, Var83]. sensor [Mil98]. Sensors [Bal94]. Separable [KKH95, Ge89c]. Separable-Denominator [KKH95]. separable-factorable [Ge89c]. Separated [TSK98, And93]. separation [HF95, Mur91b]. sequence [Wan81b]. Sequences [Fio95, Hir85, Zha99b]. Sequential [BR99a, Cro95]. Series [Boy95, Boy97a, BG94b, CS97, EHE93, KWS95, Koł94, Koł95, LSKH96, RAL95, RAL97, SS98a, Waz97a, ZB97, Ami99, AT85, BCG91, BJ85, Car88b, Cha87, Cha89b, CRS99, FMR86, GW89b, Kin94, Kum92b, Kur99, Lam99, LS81, Lon87a, Lon87b, Lon87c, Lon88, MMC98, Pal86, RRA90, Sri98, Waz98a, Wu94, WC99, YF91]. Series-Parallel [Koł94, Koł95]. series-solution [BJ85]. Server [MS95b, Mou96]. Servers [MES92]. Service [MS95b, Pat87, PM88]. services [Pat92]. Servomechanism [Rya96]. Set [Bad93, Bay93, Cro95, Ibr96, Lin96, PZ97, Sch91a, Sha97, Ahm87, BBD91, Kin91, Nie85, Pig87, Szi90]. Set-spectrum [Sch91a]. Set-Valued [Ibr96]. Sets [HK96, RK93, Yan94, Zha97c, EL77, Kha81, KB89, Sos88]. Sevcik [LE85, Mac93]. Several [AB94, SA94, Gol87, Hua99, IM80, Sca89]. sewer [BDJR98]. sex [HK81]. Sexually [MAS95]. Shading [Koz95]. shallow [HF93, Pap84, van85]. shallow-lake [van85]. shallow-water [Pap84]. Shape [Bal94, Koz95, KR81b]. shaped [ASZ94, CG88a]. Shapes [Jan95, Har82, WZ82]. share [PZZy99]. shares [ZPY99]. Sharp [AP95, Gyö90]. Sharpe [MW98a]. Sharpe-Lotka [MW98a]. shells [HF93]. Shift [DGV94]. Shift-Invariant [DGV94]. Ship [BD95, Zam88]. Shock [Dav98, HK89, Zak89]. Shocks [HLW94]. Shooting [GM94, Sel95, WJ92, Bre81, JBV98, KN89]. Short [Zha89, TT88]. Short-run [Zha89]. shortage [Sos88, Sos91]. Shortest [WLB97, Wai90]. shuffle [Lam99]. sickle [MH81]. sided [MNK99, Tag98]. Sidi [GW89b]. Sigmoid [WR96]. Sign [LP97, DLR81, Gup90, RR99]. signal [Abu99a]. signals [Abu99b]. significance [HB91]. similarity [Sai96]. Simple [HH96a, vKH95, WZ98, Abu99b, IA85, JA99, KKM90, Lui99, Pet88, Yip92].

SIMPLER [Bar91]. Simplest [KS84]. Simplexes [JkC97]. Simplicial [JCC98]. SIMPR [Sha89b]. Simulating [HY99a, Zhu94]. Simulation [Ade94, DGU96, Gre97, GMS96, JT80, Raj92, Raj94, Raj97, TH95, TPS98, ALMS90, BW88, BR99a, Cas81, Cos99, FPGV95, HFA98, JS98b, QC98, SK98a, Tur88, WN91, Zam88]. Simulations [MAS95, MPS97, RBR94, ALZ99, LF94, ZS98]. Simulator [YCZ98]. simultaneous [Li98, MW88a, MW88b, VS90, ZS99]. simultaneous-equations [VS90]. Sinc [LB96, MLB99, SS86, MSLB92]. sine [Eva90c, BT98, FV91, GPRV86]. sine-Gordon [BT98, FV91, GPRV86]. Single [Ham94b, HF95, MS95b, Mou96, Li94, RS86, ZK98]. single-equation [RS86]. Single-Phase [Ham94b, HF95]. singlets [Gre87b]. Singular [BS93, CW92, EN92, GM93, Gro93, GWL94, JR93, JR95, JNM92, JL93, LUL97, RR98, RB95, Sel94, Sem92, Sen93, WK96a, Arg99a, AG89, Boy89, CW91, CS88, Da93c, Da93d, DEV95, Den88, Gar89, Gol88b, Gol89, GB90, Ioa83, Kad86, KR87b, KR87a, KR87c, KR89, Kam88, LV87, LM94, LO99, NR99a, NR99b, RBA92, Red90a, Red90b, Red91, SR90, TT84, VC98, Waz99b, Wei98, Wei99a, Wei89]. singularities [Boy89]. Singularity [AC97, Cas87, Ioa92, KR87a, KN90, Kle82, NSE81]. Singularly [KR95a, KM95, Lop92, VB92, GP88, KR87d, KB93, LR85, NR98]. singularly-perturbed [KB93]. sink [BH93]. Sinusoid [Reg96]. Sinusoids [Reg94]. Size [Cha92b, Cha92c, Cha92d, Cha92e, GFM88, Gil89, WK98]. size-structured [WK98]. Skan [Asa97a, Asa98]. skein [GMGCL99]. skein-template [GMGCL99]. skill [ESL98]. slackness [Moh90]. slave [Ada98]. Slip [Ham92]. Sloan [Gol93]. Slope [SW92b, SW90b]. Slotted [Sch92]. slow [Mur91b]. slow-fast [Mur91b]. Slowly [SS94]. Small [Bha94, KA95, Cha96b]. Smallest [BG94c]. Smooth [Da93c, Bow78, Da93b, Sol99b, SSY99, Wat89c]. Smoothed [Sla93b]. Smoothers [SO95, Axe86, MM87a]. Smoothing [CG96, HSA97, Sla93a, Zha97a, Cao98, CK79, HH99, Jac90, KL89, Man86, Mey78, TT86, YO98, Zha98b, Zha99a]. soap [CG88b]. Sobolev [Gu99, VV80, Vau80]. social [DML98]. Software [MTP93, Ada89, Ada91, Ada93b, Die89, DF96]. Soils [Bar95]. solar [zHhT99, OL80b, TT89, Uen80]. Solid [Abd92, McW82, CK90]. solid-hydrogen-pellet [CK90]. solidification [Fre90]. Solids [LK95]. Solitons [Wer93]. Solution [ARM94, Ado97b, AK96, ADHE82, Asa97a, Béc98, Cha90b, Cha94, EHE93, Fab96, Gro93, Hau82, HLXZ97, Jam83, JNM92, KKPL96, KW96, KW97, KSZ^{+}95, MW98a, MES92, MFF92, NFJ94, NN87, Net89, OU99, RC91, SS94, SKG97, Sha97, SA96, Sim97, Siv92, SK96, Tho82d, TSK98, WHF89, Waz93, Waz95a, Waz97a, YST97, Yee93, YK97a, Ada89, AS81, AR98a, AG99, AP87, AR82, AG89, Bai81, BF83, Ber89, Bow88, Bra78, BT98, Bre81, BJ85, CAP88, Cha88a, CCSK82, CW89, Cro80, DEV95, DF96, EJNT88, Eva90b, Fei87, Gla78, GPRV86, Ioa83, JK98, JN91, JF91, JNF95, JPS86, KR84, KB84c, Kad86, KR87b, KR87c, KS84, KN90, KN79, Khu98c, Kin93, Kin94, KH78, KV84, MMP89a, MX90a, MX90b, Mey78, Mik87, MES94, NR99b, Nie85]. solution [OL80a, Ohm98, Sai99, Sas83, Sca89, SR90, Sha87, Sim99a, Sim99b, Smi84b, Ste82, SK98b, SL98, Szi99, Tau80, TT84, Uen83, UA88, WW99b, WKTM88, ZZ98, Zha91, ZL98a, ZMV99]. Solution-Adaptive [TSK98]. Solutions [AS94b, AR93, Ado97a, BD96, BG97, Boy97b, CW92, CY98, DVD95, ESS94, ES94, Ela94, Gol96, HHZ97, Jód92, JV94, JPF95, JN95, Kan97, Kar98, Kim96, Lam86, LWLL95, LWX92, LLS94, LLS96, Lu98, OT92, PJG^{+}97, RRA90, RK93, SW96, She96, Sim97, SV92b, TM92, Vau80, WA96, XZW96, Zag92, AR87, Ado88, AW99b, ASK98, Ben87, CN90, Cas94, CW91, CZ99, CC99a, DPTT98, DSK82, Eir89, Ge90b, GS94b, Ham98a, HK89, Kel81, LS87, LS91a, Lin94, Liu94, LS98b, MB88b, Mor86b, MS87a, MS92c, Mur89a, Mus84b, NJC99, Nie84, OM99, Oma86, OT93, PF99, PGHX99, QA98, RSS89, Roa82, Sai96, Sán80, Sau83, Sey81, Sha89a, She90, SL92, Waz98a, WH84, WA98, Yan99]. solvability [GK94]. Solve [Mor94, Mor95, Die89, Pap84, ZL98b]. solvents [HD88, TCS92]. solver [BB83b, Bar88, Bar89, Hol86, KSB89, Mae89, Mer86, NW90, RS86, Tha82]. solvers [Enr89, HKWdZ83, KB84a, Van83]. Solving [Abd93, BPZ96, Cha89b, Che97, CH99, FS98, Ge91, JR93, JR95, JL93, KT91, LMSS97, LP97, MBS98, NP92, PS94a, SV94, SJ96, Sel94, Sel95, SS98b, SBH96, WS87, WSL83, WK96a, WK96b, ZZZ98, bS89, AR82, BW77, Cas89, Din98, Fei85, GS99, GMS83, GT91, Gol79, Gol87, HV90, IST93, JBV98, KB85, KS93b, MM87b, MB88a, Mit90b, Mor86a, MS87b, NR98, Pes91, RC99, Red90a, Red91, SK86, SR88, VBH91, WK98, Waz98b, Waz99c, ZH93, Zha98c]. Some [Abu97, AH99b, Bel95, BB96, CW84, CSY98, CRS99, Die89, Fei87, FS94, FC93, GS97, HQU88, HSZ94, HH96a, IKP97, Koł94, Koł95, KB80, KH87, Le98, Lia97, MM94, MSZ96, Mus84b, Par95, Per96, SS98a, SM88, Sim94, SS97, Udw92b, Unb96, Vul90, ZS99, AT80, Bee87, Cas89, DPTT98, EORA99, GS86, Gol79, GB90, GB98, IST93, KVT98, KP85, Kur99, Mad90, MW88b, Sau83, TT90a, TT90b, WH84]. Soriano [Di 98]. Source [AA95, AA99b, ZZ98]. sources [AKK99]. Southern [PN97b]. Space [Arg96a, Ezz94, Ibr96, SK95, dL95, AA99c, Ahm87, Arg99d, Cic99, Fei87, Guo99, LS78, Liu94, OL80b, PV81, Tom80, Tuc97, van85]. space-based [Cic99]. Spaced [Koh92]. Spaces [Arg96b, Cha95b, CAQ93, Cro93, Arg90, AC94, Arg99e, Kha81, KCS90, Vál88]. spacing [Spi90]. span [Hof88]. Spanning [Moh94]. Sparse [Gol95, Bai98, HC85]. Spatial [SK95, ABES84, BB83a, BR99b, DS87, Puu84, Smi84b, Tur88]. Spatially [Den90]. Special [DGP93, Net89, SS98b, Sim94, SU95, SP94, UTL^{+}95, UHL96, War94, AaAaZ00b, Ham98a]. specialist [PZZy99]. Specializing [XZL97]. Species [SA96]. specification [KP86]. Specifications [NAM93, EB91, Sug83]. Specified [Sas95a, Har82]. spectra [Gre87b]. Spectral [DVD95, DLR81, nHyG97, Hei93a, Hei93b, RCG94, CyGK99, CC99b, Hei91, KS93a, Opp92, TCS92, ZH86]. Spectral-Finite [nHyG97]. Spectrophotometry [SK95]. Spectrum [HZL97, Abu99b, Sch91a]. Speed [SP95, Yse86]. SPH [HSA97]. sphere [Mus97a, Mus97b, Mus97c]. sphere-packing [Mus97c]. Spherical [IM92, AAAaZ99b, HF93, RMR86]. spherically [AAAaZ99b, AaAaZ00b]. spills [BR99b]. Spline [Beh96b, BG93, Ona97, Ste94, Wan92, Wan95b, AM98a, Beh89, Beh90, Beh96a, Ham94a, KB93, Ram90, Rob82, SW79b, WS87, Zha91]. spline-collocation [WS87]. Spline-on-Spline [Beh96b]. Splines [Beh95, SV94, HJ92, HH99]. Splitting [IKP97, NLC95, RGL98, Wei98]. splittings [Dav94]. spread [Gle88, Wat89b]. Spreading [CFT^{+}93]. sprinkler [MN91]. SQP [Sun98]. Square [LYX^{+}93, TM92, Ahm79, CCSK82, Ye98a]. squared [LMV91]. squared-magnitude [LMV91]. Squares [AB95a, Cha96a, Whi97, Ber89, CXL94, HW80, KRT83, SUJ98, Sha89a, Sug83, SSY99, Uen83, Yan98, YL98a, Ye98b]. squeezed [BKS98]. St. [DH99b]. Stability [Ahm87, AR98b, Bel95, Bil96, Bla96, Cha85, EN94, FLG92, FL94, GQS98, GF96, HH96b, HH96a, HH97, ITN98a, JNF96, Kim96, Kum95, LS95, LL89, LL96, Lop92, Mor98, PN97a, Rua92, SSW99, Sol98, SBH96, Ten90, Tom80, Tuc95, VLP95, VB92, Zag91, ZL90, Ahm79, AB89, ER88, FLZ91, FG81, FL99, GS94a, Gar89, GW89a, Gri89, IM80, JNF95, KS93b, KB89, LR85, LS99, LP89, LS91b, LG99, Mat89, Mik87, RS88, Sol99a, Sos88, Vis80, WKM84, Yan98]. stabilization [Lia98, Mat89, Zha99a]. Stabilizes [HSA97]. Stabilizing [LSVS94, TFM93, HH85]. Stable [GGS97, Lar97, Sim92, BKS98, CV89, Zho89]. Stackelberg [SMO91]. Stage [GR97, Kha96, ZA96, Bai99b, SSW99]. Stage-Structured [ZA96]. stages [Ass88]. stamps [Abd99]. Standard [BB96, WR96, AD98, TT88, VC98, Yan98]. Standardized [BD94]. start [VBH91]. starting [KK83b]. State [Bar95, BL87, Ezz94, GP97, Kim96, MS93c, MK93, Siv92, Yam95, ZA96, BKS98, EB91, HW85, NC85, PM89, Pat90]. State-Dependent [ZA96]. state-variable [Pat90]. states [Gre87b]. Static [KTW81, Sha97]. Stationary [Sug95, Han92, JCC98, MS94a, Sey81]. Statistical [Bay93, MP86, PM89, Yip93, WC99, Hir84]. Statistics [CH93, JY97, Ben87]. Steady [Bar95, BJTW98, GP97, Kim96, Siv92, AH99b, BH90, CL98a, NC85, Raj97]. steady- [NC85]. Steady-State [Bar95, GP97, Siv92]. Stefan [Asa92a, Asa97b, Mey78, RMR86, YS98a]. stencil [Li94]. stencils [YO98]. Step [Arg97b, Asa97b, RTJL92, Sim92, Sim94, Arg99c, GW89a, Man86, SSV89, Vos89, Zho89]. Stepping [Siv92, EM90, Moh99c]. stepsize [And93, Zla89]. stereo [Koz91]. Stieltjes [FT97, FT98]. Stiff [MRJ95, BH89, Cha89b, CW89, Sha89b].

Stiffness [AST86, Bur89]. Stochastic [AR90a, AR90b, Cha95b, CS96, FLG92, FL94, MBCV95, MU90a, MU90b, MS97, Moh94, Nam90a, Nam90b, Sha90a, Sha90b, Sug95, Uen89, Uen91, Yak92a, AM81, AS81, Ado81, Ami99, Cha85, CK86, Cha96b, EB91, GD87, HT90a, HT90b, HW80, Heu90a, Heu90b, JTT80, JT80, JCC98, JPS86, OU99, PM88, Pat89, Pat91, PS94b, SL98, TT90a, TT90b, Zie84]. stochastic-dynamic-programming [EB91]. Stock [GP94, PBL92, Wei84]. stock-flow [Wei84]. Stock-Holding [PBL92]. stocking [Fis85]. stocks [Den90, Puu84]. Stokes [Cha90a, Cha95a, Cha96b, CC99b, Gno82, Ham98a, nHyG97, yHyG99, KW97, Khu98a, LL95, LY98, NW90, Shy87, SJK86, VWSS94, Ye98b, Ye99, ZH86]. Stopping [DH99b]. Storage [FS94, VV85, BH89, ORSS83]. straightforward [WH87]. Strain [KD95a]. Strang [Wu97]. Strang-Fix [Wu97]. Strange [DVD95]. Strategies [BW88, CV89, DG87a, KP86, Mos88a]. strategy [Jer88, SWL98]. stratification [GG91a]. Stratified [PSMM93]. Stream [LY98, Raj94, RPT97, RC91, Ye99]. Strength [Ami97]. Stress [AA95, Ami97, AA99a, AAA99]. Stress-Strength [Ami97]. Stresses [ENS96, dL95, AAAaZ99b, AAAaZ99a, AaAaZ00b, AaAaZ00a]. Strict [Vál88, Moh90]. String [IRU83, KSZ^{+}95, IA80]. Strings [Cha93, Ebe93, RUS^{+}84]. Strokes [GP97]. Strong [Gar89]. strongly [Bar89, JC97, JN91]. Structural [Ben92, FMR86, Kum95, Pal86, HG84, Kee89, RUS^{+}84]. Structure [FHG96a, Jin95, LK95, PBL92, Wan94, Wan95b, ZH93, Baa88, DML98, KK91, Nie85, SSW99, Sch91a, TyG99, Zak89]. Structured [IKP97, TSK98, YS94, ZA96, KP98, WK98].

Structures [Mus97a, RRB95, CY99, MS87b, Mus97b, Sol99b, Van83]. studied [KK91]. Studies [AB95a, Kub94a, SKCW94]. Study [Cen97, GF96, Koč92, Kor92, KA96b, LL94a, MFGW97, Mur89a, Mus97a, SM95, Waz97b, Waz97c, AKV83, AH99b, DH99a, HK84, HC85, KP85, LS98b, Man88, Mus97b, OTA83, RR98, TW89, Waz99b]. Sturm [ABW99, Bai81, WJ92, Zha95]. Subbotin [Beh90]. Subdomain [KR95b]. subgrid [RS86]. Subject [Ano95, Ano96, Ano97]. subjected [Mos88a]. Subjects [ITN97]. Submerged [KD95a, KD95b]. Suboptimal [TFM93]. subpopulations [MH80]. subspaces [KCS90, Sen90]. Substrates [Zag92]. substructuring [Vas91]. Subsystem [Mur96]. Successive [GMS83, AM81, CK86, VR99e, VR99f]. successive-correction [VR99e]. successive-corrections [VR99f]. Sufficient [Arg98, FLZ91, Shi89]. suitability [GG91a]. sum [Gri98, MMP89a, Mit92, RR99]. sum-of-product [Mit92]. Summary [SA94, Mus80]. Summation [CS97, Mur91a, SS98a, CSZ98, Lon87a, Lon87c, Lon88]. Summation-over-paths [Mur91a]. Sums [MMP90, PY94, Kha81, WLTS99]. Sun [Tuc97]. Sun-Jupiter [Tuc97]. supercomputer [HGS88]. Supercomputers [LMM86]. superconductors [Cos99]. superconvergence [Gol88a]. supergrid [RS86]. superposition [CG98]. Supersonic [CJF93]. Support [HKLP97, KPP93, Mal93, RPT97, HKK^{+}79, HKK^{+}80, YB99]. Supported [Wu97]. Suppression [RL96]. Surface [BJTW98, KUO95, KM94, Knu93, BTW94, Cas81, MS94b, Wan80]. surface-area [Wan80]. Surfaces [BK99, CG83, HT98, SUJ98]. Survey [Cor89, Kha96, Baa88, KR89]. Survival [ADS96, MU90a, MU90b, MH80, MH81]. Sustainable [SAF^{+}97]. Sustained [KS92b, WZ95]. SWAGMAN [PMMA97].

Sweeping [Kwo93]. Sylvester [MD81]. Symbolic [GG91b, GB95]. Symmetric [BKS83, BKS84a, BKS84b, BB96, KB84c, KB84b, LSVS94, Sha94, Bra86, Hua91, MM87a, Man88, Sca89, ST96, Szi99, Udw92b, YO98]. symmetrical [Tes80]. symmetry [Yan99]. symmetry-like [Yan99]. Symplectic [TPGV97]. Symposium [Tho82d]. synopsis [War87]. Synthesis [Sor85, HH85]. System [Ado97d, AM98b, CT96, CT97, FW98, FBSS96, Gun93, GSI96, GF96, Ham97, Har97, HKLP97, Lin96, LS98a, Mal93, MZV96, Mor94, MBS98, PL93, SA96, SST95, Tuc95, VLP95, VB92, XS95, XZW96, Yil96, Abd99, Ais88, AHKK80, Ass87, Ass88, BY99, BDJR98, Bra82, CCSK82, EB99, Ge90b, HK81, IFU98, JC97, JV98, Joh82, KKU91, KS80, LRV84, Lia98, MR89, Mur91b, MR99, Pat87, PM88, PZZy99, RR98, Sag91, Sav93, SM88, Söd89, Son93, Sor85, TETH98, TGK91, Uen89, VBH91, WW98, Wei98, Zam88, Zha91, ZSL98]. Systems [Ayo93, Bai97, Bay97, Béc98, Bil96, BT97, Che95, Che96b, CUK92, FT94, FLG92, FL94, Fri97, FHG96b, GWL94, Har95, Har97, HHT98, HH96b, HH96a, HH97, Jin95, JPF95, JNF96, KTW81, KM95, Koł94, Koł95, Kon92, Kor92, Kum95, LMS92, Lar97, LFS95, LR95, Lei95, LL96, MMMK96, Mo97, MS94c, Mor95, MFGW97, MKT96, MBS98, NFJ94, OMS95, RAL97, RL96, RRB95, Rua92, Rya96, Sha97, Son96, Sug95, Tho82d, TH95, UK95, Unb96, Wam92, WK95, Ada99, AM98a, ALZ99, AS99, And93, Any82, Bai99a, Bai99b, BdA89, BL87, BH89, Cas88, Cha85, CG87, Cha89b, Chu98, Col82, CW89, Den86, ER88, EL91, FPGV95, FLZ91, GS99, GG91b, GB95, GL99, HG84, HEH81, Hau82, He96, Hil80, IM80, Jam86, JS98b, JF91, JNF95, Kee89, KM86, KB89, Kur99]. systems [LR85, LL89, LSP91, Li94, LS99, LP89, LS91b, pLR87b, MD98, MM91, Mas82, McW82, Mei89, MM87b, Mor98, Mor86a, Mor86b, MS87a, MS87b, MS92c, MR91, NA91, O’N90, RSS89, RL82, Ser79, Sol98, Sol99a, Sol99b, SK98a, SP93, Tha82, Tho82a, Tho82c, Tom80, Udw81, Udw87, Udw91, Udw92a, UK94, VBH91, VL83, VB89, mW98b, WSL83, Wei99a, Yan98, YL98a, YW85, ZH93, dAMY89, dPL84].

Table [PMMA97]. tactics

[HKK^{+}79, HKK^{+}80]. Tailed [WLB97]. Tailoring [Mei89]. taken [BD90]. Taking [LC97]. Tam [Lan89]. Tam-Zardecki [Lan89]. Tangent [Knu93]. tapered [Hof88]. Target [Ahm93, CO96, Cic99]. Targeting [HYL97]. Taylor [BR88, Cha89b, FG94a, HY99b, Waz98a]. TDM [WW99a, WW99b]. teacher [Cas91]. team [AHKK80, Lak81]. Technique [Bad93, BG94c, CN97, De 94, HZL97, JM93, LM95, RBPK96, RK93, TSK98, WK96a, BKS83, BKS84a, BKS84b, CL98a, Du83, JC97, KB84c, KB84b, Khu98b, Waz98b, WC99, Yag82, ZS98, Zha98b, Zha99a]. Techniques [Cic93a, Jac93, BBD91, Bay98b, Bay98a, BY99, DH99a, Hei93b, Kni82, KH78, MM91, Mit86, MCD89, Moo81, RGL98, Ram99d, Ram99b, Red90b, RR98, Sca89, VR99b, VR99c, VR99d, VR99e, VR99f, Vol92, Zha97b, Zha98a]. Temperature [MBG95]. template [GMGCL99]. tendency [ZPY99]. Tennessee [Tho82d]. Tensor [Jin95]. Term [Kum92a, Kin94, Pen84, TS87]. terminal [BM89]. termination [VA88]. Terms [Waz97a, FL99, LL89, RS88, Whi89, ZZ98]. Test [Yag82, GG91a, Jam85]. Testing [Bha94, KP86, NH86, TA93]. Tests [BB96, SW92b, Cha91]. Tether [Jan95]. Tethered [SST95, VLP95]. th [FLG92, SV92b]. Thailand [PN97b]. Their [Arg96a, Arg97a, Kim96, SAL96, ABT98, Arg99d, CS97, CSY98, GW89a, GT82, KCS90, Mus80, Tho82d, Yür99]. Theorem [Arg94, FG94a, Gun94, Ibr96, LS97, Zha95, Aft80, BK80, Got85, KP91, Opp92, Sas88, Zan87]. Theorems [GL92, LG94, SS97, AS90, CY80, EL77, ES90, Par87, Smi86]. Theoretic [Cha92a, EM98]. Theoretical [Bog91]. Theory [Cen97, Chu92, CH93, EHE93, FM96, KW96, LK95, MS93c, MK93, PY94, AS81, AHKK80, Alb89, ADM91, BR99a, Bra86, BKM79, Bur89, Cas87, Cas88, Cas94, CY99, DLRJ81, EE89, Eva90a, Eva90b, Eva90c, ER94b, ER94a, Fym79, HEH81, Heu90a, Heu90b, KN90, Khu98c, LS99, MM87a, Mur89a, Mur89b, Mus80, NR99a, Opp92, Pet88, RR98, Shi89, Ste84, Szi89, VA88, Zie84, Zla89]. Therapies [HD96]. Therapy [DGU96, CAP88, Rap92]. Thermal [AA95, CK95, AAAaZ99b, AAAaZ99a, AaAaZ00b, AaAaZ00a]. Thermally [Mei94, Cas81]. thermistor [KBÖ99]. Thermoelastic [AA96, AA99b]. thickness [Uen80]. Thin [KH96]. Thin-Film [KH96]. Third [EN94, KB93, OT92, Wan96a, OT93]. Third-Order [EN94, OT92, Wan96a, KB93, OT93]. Thomas [CL98b, Waz99c]. Three [CS92, nHyG97, JM93, KD95b, Lop92, LMM86, RCG94, SA96, Shi82, SSB82, Tuc95, BF83, Cau83, CG87, Dua99a, Dul82, EB99, GL99, Hol86, yHyG99, Jou82, Kap88, KW86, KM87, MB82, Mur91b, OL80b, ST96, Tha82, Tha84, Tho82a, TS87, WZ82, WM91, ZK98, ZA88]. Three-Body [Tuc95]. three-degree-of-freedom [EB99]. Three-Dimensional [CS92, nHyG97, KD95b, Shi82, SSB82, BF83, Cau83, Dul82, Hol86, yHyG99, Jou82, KM87, MB82, Mur91b, OL80b, ST96, Tha82, Tha84, Tho82a, WZ82, WM91, ZA88]. Three-Point [Lop92]. Three-Species [SA96]. three-term [TS87]. Time [AR98b, Asa97b, CT96, Dua99b, FL94, Fri97, KPP93, Kum95, LSKH96, LL96, Lia97, LZ96, Lu98, LLX96, LUL97, Moh99c, MS94c, NG94, Pap94, PS94b, PY94, RTJL92, Sug95, SO95, Tau80, Wan93, ZA96, ZZZ98, ABW99, AKK99, Ami99, AT85, BD90, BTW94, CW91, Da93c, Da93d, EM90, FMR86, GW89a, HW85, HH85, HKK^{+}79, Kum92b, LR85, LS99, Lin94, LSY98, MMC98, Mus83a, Mus83b, Mus84a, Pal86, PM89, Pat91, Ram98c, TLG90, Udw91, Udw92a, VV85, Wu94, WC99, YR84, ZH86]. Time-Delay [FL94, LL96]. Time-Dependent [MS94c, PS94b, ZH86]. time-invariant [LS99]. time-linearized [Ram98c]. Time-Periodic [LZ96, Dua99b, CW91]. Time-Reverse [NG94]. time-series [AT85]. time-varying [HKK^{+}79, LSY98]. Times [RB95]. tip [Tha82, Udw92a]. Toeplitz [GS99, Jin95, LSVS94, Mor94, Sha94]. tokamak [VCM91]. Tomographic [Kha94]. tomography [FEGL90, PZ99]. Tool [TH96, KN79, MU90a, MU90b]. Tools [Van97]. Topological [Zel87, AW99a, Sch91a]. Topology [SW92a, Rob82]. Tori [Tak98]. toroidal [Khu98a, VCM91]. Torsional [AA94]. total [JMS98, SUJ98]. Tours [War91]. toxic [Pat90]. Tracking [CO96, KK95, PS94b, SO95, Cic99, Mey78]. traction [Lee98]. Trade [Van97, Wei84]. Trade-off [Van97]. Tradeoffs [PL93]. traders [Pal78]. Trading [MPD97]. Traffic [Mil95, MFGW97, Har86, KL85]. Training [AH94]. Trajectory [SMY95]. Transfer [HM94, MA95, SKB95, DSK82, KKS87, KL99a, OL80b, Tak85]. Transfinite [GT82, Hof87]. Transform [Hai97, Kha92, CH99, Kan99, KNS81, RJY80, Sum84, Tag98, VA88, WH85]. Transformation [CH96, Cic92, FH95, FHG96a, ADHE82, CL98a, Die89, Faz90, Faz91, GD87, JC97, KN79, MD98, Mor86b]. Transformations [DGV94, Ge90a, GW89b, Kop90, KR88, LS81, Yag82]. Transforms [GM93, Gzy95, Gzy97, Jay96, KVT98, MNK99, Yür99, Zen99]. Transient [AAAaZ99b, AAAaZ99a, AA96, AaAaZ00b, AaAaZ00a, SZ97, IBS94]. Transistors [SEG96]. Transition [Kub97, NS97, VD97, Ben89]. Transmission [Alw97, LL94a]. Transmitted [MAS95, Uen80]. transonic [Cau83, Jam83, Kle82, MS83, Mel86]. transonic-flow [Mel86]. Transport [BMS92, LS98a, LZ96, PA94, AKK99, HY98, HY99b, HM99, HY99c, Hos99, HY99a, LP81, PV81, Pat90, WW98, Zia89]. Transportation [MFGW97]. Transversal [VLP95]. transversality [AW99a]. Transversely [AA95, AA96]. Trapezoidal [Yeh97]. trapping [Cos99]. Travel [Sla93a]. Traveling [Boy97b, War94, Dro90, Fei87, War91]. traveling-salesman [Fei87]. Treatment [LS93a, BR99b, CAP88, FGLF88, Got84, KR87d, MS83, PF99, Rap92, TT86]. Tree [Moh94, Pes91]. trees [ZSL98]. Trench [LSVS94]. Trends [Unb96]. tri [Sor82]. tri-element [Sor82]. triangle [Beh96a, Fun91, Kor97]. Triangular [Die96, CL98b, KKL99, Tuc97]. Triangularization [SH92]. Triangulated [JkC97]. Triangulations [Wan92, Wan95b]. tributary [Bec85]. Tridiagonal [MS97, Hil80]. trigonometric [Jia99, Ser79]. trimmed [HT98]. Truncated [Kap88, Cra79]. Trust [dYS97, SSY99]. tubes [AAAaZ99a, AaAaZ00a, CCSK82, ST96]. tubular [Faz91]. turbine [WC88]. Turbomachinery [JM93]. turbulence [DS87, SKCW94]. Turbulent [Raj92, Raj94, Raj97]. Turning [LW94, Lop92, NR98]. tutorial [Wat89a]. twin [NR98]. Two [Arg97b, Da93a, Eir89, FV91, GR97, GM94, Gra97, GL94, JY97, KBS96, KKH95, Kha96, KL89, LMS92, LL95, LY98, LS81, Mou96, MFF92, MR99, NSV92, NS97, SW96, SV94, Sim92, Smi86, TM92, WK96b, WLB97, Ye99, YK97a, YK97b, Zag92, Zha98b, Arg99c, Ass88, Bai99b, Bai81, BH93, Bar91, BS86, Cas89, CH99, CS99, DG87b, Faz90, FL99, HK81, HV90, Jam83, KR87d, Kap88, KS93b, LL89, Lak81, MNK99, MS83, Mit90b, ML90, NSE81, Ohm98, OU99, Pen84, RS88, SSW99, Sca89, Tag98, UA88, VR99f, Vos89, WS87, Waz99b, WW99a, WKTM88, Zia89, ZMV99]. Two-color [KL89]. Two-Dimensional [GL94, KKH95, NSV92, NS97, BH93, Bar91, CH99, HV90, Jam83, MS83, OU99, Zia89]. two-fluid [Ohm98]. Two-grid [Zha98b]. Two-Level [LL95, LY98, YK97b]. two-parameter [Bai81]. two-person [Lak81]. two-pipe [Sca89]. Two-Point [GM94, KBS96, LMS92, MFF92, SV94, TM92, MR99, Cas89, Faz90, KR87d, KS93b, ML90, NSE81, Pen84, UA88, VR99f, WS87, WKTM88]. two-sex [HK81]. two-sided [MNK99, Tag98]. Two-Stage [GR97, Kha96, Bai99b, SSW99]. Two-Step [Arg97b, Sim92, Arg99c, Vos89]. Two-Tailed [WLB97]. Type [Ado97c, BS97, Cen97, Cic92, Cic93a, Cic93b, EN94, EHE93, HZ95a, HHT98, MS94c, Sri95, TH94, Wan92, Wan96a, AZK96, Dul82, EORA99, Gup90, HL88b, HL88a, Ioa83, KN79, LS78, LMV91, McG79, MS94b, MB88b, VV80, Yse86, ZL90, Arg94, Cos99, MEZI93, SW92b, WK96a, He96, RMR86]. Type-2 [Wan92]. Type-II [Cos99]. Types [CB94b, MSZ96, CB94a]. Typological [Mus97a, Mus97b].

U.S. [DS88]. Ultra [TPS98].

Ultra-Relativistic [TPS98]. Unbounded [JPF95, Mo97, VB92, AA99b, Guo99]. Uncertain [Che95, Lei95, RRB95, TH95, BL87, Udw87]. Uncertainty [EFS95, Fri97, IL95, KPP93, Mal93, MB95, Sar93, SL96, DDS85, Ham98b]. Unconditionally [GGS97]. Unconstrained [Lia97, Moh90, KB80]. undersized [Fie84]. Underwater [BD95, KP96]. Unequally [Koh92]. unexpended [Pat92]. unification [Cha87]. Unified [Arg96a, CS93, Jol92, Arg99d, Con89, Tag99, YL98a, YW85]. Uniform [FL99, Ada93b, Gar89]. Uniqueness [Koz95, GS94a, Koz91, MR99, Pen84, SL83, Sev99b]. unit [Bar88]. Unitary [Boy97a]. Units [LSKH96]. Unity [Boy97a]. Universal [Chu96, Rya96]. universality [CM90]. Unknowns [GP97]. Unprogrammability [Gun93]. unstable [Bra78, Bre81, TS87]. Unsteady [BJTW98, CJF93, Ezz94, nHyG97, JFW96, Raj92, Raj94, Raj97, yHyG99]. Unstirred [BR98]. Unstructured [LS98a]. until [PY94]. unweighted [Sha89a]. updating [AB95b]. upon [MH80]. Upper [BDS96, OT92, LS87, OT93]. Upstream [Ram98d]. Uptake [OA97]. upwind [Váz98]. urine [BKM79, WM91]. Urn [MAS95]. Use [AFM97, Beh96b, HD96, Mur89b, PZ94, Tho82d, WC99, AR82, Ham94a, JB89a, ONOT98, Tur88]. Used [BG94c, TJ95, BR99b]. Useful [Par95]. Users [LL94a]. Uses [PMMA97]. Using [Bay93, BP92, Cen97, CWL92, CW97, CH92, CLC97, Hoj94, HZ96, Jay96, JY97, JL97, MBCV95, Raj97, RK87, RBR94, RK93, Sel94, Sel95, SB92, SIA96, WK95, Bar88, Bar91, BY99, BTW94, CCSK82, CDW91, Cic99, Col82, DLR81, Die89, DF96, EJNT88, ESL98, EE89, Eva90c, EHE93, FS98, GG91b, GB95, GMGCL99, GP97, HT98, Hau82, Hei93b, HC85, HL82, JC97, KT83, KL99b, Kle82, LS98a, MM91, MBS80, Moh90, MS87a, MS92c, MKT96, Nak82, NH86, Ohm98, Ona97, PF99, Pat89, Pet88, Sav93, SW90b, Shi82, SSB82, Tha82, Udw91, Udw92a, VA88, VBH91, Vol92, WN91, ZPY99]. usual [MM87a]. Utility [AAHM93, DH99b, Mal91, MS99]. Utilization [SS98b].

V [VR99e, XL99]. V-cycle [XL99]. Validated [NJC99]. Validation [Lor80, RW93]. validity [Jam85, Mik87]. Value [BKM95, Beh96b, BT97, CN97, Cha96a, GM94, Gup98, Ibr96, JNM92, KR95a, KBS96, LMS92, Mo97, MFF92, OT92, PA94, RGL97b, Sas95b, SV94, ŠNG92, SS94, Sim92, Sim94, Ste94, TM92, VS96, Wan95a, Wan96a, WCG97, YK97a, AG85, Aga86, AW99b, AG99, Als79, AK80, AY98, Bai91, Bow88, Bre81, BKM79, CN90, Cas89, Cas94, CS88, Enr89, EL90, Faz90, GZ84, Guo99, Gup90, GK94, HSZ94, HG99, HC87, HL88b, HL88a, KR84, Kad86, KR87a, KR87d, KB93, KK83a, KS93b, KT89, KH78, KM79, LP81, LRV84, LM94, pLR87a, pLR87b, Mär89, Mit86, ML90, MR99, NJC99, NSE81, Oma86, OT93, Pan88, Pen84, QA98, Sev99b, SV83, SW91, Smi84a, SW89, SV92a, SV92b, TS98, Uen83, UA88, VR99f, Wan99a, Wan99b, WS87, Wei89, WKTM88]. value [WA99]. Valued [Ibr96, Mv98, SS91, WSL83]. Values [Lan92, Waz97c, Gri98, Gyö90, Hir84, Hir85, Hol89, HC85]. valve [VWSS94]. Vandermonde [GS97]. Variable [Asa97b, BG97, Dég96, EHE93, FBSS96, SW96, Sim94, Eva90a, Eva90b, ER94b, ER94a, KB84b, KB93, Mos88a, Pat90, TT84, ZH93, Zla89]. Variable-Coefficient [BG97]. variable-mesh [KB93]. Variable-Step [Sim94]. Variables [Ami97, Hoh97, Jol92, SA94, WP97, AP95, BB83a, CS85, DG87b, EB91, KT83, Mit90b, VS90]. Variance [Jac93]. variant [AY98]. Variation [Aft80, CB94a, CB94b, GWL94, SA94, Bee87, BL78, DT87]. variation-of-constants [DT87]. Variational [CK90, Koč92, LSP91, PZ97, Zha97c, AHKK80, CSR88, Knu91, MM87a, Man86, PK98, SB82, ZZ98]. variations [KS84]. variety [KBÖ99]. Various [KSZ^{+}95, ESL98, MKJ87]. Varying [SS94, SBH96, HKK^{+}79, LSY98]. Vector [CC94, XZL97, Aft80, Axe86, CJ88, GMS83, NH86, Rug93, Vál88, ZM91, ZG90]. Vectorizable [Mel86]. vectorization [HT88]. Vectorized [BB83b, Gno82, Hol86, HV90]. Vectors [LUL97]. Vehicle [BD95, CO96, KP96, Mat93, SWL98]. Velocity [BNN95, Raj94, Raj97, MS94a, MKJ87]. velocity-current [MS94a]. velocity-pressure-gradient [MKJ87]. ventricular [YDB79]. version [Mit92]. versions [MW88b]. versus [Bur89, HV90]. vertex [Kor97]. Verticum [MS94c]. Very [KWS95, Doa90]. VI [VR99f]. via [ABT98, BK99, BG97, BS93, CS96, CS88, Die89, Ge91, HKLS93, HZ95b, Kad86, KR87c, KRT83, LLS94, LLS96, MF99, MW98a, Mil98, Mus97c, PVM88, Pap96, RC91, RRA90, RAL95, RAL97, Sen90, Sol98, Sug83, TM92, VCM91]. Vibration [Ado96d, LR95, MD98, MB95, RL96, HF93]. Vibrations [Bar95]. view [RB85]. Violin [Mit97, Mit99]. virtual [Zha98c]. viscous [ADHE82, BS86, HF95, Joh83]. viscously [HG84]. Visibility [BD94]. visual [Hof89, Baa88]. Visualization [Kar98, BK99]. VML [Zha98c]. Voigt [LW90]. Volterra [Abd93, AM98b, AR98b, Bla96, BW77, Bow78, ER88, GL94, LL96, Mor90, Rua92, SA96, Sol99a, WK96a, Waz99a, YF91, ZL90]. Volterra-series [YF91]. Volterra-Type [WK96a]. Volume [LM95, LS98a, And90, Gno82, STD98]. Volumes [Lan89]. Voronoï [HT98]. vortex [Cos99]. Vries [RC99]. vs [AB95a, RBR94].

waiting [Ass88]. Water [BDS96, HYL97, OA97, PL93, PMMA97, RPT97, YBR97, AT82, Bog91, DH99a, GGR91, Hau82, Khu98c, MN91, Pap84]. Wave [AA94, Ado87b, Boy97b, ENS96, KD95a, KD95b, OŞL80, Reg94, Reg96, Abu99a, AT82, Hau82, Jac90, KV84, Sha87, Soc88, Waz98b]. Waveform [Son96, Zha96a]. Wavelet [XS95, ZL98c]. Waves [Boy97b, AA99a, AAA99, AA99c, Ahm99, DH99a, Khu98c, Zak89]. Way [Mat93, WH87]. Weak [Abu97, Cro93, Hua91, JK98, Zak89]. Weakly [Pap92, WK96a, Bai99a, Da93c, Da93d, dAMY89]. weakly-coupled [dAMY89]. Weapon [Ahm93]. weather [AK85]. Weibull [BP92, Sos91]. weight [Da93b, GBS99]. Weighted [CLC97, RBPK96, Sha89a, Pan88, VS90, Wan87, YL98a]. Weighting [CH93, CL99, Lin94]. well [HK88, ML90]. well-conditioning [ML90]. well-posed [HK88]. which [FS90]. White [WL97]. whole [Ahm87]. whose [Net89]. Widely [HSG94, HZL97]. Wiener [NLC95]. Wigley [BJTW98]. Windowing [Zha96a]. wine [SS87]. wing [Sor82, Tha82]. wing-tip [Tha82]. wing/wing [Tha82]. wing/wing-tip [Tha82]. wishes [Kag91a]. within [Mur93, Pig87]. Work [Tes91b, Tes91c, Kag91b]. World [Mus94, Sla93b]. writing [SW79a].

X [LMM86]. X-MP [LMM86]. X-MP/24 [LMM86]. XP [LMM86]. XP-200 [LMM86].

Yamamoto [Arg94]. Yorke [Wat81].

Zakharov [Ado97b]. Zardecki [Lan89]. Zero [PY94, Sor96, Di 98]. Zeros [KS92a, Mil87b, Sor97b, XZL97, IA85, LS78, Mil88a, MW88a, MW88b, SW90a, Sor99, Wol89, ZS99].

References

Abd-Alla:1994:TWP

[AA94] Abo-El-Nour N. Abd-Alla. Torsional wave propagation in an orthotropic magneto-elastic hollow circular cylinder. Applied Mathematics and Computation, 63(2–3):281–293, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abd-Alla:1995:TST

[AA95] A. M. Abd-Alla. Thermal stress in a transversely isotropic circular cylinder due to an instantaneous heat source. Applied Mathematics and Computation, 68(2–3):113–124, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=2-3&aid=9400085.

Abd-Alla:1996:TRI

[AA96] A. M. Abd-Alla. Transient response of inhomogeneous transversely isotropic thermoelastic cylinder to a dynamic input. Applied Mathematics and Computation, 74(1):1–13, January 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=1&aid=9400246.

Abd-Alla:1999:EIS

[AA99a] A. M. Abd-Alla. The effect of initial stress and orthotropy on the propagation waves in a hollow cylinder. Applied Mathematics and Computation, 106 (2–3):237–244, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/28/
article.pdf.

Abd-Alla:1999:GTI

[AA99b] A. M. Abd-Alla. On a generalized thermoelastic interaction in an unbounded body due to a line heat source. Applied Mathematics and Computation, 100(2–3):285–295, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6216.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6216.

Abd-Alla:1999:PRW

[AA99c] A. M. Abd-Alla. Propagation of Rayleigh waves in an elastic halfspace of orthotropic material. Applied Mathematics and Computation, 99(1):61–69, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/1/6171.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
1&aid=6171.

Abd-Alla:1999:PLW

[AAA99] A. M. Abd-Alla and S. M.

Ahmed. Propagation of Love waves in a non-homogeneous orthotropic elastic layer under initial stress overlying semiinfinite medium. Applied Mathematics and Computation, 106 (2–3):265–275, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/30/
article.pdf.

Abd-All:1999:TTSb

[AAAaZ99a] A. M. Abd-All, A. N. Abdalla, and N. A. Zeidan. Transient thermal stresses in a rotation non-homogeneous cylindrically orthotropic composite tubes. Applied Mathematics and Computation, 105(2–3): 253–269, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/2-
3/6323.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=

                                        105&issue=2-3&aid=6323. See

erratum [AaAaZ00a].

Abd-All:1999:TTSa

[AAAaZ99b] A. M. Abd-All, A. N. Abd-alla, and N. A. Zeidan. Transient thermal stresses in a spherically orthotropic elastic medium with spherical cavity. Applied Mathematics and Computation, 105(2–3):231–252, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/105/2-
                                        3/6314.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        105&issue=2-3&aid=6314. See

erratum [AaAaZ00b].

Abd-alla:2000:ETTb

[AaAaZ00a] A. M. Abd-alla, A. N. Abdalla, and N. A. Ziedan. Erratum to: “Transient thermal stresses in a rotation non-homogeneous cylindrically orthotropic composite tubes” [Appl. Math. Comput. 105 (1999) 253–269]. Applied Mathematics and Computation, 112(2– 3):346, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/29/17/20/86/23/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        29/17/20/86/23/36/article.

pdf. See [AAAaZ99a].

Abd-alla:2000:ETTa

[AaAaZ00b] A. M. Abd-alla, A. N. Abd-alla, and N. A. Ziedan. Erratum to: “Transient thermal stresses in a spherically orthotropic elastic medium with special cavity” [Appl. Math. Comput. 105 (1999) 231–252]. Applied Mathematics and Computation, 112(2– 3):345, June 15, 2000. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.elsevier.
nl/gej-ng/29/17/20/86/23/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
29/17/20/86/23/35/article.

pdf. See [AAAaZ99b].

Alwasel:1998:EPB

[AAAN98] I. Alwasel, A. Alzaid, and H. Al-Nachawati. Estimating the parameters of the binomial autoregressive process of order one. Applied Mathematics and Computation, 95 (2–3):193–204, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6110.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6110.

Abdel-Aziz:1999:SAL

[AAES99] Mohammedi R. Abdel-Aziz and Salah M. El-Sayed. Sensitivity analysis of the largest dependent eigenvalue functions of eigensystems. Applied Mathematics and Computation, 100 (2–3):103–110, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/100/2-
                                        3/6215.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        100&issue=2-3&aid=6215.

Al-Alwani:1993:IIM

[AAHM93] Jumah E. Al-Alwani, Benjamin F. Hobbs, and Behnam B. Malakooti. An interactive integrated multiobjective optimization approach for quasiconcave/quasiconvex utility functions. Applied Mathematics and Computation, 54(2–3):241– 257, March 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmed:1987:AGP

[AB87] Javed Ahmed and Charles D. Bonham. Application of geometric programming to optimum allocation problems in multivariate doubling sampling. Applied Mathematics and Computation, 21(2):157–169, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ascher:1989:NCS

[AB89] U. Ascher and G. Bader. A note on conditioning, stability and collocation matrices. Applied Mathematics and Computation, 31(Special issue):437–448, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Azaiez:1994:PAB

[AB94] M. Naceur Azaiez and Vicki M.

Bier. Perfect aggregation in Bayesian estimation for several realistic reliability models. Applied Mathematics and Computation, 64(2–3):139–153, September 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Angleton:1995:LSR

[AB95a] George M. Angleton and Charles D.

Bonham. Least squares regression vs. geometric mean regression for ecotoxicology studies. Applied Mathematics and Computation, 72(1):21–32, September 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=1&aid=9400161.

Azaiez:1995:PAC

[AB95b] M. Naceur Azaiez and Vicki M.

Bier. Perfect aggregation for a class of general reliability models with Bayesian updating. Applied Mathematics and Computation, 73(2–3):281–302, December 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        73&issue=2-3&aid=9500068.

Abdalkhani:1992:NCS

[Abd92] J. Abdalkhani. The nonlinear cooling of a semi-infinite solid — implicit Runge–Kutta (IRK) methods. Applied Mathematics and Computation, 52(2–3): 233–237, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdalkhani:1993:HOM

[Abd93] J. Abdalkhani. Higher order methods for solving Volterra integrodifferential equations of the first kind. Applied Mathematics and Computation, 57(1):97– 101, August 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abdou:1999:IEC

[Abd99]  M. A. Abdou.  Integral

equation and contact problem for a system of impressing stamps. Applied Mathematics and Computation, 106 (2–3):141–148, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.elsevier.nl/gej-ng/
                                        10/9/12/71/18/19/abstract.
                                        html;   http://www.elsevier.

nl/gej-ng/10/9/12/71/18/19/
article.pdf.

Allen:1984:MES

[ABES84] P. M. Allen, F. Boon, G. Engelen, and M. Sanglier. Modeling evolving spatial choice patterns. Applied Mathematics and Computation, 14(??):97–129, 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ablow:1982:EMG

[Abl82] C. M. Ablow. Equidistant mesh for gas dynamic calculations. Applied Mathematics and Computation, 10/11(1): 859–863, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Abaffy:1998:PMA

[ABT98] Jozsef Abaffy, Marida Bertocchi, and Anna Torriero. Perturbations of M-matrices via ABS methods and their applications to input-output analysis. Applied Mathematics and Computation, 94(2–3):145–170, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/2-3/
6087.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
2-3&aid=6087.

Abuelmaatti:1997:HIP

[Abu97] Muhammad Taher Abuelma’atti. Harmonic and intermodulation performance of some abrupt and weak nonlinearities. Applied Mathematics and Computation, 83(1):87–96, April 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=1&aid=9600049.

Abuelmaatti:1999:NDF

[Abu99a] Muhammad Taher Abuelma’atti. Nonlinear distortion of an FM signal by half-wave linear rectification. Applied Mathematics and Computation, 99 (1):47–59, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/1/6180.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
1&aid=6180.

Abuelmaatti:1999:SAC

[Abu99b] Muhammad Taher Abuelma’atti. A simple algorithm for computing the Fourier spectrum of experimentally obtained signals. Applied Mathematics and Computation, 98(2– 3):229–239, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/
6162.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6162.

Agarwal:1999:SLE

[ABW99] Ravi P. Agarwal, Martin Bohner, and Patricia J. Y. Wong. Sturm–Liouville eigenvalue problems on time scales. Applied Mathematics and Computation, 99(2–3):153–166, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6188.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6188.

Argyros:1994:MMB

[AC94] Ioannis K. Argyros and Dong Chen. The midpoint method in Banach spaces and the Pták error estimates. Applied Mathematics and Computation, 62 (1):1–15, April 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anghel:1997:NMM

[AC97] V. N. P. Anghel and W.-H.

Choe. Numerical methods for magnetized plasma equations with boundary singularity. Applied Mathematics and Computation, 85(2–3):149– 164, September 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600132.

Agouzal:1998:EMA

[AD98] A. Agouzal and N. Debit. Equilibrium method to approximate elliptic problems with non-standard boundary conditions. Applied Mathematics and Computation, 95(2–3): 165–171, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6111.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6111.

Adams:1989:MMP

[Ada89] J. C. Adams. MUDPACK: multigrid portable FORTRAN software for the efficient solution of linear elliptic partial differential equations. Applied Mathematics and Computation, 34 (2 (part II)):113–146, November 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adams:1991:REM

[Ada91] John C. Adams. Recent enhancements in MUDPACK, a multigrid software package for elliptic partial differential equations. Applied Mathematics and Computation, 43(1):79–93, 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adamatzky:1993:ICA

[Ada93a] Andrew I. Adamatzky. Implantation of cellular automata. Applied Mathematics and Computation, 55(1):49–71, April 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adams:1993:MMS

[Ada93b] John C. Adams. MUDPACK-2:

Multigrid software for approximating elliptic partial differential equations on uniform grids with any resolution. Applied Mathematics and Computation, 53(2–3):235–249, March 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adamatzky:1997:HCA

[Ada97] Andrew Adamatzky. How cellular automaton plays Minesweeper. Applied Mathematics and Computation, 85(2–3):127–137, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.elsevier.com/cgi-bin/
cas/tree/store/amc/cas_sub/
browse/browse.cgi?year=1997&
volume=85&issue=2-3&aid=9600117.

Adamatzky:1998:MSA

[Ada98] Andrew Adamatzky. Masterslave algebra: On the binary compositions of distributed beliefs. Applied Mathematics and Computation, 95(2–3): 173–180, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6107.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6107.

Adamatzky:1999:NBC

[Ada99] Andrew Adamatzky. Nonconstructible blocks in 1D cellular automata: minimal generators and natural systems. Applied Mathematics and Computation, 99(1):77–91, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/1/6178.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
1&aid=6178.

Adeff:1994:SAP

[Ade94] S. E. Adeff. Simulation of axisymmetric piezoelectric devices. Applied Mathematics and Computation, 65(1–3):355–369, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anderson:1982:SVI

[ADHE82] O. L. Anderson, R. T. Davis, G. B. Hankins, and D. E. Edwards. Solution of viscous internal flows on curvilinear grids generated by the Schwarz–Christoffel transformation. Applied Mathematics and Computation, 10/11(1):507–524, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Ammar:1991:EIT

[ADM91] G. Ammar, W. Dayawansa, and C. Martin. Exponential interpolation: theory and numerical algorithms. Applied Mathematics and Computation, 41 (3):189–232, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1981:PNS

[Ado81] G. Adomian. On product nonlinearities in stochastic differential equations. Applied Mathematics and Computation, 8(1): 79–82, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1987:IAD

[Ado87a] G. Adomian. An investigation of the asymptotic decomposition method for nonlinear equations in physics. Applied Mathematics and Computation, 24(1):1–17, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1987:WPN

[Ado87b] G. Adomian. Wave propagation in nonlinear media. Applied Mathematics and Computation, 24(4):311–332, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1988:ASN

[Ado88] G. Adomian. Analytic solutions for nonlinear equations. Applied Mathematics and Computation, 26(1):77–88, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1996:BKM

[Ado96a] G. Adomian. The Burridge– Knopoff model. Applied Mathematics and Computation, 77(2– 3):131–132, ???? 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1996:CME

[Ado96b] G. Adomian. Coupled Maxwell equations for electromagnetic scattering. Applied Mathematics and Computation, 77(2–3): 133–135, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1996:KPE

[Ado96c] G. Adomian. The Kadomtsev– Petviashvili equation. Applied Mathematics and Computation, 76(1):95–97, April 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=1&aid=9500186.

Adomian:1996:NRV

[Ado96d] G. Adomian. Nonlinear random vibration. Applied Mathematics and Computation, 77(2– 3):109–112, ???? 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1997:ESN

[Ado97a] G. Adomian. Explicit solutions of nonlinear partial differential equations. Applied Mathematics and Computation, 88(2–3):117– 126, December 30, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600141.

Adomian:1997:NPS

[Ado97b] G. Adomian. Non-perturbative solution of the Klein–Gordon– Zakharov equation. Applied Mathematics and Computation, 81(1):89–92, January 1, 1997.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=1&aid=9500311.

Adomian:1997:KTE

[Ado97c] G. Adomian. On KdV type equations. Applied Mathematics and Computation, 88(2–3):131– 135, December 30, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600283.

Adomian:1997:DRD

[Ado97d]  G. Adomian.  On the dy-

namics of a reaction-diffusion system. Applied Mathematics and Computation, 81(1):93– 95, January 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=1&aid=9500312.

Adomian:1997:OPR

[Ado97e] G. Adomian. Optical propagation in random media. Applied Mathematics and Computation, 88(2–3):127–129, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/

tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600282.

Ananda:1996:ABE

[ADS96] Malwane M. Ananda, Rohan J.

Dalpatadu, and Ashok K. Singh. Adaptive Bayes estimators for parameters of the Gompertz survival model. Applied Mathematics and Computation, 75(2–3): 167–177, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500124.

Antoine:1997:MCL

[AFM97] Jacques Antoine, Günther Fischer, and Marek Makowski. Multiple criteria land use analysis. Applied Mathematics and Computation, 83(2–3):195– 215, May 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=2-3&aid=9600190.

Aftabizadeh:1980:VCV

[Aft80] A. R. Aftabizadeh. Variation of constants, vector Lyapunov functions, and comparison theorem. Applied Mathematics and Computation, 7(4):341–352, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agarwal:1985:MCM

[AG85] Ravi P. Agarwal and Ramesh C. Gupta. Method of chasing for multipoint boundary value problems. Applied Mathematics and Computation, 17(1):37–43, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Asaithambi:1989:PSB

[AG89] N. S. Asaithambi and J. B. Garner. Pointwise solution bounds for a class of singular diffusion problems in physiology. Applied Mathematics and Computation, 30(3):215–222, April 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Gahtani:1999:IBS

[AG99] H. J. Al-Gahtani. Integral-based solution for a class of second order boundary value problems. Applied Mathematics and Computation, 98(1):43–48, January 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/98/1/6154.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=98&issue=
                                        1&aid=6154.

Agarwal:1986:CMD

[Aga86]  Ravi P. Agarwal.  Compu-

tational methods for discrete boundary value problems. Applied Mathematics and Computation, 18(1):15–41, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adeli:1994:ACG

[AH94] H. Adeli and S. L. Hung.

An adaptive conjugate gradient learning algorithm for efficient training of neural networks. Applied Mathematics and Computation, 62(1):81–102, April 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Awartani:1999:NRG

[AH99a] Marwan Awartani and M. H.

Hamdan. Non-reactive gasparticulate models of flow through porous media. Applied Mathematics and Computation, 100(1):93–102, April 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/1/
6197.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=100&
issue=1&aid=6197.

Awartani:1999:SAG

[AH99b] Marwan Awartani and M. H.

Hamdan. Some admissible geometries in the study of steady plane flow of a dusty fluid through porous media. Applied Mathematics and Computation, 100(1):85–92, April 1, 1999. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/100/1/
                                        6196.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=100&
                                        issue=1&aid=6196.

Akbari:1980:ETT

[AHKK80] Alireza Akbari, James Hess, Harriet Kagiwada, and Robert Kalaba. The equivalence of team theory’s integral equations and a Cauchy system: sensitivity analysis of a variational problem. Applied Mathematics and Computation, 6(1):21–36, January 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmadi:1979:MSS

[Ahm79] Goodarz Ahmadi. On the mean square stability of linear difference equations. Applied Mathematics and Computation, 5(3): 233–241, ???? 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmad:1987:SRS

[Ahm87] Khader H. Ahmad. Stability relative to a set and to the whole space revisited. Applied Mathematics and Computation, 24(2):91–99, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmad:1993:NEC

[Ahm93] Ibrahim A. Ahmad. Nonparametric estimation of the CDF of weapon effectiveness on target performance. Applied Mathematics and Computation, 57(2– 3):289–303, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmed:1999:IGP

[Ahm99] S. M. Ahmed. Influence of gravity on the propagation of waves in granular medium. Applied Mathematics and Computation, 101(2–3):269–280, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/101/2-
3/6232.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
101&issue=2-3&aid=6232.

Aisbett:1988:CAM

[Ais88] Janet Aisbett. Cyclic audit of a multichannel system. Applied Mathematics and Computation, 25(1 (part I)):11–28, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Armstrong:1980:ALA

[AK80] Ronald D. Armstrong and Mabel Tam Kung. An algorithm for a least absolute value regression problem with bounds on the parameters. Applied Mathematics and Computation, 7(3):267–279, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alexandridis:1985:DMC

[AK85] Mark G. Alexandridis and Roman Krzysztofowicz. Decision models for categorical and probabilistic weather forecasts. Applied Mathematics and Computation, 17(3):241–266, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ali:1996:AFC

[AK96] Ismail Ali and Shyam Kalla. An application of fractional calculus to the solution of a general class of differintegral equations. Applied Mathematics and Computation, 77(2–3):137–152, ???? 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ali:1999:TDM

[AKK99] I. Ali, S. L. Kalla, and H. G. Khajah. A time dependent model for the transport of heavy pollutants from groundlevel aerial sources. Applied Mathematics and Computation, 105(1):91–99, October 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/1/
6312.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.

cgi?year=1999&volume=105&
issue=1&aid=6312.

Alvarez:1983:NSN

[AKV83] A. Alvarez, Pen Yu Kuo, and Luis Vázquez. The numerical study of a nonlinear onedimensional Dirac equation. Applied Mathematics and Computation, 13(1–2):1–15, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Albrecht:1989:EGT

[Alb89] Peter Albrecht. Elements of a general theory of composite integration methods. Applied Mathematics and Computation, 31 (Special issue):1–17, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Allen:1990:MAS

[ALMS90] Linda Allen, Truman Lewis, Clyde F. Martin, and Mark Stamp. A mathematical analysis and simulation of a localized measles epidemic. Applied Mathematics and Computation, 39(1):61–77, 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alspaugh:1979:AII

[Als79] Dale W. Alspaugh. Application of invariant imbedding to linear multipoint boundary value problems with general boundary conditions. Applied Mathematics and Computation, 5(3):213–232, ???? 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Alwash:1997:LCD

[Alw97]  M. A. M. Alwash.  Limit

cycles in a disease transmission model. Applied Mathematics and Computation, 86 (1):85–92, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1997/86/1/9600174.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1997&volume=86&issue=
1&aid=9600174.

Allegretto:1999:NAS

[ALZ99] W. Allegretto, Y. Lin, and A. Zhou. Numerical analysis for systems with memory arising from semiconductor simulations. Applied Mathematics and Computation, 105(2–3): 101–119, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/2-
3/6302.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
105&issue=2-3&aid=6302.

Adomian:1981:CIM

[AM81] G. Adomian and K. Malakian.

A comparison of the iterative method and Picard’s successive approximations for deterministic and stochastic differential equations. Applied Mathematics and Computation, 8(3):187–204, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agwu:1998:OCD

[AM98a] Nwojo N. Agwu and Clyde F.

Martin. Optimal control of dynamic systems: Application to spline approximations. Applied Mathematics and Computation, 97(2–3):99–138, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6109.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6109.

Ahmad:1998:ENP

[AM98b] Shair Ahmad and Francisco Montes De Oca. Extinction in nonautonomous T -periodic competitive Lotka–Volterra system. Applied Mathematics and Computation, 90(2– 3):155–166, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/

                                        9700396.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        90&issue=2-3&aid=9700396.

Aminzadeh:1997:ERE

[Ami97] M. S. Aminzadeh. Estimation of reliability for exponential stressstrength models with explanatory variables. Applied Mathematics and Computation, 84(2– 3):269–274, July 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        84&issue=2-3&aid=9600090.

Aminzadeh:1999:ECS

[Ami99] Mostafa S. Aminzadeh. Estimation of P (Z < Y) for correlated stochastic time series models. Applied Mathematics and Computation, 104 (2–3):179–189, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/104/2-
                                        3/6294.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        104&issue=2-3&aid=6294.

Anderson:1987:ESP

[And87] Dale A. Anderson. Equidistribution schemes, Poisson generators, and adaptive grids. Applied Mathematics and Computation, 24(3):211–227, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anderson:1990:GCV

[And90] Dale A. Anderson. Grid cell volume control with an adaptive grid generator. Applied Mathematics and Computation, 35(3): 309–317 (or 209–217??), February 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Andrus:1993:RKM

[And93]  Jan Frederick Andrus.  A

Runge–Kutta method with stepsize control for separated systems of first-order ODEs. Applied Mathematics and Computation, 59(2–3):193–214, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:1989:OC

[Ano89] Anonymous, editor. 1986 ODE Conference, volume 31 of Applied Mathematics and Computation. North-Holland, Amsterdam, The Netherlands, May 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:1995:SI

[Ano95] Anonymous. 1995 subject index.

Applied Mathematics and Computation, 73(2–3):303–??, ???? 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:1996:SI

[Ano96] Anonymous. 1996 subject index. Applied Mathematics and Computation, 80(2–3):301–??, ???? 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:1997:SI

[Ano97] Anonymous. 1998 subject index. Applied Mathematics and Computation, 87(2–3):313–??, ???? 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:1999:CP

[Ano99a] Anonymous. Conference paper(s). Applied Mathematics and Computation, 106(2– 3):AD1, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/
                                        9/12/71/18/35/;   http://
                                        www.elsevier.nl/gej-ng/10/
                                        9/12/71/18/35/article.pdf.

Anonymous:1999:I

[Ano99b] Anonymous. Index. Applied Mathematics and Computation, 106(2–3):305–306, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/
                                        9/12/71/18/34/;   http://
                                        www.elsevier.nl/gej-ng/10/
                                        9/12/71/18/34/article.pdf.

Anyiwo:1982:IDG

[Any82] Joshua C. Anyiwo. Idealized dynamic grid computation of physical systems. Applied Mathematics and Computation, 10/11 (1):837–858, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Anastasselou:1987:ASP

[AP87] E. G. Anastasselou and D. E.

Panayotounakos. Analytical solution of polynomial equations with an application to the quintic equation. Applied Mathematics and Computation, 21(2): 145–155, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agarwal:1995:SDI

[AP95] Ravi P. Agarwal and Peter Y. H. Pang. Sharp discrete inequalities in n independent variables. Applied Mathematics and Computation, 72(2–3): 97–112, October 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400130.

Anderson:1982:USA

[AR82] Dale A. Anderson and M. M.

Rai. The use of solution adaptive grids in solving partial differential equations. Applied Mathematics and Computation, 10/11(1):317–338, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Adomian:1987:ESD

[AR87] G. Adomian and R. Rach. Explicit solutions for differential equations. Applied Mathematics and Computation, 23(1):49–59, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Artzrouni:1990:SDEa

[AR90a] Marc Artzrouni and James Reneke. Stochastic differential equations in mathematical demography: a review. Applied Mathematics and Computation, 38(1 (part I)):7–21, July 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Artzrouni:1990:SDEb

[AR90b] Marc Artzrouni and James Reneke. Stochastic differential equations in mathematical demography: a review. Applied Mathematics and Computation, 39(3 (Supplement)):139s–153s, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1993:NAM

[AR93] G. Adomian and R. Rach. A new algorithm for matching boundary conditions in decomposition solutions. Applied Mathematics and Computation, 58(1):61– 68, September 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ahmad:1998:NSG

[AR98a] Falih Ahmad and Mohsen Razzaghi. A numerical solution to the Gel′fand–Levitan– Marchenko equation. Applied Mathematics and Computation, 89(1–3):31–39, January 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781646.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Ahmad:1998:SVD

[AR98b] Shair Ahmad and M. Rama Mohana Rao. Stability of Volterra diffusion equations with time delays. Applied Mathematics and Computation, 90 (2–3):143–154, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/
9700395.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=2-3&aid=9700395.

Argyros:1990:SMG

[Arg90] Ioannis K. Argyros. The secant method in generalized Banach spaces. Applied Mathematics and Computation, 39(2 (part II)):111–121, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Argyros:1994:CTN

[Arg94] Ioannis K. Argyros. A convergence theorem for Newton-Like methods under generalized Chen-Yamamoto-Type assumptions. Applied Mathematics and Computation, 61(1):25– 37, March 01, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Argyros:1996:RNMa

[Arg96a] Ioannis K. Argyros. Results on Newton methods: Part I. A unified approach for constructing perturbed Newton-Like methods in Banach space and their applications. Applied Mathematics and Computation, 74(2–3): 119–141, February 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        74&issue=2-3&aid=9500090.

Argyros:1996:RNMb

[Arg96b] Ioannis K. Argyros. Results on Newton methods: Part II. perturbed Newton-like methods in generalized Banach spaces. Applied Mathematics and Computation, 74(2–3):143– 159, February 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500118.

Argyros:1997:MIP

[Arg97a] Ioannis K. Argyros. A mesh independence principle for inexact Newton-like methods and their discretizations under generalized Lipschitz conditions. Applied Mathematics and Computation, 87(1):15–48, November 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=1&aid=9600219.

Argyros:1997:CTS

[Arg97b] Ioannis K. Argyros. On the convergence of two-step methods generated by point-to-point operators. Applied Mathematics and Computation, 82(1): 85–96, March 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
82&issue=1&aid=9600025.

Argyros:1998:SCC

[Arg98]  Ioannis K. Argyros.  Suffi-

cient conditions for constructing methods faster than Newton’s. Applied Mathematics and Computation, 93(2–3):169–181, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/2-3/
6065.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6065.

Argyros:1999:CRI

[Arg99a] Ioannis K. Argyros. Convergence rates for inexact Newtonlike methods at singular points and applications. Applied Mathematics and Computation, 102 (2–3):185–201, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6241.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6241.

Argyros:1999:NMU

[Arg99b] Ioannis K. Argyros. On Newton’s method under mild differentiability conditions and applications. Applied Mathematics and Computation, 102 (2–3):177–183, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/

store/amc/sub/1999/102/2-
3/6240.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6240.

Argyros:1999:CTS

[Arg99c] Ioannis K. Argyros. On the convergence of two-step methods generated by point-to-point operators. Applied Mathematics and Computation, 102 (2–3):165–176, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6242.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6242.

Argyros:1999:RNMa

[Arg99d] Ioannis K. Argyros. Results on Newton methods. Part 1: A unified approach for constructing perturbed Newtonlike methods in Banach space and their applications. Applied Mathematics and Computation, 102(2–3):203–222, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6243.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6243.

Argyros:1999:RNMb

[Arg99e] Ioannis K. Argyros. Results on Newton methods. Part II: Perturbed Newton-like methods in generalized Banach spaces. Applied Mathematics and Computation, 102(2–3):223–236, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/102/2-
                                        3/6244.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        102&issue=2-3&aid=6244.

Adomian:1994:SGN

[ARM94] G. Adomian, R. Rach, and R. E. Meyers. Solution of genetic nonlinear oscillators. Applied Mathematics and Computation, 64(2– 3):167–170, ???? 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Adomian:1997:NIA

[ARM97] G. Adomian, R. C. Rach, and R. E. Meyers. Numerical integration, analytic continuation, and decomposition. Applied Mathematics and Computation, 88(2–3):95–116, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        88&issue=2-3&aid=9600052.

Adomian:1981:NSD

[AS81] G. Adomian and D. Sarafyan.

Numerical solution of differential equations in the deterministic limit of stochastic theory. Applied Mathematics and Computation, 8(2):111–119, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Argyros:1990:CTA

[AS90] Ioannis K. Argyros and Ferenc Szidarovszky. Comparison theorems for algorithmic models. Applied Mathematics and Computation, 40(2 (part II)): 179–185, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Argyros:1992:MCA

[AS92] Ioannis K. Argyros and Ferenc Szidarovszky. On the monotone convergence of algorithmic models. Applied Mathematics and Computation, 48(2–3):167–176, April 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Sitta:1994:NCB

[AS94a] A. M. M. Abu-Sitta. A note on a certain boundary-layer equation. Applied Mathematics and Computation, 64(1):73–77, August 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abu-Sitta:1994:SBL

[AS94b] A. M. M. Abu-Sitta. On solutions of a boundary layer equation. Applied Mathematics and Computation, 63(1):95–100, June 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ananda:1999:PAM

[AS99] Malwane M. A. Ananda and Ashok K. Singh. Performance assessment of multiple engineered barrier systems. Applied Mathematics and Computation, 102(1):25–33, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/1/
6237.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=102&
issue=1&aid=6237.

Asaithambi:1992:GMS

[Asa92a] N. S. Asaithambi. A Galerkin method for Stefan problems. Applied Mathematics and Computation, 52(2–3):239–250, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Asaithambi:1992:NMI

[Asa92b] N. S. Asaithambi. A numerical method for an inhomogeneous nonlinear diffusion problem. Applied Mathematics and Computation, 50(2–3):307–319, August 1992. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Asaithambi:1997:NMS

[Asa97a] N. S. Asaithambi. A numerical method for the solution of the Falkner–Skan equation. Applied Mathematics and Computation, 81(2–3):259–264, February 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=2-3&aid=9500325.

Asaithambi:1997:VTS

[Asa97b] N. S. Asaithambi. A variable time step Galerkin method for a one-dimensional Stefan problem. Applied Mathematics and Computation, 81(2–3):189– 200, February 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=2-3&aid=9500329.

Asaithambi:1998:FDM

[Asa98] Asai Asaithambi. A finitedifference method for the Falkner–Skan equation. Applied Mathematics and Computation, 92(2–3):135–141, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/2-3/
9710042.pdf;   http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        92&issue=2-3&aid=9710042.

Al-Saqabi:1998:ESF

[ASK98] Bader Al-Saqabi and Virginia S. Kiryakova. Explicit solutions of fractional integral and differential equations involving Erdélyi-Kober operators. Applied Mathematics and Computation, 95(1):1–13, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/95/1/6104.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=95&issue=
                                        1&aid=6104.

Assimakopoulos:1987:MMS

[Ass87]  Nikitas Assimakopoulos.  A

medication management system. Applied Mathematics and Computation, 21(1):73–91, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Assimakopoulos:1988:WST

[Ass88] Nikitas Assimakopoulos. A waiting system with two stages. Applied Mathematics and Computation, 27(2):173–190, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Abolhassani:1986:SCA

[AST86] Jamshid S. Abolhassani, Robert E.

Smith, and Surendra N. Tiwari. Stiffness characteristic and the applications of the method of lines on nonuniform grids. Applied Mathematics and Computation, 18(4):285–311, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anuradha:1994:SBC

[ASZ94] V. Anuradha, R. Shivaji, and J. Zhu. On s-shaped bifurcation curves for multiparameter positone problems. Applied Mathematics and Computation, 65(1– 3):171–182, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Agarwal:1980:SND

[AT80] Ravi P. Agarwal and E. Thandapani. On some new discrete inequalities. Applied Mathematics and Computation, 7(3):205–224, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Aston:1982:ISW

[AT82] Martha B. Aston and J. W.

Thomas. An implicit scheme for water wave problems. Applied Mathematics and Computation, 10/11(1):809–818, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Appelbaum:1985:HME

[AT85] Mel Appelbaum and Chris P. Tsokos. A heuristic method for estimating time-series models for forecasting. I. Applied Mathematics and Computation, 16(??):265–275, 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Avgerinos:1994:OCP

[Avg94] Evgenios P. Avgerinos. Optimal control problems involving optimal economic growth with infinite horizon. Applied Mathematics and Computation, 66(2–3): 293–311, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Aftabizadeh:1987:DID

[AW87] A. R. Aftabizadeh and Joseph Wiener. Differential inequalities for delay differential equations with piecewise constant argument. Applied Mathematics and Computation, 24(3):183– 194, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Agarwal:1999:ATT

[AW99a] Ravi P. Agarwal and Fu-Hsiang Wong. An application of topological transversality to nonpositive higher order difference equations. Applied Mathematics and Computation, 99(2– 3):167–178, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6189.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6189.

Agarwal:1999:ESB

[AW99b] Ravi P. Agarwal and Fu-Hsiang Wong. Existence of solutions to (k, n − k − 2) boundary value problems. Applied Mathematics and Computation, 104 (1):33–50, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/1/
6268.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=104&
issue=1&aid=6268.

Axelsson:1986:AIM

[Axe86] O. Axelsson. Analysis of incomplete matrix factorizations as multigrid smoothers for vector and parallel computers. Applied Mathematics and Computation, 19(1–4):3–22, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Aykut:1998:BVP

[AY98] Arzu Aykut and Bünyamin Yildiz. On a boundary value problem for a differential equation with variant retarded argument. Applied Mathematics and Computation, 93(1):63–71, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/93/1/6070.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=93&issue=
                                        1&aid=6070.

Ayoubi:1993:RGF

[Ayo93] I. S. Ayoubi. The Riemann-Green function and the invariant imbedding equations for hyperbolic systems of first-order. Applied Mathematics and Computation, 55(2–3):101–129, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Al-Zamel:1996:EHE

[AZK96] A. Al-Zamel and S. Kalla. Epstein–Hubbell elliptic-type integral and its generalizations. Applied Mathematics and Computation, 77(1):9–32, June 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        77&issue=1&aid=9500147.

Baaske:1988:PPV

[Baa88] Wolfgang Baaske. Principal planes: Visual structure analysis by cluster biplot projection pursuit in a multinational survey. Applied Mathematics and Computation, 26(4):303–314, 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Badiru:1993:NCS

[Bad93] Adedeji B. Badiru. A new computational search technique for AI based on the Cantor set. Applied Mathematics and Computation, 57(2–3):255–274, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bailey:1981:AST

[Bai81] Paul B. Bailey. The automatic solution of two-parameter Sturm–Liouville eigenvalue problems in ordinary differential equations. Applied Mathematics and Computation, 8(4):251–259, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bai:1991:PCM

[Bai91] Yin Bai. A perturbed collocation method for boundaryvalue problems in differentialalgebraic equations. Applied Mathematics and Computation, 45(3):269–291, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bai:1997:PHI

[Bai97] Zhong-Zhi Bai. Parallel hybrid iteration methods for block bordered linear systems. Applied Mathematics and Computation, 86(1):37–60, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1997/86/1/9600172.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1997&volume=86&issue=
1&aid=9600172.

Bai:1998:APN

[Bai98] Zhong-Zhi Bai. Asynchronous parallel nonlinear multisplitting relaxation methods for large sparse nonlinear complementarity problems. Applied Mathematics and Computation, 92(1):85–100, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/1/6037.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=92&issue=
1&aid=6037.

Bai:1999:AMA

[Bai99a] Zhong-Zhi Bai. Asynchronous multisplitting AOR methods for a class of systems of weakly nonlinear equations. Applied Mathematics and Computation, 98(1):49–59, January 1, 1999.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/1/6157.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
1&aid=6157.

Bai:1999:CPH

[Bai99b] Zhong-Zhi Bai. A class of parallel hybrid two-stage iteration methods for block bordered linear systems. Applied Mathematics and Computation, 101(2–3):245–267, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/101/2-
3/6230.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
101&issue=2-3&aid=6230.

Balakrishnan:1994:SCP

[Bal94] A. V. Balakrishnan. Shape control of plates with piezo actuators and collocated position/ rate sensors. Applied Mathematics and Computation, 63(2– 3):213–234, July 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barros:1988:PEU

[Bar88] Saulo R. M. Barros. The Poisson equation on the unit disk: a multigrid solver using polar coordinates. Applied Mathematics and Computation, 25(2 (part II)):123–135, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barthelson:1989:MOS

[Bar89] Stephen H. Barthelson. The modified operator strongly implicit scheme, MOSI, as a fast elliptic solver. Applied Mathematics and Computation, 33(3): 185–196, October 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barthelson:1991:ATD

[Bar91] Stephen H. Barthelson. Acceleration of a two-dimensional combustion code using MOSI to find the pressure and pressurecorrection fields for a SIMPLER based algorithm. Applied Mathematics and Computation, 46(3): 211–220, December 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bardet:1995:DSP

[Bar95] J. P. Bardet. The damping of saturated poroelastic soils during steady-state vibrations. Applied Mathematics and Computation, 67(1–3):3–31, January 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/

browse.cgi?year=1995&volume=
67&issue=1-3&aid=9400052.

Bayard:1993:SPS

[Bay93] D. S. Bayard. Statistical plant set estimation using Schroeder-Phased multisinusoidal input design. Applied Mathematics and Computation, 58(2–3):169– 198, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bayram:1997:NMA

[Bay97]  Mustafa Bayram.  A novel

method for analyzing enzyme kinetic systems. Applied Mathematics and Computation, 87(2– 3):161–174, December 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=2-3&aid=9600227.

Bayram:1998:ACAb

[Bay98a] Mustafa Bayram. Application of computer algebra techniques to affinity binding equations. Applied Mathematics and Computation, 94(1):83–90, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/1/6041.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
1&aid=6041.

Bayram:1998:ACAa

[Bay98b]  Mustafa Bayram.  Applica-

tion of computer algebra techniques to enzyme kinetics. Applied Mathematics and Computation, 94(1):73–81, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/94/1/6040.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=94&issue=
                                        1&aid=6040.

Bardossy:1983:NDS

[BB83a] A. Bardossy and I. Bogardi. Network design for the spatial estimation of environmental variables. Applied Mathematics and Computation, 12(??):339– 365, 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barkai:1983:VMP

[BB83b] D. Barkai and A. Brandt. Vectorized multigrid Poisson solver for the CDC Cyber 205. Applied Mathematics and Computation, 13(3–4):215–227, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bhatti:1996:SOT

[BB96] M. I. Bhatti and A. M. Barry. Some optimal tests for the equicorrelation coefficient in standard symmetric multivariate normal distribution. Applied Mathematics and Computation, 75(2–3):269–285, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500130.

Barry:1997:OII

[BBB97] A. M. Barry, S. M. A. Burney, and M. I. Bhatti. Optimum influence of initial observations in regression’ models with AR(2) errors. Applied Mathematics and Computation, 82(1): 57–65, March 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
82&issue=1&aid=9600024.

Bardossy:1991:FSP

[BBD91] A. Bárdossy, I. Bogardi, and L. Duckstein. Fuzzy set and probabilistic techniques for health-risk analysis. Applied Mathematics and Computation, 45(3):241–268, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bunch:1991:DFA

[BCG91] K. J. Bunch, W. N. Cain, and R. W. Grow. A data fitting approach to series convergence acceleration. Applied Mathematics and Computation, 42(2 (part II)):189–195, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bulgarelli:1988:NMM

[BCZ88] U. Bulgarelli, M. M. Cerimele, and A. Zaretti. A numerical method for a mathematical model in gravitational hydrodynamics. Applied Mathematics and Computation, 27(3 (part I)): 251–264, September 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bulgarelli:1989:NMR

[BCZ89] Ulderico Bulgarelli, Maria Merceda Cerimele, and Anna Zaretti. A numerical model related to an astrophysical problem. Applied Mathematics and Computation, 29(1 (part I)):17– 38, January 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bainov:1990:PDC

[BD90] D. D. Baĭnov and A. B. Dishliev. Population dynamics control in regard to minimizing the time necessary for the regeneration of a biomass taken away from the population. Applied Mathematics and Computation, 39(1):37–48, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boreux:1994:FAD

[BD94] Jean-Jacques Boreux and Lucien Duckstein. A fuzzy approach to the definition of standardized visibility in fog. Applied Mathematics and Computation, 61(2–3):287–299, March 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Banerjee:1995:DCC

[BD95] Arun K. Banerjee and Van N.

Do. Deployment control of a cable connecting a ship to an underwater vehicle. Applied Mathematics and Computation, 70(2– 3):97–116, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400124.

Dynamics and control (Vienna, 1993).

Bainov:1996:QOS

[BD96] D. D. Bainov and M. B. Dimitrova. Quickly oscillating solutions of operator-differential equations. Applied Mathematics and Computation, 75(1): 1–11, March 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=1&aid=9500065.

Ballesteros:1989:OCP

[BdA89] F. Ballesteros and F. de Arriaga. Optimal control of power systems and the Riccati equation. Applied Mathematics and Computation, 33(1 (part I)):43– 52, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bohm:1998:MBS

[BDJR98] Michael Böhm, Joseph Devinny, Fereidoun Jahani, and Gary Rosen. On a moving-boundary system modeling corrosion in sewer pipes. Applied Mathematics and Computation, 92 (2–3):247–269, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/92/2-3/
                                        9710039.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        92&issue=2-3&aid=9710039.

Bella:1996:MAW

[BDS96] Aimee Bella, Lucien Duckstein, and Ferenc Szidarovszky. A multicriterion analysis of the water allocation conflict in the upper Rio Grande basin. Applied Mathematics and Computation, 77(2–3):245–265, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bonney:1985:IPM

[BE85] George E. Bonney and Robert C.

Elston. Integrals of products of multinormal mixtures. Applied Mathematics and Computation, 16(2):93–104, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Beck:1985:LEI

[Bec85] M. B. Beck. Lake eutrophication: identification of tributary nutrient loading and sediment resuspension dynamics. Applied Mathematics and Computation, 17(4):433–458, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). System identification and parameter estimation of environmental systems.

Becarie:1998:IPS

[Béc98] Cécile Bécarie. Intrinsically parallel solution of systems of linear partial differential equations. Applied Mathematics and Computation, 90(2– 3):205–227, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/
9700399.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=2-3&aid=9700399.

Beesack:1987:SVP

[Bee87] Paul R. Beesack. On some variation of parameter methods for integro-differential, integral, and quasilinear partial integrodifferential equations. Applied Mathematics and Computation, 22(2–3):189–215, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Behforooz:1989:ECC

[Beh89] G. Behforooz. End conditions for cubic spline interpolation derived from integration. Applied Mathematics and Computation, 29(3):231–244, February 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Behforooz:1990:MSS

[Beh90] G. Behforooz. Manifold Subbotin spline. Applied Mathematics and Computation, 40(3): 225–231, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Behforooz:1992:KPI

[Beh92]  G. Behforooz.  The not-a-

knot piecewise interpolatory cubic polynomial. Applied Mathematics and Computation, 52(1): 29–35, November 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Behforooz:1995:CKC

[Beh95] G. Hossein Behforooz. A comparison of the E(3) and not-A-knot cubic splines. Applied Mathematics and Computation, 72(2–3):219–223, October 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400185.

Behforooz:1996:CRS

[Beh96a] G. Hossein Behforooz. Consistency relations of the spline functions derived from a Pascallike triangle. Applied Mathematics and Computation, 74(2–3): 293–297, February 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500105.

Behforooz:1996:USS

[Beh96b] G. Hossein Behforooz. The use of spline-on-spline for the approximation of Cauchy principal value integrals. Applied Mathematics and Computation, 80(1): 23–32, November 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=1&aid=9500277.

Beletsky:1995:SSP

[Bel95] V. V. Beletsky. Some stability problems in applied mechanics. Applied Mathematics and Computation, 70(2–3): 117–141, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        70&issue=2-3&aid=9400123.

Dynamics and control (Vienna, 1993).

Benaroya:1987:NSI

[Ben87] Haym Benaroya. Nonrecursive statistics for integral equation solutions. Applied Mathematics and Computation, 24(??):275– 280, 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Benaroya:1989:MCT

[Ben89] Haym Benaroya. Markov chain transition probabilities and experimental data. Applied Mathematics and Computation, 29(2 (part II)):107–121, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Benaroya:1992:REA

[Ben92] Haym Benaroya. Random eigenvalues, algebraic methods and structural dynamic models. Applied Mathematics and Computation, 52(1):37–66, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Berman:1989:EPE

[Ber89] Sorrell Berman. An example of propagation of error associated with an iterated least squares solution. Applied Mathematics and Computation, 34(3):173– 193, December 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Behie:1983:MST

[BF83] A. Behie and P. A. Forsyth, Jr. Multigrid solution of threedimensional problems with discontinuous coefficients. Applied Mathematics and Computation, 13(3–4):229–240, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bunch:1993:CCS

[BG93] K. J. Bunch and R. W. Grow. A compact cubic spline algorithm for converging to function minimums. Applied Mathematics and Computation, 56(1):29–34, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bomze:1994:DSE

[BG94a] Immanuel M. Bomze and Walter Gutjahr. The dynamics of self-evaluation. Applied Mathematics and Computation, 64 (1):47–63, ???? 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bunch:1994:DEM

[BG94b] K. J. Bunch and R. W. Grow. A difference equation model of partial series. Applied Mathematics and Computation, 61(2–3):

301–305, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bunch:1994:ETI

[BG94c] K. J. Bunch and R. W. Grow. An extrapolation technique to iterate to the smallest and largest eigenvalues of an infinitedimensional normal matrix used in function fitting. Applied Mathematics and Computation, 62(1):17–28, April 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bo:1997:SVC

[BG97] Tian Bo and Yi-Tian Gao. Solutions of a variable-coefficient Kadomtsev–Petviashvili equation via computer alegebra. Applied Mathematics and Computation, 84(2–3):125–130, July 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600115.

Blount:1997:NDP

[BGY97] S. Blount, A. Galambosi, and S. Yakowitz. Nonlinear and dynamic programming for epidemic intervention. Applied Mathematics and Computation, 86(2–3):123–136, October 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
86&issue=2-3&aid=9600177.

Burns:1980:DEA

[BH80] J. A. Burns and P. D. Hirsch.

A difference equation approach to parameter estimation for differential-delay equations. Applied Mathematics and Computation, 7(4):281–311, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Brown:1989:RSM

[BH89] Peter N. Brown and Alan C.

Hindmarsh. Reduced storage matrix methods in stiff ODE systems. Applied Mathematics and Computation, 31(Special issue):40–91, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Barron:1990:SMI

[BH90] R. M. Barron and M. H. Hamdan. The steady motion of an incompressible dusty gas in porous media. Applied Mathematics and Computation, 37 (3):149–167, June 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barron:1993:DLB

[BH93] R. M. Barron and M. H. Hamdan. On the Darcy–Lapwood– Brinkman–Saffman dusty fluid flow models through porous media. II. applications to flow into a two-dimensional sink. Applied Mathematics and Computation, 54(1):81–97, March 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bhatti:1994:AHE

[Bha94] M. Ishaq Bhatti. Assessment of hepatic extraction of insulin by testing the small sample mean correlation coefficient. Applied Mathematics and Computation, 66(1):77–??, ???? 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bogaert:1998:GRP

[BI98] J. Bogaert and I. Impens. Generating random percolation clusters. Applied Mathematics and Computation, 91 (2–3):197–208, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6018.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6018.

Billard:1996:SDG

[Bil96] Edward A. Billard. Stability of dynamic groups in distributed computing systems. Applied Mathematics and Computation, 74(2–3):223–248, February 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500094.

Byrd:1985:DPP

[BJ85] J. William Byrd and Robert R.

Jones. The determination of periodicity in physical problems from series-solution representation. Applied Mathematics and Computation, 16(3):203– 211, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Beddhu:1998:CSU

[BJTW98] Murali Beddhu, Min-Yee Jiang, Lafayette K. Taylor, and David L. Whitfield. Computation of steady and unsteady flows with a free surface around the Wigley hull. Applied Mathematics and Computation, 89(1– 3):67–84, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781648.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Brown:1980:EBR

[BK80] Donald J. Brown and M. Ali Khan. An extension of the Brown-Robinson equivalence theorem. Applied Mathematics and Computation, 6(2):167– 175, March 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Beutler:1999:SRS

[BK99] R. Beutler and B. G. Konopelchenko. Surfaces of revolution via the Schrödinger equation: Construction, integrable dynamics and visualization. Applied Mathematics and Computation, 101(1):13–43, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/101/1/
                                        6220.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=101&
                                        issue=1&aid=6220.

Buell:1979:BVP

[BKM79] J. Buell, R. Kalaba, and D. Marsh. On a boundary value problem in the theory of urine formation. Applied Mathematics and Computation, 5(1):55–67, ???? 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bainov:1994:FOI

[BKM94] Drumi Baĭnov, Zdzisław Kamont, and Emil Minchev. On first order impulsive partial differential inequalities. Applied Mathematics and Computation, 61(2–3):207–230, March 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bainov:1995:PBV

[BKM95] Drumi Baĭnov, Zdzisław Kamont, and Emil Minchev. Periodic boundary value problem for impulsive hyperbolic partial differential equations of first order. Applied Mathematics and Computation, 68(2–3):95– 104, March 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=2-3&aid=9400083.

Bharadwaj:1983:SMT

[BKS83] K. K. Bharadwaj, M. K. Kadalbajoo, and R. Sankar. Symmetric marching technique for the discretized Poisson equation. Applied Mathematics and Computation, 12(2–3):187–198, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bharadwaj:1984:SMTa

[BKS84a] Kamal K. Bharadwaj, Mohan K.

Kadalbajoo, and R. Sankar. Symmetric marching technique for the Poisson equation. I. Dirichlet boundary conditions. Applied Mathematics and Computation, 15(2):137–149, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bharadwaj:1984:SMTb

[BKS84b] Kamal K. Bharadwaj, Mohan K. Kadalbajoo, and R. Sankar. Symmetric marching technique for the Poisson equation. II. mixed boundary conditions. Applied Mathematics and Computation, 15(2):151–163, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bajpai:1998:BEE

[BKS98] R. P. Bajpai, S. Kumar, and V. A. Sivadasan. Biophoton emission in the evolution of a squeezed state of frequency stable damped oscillator. Applied Mathematics and Computation, 93(2–3):277–288, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/93/2-3/
                                        6072.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=93&issue=
                                        2-3&aid=6072.

Bernfeld:1978:NVC

[BL78] S. R. Bernfeld and M. E. Lord. A nonlinear variation of constants method for integro differential and integral equations. Applied Mathematics and Computation, 4(??):1–14, 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Breinl:1987:SFU

[BL87] W. Breinl and G. Leitmann.

State feedback for uncertain dynamical systems. Applied Mathematics and Computation, 22 (1):65–87, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Breinbauer:1991:KMI

[BL91] M. Breinbauer and P. Lory.

The kidney model as an inverse problem. Applied Mathematics and Computation, 44(3):195– 223, 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Blank:1996:SRC

[Bla96] Luise Blank. Stability results for collocation methods for Volterra integral equations. Applied Mathematics and Computation, 79(2–3):267–288, October 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500270.

Bainov:1989:MPM

[BM89] D. D. Baĭnov and S. D. Milusheva. The method of partially multiplicative averaging in the problems of terminal control. Applied Mathematics and Computation, 33(1 (part I)):69– 74, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Budd:1992:MTA

[BMS92] C. Budd, S. McKee, and D. Swailes. Modeling H^{+} and K^{+} transport across cell membranes. Applied Mathematics and Computation, 50(1):33–44, July 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Balendran:1995:IIC

[BNN95] B. Balendran and Sia Nemat-Nasser. Integration of inelastic constitutive equations for constant velocity gradient with large rotation. Applied Mathematics and Computation, 67 (1–3):161–195, January 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        67&issue=1-3&aid=9400058.

Bogobowicz:1991:TAM

[Bog91] Agnieszka Bogobowicz. Theoretical aspects of modeling and control of water quality in a river section. Applied Mathematics and Computation, 41(1 (part I)):35–60, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bownds:1978:MGA

[Bow78] John M. Bownds. A modified Galerkin approximation method for Volterra equations with smooth kernels. Applied Mathematics and Computation, 4(1):67–79, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bownds:1988:FSC

[Bow88] John M. Bownds. A formal solution to a classical initialboundary value problem in groundwater hydraulics. Applied Mathematics and Computation, 26(4):333–354, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boyd:1989:ACC

[Boy89] John P. Boyd. The asymptotic Chebyshev coefficients for functions with logarithmic endpoint singularities: mappings and singular basis functions. Applied Mathematics and Computation, 29(1 (part I)):49–67, January 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boyd:1995:LAG

[Boy95] John P. Boyd. A lag-averaged generalization of Euler’s method for accelerating series. Applied Mathematics and Computation, 72(2–3):143–166, October 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400180.

Boyd:1997:CLU

[Boy97a] John P. Boyd. Construction of Lighthill’s unitary functions: The imbricate series of unity. Applied Mathematics and Computation, 86(1):1–10, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1997/86/1/9600145.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1997&volume=86&issue=
1&aid=9600145.

Boyd:1997:PCW

[Boy97b] John P. Boyd. Peakons and coshoidal waves: Traveling wave solutions of the Camassa–Holm equation. Applied Mathematics and Computation, 81(2–3): 173–187, February 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=2-3&aid=9500326.

Bradley:1992:GSD

[BP92] Charles E. Bradley and Tom Price. Graduating sample data using generalized Weibull functions. Applied Mathematics and Computation, 50(2–3):115– ??, August 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Barrio:1999:CCM

[BPE99] Roberto Barrio, Manuel Palacios, and Antonio Elipe. Chebyshev collocation methods for fast orbit determination. Applied Mathematics and Computation, 99(2–3):195–207, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6187.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6187.

Baronio:1996:MSI

[BPZ96] A. Baronio, E. Loli Piccolomini, and F. Zama. A method for solving the indirect approximation problem. Applied Mathematics and Computation, 77 (2–3):97–107, July 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bao:1988:LCD

[BR88] Paul G. Bao and Jon G. Rokne.

Low complexity k-dimensional Taylor forms. Applied Mathematics and Computation, 27 (3 (part I)):265–280, September 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Baxley:1998:CUC

[BR98] J. V. Baxley and S. B. Robinson. Coexistence in the unstirred chemostat. Applied Mathematics and Computation, 89(1–3): 41–65, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781647.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Barrett:1999:ETC

[BR99a] C. L. Barrett and C. M. Reidys. Elements of a theory of computer simulation I: Sequential CA over random graphs. Applied Mathematics and Computation, 98(2–3):241–259, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/
6167.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6167.

Bonham:1999:ISA

[BR99b] Charles D. Bonham and Robin M. Reich. Influence of spatial autocorrelation on a fixed-effect model used to evaluate treatment of oil spills. Applied Mathematics and Computation, 106(2–3):149–162, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/

10/9/12/71/18/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/20/
article.pdf.

Bramley:1978:NSU

[Bra78] J. S. Bramley. The numerical solution of unstable ordinary differential equations. Applied Mathematics and Computation, 4(4):307–316, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Brackbill:1982:CSC

[Bra82]  J. U. Brackbill.  Coordi-

nate system control: adaptive meshes. Applied Mathematics and Computation, 10/11(1): 277–294, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Brandt:1986:AMT

[Bra86] Achi Brandt. Algebraic multigrid theory: the symmetric case. Applied Mathematics and Computation, 19(1–4):23– 56, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Breitenecker:1981:IIM

[Bre81]  F. Breitenecker.  Invariant

imbedding and multiple shooting for the solution of unstable linear boundary value problems. Applied Mathematics and Computation, 9(4):235–244, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Brezinski:1991:IJ

[Bre91] Claude Brezinski. Implementing the jackknife. Applied Mathematics and Computation, 42(2 (part II)):111–119, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bulgarelli:1986:NMM

[BS86] U. Bulgarelli and R. M. Spitaleri. A numerical method for the motion of a mixture of two viscous incompressible fluids. Applied Mathematics and Computation, 18(3):223– 237, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

binSuleiman:1989:SNH

[bS89] Mohamed bin Suleiman. Solving nonstiff higher order ODEs directly by the direct integration method. Applied Mathematics and Computation, 33(3): 197–219, October 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Boglaev:1993:CMS

[BS93] I. P. Boglaev and V. V. Sirotkin. Computational method for a singular perturbation problem via domain decomposition and its parallel implementation. Applied Mathematics and Computation, 56(1):71–95, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bai:1997:CCP

[BS97] Zhongzhi Bai and Yangfeng Su.

On the convergence of a class of parallel decomposition-type relaxation methods. Applied Mathematics and Computation, 81(1):1–21, January 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=1&aid=9500297.

Buetow:1998:MAF

[BS98] Gerald W. Buetow, Jr. and James S. Sochacki. A more accurate finite difference approach to the pricing of contingent claims. Applied Mathematics and Computation, 91(2–3):111–126, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6025.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6025.

Brugnano:1997:BBV

[BT97] L. Brugnano and D. Trigiante.

Block boundary value methods for linear Hamiltonian systems. Applied Mathematics and Computation, 81(1):49– 68, January 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=1&aid=9500308.

Bratsos:1998:FPF

[BT98] A. G. Bratsos and E. H. Twizell. A family of parametric finite-difference methods for the solution of the sine-Gordon equation. Applied Mathematics and Computation, 93(2–3):117–137, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/2-3/
6066.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6066.

Beddhu:1994:TAC

[BTW94] Murali Beddhu, Lafayette K. Taylor, and David L. Whitfield. A time accurate calculation procedure for flows with a free surface using a modified artificial compressibility formulation. Applied Mathematics and Computation, 65(1–3):33–48, September/October 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Bhargava:1977:SDG

[BU77] T. N. Bhargava and V. R. R.

Uppuluri. Sampling distribution of Gini’s index of diversity. Applied Mathematics and Computation, 3(??):1–24, 1977. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Burrage:1989:DSP

[Bur89] Kevin Burrage. The dichotomy of stiffness: pragmatism versus theory. Applied Mathematics and Computation, 31(Special issue):92–111, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Bursal:1996:IBP

[Bur96] Faruk H. Bursal. On interpolating between probability distributions. Applied Mathematics and Computation, 77(2–3): 213–244, ???? 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bownds:1977:NSV

[BW77] John M. Bownds and Bruce Wood. A note on solving Volterra integral equations with convolution kernels. Applied Mathematics and Computation, 3(??):307–315, 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Band:1988:SLS

[BW88] Lawrence E. Band and Eric F. Wood. Strategies for large-scale, distributed hydrologic simulation. Applied Mathematics and Computation, 27(1):23–37, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Bayram:1999:PEE

[BY99] Mustafa Bayram and Bunyamin Yildiz. Parameter estimation of the enzyme kinetic system using computer algebra techniques. Applied Mathematics and Computation, 99(1):93–98, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/1/6179.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
1&aid=6179.

Cabada:1994:MMF

[Cab94] Alberto Cabada. The monotone method for first-order problems with linear and nonlinear boundary conditions. Applied Mathematics and Computation, 63(2–3):163–186, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [Cab95].

Cabada:1995:ERM

[Cab95]  Alberto Cabada.  Erratum:

Remarks on “The monotone method for first-order problems with linear and nonlinear boundary conditions” [Appl. Math. Comput. 63 (1994), no. 2–3, 163–186, MR 95h:34032]. Applied Mathematics and Computation, 69(2–3): 353–354, May 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=2-3&aid=9400147.

See [Cab94].

Cai:1993:NEP

[Cai93] Zhi Qiang Cai. Norm estimates of product operators with application to domain decomposition. Applied Mathematics and Computation, 53(2–3): 251–276, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:1988:CMM

[Cao88] Zhi Hao Cao. Convergence of multigrid methods for nonsymmetric, indefinite problems. Applied Mathematics and Computation, 28(4 (part II)):269– 288, December 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cao:1998:NCQ

[Cao98] Zhi-Hao Cao. A note on convergence of quasi-minimal residual smoothing. Applied Mathematics and Computation, 93 (2–3):289–297, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/93/2-3/
                                        6081.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=93&issue=
                                        2-3&aid=6081.

Censor:1988:CSI

[CAP88] Yair Censor, Martin D. Altschuler, and William D. Powlis. A computational solution of the inverse problem in radiation-therapy treatment planning. Applied Mathematics and Computation, 25(1 (part I)):57–87, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:1993:NHM

[CAQ93] Dong Chen, I. K. Argyros, and Q. S. Qian. A note on the Halley method in Banach spaces. Applied Mathematics and Computation, 58(2–3):215–224, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Carmody:1988:CHQ

[Car88a] Kevin Carmody. Circular and hyperbolic quaternions, octonions, and sedenions. Applied Mathematics and Computation, 28(1):47–72, October 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Carmody:1988:FFS

[Car88b] Kevin Carmody. Finite Fourier series. Applied Mathematics and Computation, 28(2):89– 96, November 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Carmody:1990:MPF

[Car90] Kevin Carmody. Modular parts of a function. Applied Mathematics and Computation, 36 (1 (part I)):63–74, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Carmody:1997:CHQ

[Car97] Kevin Carmody. Circular and hyperbolic quaternions, octonions, and sedenions — further results. Applied Mathematics and Computation, 84 (1):27–47, June 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=1&aid=9600051.

Casulli:1981:NSF

[Cas81] Vincenzo Casulli. Numerical simulation of free-surface thermally influenced flows for nonhomogeneous fluids. Applied Mathematics and Computation, 8(4):261–280, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Casti:1983:FMH

[Cas83] J. Casti. Forest monitoring and harvesting policies. Applied Mathematics and Computation, 12(1):19–48, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Casti:1987:STN

[Cas87]  J. Casti.  Singularity the-

ory for nonlinear optimization problems. Applied Mathematics and Computation, 23(2):137– 161, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Casti:1988:TMR

[Cas88] John L. Casti. The theory of metabolism-repair systems. Applied Mathematics and Computation, 28(2):113–154, November 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cash:1989:CSG

[Cas89] J. R. Cash. A comparison of some global methods for solving two-point boundary value problems. Applied Mathematics and Computation, 31(Special issue):449–462, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Casti:1991:FBK

[Cas91] John L. Casti. Foreword. Bob Kalaba: teacher, colleague, and friend. Applied Mathematics and Computation, 45(2 (part II)): 85–88, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Castro:1994:BTR

[Cas94] Alfonso Castro. Bifurcation theory and radial solutions to elliptic boundary value problems. Applied Mathematics and Computation, 65(1–3):223–230, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Caughey:1983:MCT

[Cau83]  D. A. Caughey.  Multigrid

calculation of three-dimensional transonic potential flows. Applied Mathematics and Computation, 13(3–4):241–260, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cicci:1994:SEKa

[CB94a] David A. Cicci and Gary H.

Ballard. Sensitivity of an extended Kalman filter. I. variation in the number of observers and types of observations. Applied Mathematics and Computation, 66(2–3):233–246, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cicci:1994:SEKb

[CB94b] David A. Cicci and Gary H. Ballard. Sensitivity of an extended Kalman filter. II. variation in the observation error levels, observation rates, and types of observations. Applied Mathematics and Computation, 66(2–3): 247–259, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cho:1999:NAN

[CBYK99] Chung-Ki Cho, Guo Ben-Yu, and YongHoon Kwon. A new approach for numerical identification of conductivity. Applied Mathematics and Computation, 100(2–3):265–283, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6211.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6211.

Cavazos-Cadena:1994:CAC

[CC94] Rolando Cavazos-Cadena. Computing the asymptotic covariance matrix of a vector of sample autocorrelations for ARMA processes. Applied Mathematics and Computation, 64(2–3):121–137, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:1999:ECS

[CC99a] C. Y. Chan and W. Y. Chan.

Existence of classical solutions for degenerate semilinear parabolic problems. Applied Mathematics and Computation, 101(2–3):125–149, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/101/2-
3/6228.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
101&issue=2-3&aid=6228.

Choe:1999:SMS

[CC99b] Hi Jun Choe and Jeong Ho Chu. A spectral method for Stokes equations. Applied Mathematics and Computation, 104 (2–3):131–151, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/2-
3/6284.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
104&issue=2-3&aid=6284.

Chen:1982:DSF

[CCSK82] B. C.-J. Chen, T. H. Chien, W. T. Sha, and J. H. Kim. 3-D solution of flow in an infinite square array of circular tubes by using boundary-fitted coordinate system. Applied Mathematics and Computation, 10/11 (1):619–632, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Chin:1991:FAP

[CDW91] C. Chin, L. Duckstein, and L. Wymore. Factory automation project selection using multicriterion Q-analysis. Applied Mathematics and Computation, 46(2 (part II)):107–126, November 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cengiz:1997:SEH

[Cen97] Ahmet Cengiz. A study on the Epstein–Hubbell generalized elliptic-type integral using residue theory. Applied Mathematics and Computation, 83 (1):19–26, April 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        83&issue=1&aid=9600041.

Comincioli:1993:MMP

[CFT^{+}93] V. Comincioli, D. Funaro, A. Torelli, E. D’Angelo, and P. Rossi. A mathematical model of potential spreading along neuron dendrites of cerebellar granule cells. Applied Mathematics and Computation, 59(1):73– 87, November 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Coppin:1983:DMM

[CG83] C. Coppin and D. Greenspan.

Discrete modeling of minimal surfaces. Applied Mathematics and Computation, 13(1–2): 17–35, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1987:FEM

[CG87] Ching-Lung Chang and Max D.

Gunzburger. A finite element method for first order elliptic systems in three dimensions. Applied Mathematics and Computation, 23(2):171–184, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cai:1988:EES

[CG88a] Zhi Qiang Cai and Benju Guo.

Error estimates for a Schwarz alternating procedure on Lshaped regions. Applied Mathematics and Computation, 28 (1):39–46, October 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Coppin:1988:CPM

[CG88b] C. Coppin and D. Greenspan.

A contribution to the particle modeling of soap films. Applied Mathematics and Computation, 26(4):315–331, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Carl:1996:SMI

[CG96] S. Carl and C. Grossmann. Smoothing and monotone iterations for elliptic differential inclusions. Applied Mathematics and Computation, 74(1):15– 35, January 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=1&aid=9500070.

Chang:1998:PES

[CG98] Carolina Chang and Henryk Gzyl. Parameter estimation in superposition of decaying exponentials. Applied Mathematics and Computation, 96(2–3): 101–116, November 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/2-3/
6120.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
2-3&aid=6120.

Chen:1992:UCS

[CH92] Zheng Chen and Ping He. Using computers to seek the optimum combination of investment. Applied Mathematics and Computation, 49(2–3):197–203, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cicci:1993:OWP

[CH93] David A. Cicci and Roger L.

Hall. Optimal weighting of a priori statistics in linear estimation theory. Applied Mathematics and Computation, 57(2– 3):221–??, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:1996:ADT

[CH96] Chaó-Kuang Chen and Shing-Huei Ho. Application of differential transformation to eigenvalue problems. Applied Mathematics and Computation, 79(2–3): 173–188, October 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500253.

Chen:1999:SPD

[CH99] Cha’o Kuang Chen and Shing Huei Ho. Solving partial differential equations by twodimensional differential transform method. Applied Mathematics and Computation, 106 (2–3):171–179, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/22/
article.pdf.

Chang:1985:SIS

[Cha85] Mou Hsiung Chang. Stability of interconnected stochastic delay systems. Applied Mathematics and Computation, 16 (4):277–295, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1987:PSU

[Cha87] Feng Cheng Chang. Power-series unification and reversion. Applied Mathematics and Computation, 23(1):7–23, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1988:FEM

[Cha88a] Ching-Lung Chang. Finite element method for the solution of Maxwell’s equations in multiple media. Applied Mathematics and Computation, 25(2 (part II)):89–99, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1988:MDO

[Cha88b] Mou-Hsiung Chang. Martingale dynamics and optimal routing in a network. Applied Mathematics and Computation, 28 (4 (part II)):309–327, December 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1989:NFI

[Cha89a] Mou-Hsiung Chang. Nonlinear filtering for image restoration. Applied Mathematics and Computation, 33(2 (part II)): 131–147, September 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1989:SSS

[Cha89b] Y. F. Chang. Solving stiff systems by Taylor series. Applied Mathematics and Computation, 31(Special issue):251–269, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Chang:1990:MFE

[Cha90a] Ching Lung Chang. A mixed finite element method for the Stokes problem: an accelerationpressure formulation. Applied Mathematics and Computation, 36(2 (part II)):135–146, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1990:SME

[Cha90b] Feng Cheng Chang. Solution of Maxwell’s equations by matrix formulation. Applied Mathematics and Computation, 40 (3):239–262, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chankong:1991:OTP

[Cha91] Vira Chankong. Optimality tests for partitioning and sectional search algorithms. Applied Mathematics and Computation, 44(2 (part II)):157–182, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chaitin:1992:ITI

[Cha92a] G. J. Chaitin. Informationtheoretic incompleteness. Applied Mathematics and Computation, 52(1):83–101, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chaitin:1992:LPSa

[Cha92b] G. J. Chaitin. LISP programsize complexity. Applied Mathematics and Computation, 49 (1):79–93, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chaitin:1992:LPSb

[Cha92c] G. J. Chaitin. LISP programsize complexity. II. Applied Mathematics and Computation, 52(1):103–126, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chaitin:1992:LPSc

[Cha92d] G. J. Chaitin. LISP programsize complexity. III. Applied Mathematics and Computation, 52(1):127–139, November 1992.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chaitin:1992:LPSd

[Cha92e] G. J. Chaitin. LISP programsize complexity. IV. Applied Mathematics and Computation, 52(1):141–147, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chaitin:1993:NNB

[Cha93] G. J. Chaitin. Note on the number of N-bit strings with maximum complexity. Applied Mathematics and Computation, 59(1):97–100, November 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1994:MFP

[Cha94] Jin Jiang Chang. A method for finding the periodic solution of autonomous ordinary differential equations. Applied Mathematics and Computation, 66(2– 3):161–173, ???? 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1995:PLA

[Cha95a] Ching Lung Chang. Piecewise linear approach to the Stokes equations in 3-D. Applied Mathematics and Computation, 72 (1):61–75, September 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=1&aid=9400176.

Chang:1995:DBE

[Cha95b] Mou-Hsiung Chang. On duality between estimation and control for linear stochastic functional evolution equations in Hilbert spaces. Applied Mathematics and Computation, 68 (1):51–70, March 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=1&aid=9400089.

Chang:1996:LSF

[Cha96a] Ching Lung Chang. Leastsquares finite element for second order boundary value problem with optimal rates of convergence. Applied Mathematics and Computation, 76(2–3): 267–284, May 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=2-3&aid=9500164.

Chang:1996:LDN

[Cha96b] Mou-Hsiung Chang. Large deviation for Navier–Stokes equations with small stochastic perturbation. Applied Mathematics and Computation, 76(1): 65–93, April 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=1&aid=9500150.

Chen:1995:RCD

[Che95] Y. H. Chen. Robust control design for a class of mismatched uncertain nonlinear systems. Applied Mathematics and Computation, 70(2–3): 155–167, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400105.

Dynamics and control (Vienna, 1993).

Chen:1996:IMC

[Che96a] Yong-Lin Chen. Iterative methods for computing the generalized inverses A_{T,S}^{(2)} of a matrix A. Applied Mathematics and Computation, 75(2–3):207– 222, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500126.

Chernousko:1996:CES

[Che96b] F. L. Chernousko. Control of elastic systems by bounded distributed forces. Applied Mathematics and Computation, 78(2– 3):103–110, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        78&issue=2-3&aid=9600001.

Chen:1997:IMS

[Che97] Yong-Lin Chen. Iterative methods for solving restricted linear equations. Applied Mathematics and Computation, 86(2–3): 171–184, October 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        86&issue=2-3&aid=9600180.

Christov:1982:OCM

[Chr82] C. I. Christov. Orthogonal coordinate meshes with manageable Jacobian. Applied Mathematics and Computation, 10/11 (1):885–894, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Chukwu:1992:MCT

[Chu92] E. N. Chukwu. Mathematical controllability theory of capital growth of nations. Applied Mathematics and Computation, 52(2–3):317–344, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chukwu:1994:OCG

[Chu94] E. N. Chukwu. Optimal control of the growth of income of nations. Applied Mathematics and Computation, 62(2–3):279–309, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chukwu:1996:ULC

[Chu96] E. N. Chukwu. Universal laws for the control of global economic growth with nonlinear hereditary dynamics. Applied Mathematics and Computation, 78(1):19–81, August 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=1&aid=9500155.

Chukwu:1997:CND

[Chu97] E. N. Chukwu. Control of nonlinear delay-differential equations in W_{2}^{(1)} with economic applications. Applied Mathematics and Computation, 85 (1):17–59, August 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1997/85/1/9600108.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1997&volume=85&issue=
1&aid=9600108.

Chukwu:1998:CNE

[Chu98] Ethelbert N. Chukwu. On the controllability of nonlinear economic systems with delay: The Italian example. Applied Mathematics and Computation, 95 (2–3):245–274, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/free/1998/95/2-
3/6112/menu.htm;   http://
www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6112.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6112.

Cicci:1992:AHT

[Cic92] David A. Cicci. Applications of the Householder transformation to ridge-type estimation methods. Applied Mathematics and Computation, 51(2–3): 159–??, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cicci:1993:CBP

[Cic93a] David A. Cicci. Comparison of biasing parameter computational techniques in ridge-type estimation. Applied Mathematics and Computation, 53(2–3): 111–??, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cicci:1993:SRT

[Cic93b] David A. Cicci. Sensitivity of ridge-type estimation methods to observation accuracy and sampling rate. Applied Mathematics and Computation, 57(2– 3):233–??, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cicci:1999:FPT

[Cic99] David A. Cicci. Filter performance in target tracking using space-based observers. Applied Mathematics and Computation, 99(2–3):275–293, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6191.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6191.

Casey:1988:PVP

[CJ88] Richard M. Casey and Donald A. Jameson. Parallel and vector processing in landscape dynamics. Applied Mathematics and Computation, 27(1):3–22, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ciucci:1993:NAU

[CJF93] A. Ciucci, Rhonald M. Jenkins, and Winfred A. Foster, Jr. Numerical analysis of an unsteady confined supersonic jet. Applied Mathematics and Computation, 57(2–3):245–254, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Calude:1994:IE

[CJZ94] C. Calude, H. Jürgensen, and M. Zimand. Is independence an exception? Applied Mathematics and Computation, 66 (1):63–76, ???? 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:1979:NSA

[CK79] W. K. Chan and K. S. P. Kumar. Nonlinear smoothing: approximate algorithms. Applied Mathematics and Computation, 5(1):1–22, ???? 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1986:SAA

[CK86] M. H. Chang and K. Krishna. A successive approximation algorithm for stochastic control problems. Applied Mathematics and Computation, 18(2):155– 165, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Choe:1990:VAP

[CK90] W. H. Choe and K. Kim. Variational analysis of a plasma-arc armature inside a railgun solidhydrogen-pellet accelerator for fueling fusion reactors. Applied Mathematics and Computation, 37(1 (part I)):37–56, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:1995:TEM

[CK95] C. Y. Chan and P. C. Kong. A thermal explosion model. Applied Mathematics and Computation, 71(2–3):201–210, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.elsevier.com/cgi-bin/
cas/tree/store/amc/cas_sub/
browse/browse.cgi?year=1995&
volume=71&issue=2-3&aid=9400154.

Chen:1999:GNR

[CK99] X. B. Chen and M. M. Kostreva.

A generalization of the normrelaxed method of feasible directions. Applied Mathematics and Computation, 102(2– 3):257–272, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6250.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6250.

Chen:1998:DTT

[CL98a] Chieh-Li Chen and Yung-Chin Liu. Differential transformation technique for steady nonlinear heat conduction problems. Applied Mathematics and Computation, 95(2–3): 155–164, September 15, 1998.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6105.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6105.

Chen:1998:PPN

[CL98b] Jinru Chen and Likang Li. Preconditioning projection nonconforming element method for the lowest-order Raviart– Thomas mixed triangular element method. Applied Mathematics and Computation, 93(1): 31–49, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.elsevier.
com/cas/tree/store/amc/sub/
1998/93/1/6067.pdf;   http:
//www.elsevier.com/cgi-bin/
cas/tree/store/amc/cas_sub/
browse/browse.cgi?year=1998&
volume=93&issue=1&aid=6067.

Cao:1999:CRP

[CL99] Zhi-Hao Cao and Zhong-Yun Liu. Convergence of relaxed parallel multisplitting methods with different weighting schemes. Applied Mathematics and Computation, 106 (2–3):181–196, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/23/abstract.
html;   http://www.elsevier.

nl/gej-ng/10/9/12/71/18/23/
article.pdf.

Chen:1997:EBE

[CLC97] Cha’o-Kuang Chen, Jin-Mu Lin, and Chieh-Li Chen. Error bounds estimate of weighted residuals method using genetic algorithms. Applied Mathematics and Computation, 81(2–3): 207–219, February 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=2-3&aid=9600017.

Conrad:1990:BCP

[CM90] M. Conrad and K. Matsuno.

The boundary condition paradox: a limit to the universality of differential equations. Applied Mathematics and Computation, 37(1 (part I)):67–74, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cordero:1996:GQP

[CMM96] M. Cordero, C. Martin, and J. Miller. Gaussian quadrature for products of exponential functions. Applied Mathematics and Computation, 79(2–3): 189–202, October 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500254.

Cabada:1990:ESS

[CN90] A. Cabada and Juan J. Nieto. Extremal solutions of second order nonlinear periodic boundary value problems. Applied Mathematics and Computation, 40(2 (part II)):135–145, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cabada:1997:RCI

[CN97] Alberto Cabada and Juan J. Nieto. Rapid convergence of the iterative technique for first order initial value problems. Applied Mathematics and Computation, 87(2–3):217–226, December 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        87&issue=2-3&aid=9600285.

Cicci:1996:FSE

[CO96] David A. Cicci and David L. Oakley. Filter sensitivity in exoatmospheric target vehicle tracking. Applied Mathematics and Computation, 79(2–3): 249–265, October 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        79&issue=2-3&aid=9500269.

Coleman:1982:GBF

[Col82] Roderick M. Coleman. Generation of boundary-fitted coordinate systems using segmented computational regions. Applied Mathematics and Computation, 10/11(1):633–651, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Conrad:1989:PBT

[Con89] Michael Conrad. Physics and biology: Towards a unified model. Applied Mathematics and Computation, 32(2–3):75–102, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Mathematical biology.

Conrad:1993:RLR

[Con93a] Deborah Conrad. Rules, laws, and reductionism. Applied Mathematics and Computation, 56(2–3):341–??, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Conrad:1993:B

[Con93b]  Michael Conrad.  Biophysi-

calism. Applied Mathematics and Computation, 56(2–3): 103–??, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Conrad:1993:FMF

[Con93c] Michael Conrad. The fluctuon model of force, life, and computation: A constructive analysis. Applied Mathematics and Computation, 56(2–3):203–259, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Corliss:1989:SIA

[Cor89] George F. Corliss. Survey of interval algorithms for ordinary differential equations. Applied Mathematics and Computation, 31(Special issue):112–120, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Coskun:1999:CSF

[Cos99]  Erhan Coskun.  Computa-

tional simulation of flux trapping and vortex pinning in Type-II superconductors. Applied Mathematics and Computation, 106(1):31–49, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/17/19/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/17/19/
article.pdf.

Crain:1979:EPT

[Cra79] Bradford R. Crain. Estimating the parameters of truncated normal distribution. Applied Mathematics and Computation, 5(2):149–156, ???? 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Crown:1980:NPA

[Cro80] J. C. Crown. A new PEC algorithm for the numerical solution of ordinary differential equations. Applied Mathematics and Computation, 6(3):189–209, April 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Crombez:1993:WNC

[Cro93] G. Crombez. Weak and norm convergence of a parallel projection method in Hilbert spaces. Applied Mathematics and Computation, 56(1):35–48, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Crombez:1995:PPM

[Cro95] G. Crombez. A parallel projection method based on sequential most remote set in convex feasibility problems. Applied Mathematics and Computation, 72 (2–3):113–124, October 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400177.

Choi:1999:SHO

[CRS99] Junesang Choi, Arjun K.

Rathie, and H. M. Srivastava. Some hypergeometric and other evaluations of ζ(2) and allied series. Applied Mathematics and Computation, 104 (2–3):101–108, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/2-
3/6303.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
104&issue=2-3&aid=6303.

Carrera:1985:NCN

[CS85] Jesus Carrera and Ferenc Szidarovszky. Numerical comparison of network design algorithms for regionalized variables. Applied Mathematics and Computation, 16(??):189–202, 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chow:1988:BSV

[CS88] Shui-Nee Chow and Yun Qiu Shen. Bifurcations via singular value decompositions. Applied Mathematics and Computation, 28(3 (part I)):231–245, December 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:1992:ETD

[CS92] William M. Chan and Joseph L. Steger. Enhancements of a three-dimensional hyperbolic grid generation scheme. Applied Mathematics and Computation, 51(2–3):181–205, October 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chatterjea:1993:UPC

[CS93] S. K. Chatterjea and H. M. Srivastava. A unified presentation of certain operational formulas for the Jacobi and related polynomials. Applied Mathematics and Computation, 57(1): 77–95, August 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:1995:ORG

[CS95] Ming Po Ch’ên and H. M.

Srivastava. Orthogonality relations and generating functions for Jacobi polynomials and related hypergeometric functions. Applied Mathematics and Computation, 68(2–3):153– 188, March 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=2-3&aid=9400092.

Chou:1996:MFE

[CS96] Chi-Chin Chou and William A.

Sethares. Multiplication-free evaluation of polynomials via a stochastic Bernstein representation. Applied Mathematics and Computation, 79(1):1– 25, September 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=1&aid=9500156.

Chen:1997:FCO

[CS97] Ming-Po Chen and H. M. Srivastava. Fractional calculus operators and their applications involving power functions and summation of series. Applied Mathematics and Computation, 81(2–3):287–304, February 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        81&issue=2-3&aid=9500310.

Chang:1998:MQE

[CS98] Feng Cheng Chang and Ching-Tzong Su. More on quick evaluation of determinants. Applied Mathematics and Computation, 93(1):97–99, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/93/1/6049.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=93&issue=
                                        1&aid=6049.

Chiue:1999:CCP

[CS99] Wen-Sz Chiue and Bih-Sheue Shieh. On connectivity of the Cartesian product of two graphs. Applied Mathematics and Computation, 102(2– 3):129–137, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6264.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6264.

Castillo:1988:PEV

[CSR88] José Castillo, Stanly Steinberg, and Patrick J. Roache. Parameter estimation in variational grid generation. Applied Mathematics and Computation, 28(2):155– 177, November 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:1998:SOF

[CSY98] Ming-Po Chen, H. M. Srivastava, and Ching-Shu Yu. Some operators of fractional calculus and their applications involving a novel class of analytic functions. Applied Mathematics and Computation, 91 (2–3):285–296, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6027.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6027.

Choi:1998:IIF

[CSZ98] Junesang Choi, H. M. Srivastava, and Nan-Yue Zhang. Integrals involving a function associated with the Euler–Maclaurin summation formula. Applied Mathematics and Computation, 93(2–3):101–116, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/2-3/
6060.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6060.

Chen:1996:NFM

[CT96] Cha’o-Kuang Chen and Tzu-Li Tien. A new forecasting method for time continuous model of dynamic system. Applied Mathematics and Computation, 80(2– 3):225–244, December 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500296.

Chen:1997:NFM

[CT97] Cha’o-Kuang Chen and Tzu-Li Tien. A new forecasting method of discrete dynamic system. Applied Mathematics and Computation, 86(1):61–84, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1997/86/1/9600173.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1997&volume=86&issue=
1&aid=9600173.

Chukwu:1992:GCN

[CUK92] E. N. Chukwu, Robert G. Underwood, and Louis D. Kovari. Global constrained nullcontrollability of nonlinear neutral systems. Applied Mathematics and Computation, 49(1): 95–110, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Cannings:1989:PIE

[CV89] C. Cannings and G. T. Vickers. Patterns and invasions of evolutionarily stable strategies. Applied Mathematics and Computation, 32(2–3):227–253, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Mathematical biology.

Casti:1984:SQR

[CW84] J. Casti and E. F. Wood.

Some questions of reachability in natural-resource management. Applied Mathematics and Computation, 15(3):185– 207, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Colquitt:1989:PRI

[CW89] S. D. Colquitt and Jack Williams. Polynomial and rational interpolation in the numerical solution of stiff systems. Applied Mathematics and Computation, 31(Special issue): 270–287, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Chan:1991:CMT

[CW91] C. Y. Chan and Benedict M. Wong. Computational methods for time-periodic solutions of singular semilinear parabolic problems. Applied Mathematics and Computation, 42(3):287– 312, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:1992:CEA

[CW92] C. Y. Chan and Benedict M. Wong. Computational error analysis of periodic solutions for singular semilinear parabolic problems. Applied Mathematics and Computation, 50(2–3): 279–306, August 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Charafi:1997:NAR

[CW97] A. Charafi and L. C. Wrobel. A new h-adaptive refinement scheme for the boundary element method using local reanalysis. Applied Mathematics and Computation, 82(2–3): 239–271, March 15, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
82&issue=2-3&aid=9600028.

Chang:1992:NIF

[CWL92] Qian Shun Chang, Yau Shu Wong, and Zheng Feng Li. New interpolation formulas of using geometric assumptions in the algebraic multigrid method. Applied Mathematics and Computation, 50(2–3):223–254, August 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chang:1994:AEC

[CXL94] Ching L. Chang, Xiaoqing Xiang, and Jianshi Li. An analysis of the eddy current problem by the least-squares finite element in 2-D. Applied Mathematics and Computation, 60(2– 3):179–191, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chen:1980:DST

[CY80] K. H. Chen and C. C. Yang.

On n-dimensional sampling theorems. Applied Mathematics and Computation, 7(3):247–257, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:1998:IEG

[CY98] C. Y. Chan and S. I. Yuen. Impulsive effects on global existence of solutions for degenerate semilinear parabolic equations. Applied Mathematics and Computation, 90(2–3):97–116, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/
9600353.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=2-3&aid=9600353.

Chang:1999:EOH

[CY99] Mou-Hsiung Chang and Roger K. Youree. The European option with hereditary price structures: Basic theory. Applied Mathematics and Computation, 102(2–3):279–296, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6258.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6258.

Cha:1999:PES

[CyGK99] Ki Hyun Cha, Ben yu Guo, and Yong Hoon Kwon. Parameter estimation by spectral approximation. Applied Mathematics and Computation, 104 (1):1–14, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/1/
6277.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=104&
issue=1&aid=6277.

Chronopoulos:1992:IMN

[CZ92] A. T. Chronopoulos and Z. Zlatev.

Iterative methods for nonlinear operator equations. Applied Mathematics and Computation, 51(2–3):167–180, October 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chan:1999:BSS

[CZ99] C. Y. Chan and J. K. Zhu.

Blow-up of solutions of semilinear Euler–Poisson–Darboux equations with nonlocal boundary conditions. Applied Mathematics and Computation, 99 (1):17–28, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/1/6172.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
1&aid=6172.

Da:1993:FEM

[Da93a] Xu Da. Finite element methods of the two nonlinear integrodifferential equations. Applied Mathematics and Computation, 58(2–3):241–273, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Da:1993:NSI

[Da93b] Xu Da. Non-smooth initial data error estimates with the weight norms for the linear finite element method of parabolic partial differential equations. Applied Mathematics and Computation, 54(1):1–24, March 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Da:1993:DTPa

[Da93c] Xu Da. On the discretization in time for a parabolic integrodifferential equation with a weakly singular kernel. I. smooth initial data. Applied Mathematics and Computation, 58(1):1– 27, September 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Da:1993:DTPb

[Da93d] Xu Da. On the discretization in time for a parabolic integrodifferential equation with a weakly singular kernel. II. nonsmooth initial data. Applied Mathematics and Computation, 58(1): 29–60, September 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

deAlmeida:1989:DMM

[dAMY89] A. R. C. de Almeida, I. M. MacLeod, and T. J. Ypma. Distributed-multirate methods for large weakly-coupled differential systems. Applied Mathematics and Computation, 31 (Special issue):18–39, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Davis:1994:FDS

[Dav94] Stephen F. Davis. Flux difference splittings and limiters for the resolution of contact discontinuities. Applied Mathematics and Computation, 65(1–3): 3–18, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Davis:1998:SCP

[Dav98] Stephen F. Davis. Shock capturing with Padé methods. Applied Mathematics and Computation, 89(1–3):85–98, January 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781649.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Duckstein:1988:OKB

[DBW88] L. Duckstein, P. H. Bartels, and J. E. Weber. Organization of a knowledge base by Q-analysis. Applied Mathematics and Computation, 26(4):289– 301, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ding:1993:HOA

[DDL93] J. Ding, Q. Du, and T. Y. Li. High order approximation of the Frobenius–Perron operator. Applied Mathematics and Computation, 53(2–3):151–171, February 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Davis:1985:OLD

[DDS85] D. Davis, L. Duckstein, and F. Szidarovszky. Opportunity losses due to uncertainty in multiobjective decision problems. Applied Mathematics and Computation, 17(1):69–91, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

DeKlerk:1994:SCA

[De 94] J. H. De Klerk. The sensitivity of a computational L_{1}- Approximation technique for integral equations. Applied Mathematics and Computation, 62(2– 3):171–182, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Degot:1996:MPC

[Dég96] Jérôme Dégot. Massively parallel computation of manyvariable polynomial interpolation. Applied Mathematics and Computation, 80(2–3):105– 125, December 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500283.

Dendy:1983:BBM

[Den83] J. E. Dendy, Jr. Black box multigrid for nonsymmetric problems. Applied Mathematics and Computation, 13(3–4):261– 283, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dendy:1986:BBM

[Den86] J. E. Dendy, Jr. Black box multigrid for systems. Applied Mathematics and Computation, 19(1– 4):57–74, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Dendy:1988:BBM

[Den88] J. E. Dendy, Jr. Black box multigrid for periodic and singular problems. Applied Mathematics and Computation, 25(1 (part I)):1–10, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dendrinos:1990:SDS

[Den90] Dimitrios S. Dendrinos. Spatially distributed stocks, deterministic evolution, and ecological balance. Applied Mathematics and Computation, 36(2 (part II)):113–133, March 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

DeKlerk:1995:AMN

[DEV95] J. H. De Klerk, D. Eyre, and L. M. Venter. L_{p}-approximation method for the numerical solution of singular integral equations. Applied Mathematics and Computation, 72(2–3):285– 300, October 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9500072.

Dolezal:1996:NSD

[DF96] Jaroslav Doležal and Jiří Fidler. Numerical solution of dynamic optimization problems using parameterization and op^{ti}a software. Applied Mathematics and Computation, 78(2–3):111– 121, September 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600002.

deFalco:1994:RRR

[dFG94] D. de Falco, M. Frontini, and L. Gotusso. Recent results on the regularization of Fourier polynomials. Applied Mathematics and Computation, 66(1):1–8, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Doedel:1994:LCO

[DFM94] Eusebius J. Doedel, Mark J.

Friedman, and Anand C. Monteiro. On locating connecting orbits. Applied Mathematics and Computation, 65(1–3):231–239, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Deckro:1987:DSF

[DG87a] Richard F. Deckro and Thomas M.

Gunther. Drilling strategies for frontier areas. Applied Mathematics and Computation, 23 (??):71–91, 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dror:1987:ISM

[DG87b] Moshe Dror and Saul I. Gass.

Interactive scheme for a MOLP problem given two partial orders: one on variables and one on objectives. Applied Mathematics and Computation, 24 (3):195–209, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duckstein:1993:PSI

[DGP93] Lucien Duckstein, Ambrose Goicoechea, and Eric Parent. Preface to the special issue on multicriterion decision making with engineering applications. Applied Mathematics and Computation, 54(2–3):99–??, March 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duchting:1996:CSA

[DGU96] Werner Duchting, Thomas Ginsberg, and Waldemar Ulmer. Computer simulation applied to radiation therapy in cancer research. Applied Mathematics and Computation, 74(2–3):191– 207, February 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500092.

Daemen:1994:ISI

[DGV94] Joan Daemen, René Govaerts, and Joos Vandewalle. Invertible shift-invariant transformations on binary arrays. Applied Mathematics and Computation, 62(2–3):259–277, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Daripa:1999:NSI

[DH99a] Prabir Daripa and Wei Hua. A numerical study of an illposed Boussinesq equation arising in water waves and nonlinear lattices: Filtering and regularization techniques. Applied Mathematics and Computation, 101(2–3):159–207, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/101/2-
3/6292.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
101&issue=2-3&aid=6292.

Dror:1999:SRU

[DH99b] Moshe Dror and Bruce C. Hartman. Stopping rules for utility functions and the St. Petersburg gamble. Applied Mathematics and Computation, 98(2– 3):279–291, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/
6170.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6170.

De:1990:AHP

[DHK90] Mitali De, Keith W. Hipel, and D. Marc Kilgour. Algorithms for hierarchical power. Applied Mathematics and Computation, 39(1):21–36, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

DiBella:1998:CPJ

[Di 98] Beatrice Di Bella. Comment on a paper by J. M. Soriano: “On the existence of zero points” [Appl. Math. Comput. 79 (1996), no. 1, 99–104, MR 97f:26015]. Applied Mathematics and Computation, 92(2–3):297– 298, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Sor96].

Dieci:1989:SAU

[Die89]  Luca Dieci.  Some aspects

of using IV software to solve BVP via a Riccati transformation. Applied Mathematics and Computation, 31(Special issue): 463–472, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Dietzenbacher:1996:AFB

[Die96] Erik Dietzenbacher. An algorithm for finding blocktriangular forms. Applied Mathematics and Computation, 76(2– 3):161–171, May 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        76&issue=2-3&aid=9500157.

Ding:1998:MEM

[Din98] J. Ding. A maximum entropy method for solving Frobenius– Perron operator equations. Applied Mathematics and Computation, 93(2–3):155–168, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/

store/amc/sub/1998/93/2-3/
6056.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6056.

Deeba:1996:DMA

[DK96] E. Y. Deeba and S. A. Khuri.

The decomposition method applied to Chandrasekhar Hequation. Applied Mathematics and Computation, 77(1): 67–78, June 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
77&issue=1&aid=9500188.

deBarros:1995:SDL

[dL95] F. C. P. de Barros and J. E.

Luco. Stresses and displacements in a layered half-space for a moving line load. Applied Mathematics and Computation, 67(1–3):103–134, January 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
67&issue=1-3&aid=9400056.

Denman:1981:SDM

[DLR81] E. D. Denman and J. Leyva-Ramos. Spectral decomposition of a matrix using the generalized sign matrix. Applied Mathematics and Computation, 8(4):237–250, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Denman:1985:CRR

[DLR85] E. D. Denman and J. Leyva-Ramos. Computation of roots of real and complex matrices. Applied Mathematics and Computation, 16(3):213–228, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Denman:1981:ATL

[DLRJ81] E. D. Denman, J. Leyva-Ramos, and G. J. Jeon. The algebraic theory of latent projectors in lambda matrices. Applied Mathematics and Computation, 9(4):273–300, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dunyak:1998:AIS

[DML98] James Dunyak, Clyde Martin, and Richard Lampe. Analysis of the influence of social structure on a measles epidemic. Applied Mathematics and Computation, 92(2–3):283–296, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/92/2-3/
                                        9710054.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        92&issue=2-3&aid=9710054.

Do:1996:CAM

[Do96] Y. Kwak Do. A p-cycle analysis of multigrid method. Applied Mathematics and Computation, 76(1):41–48, April 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=1&aid=9500148.

Doan:1990:NEV

[Doa90] Khanh Bui Doan. A numerical evaluation of very oscillatory infinite integrals by asymptotics. Applied Mathematics and Computation, 40(2 (part II)): 117–124, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Doan:1991:CFD

[Doa91] Khanh Bui Doan. A computation of Fermi–Dirac integrals by asymptotics and the Hermite corrector formula. Applied Mathematics and Computation, 41(1 (part I)):61–68, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

deOliveira:1989:CCS

[dOP89] Paula de Oliveira and Fernanda Patrício. A class of contractive second derivatives multistep methods. Applied Mathematics and Computation, 31(Special issue):170–182, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Dostal:1990:PPD

[Dos90] Zdeněk Dostál. Projector preconditioning and domain decomposition methods. Applied Mathematics and Computation, 37(2 (part II)):75–81, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

dePalma:1984:BBC

[dPL84] André de Palma and Claude Lefèvre. Bifurcation and behavior of complex systems. Applied Mathematics and Computation, 14(1):77–95, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Djidjeli:1998:PIS

[DPTT98] K. Djidjeli, W. G. Price, P. Temarel, and E. H. Twizell. Partially implicit schemes for the numerical solutions of some non-linear differential equations. Applied Mathematics and Computation, 96(2–3): 177–207, November 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/2-3/
6129.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.

cgi?year=1998&volume=96&issue=
2-3&aid=6129.

Devasahayam:1988:MME

[DR88] M. Paul Devasahayam and V. Subba Rao. A monotone method for the equation u_{xyz} = f. Applied Mathematics and Computation, 25(3): 175–186, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dror:1990:CAT

[Dro90]  Moshe Dror.  Cost alloca-

tion: The traveling salesman, bin packing, and the knapsack. Applied Mathematics and Computation, 35(2 (part II)):191– 207, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dongarra:1986:LAH

[DS86] J. J. Dongarra and D. C.

Sorensen. Linear algebra on high performance computers. Applied Mathematics and Computation, 20(1–2):57–88, ?? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dendrinos:1987:OTD

[DS87] Dimitrios S. Dendrinos and Michael Sonis. The onset of turbulence in discrete relative multiple spatial dynamics. Applied Mathematics and Computation, 22(1):25–44, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dendrionos:1988:NDR

[DS88] Dimitrios S. Dendrionos and Michael Sonis. Nonlinear discrete relative population dynamics of the U.S. regions. Applied Mathematics and Computation, 25(4):265–285, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dziok:1999:CAF

[DS99] J. Dziok and H. M. Srivastava. Classes of analytic functions associated with the generalized hypergeometric function. Applied Mathematics and Computation, 103(1):1–13, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/103/1/
6265.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=103&
issue=1&aid=6265.

Dwyer:1982:AGF

[DSK82] Harry A. Dwyer, Mitchell D. Smooke, and Robert J. Kee. Adaptive gridding for finite difference solutions to heat and mass transfer problems. Applied Mathematics and Computation, 10/11(1):339–356, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Deo:1987:GVC

[DT87] Sadashiv G. Deo and Euclides F.

Torres. Generalized variation-of-constants formula for nonlinear functional-differential equations. Applied Mathematics and Computation, 24(3):263–274, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dinkel:1988:INM

[DTW88] John J. Dinkel, Marietta Tretter, and Danny Wong. Interval Newton methods and perturbed problems. Applied Mathematics and Computation, 28(3 (part I)): 211–222, December 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Du:1983:MIT

[Du83] S. W. Du. Monotone iterative technique for delay differential equations in abstract cones. Applied Mathematics and Computation, 12(2–3):213–219, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Duan:1999:RTD

[Dua99a] Jinqiao Duan. A remark on the three dimensional baroclinic quasi-geostrophic dynamics. Applied Mathematics and Computation, 106(2–3): 285–288, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/32/
article.pdf.

Duan:1999:TPQ

[Dua99b] Jinqiao Duan. Time-periodic quasigeostrophic motion under dissipation and forcing. Applied Mathematics and Computation, 102(2–3):121–127, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6257.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6257.

Dulikravich:1982:FGT

[Dul82] Djordje S. Dulikravich. Fast generation of three-dimensional computational boundary-conforming periodic grids of C-type. Applied Mathematics and Computation, 10/11(1):563–584, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Dang-Vu:1995:HBS

[DVD95] H. Dang-Vu and C. Delcarte. Hopf bifurcation and strange attractors in Chebyshev spectral solutions of the Burgers equation. Applied Mathematics and Computation, 73(2–3): 99–113, December 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=2-3&aid=9400242.

deSampaio:1997:TRA

[dYS97] R. J. B. de Sampaio, Jin-Yun Yuan, and Wen-Yu Sun. Trust region algorithm for nonsmooth optimization. Applied Mathematics and Computation, 85(2– 3):109–116, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600112.

Ding:1999:FEM

[DZ99] Jiu Ding and Aihui Zhou.

A finite element method for the Frobenius–Perron operator equation. Applied Mathematics and Computation, 102 (2–3):155–164, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6262.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6262.

El-Attar:1994:CAP

[EA94] M. El-Attar. Convergence analysis for a parallel Jacobi algorithm. Applied Mathematics and Computation, 62(2–3):191–201, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Engle:1986:MSF

[EB86] Robert F. Engle and Scott J. Brown. Model selection for forecasting. Applied Mathematics and Computation, 20(??):313– 327, 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Estalrich:1991:ASS

[EB91] Juan Estalrich and Nathan Buras. Alternative specifications of state variables in stochasticdynamic-programming models of reservoir operation. Applied Mathematics and Computation, 44(2 (part II)):143–155, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Bassiouny:1999:RTD

[EB99] A. F. El-Bassiouny. Response of a three-degree-of-freedom system with cubic non-linearities to harmonic excitation. Applied Mathematics and Computation, 104(1):65–84, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/104/1/
                                        6273.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.

cgi?year=1999&volume=104&
issue=1&aid=6273.

Ebeling:1993:COI

[Ebe93] Werner Ebeling. Chaos, order, and information in the evolution of strings. Applied Mathematics and Computation, 56 (2–3):131–151, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eno:1985:SAE

[EBR85] Larry Eno, Johan G. B. Beumee, and Herschel Rabitz. Sensitivity analysis of experimental data. Applied Mathematics and Computation, 16(??):153–163, 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eisenfeld:1980:GMM

[EC80] Jerome Eisenfeld and Stephen W.

Cheng. General moment methods for a class of nonlinear models. Applied Mathematics and Computation, 6(4):335– 357, January 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Evans:1989:EFI

[EE89] James D. Evans and Lori M.

Evans. Evaluation of four irrational cosine definite integrals using residue theory. Applied Mathematics and Computation, 33(3):149–159, October 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eskandari:1995:UMC

[EFS95] A. Eskandari, P. Ffolliott, and F. Szidarovszky. Uncertainty and method choice in discrete multiobjective programming problems. Applied Mathematics and Computation, 69(2– 3):335–351, May 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=2-3&aid=9400145.

Ezquerro:1997:CPI

[EGH97] J. A. Ezquerro, J. M. Gutiérrez, and M. A. Hernández. A construction procedure of iterative methods with cubical convergence. Applied Mathematics and Computation, 85(2–3):181– 199, September 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600134.

Ezquerro:1998:CPI

[EGH98] J. A. Ezquerro, J. M. Gutiérrez, and M. A. Hernández. A construction procedure of iterative methods with cubical convergence. II. Another convergence approach. Applied Mathematics and Computation, 92(1):59–68, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/1/6026.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=92&issue=
1&aid=6026.

Evans:1993:ESS

[EHE93] James D. Evans, John H.

Hubbell, and Valerie D. Evans. Exact series solution to the Epstein–Hubbell generalized elliptic type integral using complex variable residue theory. Applied Mathematics and Computation, 53(2–3):173–189, February 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eirola:1989:TCN

[Eir89] Timo Eirola. Two concepts for numerical periodic solutions of ODEs. Applied Mathematics and Computation, 31(Special issue):121–131, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Eiseman:1982:AAC

[Eis82a]  Peter R. Eiseman.  Auto-

matic algebraic coordinate generation. Applied Mathematics and Computation, 10/11(1): 447–463, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Eiseman:1982:OGG

[Eis82b] Peter R. Eiseman. Orthogonal grid generation. Applied Mathematics and Computation, 10/11 (1):193–233, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Eiseman:1987:GCP

[Eis87] Peter R. Eiseman. The geometrical construction of pointwise distributions on curves. Applied Mathematics and Computation, 21(??):233–275, 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eisenfeld:1998:IIB

[Eis98] J. Eisenfeld. Identification, interval bounds, and linear algebraic estimation for a class of nonlinear problems. Applied Mathematics and Computation, 89(1–3):17–29, January 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        89&issue=1-3&aid=9781645.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Enright:1988:ESD

[EJNT88] W. H. Enright, K. R. Jackson, S. P. Nørsett, and P. G. Thomsen. Effective solution of discontinuous IVPs using a Runge– Kutta formula pair with interpolants. Applied Mathematics and Computation, 27(4 (part II)):313–335, September 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

ElAttar:1992:FEA

[El 92] M. El Attar. Finite element analysis on distributed memory architectures. Applied Mathematics and Computation, 52 (2–3):309–316, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eisenfeld:1977:FPT

[EL77] J. Eisenfeld and V. Lakshmikantham. Fixed point theorems on closed sets through abstract cones. Applied Mathematics and Computation, 3(2):155–167, ???? 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Erbe:1990:BVP

[EL90] L. H. Erbe and Xinzhi Liu.

Boundary value problems for nonlinear impulsive integrodifferential equations. Applied Mathematics and Computation, 36(1 (part I)):31–50, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Erbe:1991:MIM

[EL91] Lynn Erbe and Xinzhi Liu. Monotone iterative methods for differential systems with finite delay. Applied Mathematics and Computation, 43(1 (part I)): 43–64, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Elabbasy:1994:OBS

[Ela94] E. M. Elabbasy. Oscillatory behavior of solutions of forced second order differential equations with alternating coefficients. Applied Mathematics and Computation, 60(1):1–13, January 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eisenfeld:1979:APN

[EM79] Jerome Eisenfeld and Van C. Mow. Asymptotic properties of a nonlinear diffusion process arising in articular cartilage. Applied Mathematics and Computation, 5(3):187–198, ???? 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Eissfeller:1990:NRC

[EM90] Horst Eissfeller and Silvia Melitta Müller. A note on reducing communication costs in explicit time stepping methods on parallel computers. Applied Mathematics and Computation, 39 (3):191–197, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Egerstedt:1998:CTM

[EM98] Magnus Egerstedt and Clyde Martin. A control theoretic model of the combined planar motion of the human head and eye. Applied Mathematics and Computation, 90(1): 61–95, March 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=1&aid=9600351.

El-Misiery:1996:ACJ

[EMEH96] A. E. M. El-Misiery and El-Sayed M. El-Horbaty. An algorithm for calculating Jones polynomials. Applied Mathematics and Computation, 74(2–3):249– 259, February 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500096.

El-Namoury:1992:FEM

[EN92] A. R. M. El-Namoury. On the finite element method for a singular problem. Applied Mathematics and Computation, 49(2– 3):239–249, June 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Nahhas:1994:STO

[EN94] A. El-Nahhas. Stability of a third-order differential equation of retarded type. Applied Mathematics and Computation, 60(2– 3):147–152, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Naggar:1995:RNH

[ENAAA95] A. M. El-Naggar, A. M. Abd-Alla, and S. M. Ahmed. On the rotation of a non-homogeneous composite infinite cylinder of orthotropic material. Applied Mathematics and Computation, 69(2–3):147–157, May 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=2-3&aid=9400074.

Eno:1990:SAE

[Eno90] L. Eno. Sensitivity analysis of experimental data. II. new coefficients. Applied Mathematics and Computation, 37(1 (part I)):9–17, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Enright:1989:NEC

[Enr89] W. H. Enright. A new errorcontrol for initial value solvers. Applied Mathematics and Computation, 31(Special issue):288– 301, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

El-Naggar:1996:WPL

[ENS96] A. M. El-Naggar and M. M.

Saliem. Wave propagation in layered media under initial stresses. Applied Mathematics and Computation, 74(2–3):95– 117, February 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500064.

El-Owaidy:1999:SNI

[EORA99] H. El-Owaidy, A. Ragab, and A. Abdeldaim. On some new integral inequalities of Growall-Bellman type. Applied Mathematics and Computation, 106 (2–3):289–303, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/33/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/33/
article.pdf.

Elaydi:1988:LSN

[ER88] S. Elaydi and M. Rama Mohana Rao. Lipschitz stability for nonlinear Volterra integrodifferential systems. Applied Mathematics and Computation, 27 (3 (part I)):191–199, September 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Evans:1994:ACVb

[ER94a] James D. Evans and Thomas L. Roberson. Applications of complex variable residue theory for evaluating irrational definite integrals. IV. Applied Mathematics and Computation, 63(1): 35–57, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Evans:1994:ACVa

[ER94b] James D. Evans and Thomas L. Roberson. Applications of complex variable residue theory to the evaluation of irrational definite integrals. III. Applied Mathematics and Computation, 63 (1):1–34, June 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Erlander:1990:GRT

[ES90] S. Erlander and T. E. Smith. General representation theorems for efficient population behavior. Applied Mathematics and Computation, 36(3):173–217, April 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:1992:FDE

[ES92] A. M. A. El-Sayed. On the fractional differential equations. Applied Mathematics and Computation, 49(2–3):205–213, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:1993:LDE

[ES93a] A. M. A. El-Sayed. Linear differential equations of fractional orders. Applied Mathematics and Computation, 55(1):1–12, April 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sheikh:1993:ONC

[ES93b] M. M. A. El-Sheikh. Oscillation and nonoscillation criteria for second-order nonlinear differential equations. II. Applied Mathematics and Computation, 59(2– 3):101–115, ???? 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sheikh:1994:OBSb

[ES94] M. M. A. El-Sheikh. On oscillatory behavior of solutions of second order functional differential equations. Applied Mathematics and Computation, 62(2– 3):183–189, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

El-Sayed:1999:FOE

[ES99] Ahmed M. A. El-Sayed. Fractionalorder evolutionary integral equations. Applied Mathematics and Computation, 98(2– 3):139–146, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/

6166.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6166.

El-Sayed:1995:MFD

[ESI95] A. M. A. El-Sayed and A. G. Ibrahim. Multivalued fractional differential equations. Applied Mathematics and Computation, 68(1):15–25, March 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=1&aid=9400080.

Eom:1998:NSL

[ESL98] Tae-Dok Eom, Masanori Sugisaka, and Ju-Jang Lee. New skill learning paradigm using various kinds of neurons. Applied Mathematics and Computation, 91(1):9–22, April 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/1/6011.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
1&aid=6011.

El-Sayed:1999:BPA

[ESR99] A. M. A. El-Sayed and S. Z. Rida. Bell polynomials of arbitrary (fractional) orders+. Applied Mathematics and Computation, 106(1):51–62, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/17/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/17/20/
article.pdf.

El-Sheikh:1994:OBSa

[ESS94] M. M. A. El-Sheikh and R. A.

Sallam. On the oscillatory behavior of solutions of second order differential equations. Applied Mathematics and Computation, 61(1):1–16, March 01, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Evans:1990:ACVa

[Eva90a] James D. Evans. Applications of complex variable residue theory to the evaluation of irrational definite integrals. I. Applied Mathematics and Computation, 39(2 (part II)):145–157, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Evans:1990:ACVb

[Eva90b]  James D. Evans.  Applica-

tions of complex variable residue theory to the evaluation solution of irrational definite integrals. II. Applied Mathematics and Computation, 39(3):177– 189, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Evans:1990:EFI

[Eva90c] James D. Evans. Evaluation of four irrational definite sine integrals using residue theory. Applied Mathematics and Computation, 36(3):163–172, April 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ezzat:1994:SSA

[Ezz94] Magdy A. Ezzat. State space approach to unsteady free convection flow through a porous medium. Applied Mathematics and Computation, 64(2–3): 191–205, September 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fabien:1996:NSC

[Fab96] Brian C. Fabien. Numerical solution of constrained optimal control problems with parameters. Applied Mathematics and Computation, 80(1):43– 62, November 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        80&issue=1&aid=9500280.

Fahrig:1988:GMP

[Fah88] L. Fahrig. A general model of populations in patchy habitats. Applied Mathematics and Computation, 27(1):53–66, ???? 1988. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Fazio:1990:NTM

[Faz90] R. Fazio. A noniterative transformation method applied to two-point boundary-value problems. Applied Mathematics and Computation, 39(1):79–87, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fazio:1991:ITM

[Faz91] Riccardo Fazio. The iterative transformation method and length estimation for tubular flow reactors. Applied Mathematics and Computation, 42(2 (part II)):105–110, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fernandez-Berdaguer:1996:IPE

[FBSS96] Elena M. Fernandez-Berdaguer, Juan E. Santos, and Dongwoo Sheen. An iterative procedure for estimation of variable coefficients in a hyperbolic system. Applied Mathematics and Computation, 76(2–3): 213–250, May 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=2-3&aid=9500161.

Frankel:1993:SNO

[FC93] Jay I. Frankel and S. Roy Choudhury. Some new observations on the classical logistic equation with heredity. Applied Mathematics and Computation, 58(2–3):275–308, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fymat:1990:OCN

[FEGL90] Alain L. Fymat, Jack I. Eisenman, Moses A. Greenfield, and W. N. Paul Lee. Organ characterization with nuclear-medicine tomography. Applied Mathematics and Computation, 35(2 (part II)):159–189, January 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feiring:1985:EPM

[Fei85] Bruce R. Feiring. An exact penalty method for solving optimal control problems. Applied Mathematics and Computation, 16(2):105–114, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feiring:1987:SPS

[Fei87] Bruce R. Feiring. Some properties of the solution space of the N-city traveling-salesman problem. Applied Mathematics and Computation, 23(4):291– 340, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fromme:1981:CSC

[FG81] J. A. Fromme and M. A. Golberg. Convergence and stability of a collocation method for the generalized airfoil equation. Applied Mathematics and Computation, 8(4):281–292, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fernandez-Gaucherand:1994:NRT

[FG94a] Emmanuel Fernández-Gaucherand.

A note on the Ross–Taylor theorem. Applied Mathematics and Computation, 64(2–3):207–212, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fridrich:1994:RCO

[FG94b] Jiří Fridrich and James F. Geer.

Reconstruction of chaotic orbits under finite resolution. Applied Mathematics and Computation, 66(2–3):129–159, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fridrich:1995:RBO

[FG95] Jiří Fridrich and James F. Geer.

Reconstruction of blurred orbits under finite resolution. Applied Mathematics and Computation, 71(2–3):227–245, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400156.

Fymat:1988:ORT

[FGLF88] A. L. Fymat, M. A. Greenfield, S. C. Lo, and D. Findley. Optimization of the radiotherapy treatment plan. Applied Mathematics and Computation, 27 (2):117–146, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ford:1995:APF

[FH95] R. A. Ford and M. H. Hamdan. Analysis of the polar form of the von Mises transformation. Applied Mathematics and Computation, 72(2–3): 205–217, October 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400184.

Ford:1998:CPF

[FH98] R. A. Ford and M. H. Hamdan. Coupled parallel flow through composite porous layers. Applied Mathematics and Computation, 97(2–3):261–271, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6148.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6148.

Ford:1996:ECT

[FHG96a] R. A. Ford, M. H. Hamdan, and L. E. Garey. Effects of coordinate transformation and differencing schemes on matrix structure and convergence history. Applied Mathematics and Computation, 77(1):53–66, June 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
77&issue=1&aid=9500191.

Fritz:1996:ESE

[FHG96b] Christophe Fritz, Michel Hassenforder, and Gerard L. Gissinger. Evaluating systems of equations: Toward a formal approach. Applied Mathematics and Computation, 75(2–3): 223–237, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500127.

Fiebig:1984:MEA

[Fie84] Denzil G. Fiebig. On the maximum-entropy approach to undersized samples. Applied Mathematics and Computation, 14(3):301–312, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Finkelstein:1995:EFD

[Fin95] Mark Finkelstein. Estimating the frequency distribution of the numbers bet on the California lottery. Applied Mathematics and Computation, 69(2– 3):195–207, May 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        69&issue=2-3&aid=9400126.

Fiorito:1995:PPM

[Fio95] Giovanni Fiorito. On properties of periodically monotone sequences. Applied Mathematics and Computation, 72(2–3): 259–275, October 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        72&issue=2-3&aid=9400187.

Fisher:1985:DOS

[Fis85] I. H. Fisher. Derivation of optimal stocking policies for grazing in arid regions. I. methodology. Applied Mathematics and Computation, 17(1):1–35, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feng:1994:PSA

[FL94] Zhao Shu Feng and Yong Qing Liu. Practical stability analysis for stochastic time-delay systems. Applied Mathematics and Computation, 63(2–3):265–279, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fu:1999:UBS

[FL99] Xilin Fu and Xinzhi Liu. Uniform boundedness and stability criteria in terms of two measures for impulsive integrodifferential equations. Applied Mathematics and Computation, 102(2–3):237–255, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6260.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6260.

Flaam:1992:LCP

[Flå92]  Sjur Didrik Flåm.  Level-

constrained programming. Applied Mathematics and Computation, 47(2–3):259–265, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feng:1992:CPS

[FLG92] Zhao Shu Feng, Yong Qing Liu, and Feng Wei Guo. Criteria for practical stability in the pth mean of nonlinear stochastic systems. Applied Mathematics and Computation, 49(2–3):251–260, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feng:1991:NSC

[FLZ91] Zhao Shu Feng, Yong Qing Liu, and Huai Zhong Zhao. Necessary and sufficient conditions for delay-independent stability of multidelay linear systems. Applied Mathematics and Computation, 46(1 (part I)):23–32, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Freedman:1996:BPT

[FM96] H. I. Freedman and P. Moson. Bifurcations in persistence theory. Applied Mathematics and Computation, 79(2–3):125– 136, October 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500248.

Florens:1986:STS

[FMR86] J.-P. Florens, M. Mouchart, and J.-F. Richard. Structural time series modeling: a Bayesian approach. Applied Mathematics and Computation, 20(3–4):365– 400, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Foroutan-pour:1999:AIB

[FpDS99] K. Foroutan-pour, P. Dutilleul, and D. L. Smith. Advances in the implementation of the boxcounting method of fractal dimension estimation. Applied Mathematics and Computation, 105(2–3):195–210, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/2-
3/6317.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
105&issue=2-3&aid=6317.

Fei:1995:NSN

[FPGV95] Zhang Fei, Víctor M. Pérez-García, and Luis Vázquez. Numerical simulation of nonlinear Schrödinger systems: a new conservative scheme. Applied Mathematics and Computation, 71(2– 3):165–177, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400152.

Franses:1989:DBR

[Fra89] Philip Hans Franses. The distance between regression models and its impact on model selection. Applied Mathematics and Computation, 34(1 (part I)): 1–16, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

French:1990:CCH

[Fre90] Donald A. French. Computations on the Cahn–Hilliard model of solidification. Applied Mathematics and Computation, 40(1 (part I)):55–76, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Friedman:1992:DHC

[Fri92] Charles N. Friedman. Detecting Hamiltonian cycles. Applied Mathematics and Computation, 47(2–3):137–143, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fridrich:1997:DTD

[Fri97] Jiri Fridrich. Discrete-time dynamical systems under observational uncertainty. Applied Mathematics and Computation, 82(2–3):181–205, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        82&issue=2-3&aid=9600029.

French:1990:CFE

[FS90] Donald A. French and Jack W. Schaeffer. Continuous finite element methods which preserve energy properties for nonlinear problems. Applied Mathematics and Computation, 39(3):271– 295, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Finkelstein:1994:SCS

[FS94] Mark Finkelstein and Stephen Scheinberg. The storage complexity of some estimators of the median. Applied Mathematics and Computation, 60(1): 15–24, January 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Franke:1998:SPD

[FS98] C. Franke and R. Schaback.

Solving partial differential equations by collocation using radial basis functions. Applied Mathematics and Computation, 93(1):73–82, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/1/6064.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
1&aid=6064.

Fariabi:1994:CPS

[FT94] Said Fariabi and Betty Travis.

On conjugate points for second order systems. Applied Mathematics and Computation, 60 (1):37–42, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Frontini:1997:ECS

[FT97] Marco Frontini and Aldo Tagliani. Entropy-convergence in Stieltjes and Hamburger moment problem. Applied Mathematics and Computation, 88(1): 39–51, December 15, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=1&aid=9600305.

Frontini:1998:MEL

[FT98] Marco Frontini and Aldo Tagliani. Maximum entropy and lacunary Stieltjes moment problem. Applied Mathematics and Computation, 97(2–3): 183–196, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/97/2-3/
                                        6152.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=97&issue=
                                        2-3&aid=6152.

Funaro:1991:PAP

[Fun91] Daniele Funaro. Pseudospectral approximation of a PDE defined on a triangle. Applied Mathematics and Computation, 42(2 (part II)):121–138, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fei:1991:TEC

[FV91] Zhang Fei and Luis Vázquez. Two energy conserving numerical schemes for the sine-Gordon equation. Applied Mathematics and Computation, 45(1 (part I)):17–30, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feng:1998:APP

[FW98] Wei Feng and Fan Wang. Asymptotic periodicity and permanence in a competitiondiffusion system with discrete delays. Applied Mathematics and Computation, 89(1–3):99– 110, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781650.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Fymat:1979:GCI

[Fym79] Alain L. Fymat. A generalization of Cooke’s integral inversion formula with application to remote-sensing theory. Applied Mathematics and Computation, 5(1):23–39, ???? 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gaertner:1987:PAC

[Gae87] Wulf Gaertner. Periodic and aperiodic consumer behavior. Applied Mathematics and Computation, 22(2–3):233–254, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gartland:1989:SUS

[Gar89] Eugene C. Gartland, Jr. Strong uniform stability and exact discretizations of a model singular perturbation problem and its finite-difference approximations. Applied Mathematics and Computation, 31(Special issue):

473–485, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Garnaev:1990:OPN

[Gar90] Andrej Yu. Garnaev. Optimal policies in a nonlinear bioeconomic model of eutrophication. Applied Mathematics and Computation, 35(1 (part I)):95– 102, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Garner:1998:ERC

[Gar98] J. B. Garner. Existence results for a class of diffusion-reaction problems. Applied Mathematics and Computation, 89(1–3): 111–118, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        89&issue=1-3&aid=9781651.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Golberg:1990:CSP

[GB90] M. A. Golberg and H. Bowman. The conditioning of some projection methods for Fredholm and singular integral equations. Applied Mathematics and Computation, 40(2 (part II)):165–178, ???? 1990. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Goel:1995:SCU

[GB95] Narendra S. Goel and Shen Bo.

Symbolic computation using Lsystems. II. extensions. Applied Mathematics and Computation, 69(2–3):227–240, May 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=2-3&aid=9400129.

Golberg:1998:OCR

[GB98] Micheal Golberg and Harold Bowman. Optimal convergence rates for some discrete projection methods. Applied Mathematics and Computation, 96(2–3):237–271, November 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/2-3/
6127.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
2-3&aid=6127.

Gulliver:1999:AGC

[GBS99] T. Aaron Gulliver, Vijay K.

Bhargava, and Jack M. Stein. Q-ary Gray codes and weight distributions. Applied Mathematics and Computation, 103 (1):97–109, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/103/1/
                                        6272.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=103&
                                        issue=1&aid=6272.

Golberg:1994:MAE

[GC94] Michael A. Golberg and C. S. Chen. On a method of Atkinson for evaluating domain integrals in the boundary element method. Applied Mathematics and Computation, 60(2–3):125– 138, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Goicoechea:1987:NDE

[GD87] Ambrose Goicoechea and Lucien Duckstein. Nonnormal deterministic equivalents and a transformation in stochastic mathematical programming. Applied Mathematics and Computation, 21(1):51–72, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ge:1989:OCL

[Ge89a] Ren Pu Ge. Optimal choice of linear interval extension. Applied Mathematics and Computation, 30(2 (part II)):165– 189, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ge:1989:PCS

[Ge89b] Ren Pu Ge. A parallel computing scheme for minimizing a class of large scale functions. Applied Mathematics and Computation, 30(3):261–288, April 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ge:1989:PGO

[Ge89c] Ren Pu Ge. A parallel global optimization algorithm for rational separable-factorable functions. Applied Mathematics and Computation, 32(1):61–72, July 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ge:1990:FFT

[Ge90a] Ren Pu Ge. The filled function transformations for constrained global optimization. Applied Mathematics and Computation, 39(1):1–20, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ge:1990:FMM

[Ge90b] Ren Pu Ge. Finding more and more solutions of a system of nonlinear equations. Applied Mathematics and Computation, 36(1 (part I)):15–30, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ge:1991:SLP

[Ge91] Ren Pu Ge. Solving linear programming problems via linear minimax problems. Applied Mathematics and Computation, 46(1 (part I)):59–77, ???? 1991.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guttalu:1996:SSP

[GF96] Ramesh S. Guttalu and Henryk Flashner. Stability study of a periodic system by a period-to-period mapping. Applied Mathematics and Computation, 78(2– 3):123–135, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600003.

Gil:1988:CSS

[GFM88] M. A. Gil, M. J. Fernández, and I. Martínez. The choice of sample size in estimating mutual information. Applied Mathematics and Computation, 27(3 (part I)): 201–216, September 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gil:1991:PTS

[GG91a] Maria Angeles Gil and Pedro Gil. A procedure to test the suitability of a factor for stratification in estimating diversity. Applied Mathematics and Computation, 43(3):221–229, 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Goel:1991:SCU

[GG91b] Narendra S. Goel and Mark D. Goodwin. Symbolic computation using L-systems. Applied Mathematics and Computation, 42(3):223–253, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Garey:1996:HOM

[GG96] L. E. Garey and C. J. Gladwin. Higher-order methods for eigenvalues of second-order ordinary differential equations with boundary conditions. Applied Mathematics and Computation, 76(2–3):293–300, May 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=2-3&aid=9500187.

Gatto:1991:ORE

[GGR91] Marino Gatto, Luca Ghezzi, and Sergio Rinaldi. The optimal reclamation of eutrophic water bodies. Applied Mathematics and Computation, 43(2):105– 115, 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Garey:1997:USM

[GGS97] L. E. Garey, C. J. Gladwin, and R. E. Shaw. Unconditionally stable methods for second-order Fredholm integrodifferential equations. Applied Mathematics and Computation, 81(2–3):275–286, February 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        81&issue=2-3&aid=9500328.

Ge:1989:CAN

[GH89] Ren Pu Ge and Chang Bin Huang. A continuous approach to nonlinear integer programming. Applied Mathematics and Computation, 34(1 (part I)):39– 60, November 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ghorbanzadeh:1996:DRC

[Gho96] Dariush Ghorbanzadeh. Detection of random change point in one-parameter exponential families. Applied Mathematics and Computation, 77(2–3):167–177, ???? 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gil:1989:NCS

[Gil89] María Angeles Gil. A note on the choice of sample size in estimating the mutual information. Applied Mathematics and Computation, 30(2 (part II)): 125–132, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gupta:1994:SFO

[GK94] Chaitan P. Gupta and Ying C. Kwong. L^{p}-solvability of a fourth-order boundary-value problem at resonance. Applied Mathematics and Computation, 65(1–3):151–159, September/ October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Gunji:1991:ALA

[GKn91] Yukio Gunji and Norio Kon-no.

Artificial life with autonomously emerging boundaries. Applied Mathematics and Computation, 43(3):271–298, 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Grace:1992:OTS

[GL92] S. R. Grace and B. S. Lalli. Oscillation theorems for second order neutral functional differential equations. Applied Mathematics and Computation, 51(2– 3):119–133, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guoqiang:1994:AEE

[GL94] Han Guoqiang and Zhang Liqing. Asymptotic error expansion of two-dimensional Volterra integral equation by iterated collocation. Applied Mathematics and Computation, 61(2–3): 269–285, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gunzburger:1999:AAO

[GL99] Max D. Gunzburger and Hyung-Chun Lee. Analysis and approximation of optimal control problems for first-order elliptic systems in three dimensions. Applied Mathematics and Computation, 100(1):49–70, April 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/1/
6201.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=100&
issue=1&aid=6201.

Gladwell:1978:NSD

[Gla78] Ian Gladwell. On the numerical solution of a differential nonlinear eigenvalue problem on an infinite range. Applied Mathematics and Computation, 4(2): 121–137, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gleissner:1988:SE

[Gle88] Winfried Gleissner. The spread of epidemics. Applied Mathematics and Computation, 27 (2):167–171, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gopalsamy:1998:GHB

[GLL98] K. Gopalsamy, Issic K. C. Leung, and Pingzhou Liu. Global Hopf-bifurcation in a neural netlet. Applied Mathematics and Computation, 94(2– 3):171–192, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/2-3/
6097.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
2-3&aid=6097.

Golberg:1993:NEC

[GM93] Michael A. Golberg and Michael McKenna. The numerical evaluation of a class of logarithmically singular integral transforms. Applied Mathematics and Computation, 57(2–3):117– 129, September 01, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ghelardoni:1994:EES

[GM94] Paolo Ghelardoni and Pietro Marzulli. Error estimates for shooting methods in two-point boundary value problems for second-order equations. Applied Mathematics and Computation, 60(2–3):237–248, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gouesbet:1999:CEH

[GMGCL99] G. Gouesbet, S. Meunier-Guttin-Cluzel, and C. Letellier. Computer evaluation of Homfly polynomials by using Gauss codes, with a skeintemplate algorithm. Applied Mathematics and Computation, 105(2–3):271–289, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/2-
3/6325.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
105&issue=2-3&aid=6325.

Gary:1983:SOM

[GMS83] John Gary, Steve McCormick, and Roland Sweet. Successive overrelaxation, multigrid, and preconditioned conjugate gradients algorithms for solving a diffusion problem on a vector computer. Applied Mathematics and Computation, 13(3–4): 285–309, November 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guariso:1996:PSC

[GMS96] Giorgio Guariso, Vittorio Maniezzo, and Paola Salomoni. Parallel simulation of a cellular pollution model. Applied Mathematics and Computation, 79(1):27– 41, September 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=1&aid=9500214.

Gnoffo:1982:VFV

[Gno82] Peter A. Gnoffo. A vectorized, finite-volume, adaptive-grid algorithm for Navier–Stokes. Applied Mathematics and Computation, 10/11(1):819–835, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Golberg:1977:CFI

[Gol77] Michael A. Golberg. The conversion of Fredholm integral equations to equivalent Cauchy problems. II: Computation of resolvents. Applied Mathematics and Computation, 3(??):39–58, 1977. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Golberg:1979:MAS

[Gol79] Michael A. Golberg. A method of adjoints for solving some illposed equations of the first kind. Applied Mathematics and Computation, 5(2):123–129 (or 123– 130??), ???? 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Golberg:1987:CSA

[Gol87] Michael A. Golberg. The convergence of several algorithms for solving integral equations with finite part integrals. II. Applied Mathematics and Computation, 21(4):283–293, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Golberg:1988:NSR

[Gol88a] Michael A. Golberg. A note on a superconvergence result for the generalized airfoil equation. Applied Mathematics and Computation, 26(2 (part I)):105–117, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Golberg:1988:PGM

[Gol88b]  Michael A. Golberg.  The

perturbed Galerkin method for Cauchy singular integral equations with constant coefficients. Applied Mathematics and Computation, 26(1):1–33, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Golberg:1989:DPM

[Gol89] Michael A. Golberg. Discrete projection methods for Cauchy singular integral equations with constant coefficients. Applied Mathematics and Computation, 33(1 (part I)):1–41, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Golberg:1993:DSI

[Gol93] Michael A. Golberg. The discrete Sloan iterate for the generalized airfoil equation. Applied Mathematics and Computation, 57(2–3):103–115, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Golberg:1995:NSR

[Gol95] Michael Golberg. A note on the sparse representation of discrete integral operators. Applied Mathematics and Computation, 71(2–3):97–118, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400146.

Golberg:1996:RDN

[Gol96] Michael Golberg. Recent developments in the numerical evaluation of particular solutions in the boundary element method. Applied Mathematics and Computation, 75(1):91–101, March 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=1&aid=9500123.

Golberg:1999:NDM

[Gol99] Michael A. Golberg. A note on the decomposition method for operator equations. Applied Mathematics and Computation, 106(2–3):215–220, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/26/
article.pdf.

Gottinger:1984:ATP

[Got84] Hans-W. Gottinger. An analytical treatment of potential hazard and risk assessment for environmental chemicals. Applied Mathematics and Computation, 14(4):381–396, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gotusso:1985:ETL

[Got85] Laura Gotusso. On the energy theorem for the Lagrange equations in the discrete case. Applied Mathematics and Computation, 17(2):129–136, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Goldstein:1988:NES

[GP88] Charles I. Goldstein and Seymour V. Parter. On the norm equivalence of singularly perturbed elliptic difference operators. Applied Mathematics and Computation, 26(3 (part II)): 179–200, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ghezzi:1994:QAN

[GP94] Luca Luigi Ghezzi and Lorenzo Peccati. Qualitative analysis of a nonlinear stock price model. Applied Mathematics and Computation, 63(2–3):123–130, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Goyon:1997:PNM

[GP97] Olivier Goyon and Pascal Poullet. Preconditioned Newton methods using incremental unknowns methods for the resolution of a steady-state Navier-Strokes-like problem. Applied Mathematics and Computation, 87(2–3):289–311, December 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=2-3&aid=9600304.

Garcia-Pena:1983:MMC

[GPAB83] J. Garcia-Pena, S. P. Azen, and R. N. Bergman. On a modification of Marquardt’s compromise: Rationale and applications. Applied Mathematics and Computation, 12(??):1–17, 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guo:1986:NSS

[GPRV86] Ben Yu Guo, Pedro J. Pascual, María J. Rodriguez, and Luis Vázquez. Numerical solution of the sine-Gordon equation. Applied Mathematics and Computation, 18(1):1–14, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ge:1990:GCF

[GQ90] Ren Pu Ge and Yong Feng Qin.

The globally convexized filled functions for global optimization. Applied Mathematics and Computation, 35(2 (part II)): 131–158, January 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Graef:1998:SPM

[GQS98] J. R. Graef, C. Qian, and P. W. Spikes. Stability in a population model. Applied Mathematics and Computation, 89(1–3): 119–132, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        89&issue=1-3&aid=9781652.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Galligani:1997:TSA

[GR97] E. Galligani and V. Ruggiero. The two-stage arithmetic mean method. Applied Mathematics and Computation, 85(2–3):245– 264, September 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        85&issue=2-3&aid=9600139.

Grace:1995:OMN

[Gra95] S. R. Grace. Oscillations of mixed neutral functional differential equations. Applied Mathematics and Computation, 68 (1):1–13, March 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=1&aid=9400075.

Gragnani:1997:BAT

[Gra97] Alessandra Gragnani. Bifurcation analysis of two predatorprey models. Applied Mathematics and Computation, 85(2–3): 97–108, September 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600111.

Greenbaum:1986:MMM

[Gre86] A. Greenbaum. A multigrid method for multiprocessors. Applied Mathematics and Computation, 19(??):75–88, 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Greenspan:1987:NNC

[Gre87a] Donald Greenspan. Note on a new computer-oriented approach to molecular mechanics. Applied Mathematics and Computation, 21(??):45–50, 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Greenspan:1987:SMS

[Gre87b] Donald Greenspan. Semiclassical modeling of the states and spectra of para helium singlets with an algebraic formula for excited states. Applied Mathematics and Computation, 24 (??):229–240, 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gremaud:1997:NSP

[Gre97] Pierre-Alain Gremaud. Numerical simulation of pattern formation in grain flows. Applied Mathematics and Computation, 84(2–3):145–162, July 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600084.

Grigorieff:1989:SCC

[Gri89] Rolf Dieter Grigorieff. On stability constants and condition numbers of discretisation methods. Applied Mathematics and Computation, 31(Special issue): 486–516, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Grimm:1998:OSV

[Gri98] Stefan Grimm. On orbit sum values of elements of finite order. Applied Mathematics and Computation, 97(1):17–22, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/1/6141.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
1&aid=6141.

Gross:1993:CEM

[Gro93] R. Steven Gross. Computationally efficient method for the solution of a singular integrodifferential equation. Applied Mathematics and Computation, 57(2–3):161–175, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gilpin:1986:APP

[GS86] Michael J. Gilpin and Robert O.

Shelton. Applications of parallel processors to some hard discrete problems. Applied Mathematics and Computation, 20(??):1– 4, 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Garner:1994:EUS

[GS94a] J. B. Garner and Randolph James Schilling. Existence, uniqueness, and stability for a class of diffusion problems. Applied Mathematics and Computation, 65(1– 3):161–169, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Graef:1994:AOB

[GS94b] John R. Graef and Paul W. Spikes. Asymptotic and oscillatory behavior of solutions of nonlinear neutral delay equations of arbitrary order. Applied Mathematics and Computation, 65(1–3):209–221, September/ October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Gould:1997:SCI

[GS97] H. W. Gould and H. M. Srivastava. Some combinatorial identities associated with the Vandermonde convolution. Applied Mathematics and Computation, 84(2–3):97–102, July 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600054.

Garey:1999:PAS

[GS99] L. E. Garey and R. E. Shaw. A parallel algorithm for solving Toeplitz linear systems. Applied Mathematics and Computation, 100(2–3):241–247, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6212.pdf;   http://www.
elsevier.com/cgi-bin/cas/

tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6212.

Goodman:1995:LDE

[GSC95] T. N. T. Goodman, H. B. Said, and L. H. T. Chang. Local derivative estimation for scattered data interpolation. Applied Mathematics and Computation, 68(1):41–50, March 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=1&aid=9400086.

Gunji:1996:BSD

[GSI96] Yukio-Pegio Gunji, Hisato Sadaoka, and Keisuke Ito. Bootstrapping system defined by inconsistent relation between Boolean and Non-Boolean algebra. Applied Mathematics and Computation, 79(1):43–97, September 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=1&aid=9500221.

Gunji:1993:LPB

[GSK93] Yukio-Pegio Gunji, Shuji Shinohara, and Norio Konno. Learning processes based on incomplete identification and information generation. Applied Mathematics and Computation, 55(2– 3):219–253, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gordon:1982:TMT

[GT82] William J. Gordon and Linda C. Thiel. Transfinite mappings and their application to grid generation. Applied Mathematics and Computation, 10/11(1): 171–192, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Goh:1991:AAS

[GT91] C. J. Goh and K. L. Teo. Alternative algorithms for solving nonlinear function and functional inequalities. Applied Mathematics and Computation, 41(2 (part II)):159–177, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gu:1999:CFE

[Gu99] Haiming Gu. Characteristic finite element methods for nonlinear Sobolev equations. Applied Mathematics and Computation, 102(1):51–62, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/102/1/
                                        6245.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=102&
                                        issue=1&aid=6245.

Gunji:1992:FLI

[Gun92] Yukio-Pegio P. Gunji. The form of life. I. it is possible but not necessary. Applied Mathematics and Computation, 47(2–3): 267–288, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gunji:1993:FLU

[Gun93] Yukio-Pegio Gunji. Forms of life: Unprogrammability constitutes the outside of a system and its autonomy. Applied Mathematics and Computation, 57 (1):19–76, August 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gunji:1994:ALP

[Gun94] Yukio-Pegio Gunji. Autonomic life as the proof of incompleteness and lawvere’s theorem of fixed point. Applied Mathematics and Computation, 61(2–3): 231–267, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guo:1999:BVP

[Guo99] Dajun Guo. Boundary value problems for impulsive integrodifferential equations on unbounded domains in a Banach space. Applied Mathematics and Computation, 99 (1):1–15, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/1/6175.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
1&aid=6175.

Gupta:1990:FON

[Gup90] Chaitan P. Gupta. A fourthorder nonlinear boundary value problem and an integral-type sign condition. Applied Mathematics and Computation, 35(3): 231–242, February 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gupta:1998:GMP

[Gup98] Chaitan P. Gupta. A generalized multi-point boundary value problem for second order ordinary differential equations. Applied Mathematics and Computation, 89(1–3):133– 146, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781653.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Gear:1989:RTI

[GW89a] C. W. Gear and Dian Han Wang. Real-time integration formulas with off-step inputs and their stability. Applied Mathematics and Computation, 31(Special issue):132–147, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Gray:1989:ELS

[GW89b] Henry L. Gray and Suojin Wang. An extension of the Levin-Sidi class of nonlinear transformations for accelerating convergence of infinite integrals and series. Applied Mathematics and Computation, 33(2 (part II)):75–87, September 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Guan:1994:VPF

[GWL94] Zhihong Guan, Xiangcai Wen, and Yongqing Liu. Variation of the parameters formula and the problem of BIBO for singular measure differential systems with impulse effect. Applied Mathematics and Computation, 60(2–3):153–169, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gao:1999:NIM

[GXQ99] Weiguo Gao, Jungong Xue, and Yanyun Qu. A new implementation of EN method. Applied Mathematics and Computation, 98(2–3):199–208, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/98/2-3/
                                        6160.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=98&issue=
                                        2-3&aid=6160.

Gyori:1990:SCE

[Gyö90] István Györi. Sharp conditions for existence of nontrivial invariant cones of nonnegative initial values of difference equations. Applied Mathematics and Computation, 36(2 (part II)):89–111, March 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gregory:1984:GAI

[GZ84] John Gregory and Marvin Zeman. A Galerkin approximation for the initial value problem for linear second-order differential equations. Applied Mathematics and Computation, 15(2):93–108, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gzyl:1995:MRF

[Gzy95]  Henryk Gzyl.  Maxentropic

reconstruction of Fourier and Laplace transforms. Applied Mathematics and Computation, 73(2–3):181–189, December 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=2-3&aid=9400248.

Gzyl:1997:MRF

[Gzy97]  Henryk Gzyl.  Maxentropic

reconstruction of Fourier and Laplace transforms under nonlinear constraints. Applied Mathematics and Computation, 85(2–3):117–126, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600116.

Haber:1995:PCN

[Hab95]  R. Haber.  Predictive con-

trol of nonlinear dynamic processes. Applied Mathematics and Computation, 70(2–3): 169–184, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400122.

Dynamics and control (Vienna, 1993).

Hady:1995:MCB

[Had95] F. M. Hady. Mixed convection boundary-layer flow of Non-Newtonian fluids on a horizontal plate. Applied Mathematics and Computation, 68(2–3): 105–112, March 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        68&issue=2-3&aid=9400084.

Haidar:1997:RPI

[Hai97] Nassar H. S. Haidar. Recursive pseudo-inversion of the Laplace transform on the real line. Applied Mathematics and Computation, 84(2–3):213–220, July 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        84&issue=2-3&aid=9600087.

Halsey:1982:CGG

[Hal82]  Douglas Halsey.  Conformal

grid generation for multielement airfoils. Applied Mathematics and Computation, 10/11(1): 585–600, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Hale:1998:DNA

[Hal98] Jack K. Hale. Dynamics of numerical approximations. Applied Mathematics and Computation, 89(1–3):5–15, January 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        89&issue=1-3&aid=9781644.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Hamdan:1992:CAF

[Ham92] M. H. Hamdan. Computational aspects of flow through curved domains: The effect of no-slip. Applied Mathematics and Computation, 52(2–3):189– 206, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hamann:1994:CSA

[Ham94a] Bernd Hamann. Construction of B-spline approximations for use in numerical grid generation. Applied Mathematics and Computation, 65(1–3):295–314, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Hamdan:1994:SPF

[Ham94b] M. H. Hamdan. Single-phase flow through porous channels: A review of flow models and channel entry conditions. Applied Mathematics and Computation, 62(2–3):203–222, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hamdan:1997:NCS

[Ham97] M. H. Hamdan. Natural coordinate system approach to coupled n-phase fluid flow in curved domains. Applied Mathematics and Computation, 85(2–3):297– 304, September 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600144.

Hamdan:1998:AAE

[Ham98a] M. H. Hamdan. An alternative approach to exact solutions of a special class of Navier–Stokes flows. Applied Mathematics and Computation, 93(1):83–90, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/1/6033.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
1&aid=6033.

Hamdan:1998:NCU

[Ham98b] M. H. Hamdan. A note on computational uncertainty. Applied Mathematics and Computation, 94(2–3):285–291, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/2-3/
6098.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
2-3&aid=6098.

Han:1992:QEE

[Han92] Wei Min Han. Quantitative error estimate in modeling the laminar stationary flow of a Bingham fluid. Applied Mathematics and Computation, 47 (1):15–24, January 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Han:1998:PCM

[Han98]  Qiaoming Han.  Projection

and contraction methods for semidefinite programming. Applied Mathematics and Computation, 95(2–3):275–289, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.elsevier.
com/cas/tree/store/amc/sub/
1998/95/2-3/6068.pdf;   http:
//www.elsevier.com/cgi-bin/
cas/tree/store/amc/cas_sub/
browse/browse.cgi?year=1998&
volume=95&issue=2-3&aid=6068.

Harrington:1982:CMM

[Har82] Andrew Harrington. Conformal mappings onto multiply connected regions with specified boundary shapes: a preliminary discussion of computer implementation. Applied Mathematics and Computation, 10/11 (1):601–618, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Harker:1986:NET

[Har86] Patrick T. Harker. A note on the existence of traffic equilibria. Applied Mathematics and Computation, 18(4):277–283, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Harker:1987:DPR

[Har87] Patrick T. Harker. Derivatives of the Perron root of a positive reciprocal matrix: with application to the analytic hierarchy process. Applied Mathematics and Computation, 22(2– 3):217–232, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hartfiel:1995:IMS

[Har95] D. J. Hartfiel. Indeterminate Markov systems. Applied Mathematics and Computation, 72(1): 51–59, September 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        72&issue=1&aid=9400175.

Hartfiel:1997:SBQ

[Har97] D. J. Hartfiel. System behavior in quotient systems. Applied Mathematics and Computation, 81(1):31–48, January 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/

tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=1&aid=9500300.

Honma:1998:AEH

[HAST98] Noriyasu Honma, Kenichi Abe, Mitsuo Sato, and Hiroshi Takeda. Adaptive evolution of holon networks by an autonomous decentralized method. Applied Mathematics and Computation, 91(1):43–61, April 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/1/6008.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
1&aid=6008.

Haussling:1982:SNW

[Hau82] Henry J. Haussling. Solution of nonlinear water wave problems using boundary-fitted coordinate systems. Applied Mathematics and Computation, 10/11 (1):385–407, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Hamdan:1991:ADL

[HB91] M. H. Hamdan and R. M.

Barron. Analysis of the Darcy–Lapwood and the Darcy– Lapwood–Brinkman models: significance of the Laplacian. Applied Mathematics and Computation, 44(2 (part II)):121– 141, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hamdan:1993:DLB

[HB93] M. H. Hamdan and R. M. Barron. On the Darcy–Lapwood– Brinkman–Saffman dusty fluid flow models through porous media. I. models development. Applied Mathematics and Computation, 54(1):65–79, March 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hornberger:1985:SPV

[HC85] George M. Hornberger and Bernard J. Cosby. Selection of parameter values in environmental models using sparse data: a case study. Applied Mathematics and Computation, 17 (4):335–355, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). System identification and parameter estimation of environmental systems.

Hirtzel:1987:CEA

[HC87] C. S. Hirtzel and K. Y. Chan. Computer experiments for the analysis of extreme-value phenomena. Applied Mathematics and Computation, 21(??):125– 144, 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hariche:1988:SLI

[HD88] K. Hariche and E. D. Denman. On solvents and Lagrange interpolating λ-matrices. Applied Mathematics and Computation, 25(4):321–332, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hraba:1996:UMM

[HD96] Tomás Hraba and Jaroslav Dolezal. Use of a mathematical model for effectivity estimation of HIV infection therapies. Applied Mathematics and Computation, 78(2–3):153– 161, September 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600005.

Hiessl:1985:MQA

[HDP85] H. Hiessl, L. Duckstein, and E. J. Plate. Multiobjective Qanalysis with concordance and discordance concepts. Applied Mathematics and Computation, 17(??):107–122, 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:1996:DDM

[He96] Qiming He. Domain decomposition method for nonlinear generalized Schrödinger-type systems: semi-discrete problem. Applied Mathematics and Computation, 77(1):33–52, June 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        77&issue=1&aid=9500159.

He:1998:LCG

[He98] Jiaxing He. On the linear combination of Grünwald polynomial operator. Applied Mathematics and Computation, 96(2– 3):117–126, November 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/96/2-3/
                                        6130.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=96&issue=
                                        2-3&aid=6130.

Hassard:1981:NHB

[HEH81] B. Hassard and I. M. El-Henawy. Numerical Hopf bifurcation analysis in nonlinear ordinary and partial differential systems from chemical reactor theory. Applied Mathematics and Computation, 9(2):75–92, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Heinrichs:1988:CFM

[Hei88] Wilhelm Heinrichs. Collocation and full multigrid methods. Applied Mathematics and Computation, 26(1):35–43, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Heinrichs:1991:SMM

[Hei91] Wilhelm Heinrichs. A 3D spectral multigrid method. Applied Mathematics and Computation, 41(2 (part II)):117–128, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Heinrichs:1993:FEP

[Hei93a] Wilhelm Heinrichs. Finite element preconditioning for spectral multigrid methods. Applied Mathematics and Computation, 59(1):19–40, November 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Heinrichs:1993:SMM

[Hei93b]  Wilhelm Heinrichs.  Spec-

tral multigrid methods for domain decomposition problems using patching techniques. Applied Mathematics and Computation, 59(2–3):165–176, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hendry:1986:EMD

[Hen86] David F. Hendry. Empirical modeling in dynamic econometrics. Applied Mathematics and Computation, 20(3–4):201–236, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Heunis:1990:SDEa

[Heu90a] A. J. Heunis. On the stochastic differential equations of filtering theory. Applied Mathematics and Computation, 37 (3):185–218, June 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Heunis:1990:SDEb

[Heu90b] A. J. Heunis. On the stochastic differential equations of filtering theory. Applied Mathematics and Computation, 39(3 (Supplement)):3s–36s, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hipel:1991:CCA

[HF91] Keith W. Hipel and Niall M. Fraser. Cooperation in conflict analysis. Applied Mathematics and Computation, 43(2 (part II)):181–206, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hwang:1993:AVS

[HF93] Daniel Y. Hwang and Winfred A. Foster, Jr. The axisymmetric vibration of shallow spherical shells with fixed circular cutout. Applied Mathematics and Computation, 57(2–3):205– 220, September 01, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hamdan:1995:SPF

[HF95] M. H. Hamdan and R. A. Ford. Single-phase flow through porous channels. II. flow models, critical length, and viscous separation. Applied Mathematics and Computation, 69(2–3): 241–254, May 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=2-3&aid=9400132.

Hamdan:1998:NSG

[HFA98] M. H. Hamdan, R. A. Ford, and R. Abu Rabia. Numerical simulation of gas-particulate flow through curvilinear porous channels. Applied Mathematics and Computation, 94(2– 3):267–284, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/2-3/
6103.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
2-3&aid=6103.

Hansen:1983:INM

[HG83] E. R. Hansen and R. I. Greenberg. An interval Newton method. Applied Mathematics and Computation, 12(2–3): 89–98, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hanson:1984:EAM

[HG84] P. M. Hanson and I. Gladwell.

Error analysis of modal methods for viscously damped structural systems. Applied Mathematics and Computation, 14 (2):167–195, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:1999:PBV

[HG99] Zhimin He and Weigao Ge. Periodic boundary value problem for first order impulsive delay differential equations. Applied Mathematics and Computation, 104(1):51–63, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/104/1/
                                        6281.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=104&
                                        issue=1&aid=6281.

Herbin:1988:PIM

[HGS88] Raphaèle Herbin, Stephane Gerbi, and Vijay Sonnad. Parallel implementation of a multigrid method on the experimental ICAP supercomputer. Applied Mathematics and Computation, 27(4 (part II)):281–312, September 1, 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hermes:1985:ESS

[HH85] Henry Hermes and David Hogenson. The explicit synthesis of stabilizing (time optimal) feedback controls for the attitude control of a rotating satellite. Applied Mathematics and Computation, 16(3):229– 240, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hu:1996:SSC

[HH96a] Guang-Di Hu and Guang-Da Hu. Some simple criteria for stability of neutral delaydifferential systems. Applied Mathematics and Computation, 80(2–3):257–271, December 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500301.

Hu:1996:SDD

[HH96b] Guang-Di Hu and Guang-Da Hu. Stability of discretedelay systems: Boundary criteria. Applied Mathematics and Computation, 80(2–3):95– 104, December 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500282.

Hu:1997:SND

[HH97] Guang-Da Hu and Guang-Di Hu. Stability of neutral delaydifferential systems: Boundary criteria. Applied Mathematics and Computation, 87(2–3):247– 259, December 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=2-3&aid=9600300.

Hickernell:1999:RBF

[HH99] Fred J. Hickernell and Y. C. Hon. Radial basis function approximations as smoothing splines. Applied Mathematics and Computation, 102(1): 1–24, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.elsevier.
com/cas/tree/store/amc/sub/
1999/102/1/6238.pdf;   http:
//www.elsevier.com/cgi-bin/
cas/tree/store/amc/cas_sub/
browse/browse.cgi?year=1999&
volume=102&issue=1&aid=6238.

Hartung:1998:PIC

[HHT98] Ferenc Hartung, Terry L. Herdman, and Janos Turi. Parameter identification in classes of hereditary systems of neutral type. Applied Mathematics and Computation, 89(1–3): 147–160, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781654.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Huang:1997:ENS

[HHZ97] Hongci Huang, Weimin Han, and Jinshi Zhou. Extrapolation of numerical solutions for elliptic problems on corner domains. Applied Mathematics and Computation, 83(1):53–67, April 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=1&aid=9600046.

Hillion:1980:IIB

[Hil80] P. Hillion. Invariant imbedding for block tridiagonal systems. Applied Mathematics and Computation, 6(2):95–122, March 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hirtzel:1984:AEV

[Hir84] C. S. Hirtzel. Analysis of extreme values of natural processes: Statistical description of the maximum. Applied Mathematics and Computation, 15 (??):283–303, 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hirtzel:1985:EVA

[Hir85] C. S. Hirtzel. Extreme values of autocorrelated sequences. Applied Mathematics and Computation, 16(4):327–345, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Han:1992:CME

[HJ92] Wei Min Han and Emery Jou. On the computation of minimalenergy splines: convergence analysis. Applied Mathematics and Computation, 47(1):1– 13, January 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Haynie:1981:FND

[HK81] Susan P. Haynie and Robert E. Kalaba. Fitting a nonlinear differential system to population data: the two-sex geometricmean model. Applied Mathematics and Computation, 9(1):1–26, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Holodniok:1984:NAE

[HK84] M. Holodniok and M. Kubíček. New algorithms for the evaluation of complex bifurcation points in ordinary differential equations. A comparative numerical study. Applied Mathematics and Computation, 15 (3):261–274, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hicks:1988:CDH

[HK88] D. L. Hicks and K. L. Kuttler. Continuum and discrete hydrodynamical models and globally well-posed problems. Applied Mathematics and Computation, 25(4):299–320, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hicks:1989:CNS

[HK89] D. L. Hicks and K. L. Kuttler. The convergence of numerical solutions of hydrodynamical shock problems. Applied Mathematics and Computation, 30(2 (part II)):99–123, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Heymann:1996:FDP

[HK96] Valerii I. Heymann and Arkadii V.

Kryazhimskii. On finitedimensional parametrizations of attainability sets. Applied Mathematics and Computation, 78(2– 3):137–151, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600004.

Hess:1979:OTC

[HKK^{+}79] J. Hess, H. Kagiwada, R. Kalaba, K. Spingarn, and C. Tsokos. Optimal tactics for close support operations. II. degraded communications with linear timevarying dynamics. Applied Mathematics and Computation, 5(3):199–212, ???? 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hess:1980:OTC

[HKK^{+}80] J. Hess, Harriet Kagiwada, R. Kalaba, K. Spingarn, and C. Tsokos. Optimal tactics for close support operations. III. degraded intelligence and communications. Applied Mathematics and Computation, 6(3):217–228, April 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Helms:1978:FDE

[HKKN78] Jay Helms, Harriet Kagiwada, Robert Kalaba, and John Niedercorn. Fitting differential equation models to observed economic data — I. Quasilinearization. Applied Mathematics and Computation, 4(??):139– 145, 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hess:1979:CDP

[HKKT79] James Hess, Harriet Kagiwada, Robert Kalaba, and Chris Tsokos. Cooperative dynamic programming. Applied Mathematics and Computation, 5(1): 69–74, ???? 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hutchinson:1998:HIC

[HKL98] John Hutchinson, Mark J. Kaiser, and Hamid M. Lankarani. The Head Injury Criterion (HIC) functional. Applied Mathematics and Computation, 96(1):1–16, October 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/1/6119.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
1&aid=6119.

Hipel:1997:DSS

[HKLP97] Keith W. Hipel, D. Marc Kilgour, Fang Liping, and Xiaoyong (John) Peng. The decision support system GMCR in environmental conflict management. Applied Mathematics and Computation, 83(2–3): 117–152, May 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=2-3&aid=9600170.

Hicks:1993:ANC

[HKLS93] D. L. Hicks, K. L. Kuttler, L. M. Liebrock, and C. E. Southwell. Answering von Neumann’s conjecture on convergence of averages and analyzing the von Neumann-Richtmyer scheme via a material averaging method. Applied Mathematics and Computation, 55(2– 3):265–295, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hemker:1983:MMD

[HKWdZ83] P. W. Hemker, R. Kettler, P. Wesseling, and P. M. de Zeeuw. Multigrid methods: development of fast solvers. Applied Mathematics and Computation, 13(3–4):311–326, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hyman:1982:NDD

[HL82] James M. Hyman and Bernard Larrouturou. The numerical differentiation of discrete functions using polynomial interpolation methods. Applied Mathematics and Computation, 10/11 (1):487–506, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Hu:1988:PBVb

[HL88a] Shou Chuan Hu and S. Leela. Periodic boundary value problem for a second order integrodifferential equation of Hammerstein type. Applied Mathematics and Computation, 25 (3):219–226, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hu:1988:PBVa

[HL88b] Shou Chuan Hu and S. Leela. Periodic boundary value problem for integro-differential equations of Hammerstein type. Applied Mathematics and Computation, 25(1 (part I)):29–38, ???? 1988. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Hicks:1991:PAI

[HL91] Darrell L. Hicks and Lorie M.

Liebrock. Parallel algorithms for implicit material dynamics schemes. Applied Mathematics and Computation, 46(3):179– 209, December 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hicks:1994:EEE

[HLW94] D. L. Hicks, L. M. Liebrock, and Y. Wen. Eigenform error estimates of computational material dynamics with shocks. 1.1. Hooke’s law materials. Applied Mathematics and Computation, 66(2–3):181–216, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hon:1997:MMN

[HLXZ97] Y. C. Hon, M. W. Lu, W. M.

Xue, and Y. M. Zhu. Multiquadric method for the numerical solution of a biphasic mixture model. Applied Mathematics and Computation, 88(2–3): 153–175, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600309.

Hady:1994:MCR

[HM94] F. M. Hady and R. A. Mohamed. Mixed convectionradiation interaction heat transfer in boundary-layer over a semi-infinite flat plate embedded in porous medium with a transfer magnetic field. Applied Mathematics and Computation, 66(2–3):103–128, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hon:1998:ENS

[HM98] Y. C. Hon and X. Z. Mao. An efficient numerical scheme for Burgers’ equation. Applied Mathematics and Computation, 95(1):37–50, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/1/6091.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
1&aid=6091.

Hossain:1999:CNG

[HM99] Md. Akram Hossain and A. S. Miah. Crank–Nicolson–Galerkin model for transport in groundwater: Refined criteria for accuracy. Applied Mathematics and Computation, 105(2–3): 173–181, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/2-
3/6321.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/

browse.cgi?year=1999&volume=
105&issue=2-3&aid=6321.

Hoffman:1987:CCT

[Hof87] Gilbert H. Hoffman. A combination conformal-transfinite mapping method for grids about finafterbody combinations. Applied Mathematics and Computation, 22(??):309–331, 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hoffman:1988:AGG

[Hof88] Gilbert H. Hoffman. Algebraic grid generation for an afterbody with finite-span, tapered fins. Applied Mathematics and Computation, 28(1):15–37, October 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hoffman:1989:VCC

[Hof89] William C. Hoffman. The visual cortex is a contact bundle. Applied Mathematics and Computation, 32(2–3):137–167, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Mathematical biology.

Hogeweg:1988:CAP

[Hog88] P. Hogeweg. Cellular automata as a paradigm for ecological modeling. Applied Mathematics and Computation, 27(1):81–100, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hohl:1997:MPS

[Hoh97] Jean-Christophe Hohl. Massively parallel search for linear factors in polynomials with many variables. Applied Mathematics and Computation, 85(2– 3):227–243, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600138.

Hojati:1994:ULP

[Hoj94] M. Hojati. Using linear programming to seek the optimum combination of investment. Applied Mathematics and Computation, 66(2–3):175–179, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Holter:1986:VMS

[Hol86] William H. Holter. A vectorized multigrid solver for the three-dimensional Poisson equation. Applied Mathematics and Computation, 19(1–4):127– 144, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Holland:1989:SNF

[Hol89] John H. Holland. Searching nonlinear functions for high values. Applied Mathematics and Computation, 32(2–3):255–274, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Mathematical biology.

Hossain:1999:MAD

[Hos99] Md. Akram Hossain. Modeling advective–dispersive transport with reaction: An accurate explicit finite difference model. Applied Mathematics and Computation, 102 (2–3):101–108, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6259.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6259.

Howell:1994:ECE

[How94] Gary W. Howell. Efficient computation of eigenvalues of randomly generated matrices. Applied Mathematics and Computation, 66(1):9–24, November 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hanumara:1988:SRB

[HQU88] R. Choudary Hanumara, Eugene P. Quinn, and V. R. R. Uppuluri. Some results on the best of paired comparisons. Applied Mathematics and Computation, 28(3 (part I)):247–268, December 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hartfield:1998:CFI

[HRA98] Roy Hartfield, Steve Rose, and John Abbitt. Computational fluid imaging for iodine fluorescence in compressible flows. Applied Mathematics and Computation, 95(1):63–73, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/95/1/6099.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=95&issue=
                                        1&aid=6099.

Hosea:1994:EEC

[HS94] M. E. Hosea and L. F. Shampine. Estimating the error of the classic Runge–Kutta formula. Applied Mathematics and Computation, 66(2–3): 217–226, December 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hamdan:1996:DGF

[HS96] M. H. Hamdan and K. D. Sawalha. Dusty gas flow through porous media. Applied Mathematics and Computation, 75 (1):59–73, March 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        75&issue=1&aid=9500104.

Hernandez:1998:CMC

[HS98] M. A. Hernández and M. A.

Salanova. Chebyshev method and convexity. Applied Mathematics and Computation, 95 (1):51–62, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/1/6100.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
1&aid=6100.

Hernandez:1999:ICC

[HS99] M. A. Hernández and M. A.

Salanova. Indices of convexity and concavity. Application to Halley method. Applied Mathematics and Computation, 103(1):27–49, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/103/1/
6269.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=103&
issue=1&aid=6269.

Hicks:1997:CSS

[HSA97] D. L. Hicks, J. W. Swegle, and S. W. Attaway. Conservative smoothing stabilizes discrete-numerical instabilities in SPH material dynamics computations. Applied Mathematics and Computation, 85(2–3):209– 226, September 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        85&issue=2-3&aid=9600136.

Han:1994:WCM

[HSG94] Bo Han, Ji-Qun Shao, and Bao-Qi Guo. A widely convergent method for determining the distributed parameters of an elliptical equation. Applied Mathematics and Computation, 60(2– 3):139–146, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Harris:1994:SMR

[HSZ94] G. Harris, P. Sellers, and B. Zinner. Some multiplicity results for a class of nonlinear boundary value problems. Applied Mathematics and Computation, 65(1– 3):195–208, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Hempel:1988:NVA

[HT88] R. Hempel and C. Thompson. A note on the vectorization of algebraic multigrid algorithms. Applied Mathematics and Computation, 26(3):245–256, 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Habib:1990:ISNa

[HT90a] M. K. Habib and A. Thavaneswaran. Inference for stochastic neuronal models. Applied Mathematics and Computation, 38(1 (part I)):51–73, July 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Habib:1990:ISNb

[HT90b] M. K. Habib and A. Thavaneswaran. Inference for stochastic neuronal models. Applied Mathematics and Computation, 39(3 (Supplement)): 183s–205s, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hamann:1998:DTS

[HT98] Bernd Hamann and Po-Yu Tsai.

Decomposing trimmed surfaces using the Voronoï diagram and a scan line algorithm. Applied Mathematics and Computation, 89(1–3):327–344, January 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781666.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Huang:1991:CAM

[Hua91] Wei Zhang Huang. Convergence of algebraic multigrid methods for symmetric positive definite matrices with weak diagonal dominance. Applied Mathematics and Computation, 46(2 (part II)):145–164, November 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Huang:1999:EES

[Hua99] Zhengda Huang. On the error estimates of several Newtonlike methods. Applied Mathematics and Computation, 106 (1):1–16, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/17/17/00006306.
pdf;   http://www.elsevier.
nl/gej-ng/10/9/12/71/17/17/
abstract.html.

Huss:1990:ALA

[Hus90] Ronald E. Huss. An Ada library for automatic evaluation of derivatives. Applied Mathematics and Computation, 35(2 (part II)):103–123, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Also appeared as Working Paper, Hughes Aircraft Company, February, 1989.

Holter:1990:CVM

[HV90] Bill Holter and George Vanden-Berghe. A comparison of vectorized methods for solving the two-dimensional diffusion equation: multigrid versus polynomial preconditioned conjugate gradient. Applied Mathematics and Computation, 40(1 (part I)):77–103, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Haitovsky:1980:GRR

[HW80] Yoel Haitovsky and Yohanan Wax. Generalized ridge regression, least squares with stochastic prior information, and Bayesian estimators. Applied Mathematics and Computation, 7(2):125–154, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hebson:1985:PSP

[HW85] Charles Hebson and Eric F.

Wood. Partitioned state and parameter estimation for realtime flood forecasting. Applied Mathematics and Computation, 17(4):357–374, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). System identification and parameter estimation of environmental systems.

Han:1998:CDN

[HW98] Danfu Han and Xinghua Wang.

Convergence on a deformed Newton method. Applied Mathematics and Computation, 94 (1):65–72, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/1/6079.

pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
1&aid=6079.

Hossain:1998:MCT

[HY98] Md. Akram Hossain and David R. Yonge. Modeling contaminant transport in groundwater: an optimized finite element method. Applied Mathematics and Computation, 96 (1):89–100, October 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/1/6125.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
1&aid=6125.

Hossain:1999:SAD

[HY99a] Md. Akram Hossain and D. R. Yonge. Simulating advectivedispersive transport in groundwater: an accurate finite difference model. Applied Mathematics and Computation, 105 (2–3):221–230, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/2-
3/6324.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
105&issue=2-3&aid=6324.

Hossain:1999:ATG

[HY99b] Md. Akram Hossain and David R.

Yonge. Accuracy of the Taylor– Galerkin model for contaminant transport in groundwater. Applied Mathematics and Computation, 102(2–3):109–119, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6261.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6261.

Hossain:1999:GMT

[HY99c] Md. Akram Hossain and David R.

Yonge. On Galerkin models for transport in groundwater. Applied Mathematics and Computation, 100(2–3):249–263, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6209.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6209.

Heilman:1997:TFI

[HYL97] Philip Heilman, Diana S.

Yakowitz, and Leonard J. Lane. Targeting farms to improve water quality. Applied Mathematics and Computation, 83(2–3): 173–194, May 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        83&issue=2-3&aid=9600168.

Han:1995:AEE

[HZ95a] Guoqiang Han and Liqing Zhang. Asymptotic error expansion of a collocation-type method for Hammerstein equations. Applied Mathematics and Computation, 72(1):1–19, September 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        72&issue=1&aid=9400160.

Huang:1995:DOF

[HZ95b] Y. Huang and N. G. Zamani. Design optimization via finite elements: An elementary account. Applied Mathematics and Computation, 69(2–3): 275–298, May 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        69&issue=2-3&aid=9400135.

Huang:1996:OBD

[HZ96] Y. Huang and N. G. Zamani. Optimum beam design using finite elements. Applied Mathematics and Computation, 80(2– 3):155–179, December 1, 1996.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500285.

Huang:1998:FPR

[HZ98] Zhengda Huang and Shiming Zheng. On a family of parallel root-finding methods for generalized polynomials. Applied Mathematics and Computation, 91(2–3):221–231, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6021.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6021.

Han:1997:WCG

[HZL97] B. Han, M. L. Zhang, and J. Q. Liu. A widely convergent generalized pulse-spectrum technique for the coefficient inverse problem of differential equations. Applied Mathematics and Computation, 81(2–3): 97–112, February 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=2-3&aid=9500304.

Iyengar:1980:SSA

[IA80] S. Sitharama Iyengar and Vincent Alia. A string searching algorithm. Applied Mathematics and Computation, 6(??):123– 131, 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ioakimidis:1985:NSA

[IA85] N. I. Ioakimidis and E. G. Anastasselou. A new, simple approach to the derivation of exact analytical formulae for the zeros of analytic functions. Applied Mathematics and Computation, 17(2):123–127, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Itoh:1989:DLE

[IA89] Toshiaki Itoh and Kanji Abe. Discrete Lagrange’s equations and canonical equations based on the principle of least action. Applied Mathematics and Computation, 29(2 (part II)): 161–183, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ibrahim:1996:DSV

[Ibr96] A.-G. M. Ibrahim. On the differentiability of set-valued functions defined on a Banach space and mean value theorem. Applied Mathematics and Computation, 74(1):79–94, January 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=1&aid=9500095.

Issa:1994:AFE

[IBS94] Camille A. Issa, Krishnan Balasubramaniam, and Kanthikannan Srirengan. Adaptive hpfinite element for transient analysis. Applied Mathematics and Computation, 65(1–3):315–333, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Iwakoshi:1998:FCS

[IFU98] Yasushi Iwakoshi, Takeshi Furuhashi, and Yoshiki Uchikawa. A fuzzy classifier system for evolutionary learning of robot behavior. Applied Mathematics and Computation, 91 (1):73–81, April 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/1/6006.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
1&aid=6006.

Iannelli:1997:SMN

[IKP97] Mimmo Iannelli, Mi-Young Kim, and Eun-Jae Park. Splitting methods for the numerical approximation of some models of age-structured population dynamics and epidemiology. Applied Mathematics and Computation, 87(1):69– 93, November 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=1&aid=9600222.

Itiki:1996:ICP

[IKU96] Cinthia Itiki, Robert Kalaba, and Firdaus Udwadia. Inequality constraints in the process of jumping. Applied Mathematics and Computation, 78(2–3):163– 173, September 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600006.

Ivanenko:1995:NMU

[IL95] V. I. Ivanenko and V. A. Labkovsky. On the natural measure of uncertainty and information. Applied Mathematics and Computation, 70(2–3): 185–190, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400121.

Dynamics and control (Vienna, 1993).

Ianiro:1980:NDS

[IM80] N. Ianiro and C. Maffei. Nonautonomous differential systems: several functions in the stability problem. Applied Mathematics and Computation, 7(1):71–80, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Iakovidis:1992:IPE

[IM92] Ilias Iakovidis and Clyde F. Martin. The inverse problem in electrocardiography on a spherical model. Applied Mathematics and Computation, 48(2–3): 187–217, April 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Inohara:1999:CMR

[Ino99] Takehiro Inohara. On conditions for a meeting not to reach a recurrent argument. Applied Mathematics and Computation, 101(2–3):281–298, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/101/2-
3/6231.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
101&issue=2-3&aid=6231.

Ioakimidis:1983:NSC

[Ioa83] N. I. Ioakimidis. On the numerical solution of Cauchy type singular integral equations by the collocation method. Applied Mathematics and Computation, 12(1):49–60, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ioakimidis:1988:AQR

[Ioa88] N. I. Ioakimidis. Application of quadrature rules to the determination of plane equipotential lines and other curves defined by harmonic functions. Applied Mathematics and Computation, 27(2):147–154, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ioakimidis:1992:COS

[Ioa92] Nikolaos I. Ioakimidis. Computation of the orders of singularity of sectionally analytic functions. Applied Mathematics and Computation, 48(1):13– 19, March 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ibidapo-Obe:1995:ACB

[IOAB95] O. Ibidapo-Obe, A. B. Alonge, and Adedeji B. Badiru. On active controls for a biped mechanism. Applied Mathematics and Computation, 69(2–3): 159–183, May 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        69&issue=2-3&aid=9400090.

Ioakimidis:1989:NCQ

[IP89] N. I. Ioakimidis and K. E. Papadakis. A new class of quite elementary closed-form integral formulae for roots of nonlinear equations. Applied Mathematics and Computation, 29(2 (part II)):185–196, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ioakimidis:1991:NEC

[IPP91] N. I. Ioakimidis, E. A. Perdios, and K. E. Papadakis. Numerical estimation of the coefficient of the homogeneous Riemann– Hilbert problem on the basis of boundary data. Applied Mathematics and Computation, 41(1 (part I)):21–33, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Iyengar:1983:SPL

[IRU83] S. Sitharama Iyengar, A. K. Rajagopal, and V. R. R. Uppuluri. String patterns of leading digits. Applied Mathematics and Computation, 12(4):321– 337, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ismail:1993:EES

[IST93] Fudziah Bte Ismail, Mohamed B.

Suleiman, and Bachok B. Taib. Existence of h^{2}-expansion for some extrapolation methods for solving autonomous higher order ODEs. Applied Mathematics and Computation, 58(1):69– 83, September 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Inohara:1997:IDC

[ITN97] Takehiro Inohara, Shingo Takahashi, and Bunpei Nakano. Impossibility of deception in a conflict among subjects with interdependent preference. Applied Mathematics and Computation, 81(2–3):221–244, February 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=2-3&aid=9600018.

Inohara:1998:CSI

[ITN98a] Takehiro Inohara, Shingo Takahashi, and Bunpei Nakano. Complete stability and inside commonality of perceptions. Applied Mathematics and Computation, 90(1):11–25, March 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=1&aid=9600334.

Inohara:1998:CMR

[ITN98b] Takehiro Inohara, Shingo Takahashi, and Bunpei Nakano. On conditions for a meeting not to reach a deadlock. Applied Mathematics and Computation, 90 (1):1–9, March 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=1&aid=9600333.

Ives:1982:CGG

[Ive82] David C. Ives. Conformal grid generation. Applied Mathematics and Computation, 10/11(1): 107–135, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Jang:1999:SFC

[JA99] Sophia R.-J. Jang and Linda J. S. Allen. A simple food chain with a growth inhibiting nutrient. Applied Mathematics and Computation, 104 (2–3):277–298, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/2-
3/6297.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
104&issue=2-3&aid=6297.

Jackson:1990:EES

[Jac90]  Dennis E. Jackson.  Error

estimates for the semidiscrete finite-element approximation of a damped wave equation with a smoothing property. Applied Mathematics and Computation, 35(3):243–261, February 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jackson:1992:EES

[Jac92] D. E. Jackson. Error estimates for the semidiscrete Galerkin approximations of the FitzHugh– Nagumo equations. Applied Mathematics and Computation, 50(1):93–114, July 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jacobson:1993:VBR

[Jac93] Sheldon H. Jacobson. Variance and bias reduction techniques for the harmonic gradient estimator. Applied Mathematics and Computation, 55(2– 3):153–186, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jacobson:1995:AMM

[Jac95] Sheldon H. Jacobson. Analyzing the M/M/1 queue in frequency domain experiments. Applied Mathematics and Computation, 69(2–3):185–194, May 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        69&issue=2-3&aid=9400125.

Jameson:1983:SEE

[Jam83] Antony Jameson. Solution of the Euler equations for twodimensional transonic flow by a multigrid method. Applied Mathematics and Computation, 13(3–4):327–355, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jameson:1985:PAA

[Jam85] Donald A. Jameson. A priori analysis of allowable interval between measurements as a test of model validity. Applied Mathematics and Computation, 17(2):93–105, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jameson:1986:SIM

[Jam86] Donald A. Jameson. Sampling intensity for monitoring of environmental systems. Applied Mathematics and Computation, 18(??):71–76, 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Janssens:1995:EMS

[Jan95] F. Janssens. Exact mode shapes for a continuous tether. Applied Mathematics and Computation, 70(2–3):191–201, July 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400120.

Jayasuriya:1996:MAU

[Jay96] Kumara Jayasuriya. Multiprecision arithmetic using fast Hartley transforms. Applied Mathematics and Computation, 75 (2–3):239–251, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500128.

Jolayemi:1989:MCC

[JB89a] J. K. Jolayemi and J. N. Berrettoni. Multivariate control charts: an optimization approach to the effective use of measurement of performance. Applied Mathematics and Computation, 32(1):1–15, July 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jolayemi:1989:ODM

[JB89b] J. K. Jolayemi and J. N. Berrettoni. An optimal design of multivariate control charts in the presence of multiple assignable causes. Applied Mathematics and Computation, 32(1):17–33, July 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jimenez:1998:ESA

[JBV98] S. Jiménez, S. Bulgakov, and L. Vázquez. Efficient shooting algorithms for solving the nonlinear one-dimensional scalar Helmholtz equation. Applied Mathematics and Computation, 95(2–3):101–114, September 15, 1998. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6124.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6124.

Jang:1997:ARS

[JC97] Ming-Jyi Jang and Chieh-Li Chen. Analysis of the response of a strongly nonlinear damped system using a differential transformation technique. Applied Mathematics and Computation, 88(2–3):137–151, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600308.

Jonckheere:1998:SAC

[JCC98] Edmond A. Jonckheere, Chung-Kuang Chu, and Chih-Yung Cheng. Simplicial algorithms for computing stationary probabilities of stochastic matrices. Applied Mathematics and Computation, 93(2–3):207–217, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/2-3/
6089.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6089.

Jerrell:1988:RSS

[Jer88] Max E. Jerrell. A random search strategy for function optimization. Applied Mathematics and Computation, 28(3 (part I)): 223–229, December 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jespersen:1983:DIM

[Jes83] Dennis C. Jespersen. Design and implementation of a multigrid code for the Euler equations. Applied Mathematics and Computation, 13(3–4):357– 374, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jodar:1991:IDM

[JF91] L. Jódar and M. Legua Fernández. An implicit difference method for the numerical solution of coupled systems of partial differential equations. Applied Mathematics and Computation, 46 (2 (part II)):127–134, November 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jenkins:1996:NAU

[JFW96] Rhonald M. Jenkins, Winfred A. Foster, Jr., and Lora S. Wirth. Numerical analysis of unsteady multiple jet plume interactions. Applied Mathematics and Computation, 79(2–3): 239–247, October 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500268.

Ji:1994:ALE

[Ji94] Jun Ji. An alternative limit expression of Drazin inverse and its application. Applied Mathematics and Computation, 61(2–3): 151–156, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jiaxing:1999:LCB

[Jia99]  He Jiaxing.  On a linear

combination of S.N. Bernstein trigonometric interpolation polynomial. Applied Mathematics and Computation, 106 (2–3):197–203, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/24/
article.pdf.

Jimenez:1994:DDC

[Jim94] Salvador Jiménez. Derivation of the discrete conservation laws for a family of finite difference schemes. Applied Mathematics and Computation, 64(1):13–45, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jing:1994:ABC

[Jin94] Chen Jing. Alternating block Crank–Nicolson method for the 3-D heat equation. Applied Mathematics and Computation, 66(1):41–61, November 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jin:1995:FAB

[Jin95] Xiao-Qing Jin. A fast algorithm for block Toeplitz systems with tensor structure. Applied Mathematics and Computation, 73(2– 3):115–124, December 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        73&issue=2-3&aid=9400243.

Jin:1997:NCC

[Jin97] Xiao-Qing Jin. A note on construction of circulant preconditioners from kernels. Applied Mathematics and Computation, 83(1):3–12, April 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        83&issue=1&aid=9600332. See

addendum [JS98a].

Janicki:1998:WOS

[JK98] Ryszard Janicki and Waldemar W. Koczkodaj. A weak order solution to a group ranking and consistency-driven pairwise comparisons. Applied Mathematics and Computation, 94(2– 3):227–241, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/2-3/
6095.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
2-3&aid=6095.

Jonckheere:1997:BFT

[JkC97] Edmond A. Jonckheere and Chung kuang Chu. Bounded flatness in Q-triangulated regular N-simplexes. Applied Mathematics and Computation, 88(2– 3):177–198, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600310.

Jodar:1993:SSO

[JL93] Lucas Jódar and Matilde Legua.

Solving second-order matrix differential equations with regular singular points avoiding the increase of the problem dimension. Applied Mathematics and Computation, 53(2–3): 191–205, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jeong:1997:EMA

[JL97] Il-Kwon Jeong and Ju-Jang Lee. Evolving multi-agents using a self-organizing genetic algorithm. Applied Mathematics and Computation, 88(2–3):293– 303, December 30, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        88&issue=2-3&aid=9600337.

Jenkins:1993:ICT

[JM93] R. M. Jenkins and D. A. Moore. An inverse calculation technique for quasi-three-dimensional turbomachinery cascades. Applied Mathematics and Computation, 57(2–3):197–204, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jukic:1998:DTN

[JMS98] D. Jukić, T. Marošević, and R. Scitovski. Discrete total l_{p}- norm approximation problem for the exponential function. Applied Mathematics and Computation, 94(2–3):137–143, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/94/2-3/
                                        6080.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=94&issue=
                                        2-3&aid=6080.

Jodar:1993:NMM

[JMV93] L. Jódar, J. L. Morera, and R. J. Villanueva. Numerical multisteps matrix methods for Y′′ = f(t, Y). Applied Mathematics and Computation, 59 (2–3):257–274, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jodar:1991:ESC

[JN91] L. Jódar and E. Navarro. Exact solution of a class of strongly coupled Riccati matrix differential equations. Applied Mathematics and Computation, 43(2 (part II)):165–173, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Juang:1995:CRI

[JN95] Jonq Juang and Paul Nelson. Convergence rates of iterative solutions of algebraic matrix Riccati equations. Applied Mathematics and Computation, 72(2–3):125–142, October 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400179.

Jodar:1995:CSI

[JNF95] L. Jódar, E. Navarro, and M. V.

Ferrer. A correction on the stability of implicit higher order difference systems. erratum: “Closed form general solution of nonhomogeneous implicit higher order difference systems” [Appl. Math. Comput. 60 (1994), no. 2–3, 113–123, MR 95c:39004]. Applied Mathematics and Computation, 71(1):91–96, August 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=1&aid=9500060. See

[NFJ94].

Jodar:1996:CSI

[JNF96] L. Jódar, E. Navarro, and M. V. Ferrer. A correction on the stability of implicit higher order difference systems. Applied Mathematics and Computation, 74(2– 3):299–304, February 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500217.

Jodar:1992:CFS

[JNM92] L. Jódar, E. Navarro, and J. L. Morea. A closed-form solution of singular regular higher-order difference initial and boundary value problems. Applied Mathematics and Computation, 48(2– 3):153–166, April 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jodar:1992:RAS

[Jód92] Lucas Jódar. Rational approximate solutions and error bounds for the nonsymmetric Riccati matrix differential equation. Applied Mathematics and Computation, 50(1):23–31, July 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Johnson:1982:NME

[Joh82] Billy H. Johnson. Numerical modeling of estuarine hydrodynamics on a boundary-fitted coordinate system. Applied Mathematics and Computation, 10/11 (1):409–436, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Johnson:1983:MGC

[Joh83] Gary M. Johnson. Multiplegrid convergence acceleration of viscous and inviscid flow computations. Applied Mathematics and Computation, 13(3–4): 375–398, November 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jolayemi:1992:UAS

[Jol92] Joel K. Jolayemi. A unified approximation scheme for the convolution of independent binomial variables. Applied Mathematics and Computation, 49 (2–3):269–??, June 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jou:1982:EMG

[Jou82] Wen Huei Jou. An experience in mesh generation for threedimensional calculation of potential flow around a rotating propeller. Applied Mathematics and Computation, 10/11(1): 547–561, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Jodar:1995:SOD

[JPF95] Lucas Jódar, Enrique Ponsoda, and M. Legua Fernández. Second-order difference systems in unbounded domains: Existence conditions and construction of exact and approximate bounded solutions. Applied Mathematics and Computation, 71(2–3):271–287, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400158.

Jordan:1986:NSN

[JPS86] Kirk E. Jordan, George C. Papanicolaou, and R. Spigler. On the numerical solution of a nonlinear stochastic Helmholtz equation with a multigrid preconditioner. Applied Mathematics and Computation, 19(1–4): 145–157, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Jayakumar:1993:CMS

[JR93] J. Jayakumar and N. Ramanujam. A computational method for solving singular perturbation problems. Applied Mathematics and Computation, 55(1):31–48, April 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jayakumar:1995:CMS

[JR95] J. Jayakumar and N. Ramanujam. A computational method for solving quasilinear singular perturbation problems. Applied Mathematics and Computation, 71(1):1–14, August 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=1&aid=9400077.

Jukic:1996:EOP

[JS96] Dragan Jukić and Rudolf Scitovski. The existence of optimal parameters of the generalized logistic function. Applied Mathematics and Computation, 77 (2–3):281–294, July 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jin:1998:ANC

[JS98a] Xiao-Qing Jin and Vai-Kuong Sin. Addendum to: “A note on construction of circulant preconditioners from kernels” [Appl.

Math. Comput. 83 (1997), no. 1, 3–12; MR 97i:65077, by Jin]. Applied Mathematics and Computation, 95(1):91–99, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/1/6082.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=

1&aid=6082. See [Jin97].

Jinzenji:1998:ASS

[JS98b] T. Jinzenji and M. Sasaki. An advanced simulation scheme for electric railway power systems based on artificial life approach. Applied Mathematics and Computation, 91(1):23–32, April 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/1/6012.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
1&aid=6012.

Jernigan:1980:SNS

[JT80] Robert W. Jernigan and Chris P. Tsokos. Simulation of a nonlinear stochastic ecology model. Applied Mathematics and Computation, 7(1):9–25, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jernigan:1980:PMB

[JTT80] R. W. Jernigan, J. C. Turner, and C. P. Tsokos. Parameter and moment bounds for a nonlinear stochastic ecology model. Applied Mathematics and Computation, 7(1):27–53, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jimenez:1990:AFN

[JV90] Salvador Jiménez and Luis Vázquez. Analysis of four numerical schemes for a nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 35(1 (part I)):61–94, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jodar:1994:ASC

[JV94] Lucas Jódar and R. J. Villanueva. Analytic solutions of coupled second-order delay differential equations. Applied Mathematics and Computation, 62(2–3):235–247, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Jiang:1998:QMS

[JV98] Jiahui Jiang and A. S. Vatsala. The quasilinearization method in the system of reaction diffusion equations. Applied Mathematics and Computation, 97(2–3):223–235, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6145.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6145.

Jayawardene:1997:ULO

[JY97] Sumedha Jayawardene and Shie-Shien Yang. On using linear ordered rank statistics for detecting early differences between two distributions. Applied Mathematics and Computation, 84(2– 3):103–113, July 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600081.

Kadalbajoo:1995:BEM

[KA95] Mohan K. Kadalbajoo and K. Arumugam. Boundary element method for elliptic differential equations with a small parameter. Applied Mathematics and Computation, 71(2–3):151– 163, September 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400151.

Kumar:1996:AHM

[KA96a] Kuldeep Kumar and M. A. Alsaleh. Application of Hankel matrices in polynomial regression. Applied Mathematics and Computation, 77(2–3):205–211, ???? 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:1996:CSE

[KA96b] Kuldeep Kumar and M. A. Alsaleh. A comparative study for the estimation of parameters in nonlinear models. Applied Mathematics and Computation, 77 (2–3):179–183, July 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1986:NSS

[Kad86] Mohan K. Kadalbajoo. Numerical solution of singular integral equations via initial value methods. Applied Mathematics and Computation, 18(3):211– 222, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kagiwada:1991:BWB

[Kag91a] Harriet H. Kagiwada. Best wishes to Bob on his 65th birthday. Applied Mathematics and Computation, 45(2 (part II)): 91–92, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kagiwada:1991:KWA

[Kag91b] Harriet H. Kagiwada. Kalaba’s work in applied mathematics: modeling, optimization, and inverse problems. Applied Mathematics and Computation, 45(2 (part II)):93–97, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kailath:1991:ROD

[Kai91] T. Kailath. Remarks on the origin of the displacement-rank concept. Applied Mathematics and Computation, 45(2 (part II)):193–206, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kamowitz:1988:MAS

[Kam88] David Kamowitz. Multigrid applied to singular perturbation problems. Applied Mathematics and Computation, 25(2 (part II)):145–174, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kananthai:1997:SDD

[Kan97] A. Kananthai. On the solutions of the n-dimensional diamond operator. Applied Mathematics and Computation, 88(1): 27–37, December 15, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        88&issue=1&aid=9600303.

Kananthai:1999:FTD

[Kan99]  Amnuay Kananthai.  On

the Fourier transform of the Diamond Kernel of Marcel Riesz. Applied Mathematics and Computation, 101(2– 3):151–158, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/101/2-
3/6227.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
101&issue=2-3&aid=6227.

Kapur:1980:GPG

[Kap80] J. N. Kapur. A general population growth model with density dependence. Applied Mathematics and Computation, 6(1): 37–50, January 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kapitza:1988:TIF

[Kap88] H. Kapitza. Truncated incomplete factorizations for conjugate-gradient methods in two and three dimensions. Applied Mathematics and Computation, 28(1):73–87, October 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karsai:1998:ABV

[Kar98] János Karsai. Asymptotic behavior and its visualization of the solutions of intermittently and impulsively damped nonlinear oscillator equations. Applied Mathematics and Computation, 89(1–3):161–172, January 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        89&issue=1-3&aid=9781655.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Kazantsev:1999:LLE

[Kaz99]  E. Kazantsev.  Local Lya-

punov exponents of the quasigeostrophic ocean dynamics. Applied Mathematics and Computation, 104(2–3):217–257, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/amc/sub/
                                        1999/104/2-3/6299.pdf;   http:
                                        //www.elsevier.com/cgi-bin/
                                        cas/tree/store/amc/cas_sub/
                                        browse/browse.cgi?year=1999&
                                        volume=104&issue=2-3&aid=
                                        6299.

Kowalik:1980:SCA

[KB80] J. S. Kowalik and W. R. Boland. Some computational advances in unconstrained optimization. Applied Mathematics and Computation, 7(1):55–69, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1984:FES

[KB84a] Mohan K. Kadalbajoo and K. K. Bharadwaj. Fast elliptic solvers—an overview. Applied Mathematics and Computation, 14(4):331–355, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1984:SMTb

[KB84b] Mohan K. Kadalbajoo and K. K.

Bharadwaj. Symmetric marching technique and elliptic equations with variable coefficients involving mixed partial derivatives. Applied Mathematics and Computation, 15(4):315– 323, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1984:SMTa

[KB84c] Mohan K. Kadalbajoo and K. K.

Bharadwaj. Symmetric marching technique for the numerical solution of nonseparable elliptic equations. Applied Mathematics and Computation, 14 (2):197–206, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1985:FDM

[KB85] Mohan K. Kadalbajoo and K. K.

Bharadwaj. On fast direct methods for solving elliptic equations over nonrectangular regions. Applied Mathematics and Computation, 16(1):1–17, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kulev:1989:GSS

[KB89] G. K. Kulev and D. D. Baĭnov.

Global stability of sets for systems with impulses. Applied Mathematics and Computation, 29(3):255–270, February 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1993:TOV

[KB93] Mohan K. Kadalbajoo and Rajesh K. Bawa. Third-order variable-mesh cubic spline methods for singularly-perturbed boundary-value problems. Applied Mathematics and Computation, 59(2–3):117–129, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khan:1998:PIE

[KB98] Shahjahan Khan and M. I. Bhatti. Predictive inference on equicorrelated linear regression models. Applied Mathematics and Computation, 95 (2–3):205–217, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6108.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6108.

Kutluay:1999:VFD

[KBÖ99] S. Kutluay, A. R. Bahadir, and A. Özde. A variety of finite difference methods to the thermistor with a new modified electrical conductivity. Applied Mathematics and Computation, 106(2–3):205–213, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/25/
article.pdf.

Katti:1996:CFD

[KBS96] C. P. Katti, S. N. Baboo, and S. Sivaloganathan. On the convergence of finite difference methods for a class of twopoint boundary value problems with periodic boundary conditions. Applied Mathematics and Computation, 75(2–3):287– 302, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500144.

Kwak:1990:PNG

[KCS90] Do Young Kwak, U. Jin Choi, and S. Son. Prime number grid scheme to approximate function spaces with their lowdimensional subspaces. Applied Mathematics and Computation, 40(3):233–237, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Keshavamurthy:1995:EWSa

[KD95a] Ramdass Keshavamurthy and Marijan Dravinski. Elastic wave scattering by an inclusion in a multilayered medium submerged in fluid. I. plane strain model. Applied Mathematics and Computation, 67(1–3):33– 60, January 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        67&issue=1-3&aid=9400053.

Keshavamurthy:1995:EWSb

[KD95b] Ramdass Keshavamurthy and Marijan Dravinski. Elastic wave scattering by an inclusion in a multilayered medium submerged in fluid. II. three-dimensional model. Applied Mathematics and Computation, 67(1–3):265– 292, January 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        67&issue=1-3&aid=9400061.

Keesman:1989:DPS

[Kee89] K. Keesman. On the dominance of parameters in structural models of ill-defined systems. Applied Mathematics and Computation, 30(2 (part II)): 133–147, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kelman:1981:CSC

[Kel81] Robert B. Kelman. Convergence of solutions to the classic dual cosine equation. Applied Mathematics and Computation, 9(3):217–225, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Konno:1991:MCA

[KG91] Norio Konno and Yukio Gunji.

Mathematical construction of an autonomous artificial life. Applied Mathematics and Computation, 46(1 (part I)):33–58, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kubicek:1978:OPI

[KH78] Milan Kubíček and Vladimír Hlaváček. One-parameter imbedding techniques for the solution of nonlinear boundaryvalue problems. Applied Mathematics and Computation, 4(4): 317–357, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kuttler:1987:SPH

[KH87] K. L. Kuttler and D. L. Hicks.

Some progress on the hydrocode convergence problem. Applied Mathematics and Computation, 23(3):211–233, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kamel:1996:ATF

[KH96] M. T. Kamel and M. H. Hamdan. Aspects of thin-film polar fluid lubrication. Applied Mathematics and Computation, 80(1):

33–41, November 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        80&issue=1&aid=9500278.

Khan:1981:ACA

[Kha81] M. Ali Khan. Approximate convexity of average sums of sets in normed linear spaces. Applied Mathematics and Computation, 9(1):27–34, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khanh:1991:NRD

[Kha91] Bui Doan Khanh. A numerical resolution of differentialalgebraic equations. Applied Mathematics and Computation, 44(3):261–269, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khanh:1992:NEH

[Kha92] Bui Doan Khanh. A numerical evaluation of the Hankel transform of bandlimited functions. Applied Mathematics and Computation, 48(2–3):83–100, April 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khanh:1994:CTR

[Kha94] Bui Doan Khanh. A computation for tomographic reconstruction from projections. Applied Mathematics and Computation, 64(2–3):79–91, ???? 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khan:1996:RBP

[Kha96] Shahjahan Khan. Regressionbased prediction for two-stage survey data with correlated normal errors. Applied Mathematics and Computation, 79(2–3): 105–124, October 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500245.

Khuri:1998:BCA

[Khu98a] S. A. Khuri. Biorthogonality condition for axisymmetric Stokes flow in a toroidal region. Applied Mathematics and Computation, 97(2–3): 255–259, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6153.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6153.

Khuri:1998:NAC

[Khu98b] S. A. Khuri. A new approach to the cubic Schrödinger equation: An application of the decomposition technique. Applied Mathematics and Computation, 97(2–3):251–254, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6151.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6151.

Khuri:1998:SIE

[Khu98c] S. A. Khuri. On the solution of an integral equation arising in water waves theory. Applied Mathematics and Computation, 95(2–3):291–295, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6023.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6023.

Kim:1996:ESS

[Kim96] Mi-Young Kim. Existence of steady state solutions to an epidemic model with screening and their asymptotic stability. Applied Mathematics and Computation, 74(1):37–58, January 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=1&aid=9500071.

Kincanon:1991:OSC

[Kin91] Eric Kincanon. An orthogonal set composed from the functions e^{nx} and n ≥ 0 and x ≤ 0. Applied Mathematics and Computation, 41(1 (part I)):69–75, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kincanon:1993:ASG

[Kin93] Eric Kincanon. Approximate solution to the Gel′fand–Levitan equation. Applied Mathematics and Computation, 53(2–3): 121–128, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kincanon:1994:SST

[Kin94] Eric Kincanon. A series solution for the gvψ_{0} term of the Born series. Applied Mathematics and Computation, 66(2–3):227–232, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kincanon:1997:AAR

[Kin97] Eric Kincanon. Ambiguities and almost reflectionless potentials. Applied Mathematics and Computation, 83(1):79–85, April 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=1&aid=9600048.

Kjellstrom:1999:EB

[Kje99] Gregor Kjellström. The evolution in the brain. Applied Mathematics and Computation, 98(2–3):293–300, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/
6159.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6159.

Kerlick:1982:AQC

[KK82] G. David Kerlick and Goetz H. Klopfer. Assessing the quality of curvilinear coordinate meshes by decomposing the Jacobian matrix. Applied Mathematics and Computation, 10/11(1): 787–807, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Kagiwada:1983:IVM

[KK83a] Harriet H. Kagiwada and Robert E. Kalaba. An initial value method for the nonlinear integral equation F (u(t), t, λ) = λ ∫_{0}^{1} k(t, y, u(y))dy. Applied Mathematics and Computation, 13(1–2):117–124, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kubicek:1983:DBB

[KK83b] Milan Kubíček and Alois Klíč.

Direction of branches bifurcating at a bifurcation point. determination of starting points for a continuation algorithm. Applied Mathematics and Computation, 13(1–2):125–142, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karaa:1984:BDA

[KK84] Fadi Karaa and Roman Krzysztofowicz. Bayesian decision analysis of dam safety. Applied Mathematics and Computation, 14 (??):357–380, 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kubica:1990:ACA

[KK90] K. Kubica and J. Kuczera. An application of cellular automata to model a lipid membrane. Applied Mathematics and Computation, 39(1):49–59, September 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kubica:1991:EAC

[KK91] Krystian Kubica and Janina Kuczera. The effect of the alkyl chain length of modifier molecules on membrane structure studied by the cellularautomaton method. Applied Mathematics and Computation, 44(2 (part II)):183–193, 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kagiwada:1995:FEF

[KK95] Harriet H. Kagiwada and Robert E. Kalaba. Fuzzy evidential filter for detection and tracking of dim objects. Applied Mathematics and Computation, 69(1):75–96, April 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=1&aid=9400100.

Kawamata:1995:SDT

[KKH95] Masayuki Kawamata, Eiichiro Kawakami, and Tatsuo Higuchi. Separable-denominator twodimensional adaptive filters with application to noise reduction in images. Applied Mathematics and Computation, 69(1): 111–122, April 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=1&aid=9400102.

Kwak:1999:MAC

[KKL99] Do Y. Kwak, Hyun J. Kwon, and Sungyun Lee. Multigrid algorithm for cell centered finite difference on triangular meshes. Applied Mathematics and Computation, 105(1):77–85, October 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/

store/amc/sub/1999/105/1/
6315.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=105&
issue=1&aid=6315.

Kalaba:1990:LPS

[KKM90] Robert Kalaba, Moon Kim, and James E. Moore, II. Linear programming and simple associative memories. Applied Mathematics and Computation, 40 (3):203–214, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kappeller:1996:OEM

[KKPL96] M. Kappeller, M. Kiehl, M. Perzl, and M. Lenke. Optimized extrapolation methods for parallel solution of IVPs on different computer architectures. Applied Mathematics and Computation, 77(2–3):301–315, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kagiwada:1987:FIE

[KKS87] Harriet H. Kagiwada, Robert E.

Kalaba, and Melvin R. Scott. The b and h functions for integral equations with displacement kernels: a computational method and an application to radiative transfer. Applied Mathematics and Computation, 23(2):93–101, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kagiwada:1991:KAM

[KKU91] Harriet H. Kagiwada, Julia K. Kagiwada, and Sueo Ueno. Kalaba’s associative memories for system identification. Applied Mathematics and Computation, 45(2 (part II)):135–142, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Katehakis:1985:DRP

[KL85] Michael N. Katehakis and Alan Levine. A dynamic routing problem—numerical procedures for light traffic conditions. Applied Mathematics and Computation, 17(3):267–276, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kuo:1989:TCF

[KL89] C.-C. Jay Kuo and Bernard C. Levy. Two-color Fourier analysis of the multigrid method with red-black Gauss–Seidel smoothing. Applied Mathematics and Computation, 29(1 (part I)):69– 87, January 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kang:1999:NIT

[KL99a] Kab Seok Kang and Sung Yun Lee. New intergrid transfer operator in multigrid method for P_{1}-nonconforming finite element method. Applied Mathematics and Computation, 100 (2–3):139–149, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6217.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6217.

Kim:1999:PGS

[KL99b] Taegil Kim and Chang-Ock Lee. A parallel Gauss–Seidel method using NR data flow ordering. Applied Mathematics and Computation, 99(2– 3):209–220, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6192.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6192.

Klevenhusen:1982:DEG

[Kle82] Karl D. Klevenhusen. 2-D elliptic grid generation using a singularity method and its application to transonic interference flows. Applied Mathematics and Computation, 10/11(1): 739–760, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Kubicek:1979:ELB

[KM79] Milan Kubíček and Miloš Marek. Evaluation of limit and bifurcation points for algebraic equations and nonlinear boundary-value problems. Applied Mathematics and Computation, 5(3):253–264, ???? 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kolp:1986:EMA

[KM86] O. Kolp and H. Mierendorff. Efficient multigrid algorithms for locally constrained parallel systems. Applied Mathematics and Computation, 19(1–4):169– 200, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Kocvara:1987:MMT

[KM87] Michal Kočvara and Jan Mandel. A multigrid method for three-dimensional elasticity and algebraic convergence estimates. Applied Mathematics and Computation, 23(2):121–135, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khamayseh:1994:SGG

[KM94] Ahmed Khamayseh and C. Wayne Mastin. Surface grid generation based on elliptic PDE models. Applied Mathematics and Computation, 65(1–3):253–264, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Kalachev:1995:OSS

[KM95] L. V. Kalachëv and R. M. M.

Mattheij. On optimally scaled systems for second-order scalar singularly perturbed problems. Applied Mathematics and Computation, 68(1):71–93, March 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=1&aid=9400091.

Krohn:1991:IMA

[KMR^{+}91] I. Krohn, S. Moltzahn, J. Rosenmüller, P. Sudhölter, and H.-M. Wallmeier. Implementing the modified LH algorithm. Applied Mathematics and Computation, 45(1 (part I)):31–72, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Keilson:1979:LTT

[KN79] J. Keilson and W. R. Nunn. Laguerre transformation as a tool for the numerical solution of integral equations of convolution type. Applied Mathematics and Computation, 5(4):313–359, ???? 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Keller:1989:HIP

[KN89] Herbert B. Keller and Paul Nelson. Hypercube implementations of parallel shooting. Applied Mathematics and Computation, 31(Special issue):574– 603, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Kaneko:1990:AAT

[KN90] Hideaki Kaneko and Richard Noren. An application of approximation theory to the numerical solution of a Fredholm equation with a mild singularity. Applied Mathematics and Computation, 37(1 (part I)):57–66, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Knight:1982:ACC

[Kni82] Doyle D. Knight. Application of curvilinear coordinate generation techniques to the computation of internal flows. Applied Mathematics and Computation, 10/11(1):357–384, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Keilson:1981:BLT

[KNS81] J. Keilson, W. Nunn, and U. Sumita. The bilateral Laguerre transform. Applied Mathematics and Computation, 8(2): 137–174, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Knupp:1991:DVG

[Knu91] Patrick M. Knupp. The direct variational grid generation method extended to curves. Applied Mathematics and Computation, 43(1 (part I)):65–78, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Knupp:1993:SGG

[Knu93] Patrick M. Knupp. Surface grid generation in the tangent plane. Applied Mathematics and Computation, 59(1):41–62, November 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kocvara:1992:ASL

[Koč92]  Michal Kočvara.  An alge-

braic study of a local multigrid method for variational problems. Applied Mathematics and Computation, 51(1):17– 41, September 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kohlbeck:1992:MNI

[Koh92] F. Kohlbeck. A method for the numerical integration of unequally spaced data. Applied Mathematics and Computation, 50(1):45–49, July 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kolowrocki:1994:LRF

[Koł94] Krzysztof Kołowrocki. Limit reliability functions of some seriesparallel and parallel-series systems. Applied Mathematics and Computation, 62(2–3):129–151, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kolowrocki:1995:ARF

[Koł95] Krzysztof Kołowrocki. Asymptotic reliability functions of some nonhomogeneous seriesparallel and parallel-series systems. Applied Mathematics and Computation, 73(2–3):133– 151, December 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=2-3&aid=9400245.

Konno:1992:FML

[Kon92] Norio Konno. A formalism of multidimensional living systems based on nonstandard analysis. Applied Mathematics and Computation, 49(2–3):231–238, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kopf:1990:IMP

[Kop90] Christoph Kopf. Invariant measures for piecewise linear transformations of the interval. Applied Mathematics and Computation, 39(2 (part II)):123–144, ???? 1990. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Korman:1992:CAS

[Kor92] Philip Korman. A computerassisted study of periodic parabolic systems. Applied Mathematics and Computation, 50(2–3):203–221, August 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Koreas:1997:NCC

[Kor97] Diamantis P. Koreas. The NPcompleteness of chromatic index in triangle free graphs with maximum vertex of degree 3. Applied Mathematics and Computation, 83(1):13–17, April 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=1&aid=9600021.

Kozera:1991:EUP

[Koz91]  Ryszard Kozera.  Existence

and uniqueness in photometric stereo. Applied Mathematics and Computation, 44(1 (part I)): 1–103, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kozera:1995:CIU

[Koz95] Ryszard Kozera. On complete integrals and uniqueness in shape from shading. Applied Mathematics and Computation, 73(1):1–37, November 1, 1995.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        73&issue=1&aid=9400127.

Kamowitz:1985:SSM

[KP85] David Kamowitz and Seymour V. Parter. A study of some multigrid ideas. Applied Mathematics and Computation, 17(2):153–184, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kiviet:1986:TSM

[KP86] Jan F. Kiviet and Garry D. A. Phillips. Testing strategies for model specification. Applied Mathematics and Computation, 20(3–4):237–269, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kravvaritis:1991:ETO

[KP91] Dimitrios Kravvaritis and Nikolaos S. Papageorgiou. An existence theorem for an optimal economic growth problem with infinite planning horizon. Applied Mathematics and Computation, 44(3):249–259, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kreuzer:1996:CPR

[KP96] Edwin Kreuzer and Fernando C. Pinto. Controlling the position of a remotely operated underwater vehicle. Applied Mathematics and Computation, 78(2– 3):175–185, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600007.

Kim:1998:CFE

[KP98] Mi-Young Kim and Eun-Jae Park. Characteristic finite element methods for diffusion epidemic models with agestructured populations. Applied Mathematics and Computation, 97(1):55–70, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/1/6143.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
1&aid=6143.

Karmeshu:1983:DMC

[KPJ83] Karmeshu, R. K. Pathria, and V. P. Jain. A dynamical model of company growth. Applied Mathematics and Computation, 12(1):61–75, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kravvaritis:1993:ESP

[KPP93] Dimitrios Kravvaritis, George Pantelides, and Nikolaos S. Papageorgiou. Existence of support prices for a discrete time growth model with uncertainty. Applied Mathematics and Computation, 53(1):27–43, January 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:1987:ANM

[KPT87] Robert Kalaba, Thomas Plum, and Leigh Tesfatsion. Automation of nested matrix and derivative operations. Applied Mathematics and Computation, 23 (3):243–268, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:1981:AAC

[KR81a] Robert Kalaba and Nima Rasakhoo. Analytical aspects of computational linear algebra. Applied Mathematics and Computation, 8(2):83–109, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:1981:MRS

[KR81b] Robert Kalaba and Nima Rasakhoo. The most representative shape of a family of curves. Applied Mathematics and Computation, 8(3):205–236, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1984:DII

[KR84] Mohan K. Kadalbajoo and K. S. Raman. Discrete invariant imbedding for the numerical solution of boundary value problems over infinite intervals. Applied Mathematics and Computation, 15(4):305–314, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1987:DII

[KR87a] Mohan K. Kadalbajoo and K. Santhana Raman. Discrete invariant imbedding for singular boundary value problems with a regular singularity. Applied Mathematics and Computation, 22(4):291–307, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1987:AMN

[KR87b] Mohan K. Kadalbajoo and Y. N.

Reddy. Approximate method for the numerical solution of singular perturbation problems. Applied Mathematics and Computation, 21(3):185–199, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1987:NSS

[KR87c] Mohan K. Kadalbajoo and Y. N.

Reddy. Numerical solution of singular perturbation problems via deviating arguments. Applied Mathematics and Computation, 21(3):221–232, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1987:NTS

[KR87d] Mohan K. Kadalbajoo and Y. N. Reddy. Numerical treatment of singularly perturbed two-point boundary value problems. Applied Mathematics and Computation, 21(2):93–110, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Krikeles:1988:CPA

[KR88] Basil C. Krikeles and Richard L. Rubin. On the crowding of parameters associated with Schwarz–Christoffel transformations. Applied Mathematics and Computation, 28(4 (part II)): 297–308, December 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1989:ANA

[KR89] Mohan K. Kadalbajoo and Y. N. Reddy. Asymptotic and numerical analysis of singular perturbation problems: a survey. Applied Mathematics and Computation, 30(3):223–259, April 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kuznetsov:1991:NAF

[KR91] Yu. A. Kuznetsov and S. Rinaldi. Numerical analysis of the flip bifurcation of maps. Applied Mathematics and Computation, 43(3):231–236, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kadalbajoo:1995:AGE

[KR95a] Mohan K. Kadalbajoo and A. Appaji Rao. The alternating group explicit (AGE) method for singularly perturbed boundary value problems. Applied Mathematics and Computation, 68(2–3):125–142, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=2-3&aid=9400087.

Kaye:1995:SBF

[KR95b] M. Kaye and D. Reuster. Subdomain basis functions from entire domain basis functions for a certain class of problems. Applied Mathematics and Computation, 72(2–3):167–173, October 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400181.

Kalaba:1983:NLS

[KRT83] Robert Kalaba, Nima Rasakhoo, and Asher Tishler. Nonlinear least squares via automatic derivative evaluation. Applied Mathematics and Computation, 12(2–3):119–137, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:1983:NPD

[KRTW83] Robert Kalaba, Nima Rasakhoo, Asher Tishler, and Jia Song Wang. A new proof for Decell’s finite algorithm for generalized inverses. Applied Mathematics and Computation, 12(2– 3):199–211, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Krueger:1981:DPN

[Kru81] Robert J. Krueger. Determining the profile of a nonmonotonic hill. Applied Mathematics and Computation, 8(1):1–16, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:1978:CCG

[KS78] R. E. Kalaba and K. Spingarn. A criterion for the convergence of the Gauss–Seidel method. Applied Mathematics and Computation, 4(4):359–367, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:1980:SPE

[KS80] R. E. Kalaba and K. Spingarn. Sensitivity of parameter estimates to observations, system identification, and optimal inputs. Applied Mathematics and Computation, 7(3):225–235, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:1981:DLR

[KS81] R. Kalaba and K. Spingarn. Design of linear regulators with energy constraints. Applied Mathematics and Computation, 8(3): 175–185, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:1984:ASO

[KS84] R. Kalaba and K. Spingarn.

Automatic solution of optimalcontrol problems. I. simplest problem in the calculus of variations. Applied Mathematics and Computation, 14(2):131– 148, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khamayseh:1992:EBC

[KS92a] Ahmed K. Khamayseh and R. Shivaji. Evolution of bifurcation curves for semipositone problems when nonlinearities develop multiple zeros. Applied Mathematics and Computation, 52(2–3):173–188, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kriuchkov:1992:SSO

[KS92b] V. N. Kriuchkov and M. Solomonovich.

Self-sustained oscillation in economics. Applied Mathematics and Computation, 49(1):63–??, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karageorghis:1993:CSC

[KS93a] A. Karageorghis and S. Sivaloganathan. A Chebyshev spectral collocation method for the computation of buoyancy-induced flows. Applied Mathematics and Computation, 55(2–3):297–308, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Katti:1993:SCS

[KS93b] C. P. Katti and S. Sivaloganathan. On the stability of a classical second order method for solving a class of two-point boundary value problems. Applied Mathematics and Computation, 59(2–3):247–255, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kahaner:1989:PIS

[KSB89] David K. Kahaner, Selden L. Stewart, and David Barnett. PLOD: interactive solver for ordinary differential equations. Applied Mathematics and Computation, 31(Special issue):302– 315, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Kalaba:1980:NDE

[KST80] R. Kalaba, K. Spingarn, and L. Tesfatsion. A new differential equation method for finding the Perron root of a positive matrix. Applied Mathematics and Computation, 7(3):187–193, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kuhn:1995:CVM

[KSZ^{+}95] A. Kuhn, W. Steiner, J. Zemann, D. Dinevski, and H. Troger. A comparison of various mathematical formulations and numerical solution methods for the large amplitude oscillations of a string pendulum. Applied Mathematics and Computation, 67 (1–3):227–264, January 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
67&issue=1-3&aid=9400060.

Kalaba:1983:CPM

[KT83] R. Kalaba and A. Tishler. A computer program to minimize a function with many variables using computer evaluated exact higher-order derivatives. Applied Mathematics and Computation, 13(1–2):143–172, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khaliq:1989:GEN

[KT89] A. Q. M. Khaliq and E. H. Twizell. Global extrapolation of numerical methods for initial value problems. Applied Mathematics and Computation, 31(Special issue):148–160, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Kalaba:1991:SNE

[KT91] R. Kalaba and Leigh Tesfatsion. Solving nonlinear equations by adaptive homotopy continuation. Applied Mathematics and Computation, 41(2 (part II)):99–115, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalaba:1981:LNC

[KTW81] Robert Kalaba, Leigh Tesfatsion, and Jone-Lin L. Wang. Local and nonlocal comparative static analysis of economic systems. Applied Mathematics and Computation, 9(??):227– 234, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kamont:1996:DDI

[KTZK96] Z. Kamont, J. Turo, and B. Zubik-Kowal. Differential and difference inequalities generated by mixed problems for hyperbolic functional differential equations with impulses. Applied Mathematics and Computation, 80(2–3):127– 154, December 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/

tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500284.

Kalaba:1992:CM

[KU92] R. Kalaba and F. Udwadia. On constrained motion. Applied Mathematics and Computation, 51(1):85–86, September 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kubica:1994:CAS

[Kub94a] K. Kubica. Cellular automata studies of cholesterol effects on gramicidin ion conductivity. Applied Mathematics and Computation, 64(2–3):101–??, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kubica:1994:CAM

[Kub94b] Krystian Kubica. The cellular automata model for lipid membranes. Applied Mathematics and Computation, 63(2–3): 187–199, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kubica:1997:EAC

[Kub97] Krystian Kubica. The effect of amphiphilic counterions on the gel-fluid phase transition of the lipid bilayer. Applied Mathematics and Computation, 87(2–3): 261–270, December 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=2-3&aid=9600301.

Kumar:1992:OER

[Kum92a] Kuldeep Kumar. Optimum exponential regression with one nonlinear term. Applied Mathematics and Computation, 50 (1):51–??, July 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:1992:PAA

[Kum92b] Kuldeep Kumar. Padé approximation and its application in time series analysis. Applied Mathematics and Computation, 48(2–3):139–151, April 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kumar:1995:ETD

[Kum95] Ravi Kumar. Effects of time delays on the stability of collocated and noncollocated point control of discrete dynamic structural systems. Applied Mathematics and Computation, 70(2– 3):215–232, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400107.

Dynamics and control (Vienna, 1993).

Kawata:1995:SAR

[KUO95] Y. Kawata, S. Ueno, and A. Ohtani. The surface albedo retrieval of mountainous forest area from satellite MSS data. Applied Mathematics and Computation, 69(1):41–59, April 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=1&aid=9400098.

Kurowicka:1999:DAA

[Kur99] Dorota Kurowicka. Domains of attraction of asymptotic reliability functions of some homogeneous series–parallel systems. Applied Mathematics and Computation, 98(1):61–74, January 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/1/6128.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
1&aid=6128.

Kalaba:1995:CMR

[KUX95] Robert Kalaba, Firdaus Udwadia, and Rong Xu. Constrained motion revisited. Applied Mathematics and Computation, 70 (1):67–76, June 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=1&aid=9400143.

Kuo:1984:NSN

[KV84] Pen Yu Kuo and Luis Vázquez. Numerical solution of a nonlinear wave equation in polar coordinates. Applied Mathematics and Computation, 14(4):313– 329, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kalla:1998:SFO

[KVT98] Shyam L. Kalla, N. Virchenko, and V. Tsarenko. On some fractional order integral transforms generated by orthogonal polynomials. Applied Mathematics and Computation, 91 (2–3):209–219, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/91/2-3/
                                        6019.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=91&issue=
                                        2-3&aid=6019.

Kettler:1986:AMM

[KW86] R. Kettler and P. Wesseling. Aspects of multigrid methods for problems in three dimensions. Applied Mathematics and Computation, 19(1–4):159– 168, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Khuri:1996:SPD

[KW96] S. A. Khuri and A. M. Wazwaz.

The solution of a partial differential equation arising in fluid flow theory. Applied Mathematics and Computation, 77(2– 3):295–300, July 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Khuri:1997:SPD

[KW97] S. A. Khuri and A. M. Wazwaz.

On the solution of a partial differential equation arising in Stokes flow. Applied Mathematics and Computation, 85(2– 3):139–147, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600131.

Kwak:1997:PGM

[Kwa97] Do Y. Kwak. A preconditioned GMRES method. Applied Mathematics and Computation, 85(2– 3):201–208, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600135.

Kwong:1993:SAI

[Kwo93] Man Kam Kwong. Sweeping algorithms for inverting the discrete Ginzburg–Landau operator. Applied Mathematics and Computation, 53(2–3):129– 150, February 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kang:1995:FSP

[KWS95] Long-Chyuan Kang, Cheng-Hsiu Wu, and Charles R. Steele. Fourier series for polygonal plate bending: A very large plate element. Applied Mathematics and Computation, 67(1–3):197– 225, January 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
67&issue=1-3&aid=9400059.

Lakshmivarahan:1981:LAT

[Lak81] S. Lakshmivarahan. A learning approach to the two-person decentralized team problem with incomplete information. Applied Mathematics and Computation, 8(1):51–78, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lainiotis:1994:NDA

[LAK94] D. G. Lainiotis, N. D. Assimakis, and S. K. Katsikas. New doubling algorithm for the discrete periodic Riccati equation. Applied Mathematics and Computation, 60(2–3):265–283, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lamberti:1986:SHH

[Lam86] Lamberto Lamberti. Solutions to the Hodgkin–Huxley equations. Applied Mathematics and Computation, 18(1):43–70, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lam:1998:LDC

[Lam98] S. H. Lam. On Lagrangian dynamics and its control formulations. Applied Mathematics and Computation, 91 (2–3):259–284, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6004.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6004.

Lameche:1999:ISA

[Lam99] Khira Lameche. Inverse for the shuffle for algebraic series. Applied Mathematics and Computation, 98(1):1–27, January 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/1/6090.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
1&aid=6090.

Langerholc:1981:CMP

[Lan81] J. Langerholc. Calculation of moving phase boundaries with latent heat. Applied Mathematics and Computation, 9(1): 43–51, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Langerholc:1989:VDH

[Lan89]  John Langerholc.  Volumes

of diced hyperspheres: resumming the Tam-Zardecki formula. Applied Mathematics and Computation, 30(1 (part I)): 1–18, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Langerholc:1992:CFS

[Lan92] John Langerholc. Calculation of forward-scattering amplitudes from a focused beam: Exact values for the Gaussian parameters. Applied Mathematics and Computation, 48(2–3):101–137, April 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lapsa:1992:RLM

[Lap92] P. M. Lapsa. Resolution-limited measure and dimension. Applied Mathematics and Computation, 52(1):1–27, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Larsen:1997:SCS

[Lar97] Jens Chr. Larsen. Stable chemical systems. Applied Mathematics and Computation, 81(2–3):

245–257, February 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=2-3&aid=9600020.

Luo:1992:NNA

[LB92] Fa Long Luo and Zheng Bao.

Neural network approach to computing matrix inversion. Applied Mathematics and Computation, 47(2–3):109–120, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lybeck:1996:SMD

[LB96] Nancy J. Lybeck and Kenneth L. Bowers. Sinc methods for domain decomposition. Applied Mathematics and Computation, 75(1):13–41, March 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=1&aid=9500099.

Li:1997:PIF

[LC97] Likang Li and Jinru Chen. Preconditioning isoparametric finite element methods taking into account numerical integration. Applied Mathematics and Computation, 87(2–3):271– 288, December 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        87&issue=2-3&aid=9600302.

Lether:1985:EAC

[LCE85] Frank G. Lether, David M. Cline, and Otis Evans. An error analysis for the calculation of semi-integrals and semiderivatives by the RL algorithm. Applied Mathematics and Computation, 17(1):45–67, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lether:1985:ACR

[LE85] Frank G. Lether and Otis M. Evans. An algorithm for the computation of the reversible Randles-Sevcik function in electrochemistry. Applied Mathematics and Computation, 16(3): 253–264, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Le:1998:SGB

[Le98] Vy Khoi Le. Some global bifurcation results for elastic plates. Applied Mathematics and Computation, 89(1–3):185– 197, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        89&issue=1-3&aid=9781657.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Lee:1998:CMF

[Lee98] Chang-Ock Lee. A conforming mixed finite element method for the pure traction problem of linear elasticity. Applied Mathematics and Computation, 93(1):11–29, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/1/6062.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
1&aid=6062.

Leitmann:1995:OAC

[Lei95] G. Leitmann. One approach to the control of uncertain dynamical systems. Applied Mathematics and Computation, 70(2– 3):261–272, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400118.

Dynamics and control (Vienna, 1993).

Lessard:1989:MRA

[Les89] Sabin Lessard. The mathematics of resource allocation in population biology. I. interspecific competition. Applied Mathematics and Computation, 32(2– 3):207–226, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Mathematical biology.

Lether:1980:EEG

[Let80] Frank G. Lether. Error estimates for Gaussian quadrature. Applied Mathematics and Computation, 7(3):237–246, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lether:1997:CNM

[Let97] Frank G. Lether. Constrained near-minimax rational approximations to Dawson’s integral. Applied Mathematics and Computation, 88(2–3):267– 274, December 30, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        88&issue=2-3&aid=9600330.

Lu:1994:DNS

[LF94] Xin Lu and Wei Feng. Dynamics and numerical simulations of food-chain populations. Applied Mathematics and Computation, 65(1–3):335–344, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Liu:1999:HON

[LF99] Xinzhi Liu and Xilin Fu. High order nonlinear differential inequalities with distributed deviating arguments and applications. Applied Mathematics and Computation, 98(2– 3):147–167, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/
6144.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6144.

Lee:1995:PEN

[LFS95] F. C. Lee, H. Flashner, and M. G. Safonov. Positivity embedding for noncolocated and nonsquare flexible systems. Applied Mathematics and Computation, 70(2–3):233–246, July 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400108.

Dynamics and control (Vienna, 1993).

Lalli:1994:OTS

[LG94] B. S. Lalli and S. R. Grace.

Oscillation theorems for second order neutral difference equations. Applied Mathematics and Computation, 62(1):47–60, April 1994. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-

5649 (electronic).

Liu:1997:MCP

[LG97] Pingzhou Liu and K. Gopalsamy. On a model of competition in periodic environments. Applied Mathematics and Computation, 82(2–3):207– 238, March 15, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        82&issue=2-3&aid=9600044.

Liu:1999:GSC

[LG99] Pingzhou Liu and K. Gopalsamy. Global stability and chaos in a population model with piecewise constant arguments. Applied Mathematics and Computation, 101(1):63–88, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/101/1/
                                        6221.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=101&
                                        issue=1&aid=6221.

Liping:1997:HPA

[LHK97] Fang Liping, Keith W. Hipel, and D. Marc Kilgour. How penalty affects enforcement of environmental regulations. Applied Mathematics and Computation, 83(2–3):281–301, May 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=2-3&aid=9600189.

Li:1994:NMS

[Li94] Xuefeng Li. A numerical method for systems of hyperbolic conservation laws with single stencil reconstructions. Applied Mathematics and Computation, 65(1– 3):125–140, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Li:1998:SAR

[Li98] Xin Li. On simultaneous approximations by radial basis function neural networks. Applied Mathematics and Computation, 95(1):75–89, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/1/6088.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
1&aid=6088.

Liao:1997:SEA

[Lia97] Aiping Liao. Some efficient algorithms for unconstrained discrete-time optimal control problems. Applied Mathematics and Computation, 87(2–3):

175–198, December 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        87&issue=2-3&aid=9600228.

Liaw:1998:ACM

[Lia98] Der-Cherng Liaw. Application of center manifold reduction to nonlinear system stabilization. Applied Mathematics and Computation, 91(2–3):243–258, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/91/2-3/
                                        6020.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=91&issue=
                                        2-3&aid=6020.

Lin:1994:FDS

[Lin94] Yan Ping Lin. Finite difference solutions for parabolic equations with the time weighting initial conditions. Applied Mathematics and Computation, 65(1–3):49–61, September/October 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Lin:1996:ECO

[Lin96] Cheng Chang Lin. An existence condition for open set nonlinear complementarity problems with an application to a closed economic system. Applied Mathematics and Computation, 76(2– 3):285–292, May 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=2-3&aid=9500185.

Lin:1998:FEM

[Lin98] Tao Lin. A C^{0} finite element method for an inverse problem. Applied Mathematics and Computation, 90(2– 3):253–284, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/
9700401.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=2-3&aid=9700401.

Liu:1994:BPS

[Liu94]  James H. Liu.  Bounded

and periodic solutions of differential equations in Banach space. Applied Mathematics and Computation, 65(1–3):141–150, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Liu:1995:ICO

[Liu95] Xinzhi Liu. Impulsive control and optimization. Applied Mathematics and Computation, 73 (1):77–98, November 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        73&issue=1&aid=9400204.

Lubarda:1995:CSR

[LK95] V. A. Lubarda and D. Krajcinovic. Constitutive structure of rate theory of damage in brittle elastic solids. Applied Mathematics and Computation, 67(1– 3):81–101, January 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        67&issue=1-3&aid=9400055.

Lakshmikantham:1989:SID

[LL89] V. Lakshmikantham and Xinzhi Liu. Stability for impulsive differential systems in terms of two measures. Applied Mathematics and Computation, 29(1 (part I)):89–98, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:1994:QSH

[LL94a] Xiaodong Lin and Xinzhi Liu. Qualitative study of an HIV transmission model among intravenous drug users. Applied Mathematics and Computation, 60(2–3):215–236, February 1994.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Luo:1994:NNE

[LL94b] Fa-Long Luo and Yan-Da Li.

Neural network for the exact RLS adaptive algorithm. Applied Mathematics and Computation, 60(2–3):103–111, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Layton:1995:TLP

[LL95] W. Layton and W. Lenferink.

Two-level Picard and modified Picard methods for the Navier– Stokes equations. Applied Mathematics and Computation, 69(2– 3):263–274, May 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=2-3&aid=9400134.

Liao:1996:RIS

[LL96] Xiao-Xin Liao and Jia Li. Robust interval stability, persistence, and partial stability on Lotka–Volterra systems with time-delay. Applied Mathematics and Computation, 75(2–3): 103–115, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500103.

Lin:1997:MNC

[LLL97] R. Lin, G. Leng, and H. P. Lee. A method for the numerical computation of Hopf bifurcation. Applied Mathematics and Computation, 86(2–3): 137–156, October 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        86&issue=2-3&aid=9600178.

Lu:1994:FPS

[LLS94] Xiguan Lu, Yong Li, and Yi Su. Finding periodic solutions of ordinary differential equations via homotopy method. Applied Mathematics and Computation, 62(2–3):153–170, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:1996:FPS

[LLS96] Xiguan Lu, Yong Li, and Yi Su. Finding periodic solutions of ordinary differential equations via homotopy method. Applied Mathematics and Computation, 78(1):1–17, August 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        78&issue=1&aid=9400241.

Luo:1996:RTC

[LLX96] Fa-Long Luo, Yan-Da Li, and Hong-Qinq Xu. Real-time computation of the eigenvectors of a class of positive definite matrices. Applied Mathematics and Computation, 77(2–3):267–280, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:1996:CHI

[LLY96] Zhenghua Lin, Yong Li, and Bo Yu. A combined homotopy interior point method for general nonlinear programming problems. Applied Mathematics and Computation, 80(2–3):209– 224, December 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500295.

Lazer:1994:SEB

[LM94] A. C. Lazer and P. J. McKenna.

A singular elliptic boundary value problem. Applied Mathematics and Computation, 65(1– 3):183–194, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Lee:1995:VIE

[LM95] Jungki Lee and Ajit K. Mal.

A volume integral equation technique for multiple scattering problems in elastodynamics. Applied Mathematics and Computation, 67(1–3):135– 159, January 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        67&issue=1-3&aid=9400057.

Lopez:1998:NMI

[LM98] David J. López and Pablo Martín. A numerical method for the integration of perturbed linear problems. Applied Mathematics and Computation, 96(1):65–73, October 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/96/1/6116.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=96&issue=
                                        1&aid=6116.

Lubeck:1986:PET

[LMM86] O. Lubeck, J. Moore, and R. Mendez. A performance evaluation of three supercomputers, Fujitsu XP-200, Hitachi S810/20, CRAY X-MP/24. Applied Mathematics and Computation, 20(??):143–144, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:1992:IDS

[LMS92] V. Lakshmikantham, K. N. Murty, and S. Sivasundaram. Impulsive differential systems for two-point boundary value problems. Applied Mathematics and Computation, 50(2–3): 157–166, August 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:1997:SLP

[LMSS97] V. Lakshmikantham, A. K.

Maulloo, S. K. Sen, and S. Sivasundaram. Solving linear programming problems exactly. Applied Mathematics and Computation, 81(1):69–87, January 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=1&aid=9500309.

Lepschy:1991:HTI

[LMV91] A. Lepschy, G. A. Mian, and U. Viaro. Hermite-type interpolation of squared-magnitude functions. Applied Mathematics and Computation, 43(3):265– 270, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:1999:RPS

[LO99] K. H. Lee and E. H. Ong. A reduction principle for singular perturbation problems. Applied Mathematics and Computation, 101(1):45–62, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/

                                        store/amc/sub/1999/101/1/
                                        6226.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=101&
                                        issue=1&aid=6226.

Longman:1987:CMS

[Lon87a] I. M. Longman. Complementary moments and the summation of series. Applied Mathematics and Computation, 24 (2):139–151, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Longman:1987:ICR

[Lon87b] I. M. Longman. Increasing the convergence rate of series. Applied Mathematics and Computation, 24(1):77–89, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Longman:1987:SFC

[Lon87c] I. M. Longman. The summation of Fourier, Chebyshev, and Legendre series. Applied Mathematics and Computation, 23 (1):61–70, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Longman:1988:SAM

[Lon88] I. M. Longman. The summation of alternating and monotonic series. Applied Mathematics and Computation, 25(4): 287–298, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lopez:1992:STP

[Lop92] L. Lopez. Stability of a threepoint scheme for linear second order singularly perturbed BVPs with turning points. Applied Mathematics and Computation, 52(2–3):279–300, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lord:1980:VIE

[Lor80] M. E. Lord. Validation of an invariant embedding method for Fredholm integral equations. Applied Mathematics and Computation, 7(1):1–7, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:1981:MMN

[LP81] V. Lakshmikantham and B. G.

Pachpatte. Monotone method for nonlinear boundary value problems arising in transport processes. Applied Mathematics and Computation, 9(1):35–41, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:1989:ASI

[LP89] Xinzhi Liu and Ratnam Pirapakaran. Asymptotic stability for impulsive differential systems. Applied Mathematics and Computation, 30(2 (part II)): 149–163, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lu:1997:MSF

[LP97] Linzhang Lu and C. E. M. Pearce. On the matrixsign-function method for solving algebraic Riccati equations. Applied Mathematics and Computation, 86(2–3):157– 170, October 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
86&issue=2-3&aid=9600179.

Ladde:1985:DSM

[LR85] G. S. Ladde and S. G. Rajalakshmi. Diagonalization and stability of multi-time-scale singularly perturbed linear systems. Applied Mathematics and Computation, 16(2):115–140, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Leitmann:1995:CSB

[LR95] G. Leitmann and E. Reithmeier. A control scheme based on ER-Materials for vibration attenuation of dynamical systems. Applied Mathematics and Computation, 70(2–3):247–259, July 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400109.

Dynamics and control (Vienna, 1993).

Lakshmikantham:1984:MMS

[LRV84] V. Lakshmikantham, V. Sree Hari Rao, and A. S. Vatsala. Monotone method for a system of second order periodic boundary value problems. Applied Mathematics and Computation, 15 (1):71–83, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:1978:ZMO

[LS78] V. Lakshmikantham and Patrick Sutherland. On the zeros of monotone operators of retarded type in a Banach space. Applied Mathematics and Computation, 4(??):81–91, 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Levin:1981:TNC

[LS81] David Levin and Avram Sidi.

Two new classes of nonlinear transformations for accelerating the convergence of infinite integrals and series. Applied Mathematics and Computation, 9(3): 175–215, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:1987:MUL

[LS87] V. Lakshmikantham and S. Sivasundaram. The method of upper and lower solutions and interval method for first order differential equations. Applied Mathematics and Computation, 23(1):1–5, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:1991:AIA

[LS91a] V. Lakshmikantham and Yong Sun. Applications of interval analysis to minimal and maximal solutions of differential equations. Applied Mathematics and Computation, 41(1 (part I)):77–87, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:1991:GAS

[LS91b] Dan Liu and Ferenc Szidarovszky. Global asymptotical stability of dynamic systems with modified contractions. Applied Mathematics and Computation, 43(3):237–240, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ladde:1993:NTR

[LS93a] G. S. Ladde and M. Sambandham. Numerical treatment of random polynomials. Applied Mathematics and Computation, 55(1):13–30, April 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Liu:1993:NFM

[LS93b] K. Liu and R. Skelton. A new formulation of Q-Markov covariance equivalent realization. Applied Mathematics and Computation, 53(1):83–95, January 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ladde:1995:SAH

[LS95] G. S. Ladde and S. Sathananthan. Stability analysis of hereditary iterative processes. Applied Mathematics and Computation, 73(1):39–54, November 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=1&aid=9400202.

Lin:1997:TK

[LS97] Li Jian Lin and H. M. Srivastava. On a theorem of Kobori. Applied Mathematics and Computation, 85(2–3):287– 296, September 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600143.

Liu:1998:GCT

[LS98a] Jinglian J. Liu and Bharat K.

Soni. 2D groundwater contaminant transport modeling by using the finite volume method on an unstructured grid system. Applied Mathematics and Computation, 89(1–3):199– 211, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781658.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Lu:1998:NSR

[LS98b] Xiaowu Lu and Rudolf Schmid. A numerical study of the Riemann solutions for gasdynamic combustion. Applied Mathematics and Computation, 91 (2–3):143–160, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6015.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6015.

Li:1999:ERS

[LS99] Weiye Li and Ferenc Szidarovszky. An elementary result in the stability theory of time-invariant nonlinear discrete dynamical systems. Applied Mathematics and Computation, 102(1):35–49, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/1/
6239.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/

amc/cas_sub/browse/browse.
cgi?year=1999&volume=102&
issue=1&aid=6239.

Lehtokangas:1996:NAP

[LSKH96] Mikko Lehtokangas, Jukka Saarinen, Kimmo Kaski, and Pentti Huuhtanen. A network of autoregressive processing units for time series modeling. Applied Mathematics and Computation, 75(2–3):151– 165, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500122.

Lenhart:1991:VAC

[LSP91] S. Lenhart, S. Stojanovic, and V. Protopopescu. Variational approach for competitive systems with obstacles. Applied Mathematics and Computation, 42(1 (part I)):65–83, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:1994:CRP

[LSS94] V. Lakshmikantham, S. K. Sen, and S. Sivasundaram. Concise row-pruning algorithm to invert a matrix. Applied Mathematics and Computation, 61(1): 17–24, March 01, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lakshmikantham:1994:STA

[LSVS94] V. Lakshmikantham, S. K. Sen, V. C. Venkaiah, and S. Sivasundaram. Stabilizing Trench’s algorithm to invert symmetric Toeplitz matrices. Applied Mathematics and Computation, 60(2–3):249–263, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:1998:DNT

[LSY98] C. Lin, F. Szidarovszky, and J. Yen. Dynamic negotiation with time-varying Pareto frontier. Applied Mathematics and Computation, 91 (2–3):99–109, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6001.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6001.

Lu:1998:CIM

[Lu98] Xin Lu. Combined iterative methods for numerical solutions of parabolic problems with time delays. Applied Mathematics and Computation, 89(1–3):213– 224, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=

89&issue=1-3&aid=9781659.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Lui:1999:MBS

[Lui99] S. H. Lui. Multiple bifurcation from“simple” eigenvalues. Applied Mathematics and Computation, 100(2–3):111–130, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6200.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6200.

Luo:1997:RTC

[LUL97] Fa-Long Luo, Rolf Unbehauen, and Yan-Da Li. Real-time computation of singular vectors. Applied Mathematics and Computation, 86(2–3):197– 214, October 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
86&issue=2-3&aid=9600182.

Lakshmikantham:1987:MMM

[LV87] V. Lakshmikantham and A. S.

Vatsala. Method of mixed monotony for nonlinear equations with a singular linear part. Applied Mathematics and Computation, 23(3):235–241, ???? 1987. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Lether:1990:ANC

[LW90] F. G. Lether and P. R. Wenston. An algorithm for the numerical computation of the Voigt function. Applied Mathematics and Computation, 35(3): 277–289, February 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:1994:HOT

[LW94] T. Y. Li and Xiao Shen Wang. Higher order turning points. Applied Mathematics and Computation, 64(2–3):155–166, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Li:1995:PSF

[LWLL95] Yong Li, Huaizhong Wang, Xianrui Lü, and Xiguan Lu. Periodic solutions for functional differential equations with infinite lead and delay. Applied Mathematics and Computation, 70 (1):1–28, June 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        70&issue=1&aid=9400131.

Liu:1992:MIN

[LWX92] Xinzhi Liu, Yau Shu Wong, and Ji Xingzhi. Monotone iterations for numerical solutions of nonlinear elliptic partial differential equations. Applied Mathematics and Computation, 50(1):59–91, July 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Layton:1998:NTL

[LY98] W. Layton and X. Ye. Nonconforming two-level discretization of stream function form of the Navier–Stokes equations. Applied Mathematics and Computation, 89(1–3):173– 183, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781656.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Lalli:1993:OHE

[LYC93] B. S. Lalli, Y. H. Yu, and B. T. Cui. Oscillation of hyperbolic equations with functional arguments. Applied Mathematics and Computation, 53(2–3): 97–110, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lin:1997:CHI

[LYF97] Zhenghua Lin, Bo Yu, and Guochen Feng. A combined homotopy interior point method for convex nonlinear programming. Applied Mathematics and Computation, 84(2–3):

193–211, July 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600086.

Li:1993:LSF

[LYX^{+}93] J. S. Li, Z. Y. Yu, X. Q. Xiang, W. P. Ni, and C. L. Chang. Least-square finite element method for electromagnetic fields in 2-D. Applied Mathematics and Computation, 58(2–3):143–167, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Logan:1996:TPT

[LZ96] J. David Logan and Vitaly Zlotnik. Time-periodic transport in heterogeneous porous media. Applied Mathematics and Computation, 75(2–3):119– 138, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500120.

Marchuk:1995:ROC

[MA95] G. I. Marchuk and V. I. Agoshkov. Reflection operators and contemporary applications to radiative transfer. Applied Mathematics and Computation, 69(1):3–21, April 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=1&aid=9400149.

MacLeod:1993:NRS

[Mac93] Allan J. MacLeod. An note on the Randles-Sevcik function from electrochemistry. Applied Mathematics and Computation, 57(2–3):305–310, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Madan:1990:INM

[Mad90] Ved P. Madan. Interval Newton method: Hansen-Greenberg approach — some procedural improvements. Applied Mathematics and Computation, 35(3): 263–276, February 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maeder:1989:GPO

[Mae89] A. J. Maeder. A general purpose ODE solver implementation. Applied Mathematics and Computation, 31(Special issue): 316–327, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Malakooti:1991:GDM

[Mal91] Behnam B. Malakooti. A generalized decomposable multiattribute utility function. Applied Mathematics and Computation, 44(3):271–285, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Malakooti:1993:DSS

[Mal93] Behnam B. Malakooti. A decision support system for discrete multiple-criteria problems: Certainty, uncertainty, and hierarchical. Applied Mathematics and Computation, 54(2–3): 131–166, March 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mandel:1986:MCN

[Man86] Jan Mandel. Multigrid convergence for nonsymmetric, indefinite variational problems and one smoothing step. Applied Mathematics and Computation, 19(1–4):201–216, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Mandel:1988:ASM

[Man88] Jan Mandel. Algebraic study of multigrid methods for symmetric, definite problems. Applied Mathematics and Computation, 25(1 (part I)):39–56, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Marz:1989:BVP

[Mär89] Roswitha März. On boundary value problems in differentialalgebraic equations. Applied Mathematics and Computation, 31(Special issue):517–537, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Mastin:1982:EIC

[Mas82] C. Wayne Mastin. Error induced by coordinate systems. Applied Mathematics and Computation, 10/11(1):31–40, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Martin:1995:UMS

[MAS95] Clyde F. Martin, Linda J. S.

Allen, and Mark Stamp. Urn model simulations of a sexually transmitted disease epidemic. Applied Mathematics and Computation, 71(2–3):179– 199, September 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400153.

Matarazzo:1986:MAP

[Mat86] Benedetto Matarazzo. Multicriterion analysis of preferences by means of pairwise actions and criterion comparisons (MAPPACC). Applied Mathematics and Computation, 18 (2):119–141, ???? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mattheij:1989:CCS

[Mat89] R. M. M. Mattheij. Conditions and conditioning, stability and stabilization. Applied Mathematics and Computation, 31(Special issue):538–554, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Matsuno:1993:BFC

[Mat93] Koichiro Matsuno. Being free from Ceteris Paribus: A vehicle for founding physics on biology rather than the other way around. Applied Mathematics and Computation, 56(2– 3):261–??, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maroulas:1980:CFE

[MB80] John Maroulas and Stephen Barnett. Continued-fraction expansions for ratios of generalized polynomials. Applied Mathematics and Computation, 6(3): 229–249, April 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Manhardt:1982:ATD

[MB82] P. D. Manhardt and A. J. Baker.

Automated three-dimensional grid refinement on a minicomputer. Applied Mathematics and Computation, 10/11(1):437–446, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Milusheva:1988:PIM

[MB88a] S. D. Milusheva and D. D.

Baĭnov. A projection-iterative method for solving the periodic problem for integrodifferential equations with impulse effect. Applied Mathematics and Computation, 28(4 (part II)): 337–352, December 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mishev:1988:OSP

[MB88b] D. P. Mishev and D. D. Baĭnov.

Oscillation of the solutions of parabolic differential equations of neutral type. Applied Mathematics and Computation, 28(2): 97–111, November 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moss:1995:DIV

[MB95] Darrell G. Moss and Haym Benaroya. A discrete inverse vibration problem with parameter uncertainty. Applied Mathematics and Computation, 69(2– 3):313–333, May 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=2-3&aid=9400140.

Masri:1995:ASO

[MBCV95] S. F. Masri, G. A. Bekey, T. K. Caughey, and E. Van de Velde. Adaptive stochastic optimization using multiprocessors. Applied Mathematics and Computation, 72(2–3):225– 257, October 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400186.

Matyasovszky:1995:IGC

[MBG95] Istvan Matyasovszky, Istvan Bogardi, and Jacques Ganoulis. Impact of global climate change on temperature and precipitation in greece. Applied Mathematics and Computation, 71(2– 3):119–150, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400148.

Masri:1980:GOA

[MBS80] S. F. Masri, G. A. Bekey, and F. B. Safford. A global optimization algorithm using adaptive random search. Applied Mathematics and Computation, 7(4):353–375, ???? 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mueller:1998:CMM

[MBS98] K. M. Mueller, S. A. Burns, and M. A. Savageau. A comparison of the monomial method and the S-system method for solving systems of algebraic equations. Applied Mathematics and Computation, 90(2– 3):167–180, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/
9700397.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=2-3&aid=9700397.

Mitwasi:1989:AMD

[MCD89] M. Mitwasi, A.-M. Connolly, and L. Duckstein. Applying multicriteria decision making techniques for planning machining operations. Applied Mathematics and Computation, 29(3):197– 218, February 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

McGrath:1979:GPT

[McG79] Joseph F. McGrath. Gaussian product-type quadratures. Applied Mathematics and Computation, 5(3):265–280, ???? 1979.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

McGough:1998:NCG

[McG98] Jeff S. McGough. Numerical continuation and the Gelfand problem. Applied Mathematics and Computation, 89(1–3): 225–239, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        89&issue=1-3&aid=9781660.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

McWhorter:1982:SMA

[McW82] John C. McWhorter. Solid mechanics applications of boundary fitted coordinate systems. Applied Mathematics and Computation, 10/11(1):253–276, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Maroulas:1981:AGS

[MD81] J. Maroulas and D. Dascalopoulos. Applications of the generalized Sylvester matrix. Applied Mathematics and Computation, 8(2):121–135, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Malik:1998:VAC

[MD98] Moinuddin Malik and Hao Huy Dang. Vibration analysis of continuous systems by differential transformation. Applied Mathematics and Computation, 96(1):17–26, October 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/1/6083.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
1&aid=6083.

Malakooti:1994:SDA

[MDR94] B. Malakooti, Lucien Duckstein, and A. Ravindran. Screening discrete alternatives with imprecisely assessed additive multiattribute functions. Applied Mathematics and Computation, 60(1):69–87, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Meinhardt:1989:TCR

[Mei89] Hans Meinhardt. Tailoring and coupling of reaction-diffusion systems to obtain reproducible complex pattern formation during development of the higher organisms. Applied Mathematics and Computation, 32(2–3): 103–135, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Mathematical biology.

Meir:1993:FEA

[Mei93] A. J. Meir. Finite element analysis of magnetohydrodynamic pipe flow. Applied Mathematics and Computation, 57(2–3):177– 196, September 01, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Meir:1994:TCM

[Mei94] A. J. Meir. Thermally coupled magnetohydrodynamics flow. Applied Mathematics and Computation, 65(1–3):79–94, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Melson:1986:VMA

[Mel86] N. Duane Melson. Vectorizable multigrid algorithms for transonic-flow calculations. Applied Mathematics and Computation, 19(1–4):217–238, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Merriam:1986:ADF

[Mer86] Marshal L. Merriam. Application of data flow concepts to a multigrid solver for the Euler equations. Applied Mathematics and Computation, 19(??):239– 263, 1986. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Mostafa:1992:MGS

[MES92] Magdi S. Mostafa and Khaled M. F. El-Sayed. Matrixgeometric solution of a multiserver queue with Markovian group arrivals and Coxian servers. Applied Mathematics and Computation, 49(2–3): 177–196, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moustafa:1994:ISQ

[MES94] Magdi S. Moustafa and Khaled M. F. El-Sayed. Iterative solution of the M^{[x]}/C_{2}/S/N queues. Applied Mathematics and Computation, 66(2–3):313– 323, December 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Meyer:1978:NSS

[Mey78] Gunter H. Meyer. The numerical solution of Stefan problems with front-tracking and smoothing methods. Applied Mathematics and Computation, 4(4): 283–306, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mohsen:1993:HAF

[MEZI93] A. Mohsen, H. El-Zoheiry, and L. Iskandar. A highly accurate finite-difference scheme for a Boussinesq-Type equation. Applied Mathematics and Computation, 55(2–3):201–212, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martin:1999:INI

[MF99] Pablo Martín and José-Miguel Farto. Improved numerical integration of perturbed oscillators via average. Applied Mathematics and Computation, 99 (2–3):129–139, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6182.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6182.

Murty:1992:STP

[MFF92] K. N. Murty, D. W. Fausett, and L. V. Fausett. Solution of twopoint boundary value problems involving Kronecker products. Applied Mathematics and Computation, 50(2–3):145–156, August 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moskowitz:1997:PDB

[MFGW97] Simon Moskowitz, Emmanuel Fernández-Gaucherand, and Michael Whalen. Passagedetector-based traffic queue estimation in intelligent transportation systems: A computational study of competing algorithms. Applied Mathematics and Computation, 86(2–3):

93–113, October 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
86&issue=2-3&aid=9600175.

McCord:1993:MAA

[MFS93] Mark R. McCord, Oscar Franzese, and Xiao Duan Sun. Multicriteria analysis of aeromedical fleet expansion. Applied Mathematics and Computation, 54(2–3): 101–129, March 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martinez:1992:AMI

[MGL92] Ignacio Martínez, María Angeles Gil, and María Teresa López. Analysis of mutual information measures in cluster sampling. Applied Mathematics and Computation, 52(2–3): 389–402, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Murthy:1980:NME

[MH80] Vrudhula K. Murthy and L. Julian Haywood. New method for estimating survival curves based upon subpopulations. Applied Mathematics and Computation, 7(3):259–266, ???? 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Murthy:1981:CSC

[MH81] Vrudhula K. Murthy and L. Julian Haywood. Comparative survival curve analysis in sickle cell disease. Applied Mathematics and Computation, 9(??):143– 152, 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mikhail:1987:VSM

[Mik87] Mikhail Naguib Mikhail. On the validity and stability of the method of lines for the solution of partial differential equations. Applied Mathematics and Computation, 22(2–3): 89–98, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Miller:1987:FRE

[Mil87a] John D. Miller. Finding roots of equations in given neighborhoods. Applied Mathematics and Computation, 23(3):185– 192, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Miller:1987:ZM

[Mil87b] John D. Miller. Zeros of maps of r^{n} into r^{n}. Applied Mathematics and Computation, 21 (4):277–282, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Miller:1988:FZA

[Mil88a] John Miller. Finding the zeros of an analytic function. Applied Mathematics and Computation, 27(3 (part I)):243–249, September 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Milne:1988:MFG

[Mil88b] Bruce T. Milne. Measuring the fractal geometry of landscapes. Applied Mathematics and Computation, 27(1):67–79, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Milito:1995:ADQ

[Mil95] Rodolfo A. Milito. On arrival driven queueing models: Admission control, traffic policing, abandonments, and correlated arrivals. Applied Mathematics and Computation, 70 (1):77–94, June 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=1&aid=9400144.

Miller:1998:OSP

[Mil98] Robert E. Miller. Optimal sensor placement via Gaussian quadrature. Applied Mathematics and Computation, 97 (1):71–97, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/1/6132.
pdf;   http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=97&issue=
                                        1&aid=6132.

Mitra:1984:EPA

[Mit84] A. K. Mitra. An eigenvalue problem arising in a patch formation model. Applied Mathematics and Computation, 15 (4):275–281, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mittelmann:1986:MCT

[Mit86] H.-D. Mittelmann. Multilevel continuation techniques for nonlinear boundary value problems with parameter dependence. Applied Mathematics and Computation, 19(1–4):265– 282, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Mitra:1990:GIF

[Mit90a] A. K. Mitra. Graphing implicit functions f(x, y) = 0. Applied Mathematics and Computation, 39(3):199–205, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mitra:1990:NMM

[Mit90b] A. K. Mitra. A neighborhood magnification method for solving implicit equations in two variables. Applied Mathematics and Computation, 37(2 (part II)):97–110, May 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mitra:1992:CVM

[Mit92] A. K. Mitra. A C-version of minimal Boolean sum-of-product forms. Applied Mathematics and Computation, 47(2– 3):91–108, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mitra:1997:GFV

[Mit97] Arun K. Mitra. On a geometry of the flat violin boundary. Applied Mathematics and Computation, 81(2–3):201– 206, February 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=2-3&aid=9500330.

Mitra:1999:VA

[Mit99] Arun K. Mitra. The violin arcs.

Applied Mathematics and Computation, 98(1):91–102, January 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/1/6146.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
1&aid=6146.

Mishra:1993:TTLb

[MK93] R. K. Mishra and Satish Kumar. Toward a theory of living state — part II. Applied Mathematics and Computation, 56(2–3):167–??, July 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Miketinac:1987:APV

[MKJ87] M. J. Miketinac, A. W. Kelm, and D. S. Jayas. Airflow patterns for various approximations to velocity-pressure-gradient data. Applied Mathematics and Computation, 22 (??):45–63, 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mu:1996:OIP

[MKT96] Gu Mu, Robert E. Kalaba, and Grant A. Taylor. Obtaining initial parameter estimates for chaotic dynamical systems using linear associative memories. Applied Mathematics and Computation, 76(2–3):143–159, May 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        76&issue=2-3&aid=9500153.

Murty:1990:DWC

[ML90] K. N. Murty and P. V. S. Lakshmi. On dichotomy and well-conditioning in two-point boundary-value problems. Applied Mathematics and Computation, 38(3):179–199, August 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morlet:1999:CSO

[MLB99] Anne C. Morlet, Nancy J. Lybeck, and Kenneth L. Bowers. Convergence of the sinc overlapping domain decomposition method. Applied Mathematics and Computation, 98(2– 3):209–227, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/
6169.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6169.

Marsili-Libelli:1987:ASA

[MLC87] S. Marsili-Libelli and M. Castelli.

An adaptive search algorithm for numerical optimization. Applied Mathematics and Computation, 23(4):341–357, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Miel:1985:CNL

[MM85] George Miel and Rose Mooney.

On the condition number of Lagrangian numerical differentiation. Applied Mathematics and Computation, 16(3):241– 252, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maitre:1987:MMS

[MM87a] Jean-François Maitre and François Musy. Multigrid methods for symmetric variational problems: a general theory and convergence estimates for usual smoothers. Applied Mathematics and Computation, 21(1):21–43, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Meintjes:1987:MSC

[MM87b] Keith Meintjes and Alexander P. Morgan. A methodology for solving chemical equilibrium systems. Applied Mathematics and Computation, 22 (4):333–361, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martin:1991:DON

[MM91] C. Martin and J. Miller. Discrete observability of nonlinear systems using continuation techniques. Applied Mathematics and Computation, 44(3):241– 248, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Miller:1994:SPM

[MM94] J. Miller and C. Martin. Some properties of mappings induced by Dirichlet polynomials. Applied Mathematics and Computation, 64(1):1–11, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

McDonough:1998:DFP

[MMC98] J. M. McDonough, S. Mukerji, and S. Chung. A datafitting procedure for chaotic time series. Applied Mathematics and Computation, 95 (2–3):219–243, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6057.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6057.

Maier:1996:MLD

[MMMK96] Stefan Maier, Mario E. Magaña, Ronald Mohler, and Stephen Kaattari. Modified limiting dilution assay for cell systems. Applied Mathematics and Computation, 78(2–3):187– 195, September 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600008.

Martin:1989:NSP

[MMP89a] C. Martin, J. Miller, and K. Pearce. Numerical solution of positive sum exponential equations. Applied Mathematics and Computation, 34(2 (part II)): 89–93, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martin:1989:PIC

[MMP89b] C. Martin, J. Miller, and K. Pearce. Parameter identification by continuation methods. Applied Mathematics and Computation, 34(1 (part I)):17– 27, November 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martin:1990:SEP

[MMP90] C. Martin, J. Miller, and K. Pearce. Sums of exponentials with polynomial coefficients. Applied Mathematics and Computation, 40(1 (part I)): 33–39, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mastrangelo:1999:NGD

[MMT99] Michèle Mastrangelo, Victor Mastrangelo, and Jean-Marie Teuler. Non-Gaussian distributions. Applied Mathematics and Computation, 101 (2–3):99–124, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/101/2-
                                        3/6229.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        101&issue=2-3&aid=6229.

Murthy:1979:ANI

[MN79] V. K. Murthy and Gorti V. L. Narasimham. On the asymptotic normality and independence of the sample partial autocorrelations for an autoregressive process. Applied Mathematics and Computation, 5(4): 281–295, ???? 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martin:1991:AMW

[MN91] Clyde F. Martin and John Newman. Analytic models of water loss in sprinkler irrigation. Applied Mathematics and Computation, 43(1):19–41, 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Matsuzuka:1999:PTS

[MNK99] Isamu Matsuzuka, Koji Nagasawa, and Akihiro Kitahama. A proposal for two-sided Laplace transforms and its application to electronic circuits. Applied Mathematics and Computation, 100(1):1–11, April 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/1/
6203.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=100&
issue=1&aid=6203.

Mazza:1992:MMA

[MO92] Christian Mazza and Neil O’Connell. Microscopic and macroscopic aspects of epidemics. Applied Mathematics and Computation, 47(2–3):237– 258, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mitra:1993:MPC

[MO93] Arun K. Mitra and Justin Ortiz. A massive parallel computation in Boolean logic. Applied Mathematics and Computation, 57(2– 3):131–160, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mo:1997:NBV

[Mo97] Jiaqi Mo. The nonlocal boundary value problems of nonlinear elliptic systems in unbounded domains. Applied Mathematics and Computation, 86(2–3): 115–121, October 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        86&issue=2-3&aid=9600176.

Mohd:1990:UGO

[Moh90]  Ismail Bin Mohd.  Uncon-

strained global optimization using strict complementary slackness. Applied Mathematics and Computation, 36(2 (part II)): 75–87, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mohd:1994:IEM

[Moh94] Ismail Bin Mohd. Interval elimination method for stochastic spanning tree problem. Applied Mathematics and Computation, 66(2–3):325–341, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mohr:1999:FEM

[Moh99a] G. A. Mohr. Finite element modelling of distribution problems. Applied Mathematics and Computation, 105 (1):69–76, October 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/1/
6318.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=105&
issue=1&aid=6318.

Mohr:1999:NPI

[Moh99b] G. A. Mohr. Numerical procedures for input–output analysis. Applied Mathematics and Computation, 101(1):89–98, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/101/1/
6219.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=101&
issue=1&aid=6219.

Mohr:1999:TSM

[Moh99c] G. A. Mohr. Time stepping of macroeconomic models. Applied Mathematics and Computation, 102(2–3):273–278, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/102/2-
                                        3/6266.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        102&issue=2-3&aid=6266.

Moore:1981:EMQ

[Moo81] Jane Moore. Existence of multiple quasifixed points of mixed monotone operators by iterative techniques. Applied Mathematics and Computation, 9(2): 135–141, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morgan:1986:HSP

[Mor86a] Alexander P. Morgan. A homotopy for solving polynomial systems. Applied Mathematics and Computation, 18(1):87–92, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morgan:1986:TAS

[Mor86b] Alexander P. Morgan. A transformation to avoid solutions at infinity for polynomial systems. Applied Mathematics and Computation, 18(1):77–86, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morhac:1990:FEF

[Mor90] Miroslav Morháč. Fast errorfree algorithm for the determination of kernels of the periodic Volterra representation. Applied Mathematics and Computation, 38(2 (part II)):87–100, July 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morhac:1994:EFL

[Mor94] Miroslav Morháč. An errorfree Levison algorithm to solve integer Toeplitz system. Applied Mathematics and Computation, 61(2–3):135–149, March 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morhac:1995:ASH

[Mor95] M. Morháč. An algorithm to solve Hilbert systems of linear equations precisely. Applied Mathematics and Computation, 73(2–3):209–229, December 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=2-3&aid=9500061.

Morchalo:1998:SAN

[Mor98]  Jarosław Morchało.  Stabil-

ity analysis of nonlinear difference systems. Applied Mathematics and Computation, 95 (1):27–36, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/95/1/6101.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=95&issue=
                                        1&aid=6101.

Mosetti:1988:OSM

[Mos88a] R. Mosetti. Optimal strategies for monitoring a variable subjected to random changes. Applied Mathematics and Computation, 25(2 (part II)):137–143, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mosetti:1988:OPB

[Mos88b] Renzo Mosetti. Optimal policies in a bioeconomic model of eutrophication. Applied Mathematics and Computation, 26(3 (part II)):257–265, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mosetti:1990:PIM

[Mos90] Renzo Mosetti. On a possible interpretation of market dynamics based on the internal model principle. Applied Mathematics and Computation, 37(1 (part I)):1–8, May 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Moustafa:1995:CGM

[Mou95] Magdi. S. Moustafa. The conjugate gradient method for queueing networks. Applied Mathematics and Computation, 71(2– 3):289–294, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400159.

Moustafa:1996:OAP

[Mou96] Magdi S. Moustafa. Optimal assignment policy of a single server attended by two queues. Applied Mathematics and Computation, 80(2–3):245–255, December 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500298.

McAleer:1986:SIN

[MP86] Michael McAleer and M. Hashem Pesaran. Statistical inference in nonnested econometric models. Applied Mathematics and Computation, 20(3–4):271–311, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mandel:1990:MC

[MP90] Jan Mandel and Seymour V.

Parter. On the multigrid F - cycle. Applied Mathematics and Computation, 37(1 (part I)): 19–36, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Michalland:1997:BOD

[MPD97] B. Michalland, E. Parent, and L. Duckstein. Bi-objective dynamic programming for trading off hydropower and irrigation. Applied Mathematics and Computation, 88(1):53– 76, December 15, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=1&aid=9600306.

Monaco:1997:NSB

[MPS97] R. Monaco, M. Pandolfi Bianchi, and A. J. Soares. Numerical simulations of a Boltzmann model for reacting gases. Applied Mathematics and Computation, 85(1):61–85, August 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1997/85/1/9600113.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1997&volume=85&issue=
1&aid=9600113.

Muratori:1989:DSH

[MR89] Simona Muratori and Sergio Rinaldi. A dynamical system with Hopf bifurcations and catastrophes. Applied Mathematics and Computation, 29(1 (part I)):1– 15, January 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muratori:1991:ILS

[MR91] Simona Muratori and Sergio Rinaldi. On the invertibility of Leslie systems. Applied Mathematics and Computation, 41(2 (part II)):179–187, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Murty:1999:TPB

[MR99] K. N. Murty and C. V. Rao.

Two-point boundary value problems associated with first order non-linear difference system — existence and uniqueness. Applied Mathematics and Computation, 100(2–3):177–188, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6205.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6205.

Morera:1995:ANI

[MRJ95] J. L. Morera, G. Rubio, and L. Jódar. Accurate numerical integration of stiff differential Riccati equations. Applied Mathematics and Computation, 72 (2–3):183–203, October 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400183.

McCarthy:1983:EMM

[MS83] D. R. McCarthy and R. C. Swanson. Embedded mesh multigrid treatment of twodimensional transonic flows. Applied Mathematics and Computation, 13(3–4):399–417, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morgan:1987:CAS

[MS87a] Alexander Morgan and Andrew Sommese. Computing all solutions to polynomial systems using homotopy continuation. Applied Mathematics and Computation, 24(2):115–138, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See errata [MS92c].

Morgan:1987:HSG

[MS87b] Alexander Morgan and Andrew Sommese. A homotopy for solving general polynomial systems that respects m-homogeneous structures. Applied Mathematics and Computation, 24(2):101– 113, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Medhin:1988:MMM

[MS88] Negash G. Medhin and M. Sambandham. Mixed min-max optimization problem with restrictions. Applied Mathematics and Computation, 26(2 (part I)): 151–167, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morgan:1989:CPP

[MS89] Alexander P. Morgan and Andrew J. Sommese. Coefficientparameter polynomial continuation. Applied Mathematics and Computation, 29(2 (part II)): 123–160, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See errata [MS92b].

Martin:1992:NEG

[MS92a] Clyde Martin and Mark Stamp.

A note on the error in Gaussian quadrature. Applied Mathematics and Computation, 47 (1):25–35, January 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Morgan:1992:ECP

[MS92b] Alexander P. Morgan and Andrew J. Sommese. Errata: “Coefficient-parameter polynomial continuation” [Appl. Math. Comput. 29 (1989), no. 2, part II, 123–160, MR 90f:58018]. Applied Mathematics and Computation, 51(2–3):207, October 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [MS89].

Morgan:1992:ECA

[MS92c] Alexander P. Morgan and Andrew J. Sommese. Errata: “Computing all solutions to polynomial systems using homotopy continuation” [Appl. Math.

Comput. 24 (1987), no. 2, 115– 138, MR 89b:65126]. Applied Mathematics and Computation, 51(2–3):209, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [MS87a].

Martin:1993:NEF

[MS93a] Clyde Martin and Victor Shubov. Natural exponential families of probability distributions and exponentialpolynomial approximation. Applied Mathematics and Computation, 59(2–3):275–297, ???? 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martin:1993:PMA

[MS93b] Clyde Martin and Victor Shubov. Probability measures, Appel polynomials and polynomial approximation. Applied Mathematics and Computation, 53(2–3):277–298, February 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mishra:1993:TTLa

[MS93c] R. K. Mishra and G. K. Sen. Toward a theory of living state — part I. Applied Mathematics and Computation, 56(2–3): 153–??, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Meir:1994:VCF

[MS94a] A. J. Meir and Paul G. Schmidt. A velocity-current formulation for stationary MHD flow. Applied Mathematics and Computation, 65(1–3):95–109, September/October 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Melnikov:1994:MGT

[MS94b] Yu. A. Mel′nikov and K. L.

Shirley. Matrices of Green’s type for the potential equation on a cylindrical surface joined to a hemisphere. Applied Mathematics and Computation, 65(1– 3):241–252, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Molnar:1994:VTL

[MS94c] Sándor Molnár and Ferenc Szigeti. On “verticum”- type linear systems with timedependent linkage. Applied Mathematics and Computation, 60(1):89–102, January 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Miller:1995:FRC

[MS95a] A. R. Miller and H. M. Srivastava. Further reducible cases of certain kampé de Fériet functions associated with incomplete integrals of cylindrical functions. Applied Mathematics and Computation, 68(2–3): 199–216, March 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=2-3&aid=9400094.

Morrice:1995:AMS

[MS95b] Douglas J. Morrice and Mark Song. An approximation method for a single server queue with cyclically indexed arrival and service rates. Applied Mathematics and Computation, 71(2– 3):247–269, September 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400157.

Moghadam:1997:MFA

[MS97] M. Mohseni Moghadam and R. Sabeti. Markov finite approximation of Frobenius– Perron operator with doublystochastic tridiagonal action matrix. Applied Mathematics and Computation, 85(2–3):265– 280, September 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600140.

Malakooti:1999:GPD

[MS99] Behnam Malakooti and Sriram Subramanian. Generalized polynomial decomposable multiple attribute utility functions for ranking and rating multiple criteria discrete alternatives. Applied Mathematics and Computation, 106(1):69–102, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/17/22/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/17/22/
article.pdf.

McArthur:1992:SGM

[MSLB92] Kelly M. McArthur, Ralph C.

Smith, John Lund, and Kenneth L. Bowers. The Sinc– Galerkin method for parameterdependent self-adjoint problems. Applied Mathematics and Computation, 50(2–3):175–202, August 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Montaldi:1996:HIC

[MSZ96] E. Montaldi, H. M. Srivastava, and G. Zucchelli. Hypergeometric identities of Cayley and Orr types: Some alternative approaches. Applied Mathematics and Computation, 76 (1):1–39, April 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/

                                        browse.cgi?year=1996&volume=
                                        76&issue=1&aid=9500145.

Mercier:1993:AAP

[MTP93] P. Mercier, J. Teghem, and H. Pastijn. Assignment of available products to orders with the MCDM software oreste. Applied Mathematics and Computation, 54(2–3):183–196, March 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

McKeague:1990:SCTa

[MU90a] Ian W. McKeague and Klaus J. Utikal. Stochastic calculus as a tool in survival analysis: a review. Applied Mathematics and Computation, 38(1 (part I)):23–49, July 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

McKeague:1990:SCTb

[MU90b] Ian W. McKeague and Klaus J. Utikal. Stochastic calculus as a tool in survival analysis: a review. Applied Mathematics and Computation, 39(3 (Supplement)):155s–181s, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muraskin:1989:SDL

[Mur89a] M. Muraskin. Study of different lattice solutions in aesthetic field theory. Applied Mathematics and Computation, 30(1 (part I)):73–98, March 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muraskin:1989:UCR

[Mur89b] M. Muraskin. Use of commutation relations in no integrability aesthetic field theory. Applied Mathematics and Computation, 29(3):271–292, February 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muraskin:1991:SPD

[Mur91a] M. Muraskin. Summation-over-paths degree of freedom and the loop lattice. Applied Mathematics and Computation, 42(2 (part II)):197–207, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muratori:1991:ASP

[Mur91b] Simona Muratori. An application of the separation principle for detecting slow-fast limit cycles in a three-dimensional system. Applied Mathematics and Computation, 43(1 (part I)): 1–18, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muraskin:1992:BNL

[Mur92] M. Muraskin. Beyond Newton and Leibniz. Applied Mathematics and Computation, 52(2–3): 417–437, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muraskin:1993:OWO

[Mur93]  M. Muraskin.  Oscillations

within oscillations. Applied Mathematics and Computation, 53(1):45–82, January 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muraskin:1994:IOM

[Mur94] M. Muraskin. Irregular oscillations and mathematical aesthetics. Applied Mathematics and Computation, 61(2–3):183– 206, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muraskin:1996:ESN

[Mur96] Murray Muraskin. Equations satisfied by a nine parameter subsystem obtained from mathematical aesthetic principles. Applied Mathematics and Computation, 80(1):63– 72, November 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=1&aid=9500281.

Muses:1978:HIF

[Mus78] C. Musès. Hypernumbers. II. further concepts and computational applications. Applied Mathematics and Computation, 4(1):45–66, ???? 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muses:1980:HQF

[Mus80] C. Musès. Hypernumbers and quantum field theory with a summary of physically applicable hypernumber arithmetics and their geometries. Applied Mathematics and Computation, 6(1):63–94, January 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muses:1983:EHT

[Mus83a] C. Musès. Errata: “Hypernumbers and time operators”. Applied Mathematics and Computation, 12(4):395, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Mus83b, Mus84a].

Muses:1983:HTO

[Mus83b] C. Musès. Hypernumbers and time operators. Applied Mathematics and Computation, 12(2– 3):139–167, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See errata [Mus83a, Mus84a].

Muses:1984:EHT

[Mus84a] C. Musès. Errata: “Hypernumbers and time operators”. Applied Mathematics and Computation, 14(2):219, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Mus83b, Mus83a].

Muses:1984:SCD

[Mus84b] C. Musès. Some current dilemmas in applied physical mathematics with some solutions. Applied Mathematics and Computation, 14(2):207–217, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muses:1994:HAH

[Mus94] C. Musès. Hypernumbers applied or how they interface with the physical world. Applied Mathematics and Computation, 60(1):25–36, January 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Muses:1997:DFA

[Mus97a] C. Musès. The dimensional family approach in (hyper)sphere packing: A typological study of new patterns, structures, and interdimensional functions. Applied Mathematics and Computation, 88(1):1–26, December 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=1&aid=9700004. See

erratum [Mus97b].

Muses:1997:EDF

[Mus97b] C. Musés. Erratum: “The dimensional family approach in (hyper)sphere packing: a typological study of new patterns, structures, and interdimensional functions” [Appl. Math. Comput. 88 (1997), no. 1, 1–26; 1 473 570]. Applied Mathematics and Computation, 88(1):

i, 1997. URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=1&aid=9700402. See

[Mus97a].

Muses:1997:SPB

[Mus97c] C. Musès. A sphere-packing breakthrough via dimensional families. Applied Mathematics and Computation, 83(1): 1–2, April 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=1&aid=9600331.

Mollers:1998:GAI

[Mv98] Th. Möllers and O. van Laak.

On the global approximation and interpolation of locally described real-valued functions. Applied Mathematics and Computation, 93(1):1–10, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/1/6054.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
1&aid=6054.

Monsi:1988:ASI

[MW88a] M. Monsi and M. A. Wolfe.

An algorithm for the simultaneous inclusion of real polynomial zeros. Applied Mathematics and Computation, 25 (4):331–346, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Monsi:1988:IVS

[MW88b] M. Monsi and M. A. Wolfe. Interval versions of some procedures for the simultaneous estimation of complex polynomial zeros. Applied Mathematics and Computation, 28(3 (part I)): 191–209, December 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mirchandani:1993:RNM

[MW93] Pitu B. Mirchandani and Małgorzata M. Wiecek. Routing with nonlinear multiattribute cost functions. Applied Mathematics and Computation, 54 (2–3):215–239, March 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martin:1996:QSG

[MW96] C. F. Martin and D. I. Wallace. Quadrature schemes for a general family of functions. Applied Mathematics and Computation, 77(1):1–8, June 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        77&issue=1&aid=9500069.

Mckay:1998:SSL

[MW98a] Steven M. Mckay and Cecilia L. Weingartner. Solution of the Sharpe-Lotka population model via a modified method of characteristics. Applied Mathematics and Computation, 91(2–3):161–177, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6016.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6016.

Wang:1998:PGM

[mW98b] Yuan ming Wang. Petrov– Galerkin methods for systems of nonlinear reaction-diffusion equations. Applied Mathematics and Computation, 96(2–3): 209–236, November 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/2-3/
6123.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
2-3&aid=6123.

Martin:1990:OENa

[MX90a] Clyde Martin and Shishen Xie.

Observability of the electropotential on a nested cylindrical domain. I: The continuous solution. Applied Mathematics and Computation, 39(3):207– 244, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Martin:1990:OENb

[MX90b] Clyde Martin and Shishen Xie. Observability of the electropotential on a nested cylindrical domain. II. numerical solution. Applied Mathematics and Computation, 39(3):245–270, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Malakooti:1998:APF

[MZ98] Behnam Malakooti and YingQing Zhou. Approximating polynomial functions by feedforward artificial neural networks: Capacity analysis and design. Applied Mathematics and Computation, 90(1):27–51 (or 27– 52??), March 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        90&issue=1&aid=9600338.

Mohler:1996:NCA

[MZV96] Ronald R. Mohler, Radosław R. Zakrzewski, and Rajkumar Vedam. Nonlinear control algorithms and power system application. Applied Mathematics and Computation, 78(2–3):197– 207, September 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600009.

Nwannenna:1991:RMC

[NA91] C. C. Nwannenna and E. M.

Aghedo. Reliability monitoring and control in electric power systems. Applied Mathematics and Computation, 43(2 (part II)):117–138, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nakamura:1982:MGG

[Nak82] S. Nakamura. Marching grid generation using parabolic partial differential equations. Applied Mathematics and Computation, 10/11(1):775–786, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Namachchivaya:1990:SBa

[Nam90a] N. Sri Namachchivaya. Stochastic bifurcation. Applied Mathematics and Computation, 38(2 (part II)):101–159, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Namachchivaya:1990:SBb

[Nam90b] N. Sri Namachchivaya. Stochastic bifurcation. Applied Mathematics and Computation, 39 (3 (Supplement)):37s–95s, ????

1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nistal:1993:NIR

[NAM93] J. L. Freire Nistal, F. Alonso Amo, and F. Aguado Martín. On the naturality of the iteration and recursive specifications. Applied Mathematics and Computation, 57(1):1–18, August 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Narayaninsamy:1999:FID

[Nar99a] Tony Narayaninsamy. Fractional iterates for n-dimensional maps. Applied Mathematics and Computation, 98(2– 3):261–278, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/
6183.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6183.

Narayaninsamy:1999:BB

[Nar99b] Tony Narayaninsamy. On the basin boundaries. Applied Mathematics and Computation, 99 (2–3):261–274, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6185.pdf;   http://www.elsevier.

com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6185.

Neuman:1985:MLA

[NC85] S. P. Neuman and J. Carrera.

Maximum-likelihood adjointstate finite-element estimation of groundwater parameters under steady- and non-steady-state conditions. Applied Mathematics and Computation, 17(4):405– 432, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). System identification and parameter estimation of environmental systems.

Neta:1989:SMP

[Net89] Beny Neta. Special methods for problems whose oscillatory solution is damped. Applied Mathematics and Computation, 31(Special issue):161–169, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Navarro:1994:CFG

[NFJ94] E. Navarro, M. V. Ferrer, and L. Jódar. Closed form general solution of nonhomogeneous implicit higher order difference systems. Applied Mathematics and Computation, 60(2–3): 113–123, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [JNF95].

Nakamura:1994:APF

[NG94] Takashi Nakamura and Yukio-Pegio Gunji. Algebraic property of flow diagram in a timereverse cellular automaton. Applied Mathematics and Computation, 62(2–3):103–127, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Newbold:1986:TCU

[NH86] Paul Newbold and Steven M. Hotopp. Testing causality using efficiently parametrized vector ARMA models. Applied Mathematics and Computation, 20(3– 4):329–348, ???? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nathanson:1983:TOD

[NHG83] M. H. Nathanson, R. S. Hillman, and C. Georgakis. Towards an optimal drug-delivery regimen for methotrexate chemotherapy. Applied Mathematics and Computation, 12(??):99–117, 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

He:1997:CSF

[nHyG97] Song nian He and Ben yu Guo. Chebyshev spectral-finite element method for three-dimensional unsteady Navier–Stokes equation. Applied Mathematics and Computation, 82(2–3):97– 127, March 15, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
82&issue=2-3&aid=9600023.

Nieto:1984:EPS

[Nie84] Juan J. Nieto. Existence of periodic solutions for first order differential equations. Applied Mathematics and Computation, 15(3):221–232, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nieto:1985:SSS

[Nie85] Juan J. Nieto. On the structure of the solution set for first order differential equations. Applied Mathematics and Computation, 16(3):177–187, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nedialkov:1999:VSI

[NJC99] N. S. Nedialkov, K. R. Jackson, and G. F. Corliss. Validated solutions of initial value problems for ordinary differential equations. Applied Mathematics and Computation, 105 (1):21–68, October 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/1/

                                        6304.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=105&
                                        issue=1&aid=6304.

Ng:1995:CPW

[NLC95] Michael K. Ng, Fu-Rong Lin, and Raymond H. Chan. Construction of preconditioners for Wiener–Hopf equations by operator splitting. Applied Mathematics and Computation, 72(1): 77–96, September 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        72&issue=1&aid=9400178.

Neta:1987:AMN

[NN87] Beny Neta and Paul Nelson. An adaptive method for the numerical solution of Fredholm integral equations of the second kind. I: Regular kernels. Applied Mathematics and Computation, 21 (2):171–184, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Newman:1992:CCI

[NP92] D. J. Newman and M. E. Primak. Complexity of circumscribed and inscribed ellipsoid methods for solving equilibrium economical models. Applied Mathematics and Computation, 52(2–3):223–231, December 1992. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Natesan:1998:CMS

[NR98] S. Natesan and N. Ramanujam. A computational method for solving singularly perturbed turning point problems exhibiting twin boundary layers. Applied Mathematics and Computation, 93(2–3):259–275, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/2-3/
6076.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6076.

Natesan:1999:BMS

[NR99a] S. Natesan and N. Ramanujam.

A “booster method” for singular perturbation problems arising in chemical reactor theory. Applied Mathematics and Computation, 100(1):27–48, April 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/1/
6198.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=100&
issue=1&aid=6198.

Natesan:1999:INS

[NR99b] S. Natesan and N. Ramanujam. Improvement of numerical solution of self-adjoint singular perturbation problems by incorporation of asymptotic approximations. Applied Mathematics and Computation, 98(2– 3):119–137, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/98/2-3/
                                        6168.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=98&issue=
                                        2-3&aid=6168.

Nowak:1989:GDA

[NS89] Martin Nowak and Karl Sigmund. Game-dynamical aspects of the prisoner’s dilemma. Applied Mathematics and Computation, 30(3):191–213, April 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Navarro:1996:FPS

[NS96] Ramón Navarro and José Sarabia. Focal points and secondorder linear differential equations. Applied Mathematics and Computation, 79(2–3):203– 206, October 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        79&issue=2-3&aid=9500266.

Nelson:1997:IET

[NS97] Paul Nelson and Daniel L. Seth. Integrodifferential equations for the two-dimensional transition kernels of invariant imbedding. Applied Mathematics and Computation, 82(1):67–83, March 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
82&issue=1&aid=9500331.

Nelson:1981:IIA

[NSE81] Paul Nelson, Seok Sagong, and Ira T. Elder. Invariant imbedding applied to homogeneous two-point boundary value problems with a singularity of the first kind. Applied Mathematics and Computation, 9(2):93–110, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Nelson:1992:IET

[NSV92] Paul Nelson, D. L. Seth, and R. Vasudevan. An integrodifferential equation for the twodimensional reflection kernel. Applied Mathematics and Computation, 49(1):1–18, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Niestegge:1990:AMS

[NW90] A. Niestegge and K. Witsch.

Analysis of a multigrid Stokes solver. Applied Mathematics and Computation, 35(3):291– 303, February 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Obeyesekere:1997:MWU

[OA97] Mandri Obeyesekere and Ronald M. Anderson. A model for water uptake in plants. Applied Mathematics and Computation, 84(2– 3):163–191, July 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600085.

Ozelkan:1998:MOF

[ÖGDB98] Ertunga C. Özelkan, Ágnes Galambosi, Lucien Duckstein, and András Bárdossy. A multiobjective fuzzy classification of large scale atmospheric circulation patterns for precipitation modeling. Applied Mathematics and Computation, 91 (2–3):127–142, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6002.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6002.

Ohmori:1998:NST

[Ohm98] Katsushi Ohmori. Numerical solution of two-fluid flows using finite element method. Applied Mathematics and Computation, 92(2–3):125–133, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/2-3/
9710036.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
92&issue=2-3&aid=9710036.

Oguztoreli:1980:ANS

[OL80a] M. N. Oğuztöreli and K. V. Leung. Analysis and numerical solution of a functional differential equation. Applied Mathematics and Computation, 6(4): 289–307, January 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ou:1980:NEH

[OL80b] Szu Cheng S. Ou and Kuo Nan Liou. Numerical experiments on the Helmholtz equation derived from the solar radiation transfer equation in three-dimensional space. Applied Mathematics and Computation, 7(2):155–175, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

ORegan:1999:PAP

[OM99] D. O’Regan and M. Meehan. Periodic and almost periodic solutions of integral equations. Applied Mathematics and Computation, 105(2–3): 121–136, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/

                                        store/amc/sub/1999/105/2-
                                        3/6316.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        105&issue=2-3&aid=6316.

Omari:1986:MMC

[Oma86] Pierpaolo Omari. A monotone method for constructing extremal solutions of second order scalar boundary value problems. Applied Mathematics and Computation, 18(3):257–275, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

OReilly:1995:RDS

[OMS95] Oliver M. O’Reilly, Naresh K. Malhotra, and N. Sri Namachchivaya. Reversible dynamical systems: Dissipationinduced destabilization and follower forces. Applied Mathematics and Computation, 70(2– 3):273–282, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        70&issue=2-3&aid=9400117.

Dynamics and control (Vienna, 1993).

ONeil:1990:CMR

[O’N90] Kevin A. O’Neil. The continuation method for rational and deficient systems. Applied Mathematics and Computation, 37(2 (part II)):121–129, May 1990.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Onah:1997:DMD

[Ona97] E. S. Onah. On direct methods for the discretization of a heat-conduction equation using spline functions. Applied Mathematics and Computation, 85(1):87–96, August 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1997/85/1/9600114.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1997&volume=85&issue=
1&aid=9600114.

Okamoto:1998:NNO

[ONOT98] Masahiro Okamoto, Taisuke Nonaka, Shuichiro Ochiai, and Daisuke Tominaga. Nonlinear numerical optimization with use of a hybrid genetic algorithm incorporating the modified Powell method. Applied Mathematics and Computation, 91(1):63–72, April 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/1/6007.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
1&aid=6007.

Oppenheimer:1992:ASA

[Opp92] Seth F. Oppenheimer. An abstract spectral approximation theorem from the theory of semigroups. Applied Mathematics and Computation, 47(2–3):185– 199, ???? 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Orlovski:1983:MMA

[ORSS83] S. Orlovski, S. Rinaldi, and R. Soncini-Sessa. A min-max approach to storage control problems. Applied Mathematics and Computation, 12(2–3):237–254, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Oguztoreli:1980:NRE

[OS80] M. N. Oğuztöreli and R. B. Stein. Neuromuscular resetting and entrainment. Applied Mathematics and Computation, 7(2): 93–124, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Oguztoreli:1980:WPD

[OŞL80] M. N. Oğuztöreli, E. S. Şuhubi, and K. V. Leung. Wave propagation in dissipative or dispersive nonlinear media: a numerical approach. Applied Mathematics and Computation, 6(4): 309–334, January 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Omari:1992:RLU

[OT92] Pierpaolo Omari and Maurizio Trombetta. Remarks on the lower and upper solutions method for second- and thirdorder periodic boundary value problems. Applied Mathematics and Computation, 50(1):1–21, July 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [OT93].

Omari:1993:ERL

[OT93] Pierpaolo Omari and Maurizio Trombetta. Erratum: “Remarks on the lower and upper solutions method for second- and thirdorder periodic boundary value problems” [Appl. Math. Comput. 50 (1992), no. 1, 1–21, MR 93d:34040]. Applied Mathematics and Computation, 56(1):101, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [OT92].

Oguztoreli:1983:KSC

[OTA83] M. N. Oğuztöreli, C. P. Tsokos, and J. Akabutu. A kinetic study of chemotherapy. Applied Mathematics and Computation, 12(2– 3):255–300, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Onah:1999:STD

[OU99] S. E. Onah and O. O. Ugbebor.

Solution of a two-dimensional stochastic investment problem. Applied Mathematics and Computation, 98(1):75–82, January 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/98/1/6158.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=98&issue=
                                        1&aid=6158.

Pang:1994:DBV

[PA94] Peter Y. H. Pang and Ravi P. Agarwal. On discrete boundary value problems arising in transport phenomena. Applied Mathematics and Computation, 60(2– 3):193–213, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pachpatte:1985:NCR

[Pac85] B. G. Pachpatte. On nonlinear coupled reaction-diffusion equations. Applied Mathematics and Computation, 16(4):297– 307, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pallaschke:1978:MCI

[Pal78] Diethard Pallaschke. Markets with countably infinitely many traders. Applied Mathematics and Computation, 4(3):201–212, ???? 1978. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Palm:1986:SEM

[Pal86] F. C. Palm. Structural econometric modeling and time series analysis. Applied Mathematics and Computation, 20(3– 4):349–364, ???? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pandian:1988:WDS

[Pan88] M. C. Pandian. A weighted difference scheme and monotone iterative methods for quasilinear boundary value problems. Applied Mathematics and Computation, 25(3):187–205, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Papa:1984:ACI

[Pap84] L. Papa. Application of the Courant-Isaacson-Rees method to solve the shallow-water hydrodynamic equations. Applied Mathematics and Computation, 15(1):85–92, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Papageorgiou:1992:EWM

[Pap92] Nikolaos S. Papageorgiou. On the existence of weakly maximal programs for a multisector economy with consumption. Applied Mathematics and Computation, 48(2–3):177–186, April 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Papageorgiou:1994:SAC

[Pap94] Nikolaos S. Papageorgiou. Sensitivity analysis for a continuous time infinite horizon growth model. Applied Mathematics and Computation, 60(1):43– 53, January 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Papageorgiou:1995:MCN

[Pap95] Nikolaos S. Papageorgiou. Minimax control of nonlinear evolution equations. Applied Mathematics and Computation, 68 (2–3):217–236, March 15, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=2-3&aid=9400095.

Papamarkos:1996:MFA

[Pap96] Nikos Papamarkos. Multirational function approximation via linear programming. Applied Mathematics and Computation, 75(1):75–89, March 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=1&aid=9500119.

Parter:1985:NCM

[Par85] Seymour V. Parter. A note on convergence of the multigrid V -cycle. Applied Mathematics and Computation, 17 (2):137–151, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Parter:1987:RMC

[Par87] Seymour V. Parter. Remarks on multigrid convergence theorems. Applied Mathematics and Computation, 23(2):103– 120, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pardalos:1989:PSA

[Par89] P. M. Pardalos. Parallel search algorithms in global optimization. Applied Mathematics and Computation, 29(3):219– 229, February 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pardo:1995:SAU

[Par95] J. A. Pardo. Some applications of the useful mutual information. Applied Mathematics and Computation, 72(1):33– 50, September 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=1&aid=9400162.

Patterson:1987:REI

[Pat87] Richard L. Patterson. Regression estimates of inputs to an M(t)/G/∞ service system. Applied Mathematics and Computation, 24(1):47–63, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Patino:1988:FIS

[Pat88] Jorge G. S. Patiño. A fully implicit scheme for multiphase flow in one-dimensional porous media. Applied Mathematics and Computation, 28(4 (part II)): 289–296, December 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Patterson:1989:RES

[Pat89] Richard L. Patterson. Regression estimates of stochastic compartmental parameters using net balance and cumulative flow data. Applied Mathematics and Computation, 29(3): 245–254, February 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Patterson:1990:MRE

[Pat90]  Richard L. Patterson.  A

Markov-renewal extension of a deterministic state-variable model of the transport and fate of a toxic chemical in the environment. Applied Mathematics and Computation, 35(3):219– 230, February 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Patterson:1991:RAC

[Pat91] Richard L. Patterson. Regression approximations for continuous-time, stochastic compartmental models of discrete interacting populations. Applied Mathematics and Computation, 45(1 (part I)):73–89, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Patterson:1992:DUS

[Pat92] Richard L. Patterson. Distributions of unexpended services in a D-node Markov-renewal network distinguished by customer class. Applied Mathematics and Computation, 52(1): 67–82, November 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pattee:1993:LFM

[Pat93] H. H. Pattee. The limitations of formal models of measurement, control, and cognition. Applied Mathematics and Computation, 56(2–3):111–130, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pavani:1995:NAR

[Pav95] R. Pavani. A numerical approximation of the rotation number. Applied Mathematics and Computation, 73(2–3):191– 201, December 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=2-3&aid=9400249.

Poston:1992:BSE

[PBL92] Timothy Poston, Hyeong Ohk Bae, and Chung-Nim Lee. Bifurcation structure of the Eckalbar stock-holding model. Applied Mathematics and Computation, 48(1):21–43, March 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Peiguang:1999:OCN

[Pei99] Wang Peiguang. Oscillatory criteria of nonlinear hyperbolic equations with continuous deviating arguments. Applied Mathematics and Computation, 106 (2–3):163–169, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/21/
article.pdf.

Pennline:1984:CEU

[Pen84] James A. Pennline. Constructive existence and uniqueness for two-point boundary value problems with a linear gradient term. Applied Mathematics and Computation, 15(3):233– 260, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Peng:1993:IPN

[Pen93] Zhang Peng. An inverse problem in nonlinear optics. Applied Mathematics and Computation, 56(1):59–70, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pervozvanski:1996:SAL

[Per96] A. Pervozvanski. On some algorithms of learning control. Applied Mathematics and Computation, 78(2–3):209– 216, September 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600010.

Pesamosca:1991:FPA

[Pes91] Giancarlo Pesamosca. A fixed point algorithm for solving the Euclidean multifacility location problem in a tree. Applied Mathematics and Computation, 44(2 (part II)):105–120, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Peters:1988:PES

[Pet88] Randall D. Peters. Period estimate for the simple pendulum using optimization theory. Applied Mathematics and Computation, 28(3 (part I)):179– 190, December 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Papageorgiou:1999:UER

[PF99] G. Papageorgiou and I. Th.

Famelis. On using explicit Runge–Kutta–Nyström methods for the treatment of retarded differential equations with periodic solutions. Applied Mathematics and Computation, 102(1):63–76, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/1/
6246.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=102&
issue=1&aid=6246.

Peng:1999:COB

[PGHX99] Mingshu Peng, Weigao Ge, Lihong Huang, and Qianli Xu. A correction on the oscillatory behavior of solutions of certain second order nonlinear differential equations. Applied Mathematics and Computation, 104(2–3):207–215, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/2-
3/6295.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
104&issue=2-3&aid=6295.

Philos:1992:OCD

[Phi92] Ch. G. Philos. Oscillations in a class of difference equations. Applied Mathematics and Computation, 48(1):45–57, March 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Piccardi:1992:FHO

[Pic92] Carlo Piccardi. Finite-horizon optimal control with pointwise cost functional. Applied Mathematics and Computation, 52 (2–3):345–353, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pignatiello:1987:EMO

[Pig87] Joseph J. Pignatiello, Jr.

Enumerating maximal-order Abelian groups within a set of eligible elements. Applied Mathematics and Computation, 22 (2–3):99–114, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pottmann:1995:RBF

[PJ95] Martin Pottmann and H. Peter Jörgl. Radial basis function networks for internal model control. Applied Mathematics and Computation, 70(2–3): 283–298, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400116.

Dynamics and control (Vienna, 1993).

PenaMiralles:1997:FGM

[PJG^{+}97] Juan Peña Miralles, Pedro José Jiménez Olivo, Damián Ginestar Peiró, Gumersindo Verdú Martín, and José Luis Muñoz-Cobo González. A fast Galerkin method to obtain the periodic solutions of a nonlinear oscillator. Applied Mathematics and Computation, 86(2–3):261– 282, October 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
86&issue=2-3&aid=9600193.

Primak:1998:MCV

[PK98] M. E. Primak and B. L. Kheyfets. Methods of centers for variational inequalities and linear programming. Applied Mathematics and Computation, 96(2–3):273–293, November 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/2-3/
6131.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
2-3&aid=6131.

Parent:1993:BOW

[PL93] E. Parent and F. Lebdi. Bicriterion operation of a water resources system with reliabilitybased tradeoffs. Applied Mathematics and Computation, 54 (2–3):197–213, March 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pitanguy:1996:DMA

[PLPW96] I. Pitanguy, F. Leta, D. Pamplona, and H. I. Weber. Defining and measuring aging parameters. Applied Mathematics and Computation, 78(2–3):217– 227, September 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600016.

Luo:1987:EVAa

[pLR87a] Zi ping Luo and Herschel Rabitz. Expected value analysis: Experimental considerations and convergence properties. Applied Mathematics and Computation, 23(1):25–47, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Luo:1987:EVAb

[pLR87b] Zi ping Luo and Herschel Rabitz. Expected value analysis of multiparameter systems. Applied Mathematics and Computation, 24(2):153–181, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Patterson:1988:SAP

[PM88] Richard L. Patterson and Zhen Qui Ma. Sensitivity analyses of parameters of a M(t)/G/∞ stochastic service system. Applied Mathematics and Computation, 26(2 (part I)):169–177, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Patterson:1989:SDC

[PM89] Richard L. Patterson and Zhenkui Ma. Statistical distributions of compartmentalized populations governed by continuous time, discrete state semi-Markov processes. Applied Mathematics and Computation, 30(1 (part I)):49–71, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Prathapar:1997:SOH

[PMMA97] S. A. Prathapar, W. S. Meyer, J. C. Madden, and E. Alociljá. SWAGMAN options: A hierarchical multicriteria framework to identify profitable land uses that minimize water table rise and salinization. Applied Mathematics and Computation, 83(2– 3):217–240, May 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=2-3&aid=9600186.

Parhi:1997:GPA

[PN97a] S. Parhi and G. Nath. The growth of perturbations associated with the linear stability of inviscid fluid flows. Applied Mathematics and Computation, 81(1):23–30, January 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/

browse.cgi?year=1997&volume=
81&issue=1&aid=9500299.

Parton:1997:NPM

[PN97b] Kevin A. Parton and Ayut Nissapa. A nonlinear programming model for analyzing aquaculture policy decision making in southern thailand. Applied Mathematics and Computation, 83(2– 3):241–260, May 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=2-3&aid=9600187.

Petkovic:1993:GOR

[PP93] Lj. Petković and M. Petković.

Generalized Ostrowski rootfinding method. Applied Mathematics and Computation, 57(2– 3):275–287, September 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Philos:1994:GAN

[PPS94] Ch. G. Philos, I. K. Purnaras, and Y. G. Sficas. Global attractivity in a nonlinear difference equation. Applied Mathematics and Computation, 62(2– 3):249–258, May 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pokhariyal:1993:AEM

[PR93] G. P. Pokhariyal and A. J. Rodrigues. An accurate epidemiological model. Applied Mathematics and Computation, 53 (1):1–12, January 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Palusinski:1994:CEM

[PS94a] Olgierd A. Palusinski and Ferenc Szidarovszky. A Chebyshev expansion method for solving nonlinear circuit equations. Applied Mathematics and Computation, 60(2–3):299–310, February 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Patterson:1994:NMT

[PS94b] Richard L. Patterson and David Snyder. A numerical model for tracking populations through a time-dependent stochastic network. Applied Mathematics and Computation, 61(2–3):163– 182, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Primak:1996:PCP

[PS96] Mordukh E. Primak and Daniel B. Szyld. A projection cutting plane algorithm for convex programming problems. Applied Mathematics and Computation, 74(2–3):261–271, February 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/

browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500097.

Pardo:1993:APD

[PSMM93] L. Pardo, M. Salicrú, M. L.

Menéndez, and D. Morales. Asymptotic properties of (r, s)- directed divergences in a stratified sampling. Applied Mathematics and Computation, 55(2– 3):131–152, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Puu:1984:MSF

[Puu84] Toenu Puu. A model of spatial flows and growth of capital and labor stocks. Applied Mathematics and Computation, 14(??):3– 9, 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pachpatte:1981:MME

[PV81] B. G. Pachpatte and A. S.

Vatsala. Monotone method for equations describing transport phenomena in a Banach space. Applied Mathematics and Computation, 9(3):165–174, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Papamarkos:1988:OAR

[PVM88] N. Papamarkos, G. Vachtsevanos, and B. Mertzios. On the optimum approximation of real rational functions via linear programming. Applied Mathematics and Computation, 26 (4):267–287, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Pinelis:1994:TUF

[PY94] I. Pinelis and S. Yakowitz. The time until the final zero crossing of random sums with application to nonparametric bandit theory. Applied Mathematics and Computation, 63(2–3):235–263, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ping:1994:MPB

[PZ94] He Ping and Chen Zheng. A method putting bar materials to rational use and a computer program. Applied Mathematics and Computation, 63(1):59–73, June 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Primak:1997:SVI

[PZ97] M. E. Primak and A. Zeleke. Set variational inequalities. Applied Mathematics and Computation, 85(2–3):165–180, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
85&issue=2-3&aid=9600133.

Piccolomini:1999:CGR

[PZ99] Elena Loli Piccolomini and Fabiana Zama. The conjugate gradient regularization method in Computed Tomography problems. Applied Mathematics and Computation, 102(1):87–99, July 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/1/
6233.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=102&
issue=1&aid=6233.

Ping:1999:ANN

[PZZy99] He Ping, Chen Zheng, and Lu Zhi-yong. Artificial neural networks model of share analysis specialist system. Applied Mathematics and Computation, 106(2–3):277–283, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/31/
article.pdf.

Qu:1998:IEB

[QA98] Ruibin Qu and Ravi P. Agarwal. Improved error bounds for freezing solutions of linear boundary value problems. Applied Mathematics and Computation, 94(2–3):97–112, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/2-3/
6092.pdf;   http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=94&issue=
                                        2-3&aid=6092.

Qiang:1998:CCM

[QC98] J. Qiang and W. H. Choe. A combination of characteristics and Monte Carlo methods for etch profile simulation. Applied Mathematics and Computation, 91(2–3):179–195, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/91/2-3/
                                        6017.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=91&issue=
                                        2-3&aid=6017.

Quinnez:1978:IMI

[QHN78] S. Quinnez, P. Hillion, and G. Nurdin. Iterative method for invariant imbedding. Applied Mathematics and Computation, 4(??):213–337, 1978. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rajendran:1992:SUT

[Raj92] K. Rajendran. Simulation of unsteady turbulent flows by quasisteady models. Applied Mathematics and Computation, 48(1): 1–11, March 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rajendran:1994:CSU

[Raj94] K. Rajendran. Computer simulation of unsteady turbulent boundary layer flow with the effect of free stream velocity. Applied Mathematics and Computation, 63(2–3):151–162, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rajendran:1997:SUT

[Raj97] K. Rajendran. Simulation of unsteady turbulent flows with the effect of fluctuating external velocity by using quasi-steady models. Applied Mathematics and Computation, 87(2–3):95– 110, December 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=2-3&aid=9600221.

Razzaghi:1995:IND

[RAL95] M. Razzaghi, A. Arabshahi, and S. D. Lin. Identification of nonlinear differential equations via Fourier series operational matrix for repeated integration. Applied Mathematics and Computation, 68(2–3): 189–198, March 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=2-3&aid=9400093.

Razzaghi:1997:ALD

[RAL97] Mohsen Razzaghi, Abdollah Arabshahi, and Shin-Feng D. Lin. Analysis of linear distributed parameter systems via double Fourier series. Applied Mathematics and Computation, 87(2–3):205–215, December 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        87&issue=2-3&aid=9600284.

Ramachandran:1990:ECC

[Ram90] M. P. Ramachandran. End conditions for cubic spline interpolation. Applied Mathematics and Computation, 40(2 (part II)):105–116, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ramos:1997:LMRa

[Ram97a] J. I. Ramos. Linearization methods for reaction-diffusion equations: 1-D problems. Applied Mathematics and Computation, 88(2–3):199–224, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        88&issue=2-3&aid=9600328.

Ramos:1997:LMRb

[Ram97b] J. I. Ramos. Linearization methods for reaction-diffusion equations: Multidimensional problems. Applied Mathematics and Computation, 88(2–3):225– 254, December 30, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600327.

Ramos:1998:ICL

[Ram98a] J. I. Ramos. Implicit, compact, linearized θ-methods with factorization for multidimensional reaction-diffusion equations. Applied Mathematics and Computation, 94(1):17–43, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/1/6074.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
1&aid=6074.

Ramos:1998:MIL

[Ram98b] J. I. Ramos. Maps of implicit, linearized θ-methods for the logistic differential equation. Applied Mathematics and Computation, 94(1):1–15, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/1/6073.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.

                                        cgi?year=1998&volume=94&issue=
                                        1&aid=6073.

Ramos:1998:PTL

[Ram98c] J. I. Ramos. A piecewise timelinearized method for the logistic differential equation. Applied Mathematics and Computation, 93(2–3):139–148, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/93/2-3/
                                        6052.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=93&issue=
                                        2-3&aid=6052.

Ramos:1998:UBC

[Ram98d] J. I. Ramos. Upstream boundary conditions for flows in porous channels. Applied Mathematics and Computation, 93 (2–3):149–154, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/93/2-3/
                                        6058.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=93&issue=
                                        2-3&aid=6058.

Ramos:1999:AAC

[Ram99a] J. I. Ramos. Asymptotic analysis of compound liquid jets at low Reynolds numbers. Applied Mathematics and Computation, 100(2–3):223–240, May 1, 1999. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6206.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6206.

Ramos:1999:LFT

[Ram99b] J. I. Ramos. Linearized factorization techniques for multidimensional reaction–diffusion equations. Applied Mathematics and Computation, 100 (2–3):201–222, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6207.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6207.

Ramos:1999:LMO

[Ram99c] J. I. Ramos. Linearized methods for ordinary differential equations. Applied Mathematics and Computation, 104(2– 3):109–129, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/2-
3/6278.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
104&issue=2-3&aid=6278.

Ramos:1999:DME

[Ram99d] J. I. Ramos. On diffusive methods and exponentially fitted techniques. Applied Mathematics and Computation, 103 (1):69–96, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/103/1/
                                        6263.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=103&
                                        issue=1&aid=6263.

Raphael:1992:RTT

[Rap92] Christopher Raphael. Radiation therapy treatment planning: an l^{2} approach. Applied Mathematics and Computation, 52(2–3): 251–277, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ravn:1987:FMP

[Rav87] Hans F. Ravn. A forward maximum principle algorithm with decision horizon results. Applied Mathematics and Computation, 24(1):65–75, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Restrepo:1985:VML

[RB85] Pedro J. Restrepo and Rafael L. Bras. A view of maximumlikelihood estimation with large conceptual hydrologic models.

Applied Mathematics and Computation, 17(4):375–403, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). System identification and parameter estimation of environmental systems.

Ruxton:1995:JTS

[RB95] David J. W. Ruxton and David J. Bell. Junction times in singular optimal control. Applied Mathematics and Computation, 70(2–3):143–154, July 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400115.

Dynamics and control (Vienna, 1993).

Rach:1992:DES

[RBA92] R. Rach, A. Baghdasarian, and G. Adomian. Differential equations with singular coefficients. Applied Mathematics and Computation, 47(2–3):179–184, ???? 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ravindran:1996:MBC

[RBPK96] Varadarajan Ravindran, Badri N.

Badriyha, Massoud Pirbazari, and Sung-Hyun Kim. Modeling of bioactive carbon adsorbers: A hybrid weighted residual-finite difference numerical technique. Applied Mathematics and Computation, 76(2– 3):99–131, May 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        76&issue=2-3&aid=9500151.

Reich:1994:CCV

[RBR94] R. M. Reich, C. D. Bonham, and K. K. Remington. Comparison of classical vs. inverse sampling for estimating plant cover using Monte Carlo simulations. Applied Mathematics and Computation, 64(2–3): 171–??, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rogers:1981:CPM

[RC81] Thomas D. Rogers and Bruce L. Clarke. A continuous planar map with many periodic points. Applied Mathematics and Computation, 8(1):17–33, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ranasinghe:1991:SBE

[RC91] A. I. Ranasinghe and M.-H. Chang. Solution of the Burgers equation on semi-infinite and finite intervals via a stream function. Applied Mathematics and Computation, 41(2 (part II)): 145–158, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rasulov:1999:ENM

[RC99] Mahir Rasulov and Erhan Coskun. An efficient numerical method for solving the Korteweg–de Vries equation in a class of discontinuous functions. Applied Mathematics and Computation, 102 (2–3):139–154, July 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/2-
3/6247.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
102&issue=2-3&aid=6247.

Rashid:1994:TLF

[RCG94] Abdur Rashid, Weiming Cao, and Benyu Guo. Three level Fourier spectral approximations for fluid flow with low Mach number. Applied Mathematics and Computation, 63(2–3): 131–149, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reddy:1990:NIM

[Red90a] Y. N. Reddy. A numerical integration method for solving singular perturbation problems. Applied Mathematics and Computation, 37(2 (part II)):83–95, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reddy:1990:NTC

[Red90b] Y. N. Reddy. Numerical techniques for a class of singular perturbation problems. Applied Mathematics and Computation, 35(1 (part I)):1–11, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reddy:1991:NNI

[Red91] Y. N. Reddy. A note on the numerical-integration method for solving singular perturbation problems. Applied Mathematics and Computation, 43(2 (part II)):175–180, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reed:1987:OPM

[Ree87] William J. Reed. Optimal preventive maintenance, protection, and replacement of a revenueearning asset. Applied Mathematics and Computation, 24 (3):241–261, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Regan:1994:LHW

[Reg94] Marian P. Regan. Linear halfwave rectification of modulated sinusoids. Applied Mathematics and Computation, 62(1):61–79, April 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Regan:1996:HWL

[Reg96] Marian P. Regan. Half-wave linear rectification of a frequency modulated sinusoid. Applied Mathematics and Computation, 79(2–3):137–162, October 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500249.

Rivera-Gallego:1998:GAC

[RG98]  Wilson Rivera-Gallego.  A

genetic algorithm for circulant Euclidean distance matrices. Applied Mathematics and Computation, 97(2–3): 197–208, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6150.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6150.

Ramos:1997:NFD

[RGL97a] J. I. Ramos and C. M. García-López. Nonstandard finite difference equations for ODEs and 1-D PDEs based on piecewise linearization. Applied Mathematics and Computation, 86 (1):11–36, September 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1997/86/1/9600146.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.

                                        cgi?year=1997&volume=86&issue=
                                        1&aid=9600146.

Ramos:1997:PLM

[RGL97b] J. I. Ramos and C. M. García-López. Piecewise-linearized methods for initial-value problems. Applied Mathematics and Computation, 82(2–3):273– 302, March 15, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        82&issue=2-3&aid=9600040.

Ramos:1998:IBC

[RGL98] J. I. Ramos and C. M. García-López. Intermediate boundary conditions in operatorsplitting techniques and linearization methods. Applied Mathematics and Computation, 94(2–3):113–136, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/94/2-3/
                                        6094.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=94&issue=
                                        2-3&aid=6094.

Rinaldi:1998:LDC

[Rin98]  Sergio Rinaldi.  Love dy-

namics: The case of linear couples. Applied Mathematics and Computation, 95(2–3): 181–192, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6113.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6113.

Rao:1980:RHH

[RJY80] K. R. Rao, A. Jalali, and P. Yip. Rationalized Hadamard-Haar transform. Applied Mathematics and Computation, 6(3): 263–281, April 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rand:1987:DDE

[RK87] R. H. Rand and W. L. Keith.

Determinacy of degenerate equilibria with linear part x′ = y and y′ = 0 using Macsyma. Applied Mathematics and Computation, 21(1):1–19, January 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Reuster:1993:TDC

[RK93] D. Reuster and M. Kaye. A technique for developing complete orthonormal basis sets using general solutions of bessel’s differential equation. Applied Mathematics and Computation, 55(2–3):255–264, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rubbert:1982:PCS

[RL82] P. E. Rubbert and K. D. Lee. Patched coordinate systems. Applied Mathematics and Computation, 10/11(1):235–252, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Reithmeier:1996:RCC

[RL96] E. Reithmeier and G. Leitmann. Robust constrained control for vibration suppression of mismatched systems. Applied Mathematics and Computation, 78(2– 3):245–257, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        78&issue=2-3&aid=9600012.

Rogers:1986:SPI

[RMR86] C. Rogers, C. McInnis, and H. Rasmussen. On a Stefan problem involving a spherical region: application of a Bergmantype expansion. Applied Mathematics and Computation, 18 (2):143–154, ???? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Roache:1982:SMS

[Roa82] Patrick J. Roache. Semidirect/marching solutions and elliptic grid generation. Applied Mathematics and Computation, 10/11(1):729–737, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Roberts:1982:ATG

[Rob82] A. Roberts. Automatic topology generation and generalised B spline mapping. Applied Mathematics and Computation, 10/11 (1):465–486, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Rolle:1998:MPP

[Rol98] Jean-Daniel Rolle. A model for perturbed production or measurement processes involving compound normal distributions. Applied Mathematics and Computation, 96(1):75–88, October 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/1/6114.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
1&aid=6114.

Rosen:1993:MM

[Ros93] Robert Rosen. On models and modeling. Applied Mathematics and Computation, 56(2–3): 359–372, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ridgley:1997:MDS

[RPT97] Mark A. Ridgley, David C. Penn, and Liem Tran. Multicriterion decision support for a conflict over stream diversion and land-water reallocation in hawaii. Applied Mathematics and Computation, 83(2–3): 153–172, May 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=2-3&aid=9600169.

Rossler:1993:EA

[RR93] Otto E. Rossler and Jonas O. Rossler. The endo approach. Applied Mathematics and Computation, 56(2–3):281–??, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Riganti:1998:SPT

[RR98] R. Riganti and A. Rossani. Singular perturbation techniques in the study of a dynamical system arising from the kinetic theory of atoms and photons. Applied Mathematics and Computation, 96(1):47–63, October 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/1/6117.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.

cgi?year=1998&volume=96&issue=
1&aid=6117.

Ratschek:1999:ECS

[RR99] Helmut Ratschek and Jon Rokne. Exact computation of the sign of a finite sum. Applied Mathematics and Computation, 99(2–3):99–127, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6194.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6194.

Razzaghi:1990:SCI

[RRA90] Mohsen Razzaghi, Mehdi Razzaghi, and Abdollah Arabshahi. Solutions of convolution integral and Fredholm integral equations via double Fourier series. Applied Mathematics and Computation, 40(3):215–224, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rodellar:1995:ACU

[RRB95] J. Rodellar, E. P. Ryan, and A. H. Barbat. Adaptive control of uncertain coupled mechanical systems with application to base-isolated structures. Applied Mathematics and Computation, 70(2–3):299–314, July 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/

                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        70&issue=2-3&aid=9400110.

Dynamics and control (Vienna, 1993).

Roache:1986:ASE

[RS86] Patrick J. Roache and Stanly Steinberg. Application of a single-equation MG-FAS solver to elliptic grid generation equations (subgrid and supergrid coefficient generation). Applied Mathematics and Computation, 19(??):283–292, 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rao:1988:PLF

[RS88] M. Rama Mohana Rao and S. Sivasundaram. Perturbing Liapunov functions and Lagrange stability in terms of two measures. Applied Mathematics and Computation, 28(4 (part II)): 329–336, December 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rassias:1992:OPC

[RS92] Themistocles M. Rassias and H. M. Srivastava. The orthogonality property of the classical Laguerre polynomials. Applied Mathematics and Computation, 50(2–3):167–173, August 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rao:1989:ABS

[RSS89] M. Rama Mohana Rao, S. Sathananthan, and S. Sivasundaram. Asymptotic behavior of solutions of impulsive integrodifferential systems. Applied Mathematics and Computation, 34(3): 195–211, December 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rehbock:1992:EPF

[RTJL92] V. Rehbock, K. L. Teo, L. S.

Jennings, and C. S. Lee. An exact penalty function approach to all-time-step constrained discrete-time optimal control problems. Applied Mathematics and Computation, 49 (2–3):215–230, June 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ruan:1992:ASV

[Rua92] Shi Gui Ruan. Asymptotic stability for Volterra integrodifferential systems. Applied Mathematics and Computation, 52 (2–3):207–222, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ruggiero:1993:PPV

[Rug93] Valeria Ruggiero. Polynomial preconditioning on vector computers. Applied Mathematics and Computation, 59(2–3):131– 150, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Rajagopal:1984:NSP

[RUS^{+}84] A. K. Rajagopal, V. R. R. Uppuluri, David S. Scott, S. Sitharama Iyengar, and Mohan Yellayi. New structural properties of strings generated by leading digits of 2^{n}. Applied Mathematics and Computation, 14(3):221–244, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Russell:1993:FVD

[RW93] David L. Russell and Luther W. White. Formulation and validation of dynamical models for narrow plate motion. Applied Mathematics and Computation, 58(2–3):103–141, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ryan:1996:USC

[Rya96] E. P. Ryan. A universal servomechanism for a class of nonlinear planar systems. Applied Mathematics and Computation, 78(2–3):229–243, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        78&issue=2-3&aid=9600011.

Sheng:1994:NVP

[SA94] Qin Sheng and Ravi P. Agarwal. Nonlinear variation of parameter methods for summary difference equations in several independent variables. Applied Mathematics and Computation, 61(1):39–60, March 01, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shawagfeh:1996:NPA

[SA96] N. T. Shawagfeh and G. Adomian. Non-perturbative analytical solution of the general Lotka–Volterra three-species system. Applied Mathematics and Computation, 76(2–3): 251–266, May 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=2-3&aid=9500162.

Solomonovich:1997:DEM

[SAF^{+}97] M. Solomonovich, L. P. Apedaile, H. I. Freedman, A. H. Gebremedihen, S. G. M. Schilizzi, and L. Belostotski. A dynamical economic model of sustainable agriculture and the ecosphere. Applied Mathematics and Computation, 84(2–3): 221–246, July 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600088.

Sage:1991:IIC

[Sag91] Andrew P. Sage. Invariant imbedding in control, estimation, and system identification. Applied Mathematics and Computation, 45(2 (part II)):99–119, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Said:1988:IP

[Sai88] H. B. Said. C^{1} interpolation on a polygon. Applied Mathematics and Computation, 27(3 (part I)): 217–229, September 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saied:1996:SSF

[Sai96] Effat A. Saied. On the similarity solutions for the free Kramer equation. I. Applied Mathematics and Computation, 74(1): 59–63, January 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        74&issue=1&aid=9500088.

Saied:1999:NCS

[Sai99]  Effat A. Saied.  The non-

classical solution of the inhomogeneous non-linear diffusion equation. Applied Mathematics and Computation, 98(2– 3):103–108, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/98/2-3/
                                        6161.pdf;   http://www.elsevier.

com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6161.

Soubeh:1996:CFR

[SAL96] Mohammed Soubeh and Mhd. Jamal Al-Laban. Context-free relations and their characteristics. Applied Mathematics and Computation, 79(2–3):163– 172, October 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500252.

Sanchez:1980:CPS

[Sán80] David A. Sánchez. Computing periodic solutions of Riccati differential equations. Applied Mathematics and Computation, 6(4):283–287, January 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sarper:1993:EAN

[Sar93] Huseyin Sarper. Evaluation of the accuracy of Naslund’s approximation for managerial decision making under uncertainty. Applied Mathematics and Computation, 55(1):73–87, April 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sasagawa:1983:ACN

[Sas83] T. Sasagawa. Analytical construction of a negative semidefinite solution of algebraic Riccati equations and its application. Applied Mathematics and Computation, 13(1–2):37– 54, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sasagawa:1985:ASO

[Sas85]  T. Sasagawa.  Analysis of

the second order matrix Riccati equations. Applied Mathematics and Computation, 16 (2):141–152, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sasagawa:1988:ETD

[Sas88] T. Sasagawa. Existence theorem for the difference Riccati equations. Applied Mathematics and Computation, 26(2 (part I)):89–103, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sasagawa:1995:CCF

[Sas95a] T. Sasagawa. Constructions of C^{∞}-class functions having specified asymptotic properties. Applied Mathematics and Computation, 73(2–3):125– 132, December 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=2-3&aid=9400244.

Sasagawa:1995:IPA

[Sas95b] T. Sasagawa. An infinite product arising from an elementary geometric problem and the estimate of its value. Applied Mathematics and Computation, 73(2– 3):271–279, December 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=2-3&aid=9500067.

Saupe:1983:GBP

[Sau83] Dietmar Saupe. Global bifurcation of periodic solutions to some autonomous differential delay equations. Applied Mathematics and Computation, 13(1– 2):185–211, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Savageau:1993:FMR

[Sav93] Michael A. Savageau. Finding multiple roots of nonlinear algebraic equations using S-system methodology. Applied Mathematics and Computation, 55(2– 3):187–199, May 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saltzman:1982:AGV

[SB82] Jeffrey Saltzman and Jeremiah Brackbill. Applications and generalizations of variational methods for generating adaptive meshes. Applied Mathematics and Computation, 10/11 (1):865–884, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Suleiman:1992:UNI

[SB92] Mohamed Suleiman and Saiman Baok. Using nonconvergence of iteration to partition ODEs. Applied Mathematics and Computation, 49(2–3):111–140, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Suleiman:1996:VCO

[SBH96] Mohamed Bin Suleiman, Saiman Baok, and George Hall. Varying the componentwise order of the multistep methods in solving ODEs and its absolute stability. Applied Mathematics and Computation, 74(2–3):209– 222, February 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500093.

Su:1996:QED

[SC96] Ching-Tzong Su and Feng Cheng Chang. Quick evaluation of determinants. Applied Mathematics and Computation, 75(2–3): 117–118, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500106.

Scandrett:1989:CSI

[Sca89]  C. Scandrett.  Comparison

of several iterative techniques in the solution of symmetric banded equations on a twopipe Cyber 205. Applied Mathematics and Computation, 34 (2 (part II)):95–112, November 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Schulz:1991:SSM

[Sch91a]  Lech Schulz.  Set-spectrum

method in structure comparison and the concept of topological kinship. Applied Mathematics and Computation, 41(1 (part I)):1–20, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Schumitzky:1991:NEA

[Sch91b] Alan Schumitzky. Nonparametric EM algorithms for estimating prior distributions. Applied Mathematics and Computation, 45(2 (part II)):143–157, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Schmid:1992:AAC

[Sch92] Ulrich Schmid. On the asymptotics of the average CRI length of the slotted ALOHA collision resolution algorithm. Applied Mathematics and Computation, 48(1):59–70, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Scitovski:1997:API

[Sci97]  Rudolf Scitovski.  Analysis

of a parameter identification problem. Applied Mathematics and Computation, 82(1): 39–55, March 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
82&issue=1&aid=9600022.

Scott:1980:GFO

[Sco80] Edward J. Scott. Green’s functions in an octant for mixed problems relative to the Laplace equation. I. Applied Mathematics and Computation, 7(4): 321–339, ???? 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Scott:1981:GFO

[Sco81] Edward J. Scott. Green’s functions in an octant for mixed problems relative to the Laplace equation. II. Applied Mathematics and Computation, 8(4): 293–301, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sacco:1996:PAL

[SEG96] Riccardo Sacco, Danilo Erricolo, and Emilio Gatti. A perturbation approach for low frequency noise in junction field effect transistors. Applied Mathematics and Computation, 74(2–3):161– 190, February 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500091.

Selvakumar:1994:CMS

[Sel94] K. Selvakumar. A computational method for solving singular perturbation problems using exponentially fitted finite difference schemes. Applied Mathematics and Computation, 66(2– 3):277–292, ???? 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Selvakumar:1995:CPS

[Sel95] K. Selvakumar. A computational procedure for solving a chemical flow reactor problem using shooting method. Applied Mathematics and Computation, 68(1):27–40, March 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=1&aid=9400082.

Semper:1992:NCO

[Sem92] Bill Semper. A note on the computation of the outflow derivative for singular perturbation problems. Applied Mathematics and Computation, 51(2–3): 135–141, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sen:1986:CAL

[Sen86] S. Sen. A class of algorithms for large scale nonlinear minimax optimization. Applied Mathematics and Computation, 18(4): 355–361, ???? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sengupta:1990:JIS

[Sen90] A. Sengupta. Jordan invariant subspaces via algebraic generalized inverse. Applied Mathematics and Computation, 38(2 (part II)):75–85, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sengupta:1993:NRC

[Sen93] A. Sengupta. A note on a reduction of Cauchy singular integral equation to Fredholm equation in L_{p}. Applied Mathematics and Computation, 56(1):97–100, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Serbin:1979:RAT

[Ser79] Steven M. Serbin. Rational approximations of trigonometric matrices with application to second-order systems of differential equations. Applied Mathematics and Computation, 5(1): 75–92, ???? 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sever:1999:CMH

[Sev99a]  Ali Sever.  A computa-

tional method for hyperbolic inverse problem. Applied Mathematics and Computation, 101(1):5–12, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/101/1/
6224.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=101&
issue=1&aid=6224.

Sever:1999:UIB

[Sev99b] Ali Sever. On uniqueness for an inverse boundary value problem in electrical prospection. Applied Mathematics and Computation, 106(2–3):103–115, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/17/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/17/
article.pdf.

Seydel:1981:NCP

[Sey81] R. Seydel. Numerical computation of periodic orbits that bifurcate from stationary solutions of ordinary differential equations. Applied Mathematics and Computation, 9(4):257–271, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1987:AMD

[SGB87] Ferenc Szidarovszky, Mark Gershon, and Andras Bárdossy. Application of multiobjective dynamic programming to regional natural resource management. Applied Mathematics and Computation, 24(4):281–301, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sen:1992:DFP

[SH92] S. K. Sen and Gary W. Howell. Direct fail-proof triangularization algorithms for AX + XB = C with error-free and parallel implementations. Applied Mathematics and Computation, 50(2– 3):255–278, August 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shabahang:1987:MWS

[Sha87] R. Shabahang. Mean wave solution: bounded medium— discrete scatterers. Applied Mathematics and Computation, 21(2):111–123, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shalaby:1989:WUL

[Sha89a] Mohamed Ahmed Shalaby.

Weighted and unweighted leastsquares solutions of linear matrix equations. Applied Mathematics and Computation, 34(1 (part I)):61–71, November 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shampine:1989:PIS

[Sha89b] L. F. Shampine. Practical improvements to SIMPR codes for stiff ODEs. Applied Mathematics and Computation, 31(Special issue):328–345, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Sharp:1990:SDEa

[Sha90a] Keith P. Sharp. Stochastic differential equations in finance. Applied Mathematics and Computation, 37(2 (part II)):131– 148, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sharp:1990:SDEb

[Sha90b] Keith P. Sharp. Stochastic differential equations in finance. Applied Mathematics and Computation, 39(3 (Supplement)): 207s–224s, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shalaby:1994:MER

[Sha94] M. A. Shalaby. Minimal eigenvalue of a real symmetric positive definite Toeplitz matrix. Applied Mathematics and Computation, 64(2–3):93–99, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shary:1997:CSS

[Sha97] Sergey P. Shary. Controllable solution set to interval static systems. Applied Mathematics and Computation, 86(2–3): 185–196, October 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        86&issue=2-3&aid=9600181.

Shetty:1990:ASL

[She90] Bala Shetty. Approximate solutions to large scale capacitated facility location problems. Applied Mathematics and Computation, 39(2 (part II)):159–175, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shen:1996:NSD

[She96] J. H. Shen. The nonoscillatory solutions of delay differential equations with impulses. Applied Mathematics and Computation, 77(2–3):153–165, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shieh:1982:TDG

[Shi82] C. F. Shieh. Three-dimensional grid generation using Poisson equations. Applied Mathematics and Computation, 10/11(1): 687–694, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Shivaji:1989:RSC

[Shi89] Ratnasingham Shivaji. Remark on a sufficient condition for multiplicity in a model in chemical reactor theory. Applied Mathematics and Computation, 29(2 (part II)):99–105, January 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shimizu:1993:BAS

[Shi93] H. Shimizu. Biological autonomy: The self-creation of constraints. Applied Mathematics and Computation, 56(2–3): 177–??, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Stokes:1994:PMM

[SHR94] M. L. Stokes, D. H. Huddleston, and M. G. Remotigue. A practical model for multidisciplinary analysis data and algorithm abstraction. Applied Mathematics and Computation, 65(1– 3):111–123, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shen:1997:IME

[SHW97] Zuhe Shen, Zhen Yu Huang, and M. A. Wolfe. An interval maximum entropy method for a discrete minimax problem. Applied Mathematics and Computation, 87(1):49–68, November 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=1&aid=9600220.

Shyy:1987:AGM

[Shy87] Wei Shyy. An adaptive grid method for Navier–Stokes flow computation. Applied Mathematics and Computation, 21(3): 201–219, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Suleiman:1996:POD

[SIA96] Mohamed B. Suleiman, Fudziah Bt. Ismail, and Kamel Ariffin B. M. Atan. Partitioning ordinary differential equations using Runge– Kutta methods. Applied Mathematics and Computation, 79 (2–3):289–309, October 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500247.

Simos:1992:EAS

[Sim92] T. E. Simos. An explicit almost P -stable two-step method with phase-lag of order infinity for the numerical integration of secondorder periodic initial-value problems. Applied Mathematics and Computation, 49(2–3):261–268, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Simos:1994:SNV

[Sim94]  T. E. Simos.  Some new

variable-step methods with minimal phase lag for the numerical integration of special second-order initial-value problem. Applied Mathematics and Computation, 64(1):65–72, August 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Simos:1997:MRK

[Sim97] T. E. Simos. Modified Runge– Kutta methods for the numerical solution of ODEs with oscillating solutions. Applied Mathematics and Computation, 84(2– 3):131–143, July 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600083.

Simos:1999:EEF

[Sim99a] T. E. Simos. Explicit exponentially fitted methods for the numerical solution of the Schrödinger equation. Applied Mathematics and Computation, 98(2–3):185–198, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/
6165.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6165.

Simos:1999:NFD

[Sim99b] T. E. Simos. A new finite difference scheme with minimal phase-lag for the numerical solution of the Schrödinger equation. Applied Mathematics and Computation, 106 (2–3):245–264, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/29/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/29/
article.pdf.

Sivaloganathan:1992:FEP

[Siv92] S. Sivaloganathan. A finiteelement, parameter stepping solution procedure for the steadystate Boussinesq equations. Applied Mathematics and Computation, 51(1):43–83, September 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Scitovski:1996:MSP

[SJ96] Rudolf Scitovski and Dragan Jukić. A method for solving the parameter identification problem for ordinary differential equations of the second order. Applied Mathematics and Computation, 74(2–3): 273–291, February 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=2-3&aid=9500098.

Swisshelm:1986:PCE

[SJK86] J. M. Swisshelm, G. M. Johnson, and S. P. Kumar. Parallel computation of Euler and Navier–Stokes flows. Applied Mathematics and Computation, 19(??):321–331, 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Schweitzer:1986:IAD

[SK86] Paul J. Schweitzer and Kyle W.

Kindle. An iterative aggregationdisaggregation algorithm for solving linear equations. Applied Mathematics and Computation, 18(4):313–353, ???? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Smokty:1995:AAS

[SK95] O. I. Smokty and K. Y. Kondratyev. The analytical approximation of the spatial frequency filtration of radiation fields for the space spectrophotometry problems. Applied Mathematics and Computation, 69(1):23–39, April 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=1&aid=9400097.

Sivaloganathan:1996:NSD

[SK96] S. Sivaloganathan and A. Karageorghis. Numerical solution of the 3-D buoyancy driven cavity problem by a pseudospectral method. Applied Mathematics and Computation, 75(1): 43–58, March 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=1&aid=9500102.

Sudarsan:1998:EAN

[SK98a] R. Sudarsan and S. Sathiya Keerthi. An efficient approach for the numerical simulation of multibody systems. Applied Mathematics and Computation, 92(2–3):195–218, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/2-3/
9710041.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
92&issue=2-3&aid=9710041.

Sudarsan:1998:NAS

[SK98b] R. Sudarsan and S. Sathiya Keerthi. Numerical approaches for solution of differential equations on manifolds. Applied Mathematics and Computation, 92(2–3):153–193, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/2-3/
9710065.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
92&issue=2-3&aid=9710065.

Sahin:1995:MRN

[SKB95] Ahmet Z. Sahin, Davut Kavranoglu, and Maamar Bettayeb. Model reduction in numerical heat transfer problems. Applied Mathematics and Computation, 69(2–3):209–225, May 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=2-3&aid=9400128.

Shang:1994:NSD

[SKCW94] H. M. Shang, Y. M. Kim, C. P. Chen, and T. S. Wang. Numerical studies of dropletturbulence interactions. Applied Mathematics and Computation, 65(1–3):63–78, September/October 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Skeel:1989:GEE

[Ske89] Robert D. Skeel. Global error estimation and the backward differentiation formulas. Applied Mathematics and Computation, 31(Special issue):197–208, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Sharan:1997:AMM

[SKG97] Maithili Sharan, E. J. Kansa, and Suman Gupta. Application of the multiquadric method for numerical solution of elliptic partial differential equations. Applied Mathematics and Computation, 84(2–3):275– 302, July 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        84&issue=2-3&aid=9600109.

Samimi:1983:GUC

[SL83] Mansour Samimi and V. Lakshmikantham. General uniqueness criteria for ordinary differential equations. Applied Mathematics and Computation, 12(1):77–88, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:1992:IMC

[SL92] Jing Xian Sun and Li Shan Liu.

Iterative method for coupled quasi-solutions of mixed monotone operator equations. Applied Mathematics and Computation, 52(2–3):301–308, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1996:AOB

[SL96] Ferenc Szidarovszky and Chia-Hung Lin. The alternating offer bargaining method under uncertainty. Applied Mathematics and Computation, 76(2–3): 133–141, May 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=2-3&aid=9500152.

Szidarovszky:1998:SBP

[SL98] Ferenc Szidarovszky and Shen Ling. A stochastic bargaining process and solution concept in the discrete case. Applied Mathematics and Computation, 92(2–3):219–227, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/2-3/
9710032.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
92&issue=2-3&aid=9710032.

Sun:1999:AQN

[SL99] Fangyu Sun and Xiangfang Li. On an accelerating quasi-Newton circular iteration. Applied Mathematics and Computation, 106(1):17–29, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.elsevier.nl/gej-ng/
                                        10/9/12/71/17/18/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej-ng/10/9/12/71/17/18/
                                        article.pdf.

Slater:1991:PBM

[Sla91] Paul B. Slater. Physical-based models of internal migration. Applied Mathematics and Computation, 43(1 (part I)):95–103, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Slater:1993:IMA

[Sla93a] Paul B. Slater. International migration and air travel: Global smoothing and estimation. Applied Mathematics and Computation, 53(2–3):225–234, February 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Slater:1993:WPD

[Sla93b] Paul B. Slater. World population distribution: Smoothed representations. Applied Mathematics and Computation, 53(2– 3):207–223, February 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Scott:1988:SEM

[SM88] Melvin R. Scott and Gary R.

Montry. Some experiments in multitasking on an ELXSI system 6400. Applied Mathematics and Computation, 26(2):135– 150, 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1994:LDP

[SM94] Ferenc Szidarovszky and Sándor Molnár. Learning in a dynamic producer-consumer market. Applied Mathematics and Computation, 62(2–3):223–233, May 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shen:1995:SDP

[SM95] Zhizhang Shen and Christian M.

Marston. A study of a dice problem. Applied Mathematics and Computation, 73(2–3):231– 247, December 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=2-3&aid=9500062.

Smith:1982:AGG

[Smi82] Robert E. Smith. Algebraic grid generation. Applied Mathematics and Computation, 10/11 (1):137–170, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Smith:1984:CPC

[Smi84a]  Tony E. Smith.  A choice

probability characterization of generalized extreme value models. Applied Mathematics and Computation, 14(1):35–62, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Smith:1984:SCC

[Smi84b] Tony E. Smith. A solution condition for complementarity problems: with an application to spatial price equilibrium. Applied Mathematics and Computation, 15(1):61–69, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Smith:1986:TET

[Smi86] Tony E. Smith. Two extension theorems for log-linear representations of positive probabilities. Applied Mathematics and Computation, 18(3):239– 255, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1991:PSL

[SMO91] Ferenc Szidarovszky, Sándor Molnár, and Koji Okuguchi. An N-person Stackelberg leaderleader model. Applied Mathematics and Computation, 46(3): 221–232, December 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1995:PTC

[SMY95] Ferenc Szidarovszky, Sándor Molnár, and Jerome Yen. On the price trajectory control of a discrete dynamic producerconsumer market. Applied Mathematics and Computation, 73(2–3):249–256, December 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
73&issue=2-3&aid=9500063.

Seda:1992:PBV

[ŠNG92] V. Šeda, Juan J. Nieto, and M. Gera. Periodic boundary value problems for nonlinear higher order ordinary differential equations. Applied Mathematics and Computation, 48 (1):71–82, March 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1990:DOM

[SO90] Ferenc Szidarovszky and Koji Okuguchi. Dynamic oligopoly: Models with incomplete information. Applied Mathematics and Computation, 38(2 (part II)):161–177, July 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sugisaka:1995:CTC

[SO95] Masanori Sugisaka and Takemasu Okada. Continuous-time Chandrasekhar smoothers in detection and tracking. Applied Mathematics and Computation, 69(1):123–136, April 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=1&aid=9400103.

Sochacki:1988:ABC

[Soc88] James Sochacki. Absorbing boundary conditions for the elastic wave equations. Applied Mathematics and Computation, 28(1):1–14, October 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soderlind:1989:MPS

[Söd89] Gustaf Söderlind. A multipurpose system for the numerical integration of ODEs. Applied Mathematics and Computation, 31(Special issue):346–360, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Soliman:1998:SCD

[Sol98] A. A. Soliman. Stability criteria of differential systems via Liapunov’s second method. Applied Mathematics and Computation, 92(2–3):143–152, June 15, 1998. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/2-3/
9710045.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
92&issue=2-3&aid=9710045.

Soliman:1999:SNV

[Sol99a] A. A. Soliman. On ϕ_{0}-stability for nonlinear Volterra integrodifferential systems. Applied Mathematics and Computation, 100(1):71–83, April 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/1/
6199.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=100&
issue=1&aid=6199.

Solis:1999:DLS

[Sol99b] Francisco J. Solis. Dimension and local structures of attracting manifolds of smooth dynamical systems. Applied Mathematics and Computation, 100 (2–3):169–175, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6204.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6204.

Soni:1993:EGG

[Son93] Bharat K. Soni. Elliptic grid generation system: control functions revisited. I. Applied Mathematics and Computation, 59 (2–3):151–163, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Song:1996:CWR

[Son96] Yongzhong Song. Chaotic waveform relaxation methods for dynamical systems. Applied Mathematics and Computation, 78 (1):83–100, August 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        78&issue=1&aid=9500250.

Sorenson:1982:GGE

[Sor82] Reese L. Sorenson. Grid generation by elliptic partial differential equations for a tri-element augmentor-wing airfoil. Applied Mathematics and Computation, 10/11(1):653–665, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Sorooshian:1985:SHS

[Sor85] Soroosh Sorooshian. Synthesis of hydrologic and system sciences in the development of rainfall-runoff models. Applied Mathematics and Computation, 17(4):279–298, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). System identification and parameter estimation of environmental systems.

Soriano:1993:GMP

[Sor93] J. M. Soriano. Global minimum point of a convex function. Applied Mathematics and Computation, 55(2–3):213–218, May 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Soriano:1994:EPC

[Sor94] J. M. Soriano. Extremum points of a convex function. Applied Mathematics and Computation, 66(2–3):261–266, ???? 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [Sor97a].

Soriano:1996:EZP

[Sor96] J. M. Soriano. On the existence of zero points. Applied Mathematics and Computation, 79 (1):99–104, September 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=1&aid=9500246. See

comment [Di 98].

Soriano:1997:EEP

[Sor97a] J. M. Soriano. Erratum: “Extremum points of a convex function” [Appl. Math. Comput. 66 (1994), no. 2–3, 261–266]. Applied Mathematics and Computation, 84(1):95, June 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=1&aid=9700210. See

[Sor94].

Soriano:1997:NZM

[Sor97b] J. M. Soriano. On the number of zeros of a mapping. Applied Mathematics and Computation, 88(2–3):287–291, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600336.

Soriano:1999:EZB

[Sor99] J. M. Soriano. Existence of zeros for bounded perturbations of proper mappings. Applied Mathematics and Computation, 99(2–3):255–259, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6184.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6184.

Sosnowska:1988:GSB

[Sos88] Honorata Sosnowska. Global stability in bounded sets: application to a queueing model of a market in a shortage economy. Applied Mathematics and Computation, 27(2):155–165, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sosnowska:1991:CKW

[Sos91] Honorata Sosnowska. Controllability in the Kornai-Weibull queueing model of a market in a shortage economy. Applied Mathematics and Computation, 43(3):207–220, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szpiro:1993:PSC

[SP93] A. A. Szpiro and V. Protopopescu. Parabolic systems with competitive interactions and control on initial conditions. Applied Mathematics and Computation, 59(2–3):215– 245, December 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1994:SME

[SP94] Ferenc Szidarovszky and Olgierd A. Palusinski. A special matrix equation and its application in microelectronics. Applied Mathematics and Computation, 64(2–3):115–119, September 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1995:IMD

[SP95] Ferenc Szidarovszky and Olgierd Palusinski. An inverse method to determine parasitics of power interconnections in high speed electronics. Applied Mathematics and Computation, 73(2–3): 203–208, December 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        73&issue=2-3&aid=9400250.

Soyeux:1993:BQA

[SPD93] A. Soyeux, E. Parent, and L. Duckstein. A biattribute quality analysis of the food industry. Applied Mathematics and Computation, 58(2–3):199– ??, October 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Spitaleri:1990:MMG

[Spi90] R. M. Spitaleri. A multigrid method for grid generation with line-spacing control. Applied Mathematics and Computation, 40(2 (part II)):125–134, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shi:1988:OMS

[SR88] Shenghua Shi and Herschel Rabitz. An operator method for solving the determining equations for the group analysis of differential equations. Applied Mathematics and Computation, 25(2 (part II)):101–113, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sengupta:1990:ASL

[SR90] A. Sengupta and M. Rama Mohana Rao. On approximate solution of linear singular integral equations. Applied Mathematics and Computation, 36 (3):147–161, April 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Spitaleri:1997:MMG

[SR97] Rosa Maria Spitaleri and Vania Regolo. Multiblock multigrid grid generation algorithms: Overcoming multigrid anisotropy. Applied Mathematics and Computation, 84(2–3): 247–267, July 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600089.

Srivastava:1994:ORG

[Sri94] H. M. Srivastava. Orthogonality relations and generating functions for the generalized Bessel polynomials. Applied Mathematics and Computation, 61(2–3): 99–134, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Srivastava:1995:NEH

[Sri95] Rekha Srivastava. A note on the Epstein–Hubbell generalized elliptic-type integral. Applied Mathematics and Computation, 69(2–3):255–262, May 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        69&issue=2-3&aid=9400133.

Srivastava:1998:FSR

[Sri98] H. M. Srivastava. Further series representations for ζ(2n+1). Applied Mathematics and Computation, 97(1):1–15, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/97/1/6138.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=97&issue=
                                        1&aid=6138.

Schaffer:1986:MSM

[SS86] Steve Schaffer and Frank Stenger. Multigrid-sinc methods. Applied Mathematics and Computation, 19(1–4):311– 319, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Szidarovszky:1987:MOM

[SS87] Ferenc Szidarovszky and Károly Szenteleki. A multiobjective optimization model for wine production. Applied Mathematics and Computation, 22(2–3): 255–275, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sanchez:1991:DFD

[SS91] Paul J. Sanchez and Susan M.

Sanchez. Design of frequencydomain experiments for discretevalued factors. Applied Mathematics and Computation, 42(1): 1–21, 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sivasundaram:1992:AIA

[SS92] S. Sivasundaram and Yong Sun.

Application of interval analysis to impulsive differential equations. Applied Mathematics and Computation, 47(2–3):201–210, ???? 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shahruz:1994:ASL

[SS94] S. M. Shahruz and A. L.

Schwartz. An approximate solution for linear boundary-value problems with slowly varying coefficients. Applied Mathematics and Computation, 60(2–3): 285–298, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szafranski:1997:OTS

[SS97] Zdzisław Szafrański and Błażej Szmanda. Oscillation theorems for some nonlinear difference equations. Applied Mathematics and Computation, 83 (1):43–52, April 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=1&aid=9600045.

SalinasDeRomero:1998:SAF

[SS98a] Susana Salinas De Romero and H. M. Srivastava. Some applications of fractional calculus involving summation of infinite series. Applied Mathematics and Computation, 90 (2–3):129–142, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/
9700394.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=2-3&aid=9700394.

Shen:1998:GCA

[SS98b] Ling Shen and Ferenc Szidarovszky. A globally convergent algorithm for solving special utilization equations. Applied Mathematics and Computation, 90(1):53–60, March 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=1&aid=9600339.

Shubin:1982:TDG

[SSB82] G. R. Shubin, A. B. Stephens, and J. B. Bell. Threedimensional grid generation using biharmonics. Applied Mathematics and Computation, 10/11 (1):761–774, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Steiner:1995:CMA

[SST95] W. Steiner, A. Steindl, and H. Troger. Center manifold approach to the control of a tethered satellite system. Applied Mathematics and Computation, 70(2–3):315–327, July 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400111.

Dynamics and control (Vienna, 1993).

Sanz-Serna:1989:CAO

[SSV89] J. M. Sanz-Serna and J. G. Verwer. Convergence analysis of one-step schemes in the method of lines. Applied Mathematics and Computation, 31(Special issue):183–196, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Salemi:1999:SCM

[SSW99] F. Salemi, V. Salone, and Wang Wendi. Stability of a competition model with two-stage structure. Applied Mathematics and Computation, 99(2– 3):221–231, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6181.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6181.

Sun:1999:QNT

[SSY99] W. Sun, R. J. B. Sampaio, and J. Yuan. Quasi-Newton trust region algorithm for non-smooth least squares problems. Applied Mathematics and Computation, 105(2–3):183–194, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/2-
3/6322.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
105&issue=2-3&aid=6322.

Steindl:1996:HCT

[ST96] A. Steindl and H. Troger. Heteroclinic cycles in the threedimensional postbifurcation motion of O(2)-symmetric fluid conveying tubes. Applied Mathematics and Computation, 78(2– 3):269–277, September 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600014.

Stanimirovic:1999:LRG

[Sta99] Predrag S. Stanimirović. Limit representations of generalized inverses and related methods. Applied Mathematics and Computation, 103(1):51–68, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/103/1/
6270.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=103&
issue=1&aid=6270.

Sivaloganathan:1998:MPV

[STD98] S. Sivaloganathan, G. Tenti, and J. M. Drake. Mathematical pressure volume models of the cerebrospinal fluid. Applied Mathematics and Computation, 94(2–3):243–266, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/2-3/
6102.pdf;   http://www.elsevier.

                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=94&issue=
                                        2-3&aid=6102.

Steger:1982:ABC

[Ste82] Joseph L. Steger. On application of body conforming curvilinear grids for finite difference solution of external flow. Applied Mathematics and Computation, 10/11 (1):295–316, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Stewart:1984:BEC

[Ste84] Ian Stewart. Beyond elementary catastrophe theory. Applied Mathematics and Computation, 14(1):25–31, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Steit:1994:CSQ

[Ste94] H. A. H. Abou Steit. Cubic spline on quintic spline method for nonlinear boundary value problems. Applied Mathematics and Computation, 66(2–3): 267–275, ???? 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Stuben:1983:AMA

[Stü83] Klaus Stüben. Algebraic multigrid (AMG): experiences and comparisons. Applied Mathematics and Computation, 13 (3–4):419–451, November 1983.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Serrano:1990:REEa

[SU90a] Sergio E. Serrano and T. E.

Unny. Random evolution equations in hydrology. Applied Mathematics and Computation, 38(3):201–226, August 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Serrano:1990:REEb

[SU90b] Sergio E. Serrano and T. E.

Unny. Random evolution equations in hydrology. Applied Mathematics and Computation, 39(3 (Supplement)):97s–122s, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sugisaka:1995:ISI

[SU95] Masanori Sugisaka and Sueo Ueno. Introduction: Special issue on adaptive filtering and estimation. Applied Mathematics and Computation, 69 (1):1–2, April 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=1&aid=9400096.

Sugisaka:1983:DLL

[Sug83] Masanori Sugisaka. The design of on-line linear leastsquares estimators given covariance specifications via an imbedding method. Applied Mathematics and Computation, 13(1– 2):55–85, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sugisaka:1987:DLR

[Sug87] Masanori Sugisaka. The design of a linear regulator based on an integral-equation representation. Applied Mathematics and Computation, 23(2):163– 169, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sugisaka:1995:ACF

[Sug95] Masanori Sugisaka. Adaptive Chandrasekhar filter for linear discrete-time stationary stochastic systems. Applied Mathematics and Computation, 69(1): 137–145, April 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=1&aid=9400104.

Scitovski:1998:ASM

[SUJ98] R. Scitovski, Š. Ungar, and D. Jukić. Approximating surfaces by moving total least squares method. Applied Mathematics and Computation, 93 (2–3):219–232, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/2-3/
6084.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6084.

Sumita:1984:MLT

[Sum84] Ushio Sumita. The matrix Laguerre transform. Applied Mathematics and Computation, 15 (1):1–28, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sun:1998:NMQ

[Sun98] Wenyu Sun. Newton’s method and quasi-Newton-SQP method for general LC^{1} constrained optimization. Applied Mathematics and Computation, 92(1):69– 84, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.elsevier.
com/cas/tree/store/amc/sub/
1998/92/1/6035.pdf;   http:
//www.elsevier.com/cgi-bin/
cas/tree/store/amc/cas_sub/
browse/browse.cgi?year=1998&
volume=92&issue=1&aid=6035.

Shendge:1983:CRF

[SV83] G. R. Shendge and A. S. Vatsala.

Comparison results for first and second order boundary value problems at resonance. Applied Mathematics and Computation, 12(4):367–380, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shendge:1984:CRP

[SV84] G. R. Shendge and A. S. Vatsala. Comparison results for parabolic differential equations at resonance. Applied Mathematics and Computation, 15 (3):209–219, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saaty:1985:MBC

[SV85] Thomas L. Saaty and Luis G. Vargas. Modeling behavior in competition: The analytic hierarchy process. Applied Mathematics and Computation, 16 (??):49–92, 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Srinivasu:1992:QCN

[SV92a] P. D. N. Srinivasu and M. Venkatesulu. Quadratically convergent numerical schemes for nonstandard initial value problems. Applied Mathematics and Computation, 47(2–3):145–154, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Srinivasu:1992:SNO

[SV92b] P. D. N. Srinivasu and M. Venkatesulu. Solutions of nonstandard nth order initial value problems. Applied Mathematics and Computation, 47(2–3): 155–178, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saxena:1994:STP

[SV94] Anjula Saxena and Ezio Venturino. Solving two-point boundary value problems by means of deficient quartic splines. Applied Mathematics and Computation, 66(1):25– 40, November 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shampine:1979:AWR

[SW79a] Lawrence F. Shampine and Herman A. Watts. The art of writing a Runge–Kutta code. II. Applied Mathematics and Computation, 5(??):93–121, 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sidhu:1979:DRA

[SW79b] Gursharan S. Sidhu and Howard L.

Weinert. Dynamical recursive algorithms for lg-spline interpolation of EHB data. Applied Mathematics and Computation, 5(2):157–185, ???? 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shearer:1985:IFK

[SW85] J. M. Shearer and M. A.

Wolfe. An improved form of the Krawczyk-Moore algorithm. Applied Mathematics and Computation, 17(3):229–239, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

So:1989:NBV

[SW89] Joseph W.-H. So and Paul Waltman. A nonlinear boundary value problem arising from competition in the chemostat. Applied Mathematics and Computation, 32(2–3):169–183, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Mathematical biology.

Shen:1990:BNP

[SW90a] Zu He Shen and M. A. Wolfe. A ball Newton point algorithm for bounding zeros of analytic functions. Applied Mathematics and Computation, 36(1 (part I)):1–14, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shen:1990:IEU

[SW90b] Zu He Shen and M. A. Wolfe. On interval enclosures using slope arithmetic. Applied Mathematics and Computation, 39(1):89– 105, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Silliman:1991:GRF

[SW91] S. E. Silliman and A. L. Wright. Generation of random fields characterized by discrete regions of constant value. Applied Mathematics and Computation, 45 (3):293–311, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shaw:1992:ATG

[SW92a] J. A. Shaw and N. P. Weatherill. Automatic topology generation for multiblock grids. Applied Mathematics and Computation, 52(2–3):355–388, December 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Shen:1992:STN

[SW92b] Zuhe Shen and M. A. Wolfe.

Slope tests for Newton-Type methods. Applied Mathematics and Computation, 52(2–3): 403–416, December 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Sheng:1994:IAN

[SW94] Zhang Lian Sheng and M. A.

Wolfe. An interval algorithm for nondifferentiable global optimization. Applied Mathematics and Computation, 63(2–3): 101–122, July 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Saha:1996:STM

[SW96] Arindam Saha and Meghanad D.

Wagh. Solutions of two minmax recurrences in parallel processing with variable recombination overhead. Applied Mathematics and Computation, 76(2– 3):173–211, May 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=2-3&aid=9500158.

Sugisaka:1998:ICS

[SWL98] Masanori Sugisaka, Xin Wang, and Ju-Jung Lee. Intelligent control strategy for a mobile vehicle. Applied Mathematics and Computation, 91(1):91–98, April 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/1/6003.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
1&aid=6003.

Sworder:1991:HAC

[Swo91] D. D. Sworder. Hybrid adaptive control. Applied Mathematics and Computation, 45(2 (part II)):173–192, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1993:CDD

[SY93] Ferenc Szidarovszky and Jerome Yen. On the controllability of discrete dynamic oligopolies under adaptive expectations. Applied Mathematics and Computation, 56(1):49–57, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Song:1998:CRN

[SY98] Jingyan Song and Yeung Yam. Complex recurrent neural network for computing the inverse and pseudo-inverse of the complex matrix. Applied Mathematics and Computation, 93 (2–3):195–205, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/2-3/
6078.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6078.

Shih:1994:IGG

[SYS94] Ming H. Shih, Tzu-Yi Yu, and Bharat K. Soni. Interactive grid generation and NURBS applications. Applied Mathematics and Computation, 65(1–3):279–294, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Satravaha:1997:ALT

[SZ97] Pornchai Satravaha and Songping Zhu. An application of the LTDRM to transient diffusion problems with nonlinear material properties and nonlinear boundary conditions. Applied Mathematics and Computation, 87(2–3):127–160, December 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/

                                        browse.cgi?year=1997&volume=
                                        87&issue=2-3&aid=9600226.

Szidarovszky:1989:GST

[Szi89] Ferenc Szidarovszky. On a general scheme in the theory of conflicts. Applied Mathematics and Computation, 34(1 (part I)): 29–37, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1990:CNA

[Szi90] Ferenc Szidarovszky. On the convergence of nonstationary algorithms modeled by point-to-set maps. Applied Mathematics and Computation, 39(2 (part II)):107–110, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Szidarovszky:1997:NAO

[Szi97] F. Szidarovszky. A note on the alternating offer method. Applied Mathematics and Computation, 82(2–3):129–135, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.elsevier.com/cgi-bin/
                                        cas/tree/store/amc/cas_sub/
                                        browse/browse.cgi?year=1997&
                                        volume=82&issue=2-3&aid=9600026.

Szidarovszky:1998:BOD

[Szi98]  Ferenc Szidarovszky.  Bar-

gaining with offer dependent break-down probabilities. Applied Mathematics and Computation, 90(2–3):117–127, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/
9600354.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=2-3&aid=9600354.

Szidarovszky:1999:NCN

[Szi99] Ferenc Szidarovszky. A new Characterization of the nonsymmetric Nash solution. Applied Mathematics and Computation, 106(1):63–68, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/17/21/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/17/21/
article.pdf.

Tabatabai:1993:RET

[TA93] M. A. Tabatabai and I. K. Argyros. Robust estimation and testing for general nonlinear regression models. Applied Mathematics and Computation, 58(1):85– 101, September 15, 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tagliani:1998:ITS

[Tag98] Aldo Tagliani. Inverse twosided z transform and moment problem. Applied Mathematics and Computation, 95 (2–3):125–138, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6055.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6055.

Tagliani:1999:HMP

[Tag99] A. Tagliani. Hausdorff moment problem and maximum entropy: A unified approach. Applied Mathematics and Computation, 105(2–3):291–305, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/2-
3/6305.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
105&issue=2-3&aid=6305.

Takashima:1985:PER

[Tak85] Tsutomu Takashima. Polarization effect on radiative transfer in Chandrasekhar’s planetary problem. Applied Mathematics and Computation, 17 (3):185–227, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Takac:1998:BIT

[Tak98] Peter Takáč. Bifurcations to invariant 2-tori for the complex Ginzberg–Landau equation. Applied Mathematics and Computation, 89(1–3):241– 257, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781661.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Tauber:1980:TDS

[Tau80] Selmo Tauber. Time dependent solution of a pollution problem. Applied Mathematics and Computation, 6(4):359– 367, January 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tsai:1992:CAM

[TCS92] J. S. H. Tsai, C. M. Chen, and L. S. Shieh. A computeraided method for solvents and spectral factors of matrix polynomials. Applied Mathematics and Computation, 47(2–3): 211–235, February 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tecle:1994:IMP

[TDK94] Aregai Tecle, Lucien Duckstein, and Pekka Korhonen. Interactive, multiobjective programming for forest resources management. Applied Mathematics and Computation, 63(1):75–93, June 1994. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Tennakone:1990:SBC

[Ten90] K. Tennakone. Stability of the biomass-carbon dioxide equilibrium in the atmosphere: mathematical model. Applied Mathematics and Computation, 35 (2 (part II)):125–130, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tennakone:1991:KGA

[Ten91] K. Tennakone. The kinetics of the growth of an accidentally created chiral biomolecule in the racemic prebiotic medium. Applied Mathematics and Computation, 43(2):139–144, 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tesfatsion:1980:MSR

[Tes80] Leigh Tesfatsion. C^{3} modeling with symmetrical rationality. Applied Mathematics and Computation, 6(1):51–61, January 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tesfatsion:1991:I

[Tes91a] Leigh Tesfatsion. Introduction. Applied Mathematics and Computation, 45(2 (part II)):89–90, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tesfatsion:1991:WRKa

[Tes91b] Leigh Tesfatsion. Work by Robert Kalaba on automated sensitivity analysis. Applied Mathematics and Computation, 45(2 (part II)):121–133, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tesfatsion:1991:WRKb

[Tes91c] Leigh Tesfatsion. Work by Robert Kalaba on multicriteria estimation. Applied Mathematics and Computation, 45(2 (part II)):159–171, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tabata:1999:NIP

[TET99] Minoru Tabata, Nobuoki Eshima, and Ichiro Takagi. The nonlinear integro-partial differential equation describing the logistic growth of human population with migration. Applied Mathematics and Computation, 98(2–3):169–183, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/2-3/
6173.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
2-3&aid=6173.

Tabata:1998:CPS

[TETH98] Minoru Tabata, Nobuoki Eshima, Ichiro Takagi, and Takashi Hiroyama. The Cauchy problem for the system of equations describing migration motivated by regional economic disparity. Applied Mathematics and Computation, 94(1):45–64, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/1/6071.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
1&aid=6071.

Teo:1993:SFS

[TFM93] K. L. Teo, M. E. Fisher, and J. B. Moore. A suboptimal feedback stabilizing controller for a class of nonlinear regulator problems. Applied Mathematics and Computation, 59(1):1– 17, November 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tapia:1991:CML

[TGK91] M. A. Tapia, T. A. Guima, and A. Katbab. Calculus for a multivalued-logic algebraic system. Applied Mathematics and Computation, 42(3):255– 285, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Todorov:1994:EHG

[TH94] Pavel G. Todorov and John H. Hubbell. On the Epstein– Hubbell generalized elliptic-type integral. Applied Mathematics and Computation, 61(2–3):157– 161, March 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tibken:1995:SCU

[TH95] B. Tibken and E. P. Hofer.

Simulation of controlled uncertain nonlinear systems. Applied Mathematics and Computation, 70(2–3):329–338, July 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400112.

Dynamics and control (Vienna, 1993).

Tibken:1996:IAT

[TH96] B. Tibken and E. P. Hofer. Interval analysis as a tool for sensitivity analysis of a hemopoietic model. Applied Mathematics and Computation, 78(2–3):259– 267, September 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600013.

Thames:1982:GTD

[Tha82] Frank C. Thames. Generation of three-dimensional boundaryfitted curvilinear coordinate systems for wing/wing-tip geometries using the elliptic solver method. Applied Mathematics and Computation, 10/11(1): 695–716, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Thames:1984:MAT

[Tha84] Frank C. Thames. Multigrid applications to three-dimensional elliptic coordinate generation. Applied Mathematics and Computation, 15(4):325–342, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Thomas:1982:NGC

[Tho82a]  P. D. Thomas.  Numerical

generation of composite threedimensional grids by quasilinear elliptic systems. Applied Mathematics and Computation, 10/11 (1):667–686, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Thompson:1982:EGG

[Tho82b] Joe F. Thompson. Elliptic grid generation. Applied Mathematics and Computation, 10/11(1): 79–105, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Thompson:1982:GCC

[Tho82c] Joe F. Thompson. General curvilinear coordinate systems.

Applied Mathematics and Computation, 10/11(1):1–30, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Thompson:1982:NGG

[Tho82d] Joe F. Thompson, editor. Numerical grid generation: proceedings of a Symposium on the Numerical Generation of Curvilinear Coordinate Systems and their Use in the Numerical Solution of Partial Differential Equations: April 1982, Nashville, Tennessee, volume 10/11 of Applied Mathematics and Computation. North-Hol-land, Amsterdam, The Netherlands, 1982. ISBN 0-444-00757-1. LCCN QA377 .S966 1982.

Tuckness:1995:CAN

[TJ95] Dan G. Tuckness and Bertrand Jost. A critical analysis of the numerical and analytical methods used in the construction of the lunar gravity potential model. Applied Mathematics and Computation, 71(1): 69–90, August 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=1&aid=9400141.

Teo:1990:NCD

[TLG90] K. L. Teo, Y. Liu, and C. J.

Goh. Nonlinearly constrained discrete-time optimal-control problems. Applied Mathematics and Computation, 38(3):227– 248, August 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Trigiante:1992:BLS

[TM92] D. Trigiante and K. N. Murty. On best least square solutions of two-point boundary value problems via modified Gram– Schmidt process. Applied Mathematics and Computation, 51(2– 3):105–118, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tomoeda:1980:SFS

[Tom80]  K. Tomoeda.  Stability of

Friedrichs’s scheme in the maximum norm for hyperbolic systems in one space dimension. Applied Mathematics and Computation, 7(4):313–320, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tang:1997:SMA

[TPGV97] Yi-Fa Tang, Víctor M. Pérez-García, and Luis Vázquez. Symplectic methods for the Ablowitz–Ladik model. Applied Mathematics and Computation, 82(1):17–38, March 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/

browse.cgi?year=1997&volume=
82&issue=1&aid=9600019.

Thandapani:1995:ONB

[TPL95] E. Thandapani, S. Pandian, and B. S. Lalli. Oscillatory and nonoscillatory behavior of second-order functional difference equations. Applied Mathematics and Computation, 70 (1):53–66, June 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=1&aid=9400142.

Tleimat:1998:NSU

[TPS98] Hazem Tleimat, R. B. Piercey, and B. K. Soni. Numerical simulation of ultra-relativistic, heavy-ion collisions. Applied Mathematics and Computation, 89(1–3):275–293, January 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781663.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Trigiante:1987:NAU

[TS87] D. Trigiante and S. Sivasundaram. A new algorithm for unstable three-term recurrence relations. Applied Mathematics and Computation, 22(4):277– 289, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tsitouras:1998:EHO

[TS98] Ch. Tsitouras and T. E. Simos. Explicit high order methods for the numerical integration of periodic initial-value problems. Applied Mathematics and Computation, 95(1):15–26, September 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/95/1/6096.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=95&issue=
                                        1&aid=6096.

Thornburg:1998:SGB

[TSK98] Hugh Thornburg, Bharat K. Soni, and Boyalakuntla Kishore. A structured grid based solutionadaptive technique for complex separated flows. Applied Mathematics and Computation, 89 (1–3):259–273, January 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1998&volume=
                                        89&issue=1-3&aid=9781662.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Theocaris:1984:NSS

[TT84] P. S. Theocaris and G. Tsamasphyros. A numerical solution of singular integro-differential equations with variable coefficients. Applied Mathematics and Computation, 15(1):47–50, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Thole:1986:BSP

[TT86] Clemens-August Thole and Ulrich Trottenberg. Basic smoothing procedures for the multigrid treatment of elliptic 3D operators. Applied Mathematics and Computation, 19(1–4): 333–345, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Thole:1988:SNS

[TT88] Clemens-August Thole and Ulrich Trottenberg. A short note on standard parallel multigrid algorithms for 3D problems. Applied Mathematics and Computation, 27(2):101–115, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tennakone:1989:PIF

[TT89] K. Tennakone and D. Tantirigoda. Periodicity in the incidence of Filariasis and the solar cycle. Applied Mathematics and Computation, 32(1):35–43, July 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Thompson:1990:ONEa

[TT90a] M. E. Thompson and A. Thavaneswaran. Optimal nonparametric estimation for some semimartingale stochastic differential equations. Applied Mathematics and Computation, 37 (3):169–183, June 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Thompson:1990:ONEb

[TT90b] M. E. Thompson and A. Thavaneswaran. Optimal nonparametric estimation for some semimartingale stochastic differential equations. Applied Mathematics and Computation, 39(3 (Supplement)):123s–137s, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tucker:1988:BIC

[TTB88] Jerry H. Tucker, Moiez A. Tapia, and A. Wayne Bennett. Boolean integral calculus. Applied Mathematics and Computation, 26(3 (part II)):201–236, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tuckness:1995:FMQ

[Tuc95] Dan G. Tuckness, Jr. A fast method for quantitatively measuring stability in a three-body dynamical system. Applied Mathematics and Computation, 70(1):29–51, June 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=1&aid=9400136.

Tuckness:1996:AON

[Tuc96] Dan G. Tuckness. Analysis of optical navigation error during Mars entry. Applied Mathematics and Computation, 80(1):1– 22, November 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=1&aid=9500275.

Tuckness:1997:PSS

[Tuc97]  Dan G. Tuckness.  Phase

space sensitivities of the Sun-Jupiter triangular libration point. Applied Mathematics and Computation, 88(2–3):255– 265, December 30, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600329.

Turner:1988:SSM

[Tur88] Monica Goigel Turner. A spatial simulation model of land use changes in a Piedmont County in Georgia. Applied Mathematics and Computation, 27(1): 39–51, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tsokos:1979:RBO

[TW79] Chris P. Tsokos and R. L. W. Welch. Robustness of the Bayes optimal discriminant procedure with O-I loss. Applied Mathematics and Computation, 5 (??):131–148, 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tennakone:1989:CSE

[TW89] K. Tennakone and S. Wickremanayake. Computer study of the effect of noise on an epidemiological model. Applied Mathematics and Computation, 34(2 (part II)):73–87, November 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tennakone:1990:BCS

[TW90] K. Tennakone and S. Wickramanayaka. Biochemical chiral selection in the presence of instabilities, chaos, and noise. Applied Mathematics and Computation, 38(1 (part I)):1–6, July 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Toyoda:1999:AHS

[TyG99] Shin’ichi Toyoda and Pegio yukio Gunji. Appearance of hierarchical structure in hyperdilation model: model of generalized measurement process. Applied Mathematics and Computation, 104(2–3):

153–178, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/2-
3/6285.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
104&issue=2-3&aid=6285.

Tishler:1983:MMA

[TZ83] A. Tishler and I. Zang. A minmax algorithm for nonlinear regression models. Applied Mathematics and Computation, 13(1– 2):95–115, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Usmani:1988:NST

[UA88] Riaz A. Usmani and Ravi P.

Agarwal. On the numerical solution of two-point discrete boundary value problems. Applied Mathematics and Computation, 25(3):247–264, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Udwadia:1981:NRR

[Udw81] F. E. Udwadia. A note on the reliability of redundant systems. Applied Mathematics and Computation, 8(1):35–49, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Udwadia:1987:RUD

[Udw87] F. E. Udwadia. Response of uncertain dynamic systems. I. Applied Mathematics and Computation, 22(??):115–150, 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Udwadia:1991:NPC

[Udw91] Firdaus E. Udwadia. Noncollocated point control of nondispersive distributed-parameter systems using time delays. Applied Mathematics and Computation, 42(1 (part I)):23–63, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Udwadia:1992:NCD

[Udw92a] Firdaus E. Udwadia. Noncollocated control of distributedparameter nondispersive systems with tip inertias using time delays. Applied Mathematics and Computation, 47(1): 47–75, January 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Udwadia:1992:SCR

[Udw92b] Firdaus E. Udwadia. Some convergence results related to the JOR iterative method for symmetric, positive-definite matrices. Applied Mathematics and Computation, 47(1):37–45, January 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ueno:1980:IST

[Uen80] Sueo Ueno. Inversion of solar transmitted radiance profiles for the atmospheric optical thickness. Applied Mathematics and Computation, 7(2):177–186, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ueno:1983:IVS

[Uen83] Sueo Ueno. An initial value solution of the least-squares estimation problem with degenerate covariance. Applied Mathematics and Computation, 13 (1–2):87–94, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ueno:1989:SAC

[Uen89] S. Ueno. Stochastic approach to the Cauchy system for the scattering function with hybrid reflector. Applied Mathematics and Computation, 29(1 (part I)):39–47, January 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ueno:1991:SAC

[Uen91] Sueo Ueno. Stochastic approach to the Cauchy problem of linearized Couette flow. Applied Mathematics and Computation, 45(2 (part II)):207–217, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ueno:1995:PAR

[Uen95] Sueo Ueno. Probabilistic approach for rarefied gas dynamics: Linearized Couette flow. Applied Mathematics and Computation, 69(1):61–73, April 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=1&aid=9400099.

Udwadia:1996:ISI

[UHL96] F. E. Udwadia, E. Hofer, and G. Leitmann. Introduction to the special issue on dynamics and control. Applied Mathematics and Computation, 78(2– 3):101–??, ???? 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Udwadia:1994:CIM

[UK94] Firdaus E. Udwadia and Ravi Kumar. Convergence of iterative methods for non-classically damped dynamic systems. Applied Mathematics and Computation, 61(1):61–97, March 01, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Udwadia:1995:APE

[UK95] Firdaus E. Udwadia and Robert E. Kalaba. An alternate proof for the equation of motion for constrained mechanical systems. Applied Mathematics and Computation, 70(2–3):339– 342, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400113.

Dynamics and control (Vienna, 1993).

Unbehauen:1996:SNT

[Unb96]  H. Unbehauen.  Some new

trends in identification and modeling of nonlinear dynamical systems. Applied Mathematics and Computation, 78(2–3):279– 297, September 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
78&issue=2-3&aid=9600015.

Udwadia:1997:DCN

[UR97] F. E. Udwadia and N. Raju.

Dynamics of coupled nonlinear maps and its application to ecological modeling. Applied Mathematics and Computation, 82 (2–3):137–179, March 15, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
82&issue=2-3&aid=9600027.

Udwadia:1995:ISI

[UTL^{+}95] F. E. Udwadia, H. Troger, G. Leitmann, H. P. Joergl, and H. Flashner. Introduction to the special issue on dynamics and control. Applied Mathematics and Computation, 70(2– 3):95–??, July 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Venkatachalam:1988:PTU

[VA88] P. A. Venkatachalam and S. Arumugam. Program termination using Z-transform theory. Applied Mathematics and Computation, 27(3 (part I)): 231–241, September 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Valyi:1988:SAD

[Vál88] István Vályi. Strict approximate duality in vector spaces. Applied Mathematics and Computation, 25(3):227–246, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

VanRosendale:1983:ADS

[Van83]  John Van Rosendale.  Al-

gorithms and data structures for adaptive multigrid elliptic solvers. Applied Mathematics and Computation, 13(3–4):453– 470, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

vanStraten:1985:AMP

[van85] Gerrit van Straten. Analytical methods for parameter-space delimitation and application to shallow-lake phytoplanktondynamics modeling. Applied Mathematics and Computation, 17(4):459–482, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). System identification and parameter estimation of environmental systems.

Vanmaele:1995:EFE

[Van95] M. Vanmaele. On an external finite element method for a second-order eigenvalue problem on a concave 2D-domain with Dirichlet boundary conditions. Applied Mathematics and Computation, 71(2–3):211– 226, September 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=2-3&aid=9400155.

VanHuylenbroeck:1997:MTT

[Van97] G. Van Huylenbroeck. Multicriteria tools for the trade-off analysis in rural planning between economic and environmental objectives. Applied Mathematics and Computation, 83(2–3): 261–280, May 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=2-3&aid=9600188.

VanSiclen:1998:IML

[Van98] Clinton DeW. Van Siclen. Information measure of location precision. Applied Mathematics and Computation, 97(2–3): 287–293, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/97/2-3/
                                        6149.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=97&issue=
                                        2-3&aid=6149.

Vargas:1980:NEC

[Var80]  Luis G. Vargas.  A note

on the eigenvalue consistency index. Applied Mathematics and Computation, 7(3):195–203, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vargas:1983:ASR

[Var83] Luis G. Vargas. Analysis of sensitivity of reciprocal matrices. Applied Mathematics and Computation, 12(4):301–320, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vassilevski:1991:MPE

[Vas91] P. S. Vassilevski. Multilevel preconditioners for elliptic problems by substructuring. Applied Mathematics and Computation, 46(2 (part II)):79–106, ???? 1991. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Vaughn:1980:SLS

[Vau80] R. L. Vaughn. Solutions of linear Sobolev differential equations. Applied Mathematics and Computation, 6(3):211–216, April 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vazquez:1998:UNA

[Váz98] C. Vázquez. An upwind numerical approach for an American and European option pricing model. Applied Mathematics and Computation, 97 (2–3):273–286, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6133.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6133.

Vincent:1989:ERS

[VB89] Thomas L. Vincent and Joel S.

Brown. The evolutionary response of systems to a changing environment. Applied Mathematics and Computation, 32(2– 3):185–206, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Mathematical biology.

Voulov:1992:ASH

[VB92] H. D. Voulov and D. D. Baĭnov.

Asymptotic stability for a homogeneous singularly perturbed system of differential equations with unbounded delay. Applied Mathematics and Computation, 51(1):1–16, September 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Verschelde:1991:NSS

[VBH91] Jan Verschelde, Marc Beckers, and Ann Haegemans. A new start system for solving deficient polynomial systems using continuation. Applied Mathematics and Computation, 44(3): 225–239, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Varner:1998:NSD

[VC98] T. N. Varner and S. Roy Choudhury. Non-standard difference schemes for singular perturbation problems revisited. Applied Mathematics and Computation, 92(2–3):101–123, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/2-3/
9710025.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
92&issue=2-3&aid=9710025.

Varadarajan:1991:RCT

[VCM91] V. Varadarajan, W. H. Choe, and G. H. Miley. Rapid calculation of tokamak equilibria with toroidal rotation via minimization of a functional. Applied Mathematics and Computation, 43(2 (part II)):145–163, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

VanKeer:1997:FEA

[VD97] R. Van Keer and H. De Schepper. Finite element approximation for 2nd order elliptic eigenvalue problems with nonlocal boundary or transition conditions. Applied Mathematics and Computation, 82 (1):1–16, March 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
82&issue=1&aid=9600091.

Veverka:1992:MRM

[Vev92] Vladimír Veverka. A method of reconciliation of measured data with nonlinear constraints. Applied Mathematics and Computation, 49(2–3):141–176, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

vanGoethem:1979:AFC

[vGG79] P. van Goethem and M. J.

Goovaerts. Approximation formulae for compound Poisson processes for a kind of claim distributions having a prescribed asymptotic behavior. Applied Mathematics and Computation, 5(3):243–251, ???? 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Visentin:1980:LPS

[Vis80] F. Visentin. Limiting processes in the stability problem. Applied Mathematics and Computation, 7(1):81–91, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vargas-Jarillo:1986:DMR

[VJ86] Cristóbal Vargas-Jarillo. A discrete model for the recovery of oil from a reservoir. Applied Mathematics and Computation, 18(2):93–117, ???? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vazquez:1988:MCM

[VJ88] Luis Vázquez and Salvador Jiménez. Motion of a charge in a magnetic dipole field. I. Painlevé analysis and a conservative numerical scheme. Applied Mathematics and Computation, 25 (3):207–217, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kozak:1995:SRS

[vKH95] Š. Kozák and J. Hejdiš. Simple, robust, selftuning controller. Applied Mathematics and Computation, 70(2–3): 203–214, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400119.

Dynamics and control (Vienna, 1993).

Vincent:1983:POD

[VL83] T. L. Vincent and C. S. Lee.

Parametric optimization with dynamic systems. Applied Mathematics and Computation, 12(2– 3):169–185, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vestroni:1995:SCT

[VLP95] Fabrizio Vestroni, Angelo Luongo, and Monica Pasca. Stability and control of transversal oscillations of a tethered satellite system. Applied Mathematics and Computation, 70(2–3): 343–360, July 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
70&issue=2-3&aid=9400114.

Dynamics and control (Vienna, 1993).

Voller:1992:MEN

[Vol92] Rudolf L. Voller. Monotone enclosure for nonlinear PDEs using multigrid techniques. Applied Mathematics and Computation, 47(2–3):121–135, February 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Voss:1989:FTS

[Vos89] David A. Voss. Factored twostep Runge–Kutta methods. Applied Mathematics and Computation, 31(Special issue):361– 368, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Villatoro:1999:MMEa

[VR99a] F. R. Villatoro and J. I. Ramos. On the method of modified equations. I: Asymptotic analysis of the Euler forward difference method. Applied Mathematics and Computation, 103 (2–3):111–139, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/103/2-
                                        3/6256.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        103&issue=2-3&aid=6256.

Villatoro:1999:MMEb

[VR99b] F. R. Villatoro and J. I. Ramos. On the method of modified equations. II: Numerical techniques based on the equivalent equation for the Euler forward difference method. Applied Mathematics and Computation, 103(2–3):141–177, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/103/2-
3/6255.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
103&issue=2-3&aid=6255.

Villatoro:1999:MMEc

[VR99c] F. R. Villatoro and J. I.

Ramos. On the method of modified equations. III. Numerical techniques based on the second equivalent equation for the Euler forward difference method. Applied Mathematics and Computation, 103(2– 3):179–212, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/103/2-
3/6254.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
103&issue=2-3&aid=6254.

Villatoro:1999:MMEd

[VR99d] F. R. Villatoro and J. I. Ramos.

On the method of modified equations. IV. Numerical techniques based on the modified equation for the Euler forward difference method. Applied Mathematics and Computation, 103(2–3):213–240, August 15, 1999. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/103/2-
                                        3/6253.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        103&issue=2-3&aid=6253.

Villatoro:1999:MMEe

[VR99e] F. R. Villatoro and J. I. Ramos. On the method of modified equations. V: Asymptotic analysis of and direct-correction and asymptotic successive-correction techniques for the implicit midpoint method. Applied Mathematics and Computation, 103 (2–3):241–285, August 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/103/2-
                                        3/6252.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        103&issue=2-3&aid=6252.

Villatoro:1999:MMEf

[VR99f] F. R. Villatoro and J. I. Ramos. On the method of modified equations. VI: Asymptotic analysis of and asymptotic successive-corrections techniques for two-point, boundaryvalue problems in ODE’s. Applied Mathematics and Computation, 105(2–3):137–171, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.elsevier.
com/cas/tree/store/amc/sub/
1999/105/2-3/6320.pdf;   http:
//www.elsevier.com/cgi-bin/
cas/tree/store/amc/cas_sub/
browse/browse.cgi?year=1999&
volume=105&issue=2-3&aid=
6320.

Vazirinejad:1990:CSE

[VS90] Shamsedin Vazirinejad and Soroosh Sorooshian. A comparison of simultaneous-equations, weighted regression, and noisein-variables models. Applied Mathematics and Computation, 40(3):187–202, ???? 1990. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vatsala:1996:GQF

[VS96] A. S. Vatsala and Donna Stutson. Generalized quasilinearization and first-order periodic boundary value problem. Applied Mathematics and Computation, 77(2–3):113–129, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vulanovic:1990:SIN

[Vul90] Relja Vulanović. Some improvements of the nonequidistant Engquist–Osher scheme. Applied Mathematics and Computation, 40(2 (part II)):147–164, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vatsala:1980:ECR

[VV80] A. S. Vatsala and R. L. Vaughn. Existence and comparison results for differential equations of Sobolev type. Applied Mathematics and Computation, 6(2): 177–187, March 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vasudevan:1985:SPC

[VV85] R. Vasudevan and P. R. Vittal. Storage problems in continuous time with random inputs, random outputs, and deterministic release. Applied Mathematics and Computation, 16(4): 309–326, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Vu:1994:NSF

[VWSS94] Bruce Vu, Ten-See Wang, Ming-Hsin Shih, and Bharat Soni. Navier–Stokes flow field analysis of compressible flow in a high pressure safety relief valve. Applied Mathematics and Computation, 65(1–3):345–353, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wong:1996:OAM

[WA96] Patricia J. Y. Wong and Ravi P. Agarwal. The oscillation and asymptotically monotone solutions of second-order quasilinear differential equations. Applied Mathematics and Computation, 79(2–3):207–237, October 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
79&issue=2-3&aid=9500267.

Wong:1998:ABS

[WA98] Patricia J. Y. Wong and Ravi P.

Agarwal. Asymptotic behaviour of solutions of higher order difference and partial difference equations with distributed deviating arguments. Applied Mathematics and Computation, 97 (2–3):139–164, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6134.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6134.

Wong:1999:ELB

[WA99] Patricia J. Y. Wong and Ravi P.

Agarwal. Eigenvalues of Lidstone boundary value problems. Applied Mathematics and Computation, 104(1):15–31, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/1/
6267.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=104&
issue=1&aid=6267.

Waissi:1990:NSC

[Wai90] Gary R. Waissi. A new O(n^{2}) shortest chain algorithm. Applied Mathematics and Computation, 37(2 (part II)):111–120, May 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wampler:1992:BNC

[Wam92]  Charles W. Wampler.  Be-

zout number calculations for multi-homogeneous polynomial systems. Applied Mathematics and Computation, 51(2–3): 143–157, October 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1980:SAI

[Wan80]  A. P. Wang.  On surface-

area integration and the related mapping. Applied Mathematics and Computation, 6(3):251–262, April 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1981:HARa

[Wan81a] Zhong De Wang. Harmonic analysis with a real frequency function. I. aperiodic case. Applied Mathematics and Computation, 9(1):53–73, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1981:HARc

[Wan81b] Zhong De Wang. Harmonic analysis with a real frequency function. III. data sequence. Applied Mathematics and Computation, 9(4):245–255, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1981:HARb

[Wan81c] Zhong De Wang. Harmonic analysis with a real function of frequency. II. periodic and bounded cases. Applied Mathematics and Computation, 9(3): 153–163, ???? 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1987:FAC

[Wan87] Guo Rong Wang. A finite algorithm for computing the weighted Moore–Penrose inverse A_{MN}^{+}. Applied Mathematics and Computation, 23(4):277– 289, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1992:SIT

[Wan92] Shao Ming Wang. Spline interpolations over type-2 triangulations. Applied Mathematics and Computation, 49(2–3):299–313, June 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1993:RNN

[Wan93] Jun Wang. A recurrent neural network for real-time matrix inversion. Applied Mathematics and Computation, 55(1):89–100, April 1993. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1994:CSA

[Wan94] Shao Ming Wang. A cubic structure on an arbitrary quadrilateral. Applied Mathematics and Computation, 60(1):55–68, January 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1995:IMB

[Wan95a] Peiguang Wang. Iterative methods for the boundary value problem of a fourth-order differential-difference equation. Applied Mathematics and Computation, 73(2–3):257–270, December 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1995&volume=73&issue=
2-3&aid=9500066.

Wang:1995:QSS

[Wan95b] Shao Ming Wang. A quadratic spline structure over triangulations. Applied Mathematics and Computation, 68(2–3):143– 152, March 15, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
68&issue=2-3&aid=9400088.

Wang:1996:BVP

[Wan96a] Peiguang Wang. Boundary value problem of a third-order mixed type differential-difference equation. Applied Mathematics and Computation, 80(2–3):273– 286, December 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500303.

Wang:1996:NCK

[Wan96b] Zeke Wang. Notes on computation of Kakutani fixed points. Applied Mathematics and Computation, 75(2–3):139– 150, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500121.

Wang:1999:FOC

[Wan99a] Peiguang Wang. Forced oscillation of a class of delay hyperbolic equation boundary value problem. Applied Mathematics and Computation, 103 (1):15–25, August 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/103/1/
6283.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/

                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=103&
                                        issue=1&aid=6283.

Wang:1999:OCC

[Wan99b] Peiguang Wang. Oscillatory criteria for a class of delay hyperbolic equations boundary value problem (II). Applied Mathematics and Computation, 100(2–3):189–199, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/100/2-
                                        3/6210.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        100&issue=2-3&aid=6210.

Warsi:1982:BDM

[War82] Z. U. A. Warsi. Basic differential model for coordinate generation. Applied Mathematics and Computation, 10/11 (1):41–77, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Warsi:1987:SEP

[War87] Z. U. A. Warsi. A synopsis of elliptic PDE models for grid generation. Applied Mathematics and Computation, 21 (4):295–311, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Warren:1991:TTS

[War91] Richard H. Warren. Tours for the traveling salesman. Applied Mathematics and Computation, 46(3):167–177, December 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Warren:1994:SCT

[War94] Richard H. Warren. Special cases of the traveling salesman problem. Applied Mathematics and Computation, 60(2–3): 171–177, February 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Watson:1979:GCA

[Wat79] Layne T. Watson. A globally convergent algorithm for computing fixed points of C^{2} maps. Applied Mathematics and Computation, 5(4):297–311, ???? 1979. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Watson:1981:EAC

[Wat81] Layne T. Watson. Engineering applications of the Chow-Yorke algorithm. Applied Mathematics and Computation, 9(2):111–133, ???? 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Watson:1989:GCH

[Wat89a] Layne T. Watson. Globally convergent homotopy methods: a tutorial. Applied Mathematics and Computation, 31(Special issue):369–396, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Watson:1989:SFD

[Wat89b] Ray Watson. On the spread of a fatal disease. Applied Mathematics and Computation, 33(2 (part II)):123–129, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Watts:1989:SOI

[Wat89c] H. A. Watts. A smooth output interpolation process for BDF codes. Applied Mathematics and Computation, 31(Special issue): 397–418, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Wazwaz:1993:NSG

[Waz93] Abdul-Majid Wazwaz. On the numerical solution of the Goursat problem. Applied Mathematics and Computation, 59(1): 89–95, November 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wazwaz:1995:DMA

[Waz95a] Abdul-Majid Wazwaz. The decomposition method for approximate solution of the Goursat problem. Applied Mathematics and Computation, 69(2–3): 299–311, May 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
69&issue=2-3&aid=9400137.

Wazwaz:1995:NAN

[Waz95b] Abdul-Majid Wazwaz. A new approach to the nonlinear advection problem: An application of the decomposition method. Applied Mathematics and Computation, 72(2–3):175– 181, October 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400182.

Wazwaz:1997:NCA

[Waz97a] Abdul-Majid Wazwaz. Necessary conditions for the appearance of noise terms in decomposition solution series. Applied Mathematics and Computation, 81(2–3):265–274, February 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/

                                        browse.cgi?year=1997&volume=
                                        81&issue=2-3&aid=9500327.

Wazwaz:1997:SBL

[Waz97b] Abdul-Majid Wazwaz. A study on a boundary-layer equation arising in an incompressible fluid. Applied Mathematics and Computation, 87(2–3):199– 204, December 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        87&issue=2-3&aid=9600281.

Wazwaz:1997:SCI

[Waz97c] Abdul-Majid Wazwaz. A study on complex integrals involving absolute values. Applied Mathematics and Computation, 85(2– 3):281–286, September 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        85&issue=2-3&aid=9600142.

Wazwaz:1998:CBA

[Waz98a] Abdul-Majid Wazwaz. A comparison between Adomian decomposition method and Taylor series method in the series solutions. Applied Mathematics and Computation, 97 (1):37–44, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/

store/amc/sub/1998/97/1/6135.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
1&aid=6135.

Wazwaz:1998:RTS

[Waz98b] Abdul-Majid Wazwaz. A reliable technique for solving the wave equation in an infinite one-dimensional medium. Applied Mathematics and Computation, 92(1):1–7, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/1/6028.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=92&issue=
1&aid=6028.

Wazwaz:1999:AAP

[Waz99a] Abdul-Majid Wazwaz. Analytical approximations and Padé approximants for Volterra’s population model. Applied Mathematics and Computation, 100(1):13–25, April 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/1/
6202.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=100&
issue=1&aid=6202.

Wazwaz:1999:CSS

[Waz99b] Abdul-Majid Wazwaz. A comparative study on a singular perturbation problem with two singular boundary points. Applied Mathematics and Computation, 99(2–3):179–193, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/99/2-3/
                                        6190.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=99&issue=
                                        2-3&aid=6190.

Wazwaz:1999:MDM

[Waz99c] Abdul-Majid Wazwaz. The modified decomposition method and Padé approximants for solving the Thomas–Fermi equation. Applied Mathematics and Computation, 105(1):11–19, October 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/105/1/
                                        6311.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=105&
                                        issue=1&aid=6311.

Wazwaz:1999:RMA

[Waz99d] Abdul-Majid Wazwaz. A reliable modification of Adomian decomposition method. Applied Mathematics and Computation, 102(1):77–86, July 1, 1999. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/102/1/
6249.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=102&
issue=1&aid=6249.

Wang:1988:AMQ

[WC88] H. Wang and A. J. Chapman.

Applications of the modified quasilinearization algorithm to the construction of the contour of turbine blades. Applied Mathematics and Computation, 25(2 (part II)):115–122, 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:1999:UFS

[WC99] Berlin Wu and Mei-Hui Chen.

Use of fuzzy statistical technique in change periods detection of nonlinear time series. Applied Mathematics and Computation, 99(2–3):241–254, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6195.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6195.

Wong:1997:IBV

[WCG97] Yau Shu Wong, Qianshun Chang, and Lianger Gong. An initial-boundary value problem of a nonlinear Klein–Gordon equation. Applied Mathematics and Computation, 84(1): 77–93, June 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        84&issue=1&aid=9600065.

Weibull:1984:SFA

[Wei84] Jörgen W. Weibull. A stockflow approach to general equilibrium with trade frictions. Applied Mathematics and Computation, 14(1):63–76, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Weinmuller:1989:CSB

[Wei89] Ewa B. Weinmüller. On the condition of singular boundary value problems. Applied Mathematics and Computation, 31(Special issue):555–573, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Wei:1998:ISD

[Wei98] Yimin Wei. Index splitting for the Drazin inverse and the singular linear system. Applied Mathematics and Computation, 95(2–3):115–124, September 15, 1998. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6106.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6106.

Wei:1999:PBS

[Wei99a] Y. Wei. Perturbation bound of singular linear systems. Applied Mathematics and Computation, 105(2–3):211–220, November 12, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/2-
3/6479.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
105&issue=2-3&aid=6479.

Wei:1999:PGI

[Wei99b] Yimin Wei. On the perturbation of the group inverse and oblique projection. Applied Mathematics and Computation, 98(1):29–42, January 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/98/1/6155.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=98&issue=
1&aid=6155.

Werbos:1993:CSF

[Wer93] Paul J. Werbos. Chaotic solitons and the foundations of physics: A potential revolution. Applied Mathematics and Computation, 56(2–3):289–339, July 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wexler:1987:AED

[Wex87] Anthony S. Wexler. Automatic evaluation of derivatives. Applied Mathematics and Computation, 24(1):19–46, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wilcox:1984:MGC

[WH84] Ralph M. Wilcox and Leo P. Harten. MACSYMA-generated closed-form solutions to some matrix Riccati equations. Applied Mathematics and Computation, 14(2):149–166, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1985:DT

[WH85] Zhong De Wang and B. R. Hunt. The discrete W transform. Applied Mathematics and Computation, 16(1):19–48, ???? 1985. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Watanabe:1987:SWD

[WH87] K. Watanabe and D. M. Himmelblau. A straightforward way to design nonlinear controllers for nonlinear processes by means of process discretization. Applied Mathematics and Computation, 21(4):313–349, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1989:SCH

[WHF89] Muhong Wang, Keith W. Hipel, and Niall M. Fraser. Solution concepts in hypergames. Applied Mathematics and Computation, 34(3):147–171, December 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

White:1988:EFR

[Whi88] L. W. White. Estimation of flexural rigidity in a Kirchhoff plate model. Applied Mathematics and Computation, 27(4 (part II)):337–359, September 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

White:1989:EHO

[Whi89]  L. W. White.  Estimation

of higher-order damping terms in linear plate models. Applied Mathematics and Computation, 33(2 (part II)):89–122, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

White:1991:EEP

[Whi91] L. W. White. Estimation of elastic parameters in a nonlinear elliptic model of a plate. Applied Mathematics and Computation, 42(2 (part II)):139–187, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

White:1993:EMP

[Whi93] L. W. White. Estimation of material parameters in a dynamic nonlinear plate model. Applied Mathematics and Computation, 54(1):25–63, March 01, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

White:1997:RRO

[Whi97] Luther W. White. Resolution of regularized output least squares estimation procedures. Applied Mathematics and Computation, 81(2–3):139–172, February 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        81&issue=2-3&aid=9500324.

Wong:1980:EAP

[WI80] K. C. Wong and S. Sitharama Iyengar. An efficient algorithm for product computations on computer. Applied Mathematics and Computation, 6(??):1– 5, 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wong:1992:SAS

[WJ92] Yau Shu Wong and Xing Zhi Ji. On shooting algorithm for Sturm–Liouville eigenvalue problems with periodic and semi-periodic boundary conditions. Applied Mathematics and Computation, 51(2–3):87– 104, October 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

White:1997:EDE

[WJO97] Luther W. White, Ying-Jun Jin, and Daniel J. O’Meara. Estimation of discontinuous elliptic coefficients. Applied Mathematics and Computation, 81(2–3): 113–138, February 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
81&issue=2-3&aid=9500323.

Watson:1995:AMC

[WK95] Brian C. Watson and Manohar P.

Kamat. Analysis of mistuned cyclic systems using mistune modes. Applied Mathematics and Computation, 67(1–3):61– 79, January 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
67&issue=1-3&aid=9400054.

Wazwaz:1996:RTS

[WK96a] A. M. Wazwaz and S. A. Khuri.

A reliable technique for solving the weakly singular second-kind Volterra-Type integral equations. Applied Mathematics and Computation, 80(2–3):287– 299, December 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500279.

Wazwaz:1996:TMS

[WK96b] A. M. Wazwaz and S. A. Khuri. Two methods for solving integral equations. Applied Mathematics and Computation, 77 (1):79–89, June 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
77&issue=1&aid=9500189.

Wazwaz:1998:NIS

[WK98] A. M. Wazwaz and S. A. Khuri. New ideas for solving size-structured population models. Applied Mathematics and Computation, 93(1):91–96, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/1/6061.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
1&aid=6061.

Wood:1984:BPB

[WKM84] Eric F. Wood, James A. Kempf, and Raman K. Mehra. BISTAB: a portable bifurcation and stability analysis package. Applied Mathematics and Computation, 15(4):343–355, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wexler:1988:IIS

[WKTM88] A. S. Wexler, R. E. Kalaba, L. S. Tesfatsion, and D. J. Marsh. An invariant-imbedding solution of general linear twopoint boundary-value problems. Applied Mathematics and Computation, 26(3 (part II)):237– 244, ???? 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wei:1997:LEC

[WL97] J. G. Wei and G. Leng. Lyapunov exponent and chaos of Duffing’s equation perturbed by white noise. Applied Mathematics and Computation, 88(1):77– 93, December 15, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=1&aid=9600307.

Woodward:1997:STT

[WLB97] J. Arthur Woodward, Wan Ching Liu, and Douglas G. Bonett. Shortest two-tailed confidence intervals. Applied Mathematics and Computation, 84(1): 65–76, June 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=1&aid=9600080.

Wu:1999:CCI

[WLTS99] Tsu-Chen Wu, Shinq-Hwa Leu, Shih-Tong Tu, and H. M. Srivastava. A certain class of infinite sums associated with Digamma functions. Applied Mathematics and Computation, 105(1):1–9, October 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/1/
6319.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=105&
issue=1&aid=6319.

Weinberg:1987:NR

[WM87] David Weinberg and Clyde Martin. A note on resultants. Applied Mathematics and Computation, 24(4):303–309, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wexler:1991:NMT

[WM91] Anthony S. Wexler and Donald J. Marsh. Numerical methods for three-dimensional models of the urine concentrating mechanism. Applied Mathematics and Computation, 45(2 (part II)):219–240, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Weatherill:1991:SPF

[WN91] N. P. Weatherill and D. K.

Natakusumah. The simulation of potential flow around multiple bodies using overlapping connected meshes. Applied Mathematics and Computation, 46(1 (part I)):1–21, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wolfe:1989:OIM

[Wol89] M. A. Wolfe. On order-interval methods for bounding zeros of order convex operators. Applied Mathematics and Computation, 32(1):45–59, July 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wolfe:1996:IMG

[Wol96] M. A. Wolfe. Interval methods for global optimization. Applied Mathematics and Computation, 75(2–3):179–206, March 15, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500125.

Wolfe:1998:IEC

[Wol98] M. A. Wolfe. Interval enclosures for a certain class of multiple integrals. Applied Mathematics and Computation, 96(2– 3):145–159, November 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/2-3/
6115.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
2-3&aid=6115.

Wong:1998:GDD

[Won98] Tzu-Tsung Wong. Generalized Dirichlet distribution in Bayesian analysis. Applied Mathematics and Computation, 97(2–3):165–181, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6147.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6147.

Woodward:1997:ECD

[WP97] J. Arthur Woodward and Christina G. S. Palmer. On the exact convolution of discrete random variables. Applied Mathematics and Computation, 83 (1):69–77, April 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=1&aid=9600047.

Waissi:1996:SAS

[WR96] Gary R. Waissi and Donald F.

Rossin. A sigmoid approximation of the standard normal integral. Applied Mathematics and Computation, 77 (1):91–95, June 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
77&issue=1&aid=9500190.

Watson:1987:SSC

[WS87] Layne T. Watson and Melvin R.

Scott. Solving spline-collocation approximations to nonlinear two-point boundary-value problems by a homotopy method. Applied Mathematics and Computation, 24(4):333–357, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Watts:1983:SCV

[WSL83] H. A. Watts, M. R. Scott, and M. E. Lord. Solving complexvalued differential systems. Applied Mathematics and Computation, 12(4):381–394, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wang:1999:COF

[WSL99] Feng-Sheng Wang, Ching-Tzong Su, and Yung-Cheng Liu. Computation of optimal feedforward and feedback control by a modified reduced gradient method. Applied Mathematics and Computation, 104(1): 85–100, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/1/
6276.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=104&
issue=1&aid=6276.

Wu:1994:PRC

[Wu94] Berlin Wu. Pattern recognition and classification in time series analysis. Applied Mathematics and Computation, 62(1):29–45, April 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Wu:1997:CSR

[Wu97] Zongmin Wu. Compactly supported radial functions and the Strang-Fix condition. Applied Mathematics and Computation, 84(2–3):115–124, July 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=2-3&aid=9600110.

Wang:1998:IDT

[WW98] Alan P. Wang and Cecilia Y.

Wang. Identification of a dissipative transport system. Applied Mathematics and Computation, 91(1):33–41, April 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/1/6013.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
1&aid=6013.

Wen:1999:OPT

[WW99a] Jyh-Horng Wen and Jee-Wey Wang. An occupancy problem with two groups of balls resulting from TDM radio communication application. Applied Mathematics and Computation, 105(1):87–90, October 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/105/1/
6307.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=105&
issue=1&aid=6307.

Wen:1999:RSO

[WW99b] Jyh-Horng Wen and Jee-Wey Wang. A recursive solution to an occupancy problem resulting from TDM radio communication application. Applied Mathematics and Computation, 101(1):1–3, June 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/101/1/
                                        6223.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=101&
                                        issue=1&aid=6223.

Warsi:1982:NGT

[WZ82] Z. U. A. Warsi and J. P. Ziebarth. Numerical generation of three-dimensional coordinates between bodies of arbitrary shapes. Applied Mathematics and Computation, 10/11 (1):717–728, ???? 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Wu:1995:PSO

[WZ95] J. Wu and X. Zou. Patterns of sustained oscillations in neural networks with delayed interactions. Applied Mathematics and Computation, 73(1):55– 75, November 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        73&issue=1&aid=9400203.

White:1998:SMF

[WZ98] Luther W. White and Musharraf Zaman. A simple model for fluid accumulation and flow in a porous medium. Applied Mathematics and Computation, 90(2–3):181–203, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/
9700398.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=2-3&aid=9700398.

Xu:1998:CMM

[XL98] Xuejun Xu and Likang Li.

A V -cycle multigrid method and additive multilevel preconditioners for the plate bending problem discretized by conforming finite elements. Applied Mathematics and Computation, 93(2–3):233–258, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/2-3/
6086.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6086.

Xu:1999:NCV

[XL99] Xuejun Xu and Likang Li.

A note on convergence of Vcycle nonconforming multigrid methods. Applied Mathematics and Computation, 104(2– 3):191–206, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/104/2-
                                        3/6300.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1999&volume=
                                        104&issue=2-3&aid=6300.

Xin:1995:ASI

[XS95] Jingmin Xin and Akira Sano. Adaptive system identification based on generalized wavelet decomposition. Applied Mathematics and Computation, 69(1): 97–109, April 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        69&issue=1&aid=9400101.

Xu:1997:GCF

[XZL97] Zong-Ben Xu, Jiang-She Zhang, and Yiu-Wing Leung. A general CDC formulation for specializing the cell exclusion algorithms of finding all zeros of vector functions. Applied Mathematics and Computation, 86 (2–3):235–259, October 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1997&volume=
                                        86&issue=2-3&aid=9600191.

Xu:1998:AAC

[XZL98] Zong-Ben Xu, Jiang-She Zhang, and Yiu-Wing Leung. An approximate algorithm for computing multidimensional convex hulls. Applied Mathematics and Computation, 94(2– 3):193–226, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/2-3/
6031.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
2-3&aid=6031.

Xu:1996:CEA

[XZW96] Zong-Ben Xu, Jiang-She Zhang, and Wei Wang. A cell exclusion algorithm for determining all the solutions of a nonlinear system of equations. Applied Mathematics and Computation, 80(2–3):181–208, December 1, 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=2-3&aid=9500286.

Yagla:1982:TPC

[Yag82] Jon J. Yagla. Test problems, coordinate transformations, and technique for nonsteady compressible flow analysis. Applied Mathematics and Computation, 10/11(1):525–546, ???? 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Numerical grid generation (Nashville, Tenn., 1982).

Yakowitz:1992:EPA

[Yak92a] Diana Yakowitz. An exact penalty algorithm for recourseconstrained stochastic linear programs. Applied Mathematics and Computation, 49(1):39–62, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yakowitz:1992:DMM

[Yak92b] Sid Yakowitz. A decision model and methodology for the AIDS epidemic. Applied Mathematics and Computation, 52(1):149– 172, November 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yamamoto:1995:CSB

[Yam95] Tokufumi Yamamoto. Critical state between local and global interaction. Applied Mathematics and Computation, 73(2–3): 153–180, December 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/
                                        browse.cgi?year=1995&volume=
                                        73&issue=2-3&aid=9400247.

Yang:1987:DCM

[Yan87]  Hai Gang Yang.  Defect-

correction multigrid methods for nonlinear problems. Applied Mathematics and Computation, 23(4):359–364, ???? 1987. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:1994:PAA

[Yan94] Hai Lang Yang. A projective algorithm for approximation of fractal sets. Applied Mathematics and Computation, 63(2– 3):201–212, July 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yan:1997:OSO

[Yan97] Jurang Yan. Oscillations of second order neutral functional differential equations. Applied Mathematics and Computation, 83(1):27–41, April 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=1&aid=9600043.

Yang:1998:CSS

[Yan98]  Suh-Yuh Yang.  On the

convergence and stability of the standard least squares finite element method for firstorder elliptic systems. Applied Mathematics and Computation, 93(1):51–62, July 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/1/6053.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
1&aid=6053.

Yang:1999:CSL

[Yan99]  Z. J. Yang.  Cylindrical

symmetry-like solutions of Laplace equation ∇^{2}V (x_{j}) = 0. Applied Mathematics and Computation, 99(1):29–34, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/1/6164.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
1&aid=6164.

Yen:1999:NAI

[YB99] Jerome Yen and Tung X. Bui. The negotiable alternatives identifier for group negotiation support. Applied Mathematics and Computation, 104 (2–3):259–276, September 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/104/2-
3/6291.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
104&issue=2-3&aid=6291.

Yildiz:1997:MCF

[YBR97] Bünyamin Yildiz, Mustafa Bayram, and Mahir A. Rasulov. On a method of calculation of the first phase saturation during the process of displacement of oil by water from porous media. Applied Mathematics and Computation, 85(1):1–16, August 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1997/85/1/9600082.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1997&volume=85&issue=
1&aid=9600082.

Yevl:1998:PGP

[YCZ98] G. Yevl, P. Cinnella, and X. Zhuang. On parallelizing a groundwater pollution simulator. Applied Mathematics and Computation, 89(1–3):313– 325, January 1, 1998. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781665.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).

Yamashiro:1979:ORL

[YDB79] S. M. Yamashiro, J. A. Daubenspeck, and F. M. Bennett. Optimal regulation of left ventricular ejection pattern. Applied Mathematics and Computation, 5(1):41–54, ???? 1979. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ye:1998:DDLa

[Ye98a] X. Ye. Domain decomposition for a least-square finite element method for second order elliptic problem. Applied Mathematics and Computation, 91(2–3):233–242, May 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/2-3/
6014.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
2-3&aid=6014.

Ye:1998:DDLb

[Ye98b] X. Ye. Domain decomposition for least-squares finite element methods for the Stokes equations. Applied Mathematics and Computation, 97 (1):45–53, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/1/6142.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
1&aid=6142.

Ye:1999:TGD

[Ye99] X. Ye. Two grid discretizations with backtracking of the stream function form of the Navier–Stokes equations. Applied Mathematics and Computation, 100(2–3):131–138, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/

store/amc/sub/1999/100/2-
3/6208.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6208.

Yee:1993:ADM

[Yee93] Eugene Yee. Application of the decomposition method to the solution of the reactionconvection-diffusion equation. Applied Mathematics and Computation, 56(1):1–27, June 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yeh:1997:TRM

[Yeh97] Tyan Yeh. Trapezoidal rule for multiple integrals over hyperquadrilaterals. Applied Mathematics and Computation, 87(2– 3):227–246, December 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=2-3&aid=9600286.

Yang:1991:MNA

[YF91] Jun Yang and Chang Xin Fan.

Modeling of nonlinear and active cochlear mechanics and its Volterra-series analysis. Applied Mathematics and Computation, 43(3):241–263, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yakowitz:1990:CAM

[YGH90] S. Yakowitz, J. Gani, and R. Hayes. Cellular automaton modeling of epidemics. Applied Mathematics and Computation, 40(1 (part I)):41–54, ???? 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yakowitz:1997:MOD

[YH97] Diana S. Yakowitz and Keith W. Hipel. Multiple objective decision making for “Lokahi” (balance) in environmental management. Applied Mathematics and Computation, 83(2–3): 97–115, May 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
83&issue=2-3&aid=9600171.

Hou:1999:CPF

[yHyG99] Jing yu Hou and Ben yu Guo. Chebyshev pseudospectral-finite element method for the threedimensional unsteady Navier– Stokes equation. Applied Mathematics and Computation, 101 (2–3):209–244, June 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/101/2-
3/6225.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/

browse.cgi?year=1999&volume=
101&issue=2-3&aid=6225.

Yildiz:1996:MIP

[Yil96] Bünyamin Yildiz. A method increasing preciseness of finite differences system and its application. Applied Mathematics and Computation, 75(2–3): 253–267, March 15, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
75&issue=2-5&aid=9500129.

Yip:1991:IFD

[Yip91] Paul Yip. Inference on a fatal disease. Applied Mathematics and Computation, 41(1 (part I)):89–97, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yip:1992:IPS

[Yip92] P. Yip. An inference procedure for a simple epidemic model: a martingale approach. Applied Mathematics and Computation, 47(1):77–89, January 1992. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yip:1993:SIP

[Yip93] Paul Yip. Statistical inference procedure for a hypergeometric model for capture-recapture experiments. Applied Mathematics and Computation, 58(2–3):

225–240, October 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yuan:1999:CGA

[YJ99] Jin-Yun Yuan and Xiao-Qing Jin. Convergence of the generalized AOR method. Applied Mathematics and Computation, 99(1):35–46, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/1/6176.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
1&aid=6176.

Yildiz:1997:TMS

[YK97a] Bünyamin Yildiz and Ahmet Kacar. Two methods for solution of a linear boundary value problem of order 2n. Applied Mathematics and Computation, 86(2–3):215–233, October 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
86&issue=2-3&aid=9600183.

Yon:1997:TLA

[YK97b] Yoon J. Yon and Do Y. Kwak. Two-level additive Schwarz preconditioners for P 1 nonconforming finite elements for nonsymmetric and indefinite problems. Applied Mathematics and Computation, 87(1):1– 14, November 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=1&aid=9600192.

Yu:1996:HMC

[YL96] Bo Yu and Zhenghua Lin. Homotopy method for a class of nonconvex Brouwer fixed-point problems. Applied Mathematics and Computation, 74(1):65– 77, January 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
74&issue=1&aid=9500089.

Yang:1998:UAW

[YL98a] Suh-Yuh Yang and Jinn-Liang Liu. A unified analysis of a weighted least squares method for first-order systems. Applied Mathematics and Computation, 92(1):9–27, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/1/6032.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=92&issue=
1&aid=6032.

Yao:1998:TDA

[YL98b] Xin Yao and Yong Liu. Towards designing artificial neural networks by evolution. Applied Mathematics and Computation, 91(1):83–90, April 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/91/1/6005.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=91&issue=
                                        1&aid=6005.

Yakowitz:1993:MAD

[YLS93] D. S. Yakowitz, L. J. Lane, and F. Szidarovszky. Multi-attribute decision making: Dominance with respect to an importance order of the attributes. Applied Mathematics and Computation, 54(2–3):167–181, March 15, 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yavneh:1998:MSS

[YO98] Irad Yavneh and Elena Olvovsky. Multigrid smoothing for symmetric nine-point stencils. Applied Mathematics and Computation, 92(2–3):229–246, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/92/2-3/
                                        9710035.pdf;   http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/amc/cas_sub/browse/

browse.cgi?year=1998&volume=
92&issue=2-3&aid=9710035.

Yakowitz:1984:CAD

[YR84] Sidney Yakowitz and Brian Rutherford. Computational aspects of discrete-time optimal control. Applied Mathematics and Computation, 15(1):29–45, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:1994:SAG

[YS94] Jiann Cherng Yang and Bharat K.

Soni. Structured adaptive grid generation. Applied Mathematics and Computation, 65(1– 3):265–278, September/October 15, 1994. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Differential equations and computational simulations, I (Mississippi State, MS, 1993).

Yi:1998:SPC

[YS98a] Fahuai Yi and T. M. Shih.

Stefan problem with convection. Applied Mathematics and Computation, 95(2–3): 139–154, September 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/95/2-3/
6022.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=95&issue=
2-3&aid=6022.

Yildiz:1998:ICP

[YS98b] Bünyamin Yildiz and Ali Sever. On the inverse conductivity problem. Applied Mathematics and Computation, 94 (1):91–96, August 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/94/1/6043.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=94&issue=
1&aid=6043.

Yserentant:1986:HBG

[Yse86] Harry Yserentant. Hierarchical bases give conjugate gradient type methods a multigrid speed of convergence. Applied Mathematics and Computation, 19(1– 4):347–358, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Yang:1997:ASH

[YST97] Xiaofeng Yang, Tielong Shen, and Katsutoshi Tamura. Approximate solution of Hamilton– Jacobi inequality by neural networks. Applied Mathematics and Computation, 84(1): 49–64, June 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/

browse.cgi?year=1997&volume=
84&issue=1&aid=9600053.

Yurekli:1999:NII

[Yür99] Osman Yürekli. New identities involving the Laplace and the L_{2}-transforms and their applications. Applied Mathematics and Computation, 99 (2–3):141–151, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6186.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6186.

Young:1985:REU

[YW85] Peter Young and Stephen Wallis. Recursive estimation: a unified approach to the identification, estimation, and forecasting of hydrological systems. Applied Mathematics and Computation, 17(4):299–334, ???? 1985. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). System identification and parameter estimation of environmental systems.

Zemanian:1988:FDA

[ZA88] A. H. Zemanian and Hyeong Keon An. Finite-difference analysis of borehole flows involving domain contractions around threedimensional anomalies. Applied Mathematics and Computation, 26(1):45–75, ???? 1988. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zaghrout:1991:OSO

[ZA91] A. A. S. Zaghrout and Z. Ali. On oscillations of second order linear differential equations. Applied Mathematics and Computation, 46(2 (part II)):135–144, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zaghrout:1996:AMS

[ZA96] A. A. S. Zaghrout and S. H. Attalah. Analysis of a model of stage-structured population dynamics growth with time statedependent time delay. Applied Mathematics and Computation, 77(2–3):185–194, July 1996. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zaghrout:1996:OGA

[ZAES96] A. Zaghrout, A. Ammar, and M. M. A. El-Sheikh. Oscillations and global attractivity in delay differential equations of population dynamics. Applied Mathematics and Computation, 77 (2–3):195–204, July 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zaghrout:1991:SPF

[Zag91] Afaf A. S. Zaghrout. Stability and persistence of facultative mutualism with populations interacting in a food chain.

Applied Mathematics and Computation, 45(1 (part I)):1–15, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zaghrout:1992:ABS

[Zag92] Afaf A. S. Zaghrout. Asymptotic behavior of solutions of competition in gradostat with two limiting complementary substrates. Applied Mathematics and Computation, 49(1):19–37, May 1992. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zakeri:1989:GSW

[Zak89] Gholam-Ali Zakeri. Geometric structure of 2D weak shock waves. Applied Mathematics and Computation, 33(3):161– 183, October 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zamani:1984:CFE

[Zam84] Nader G. Zamani. A collocation finite element method for integration of the boundary layer equations. Applied Mathematics and Computation, 15(2):109– 122, ???? 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zamani:1988:BES

[Zam88] N. G. Zamani. Boundary element simulation of the cathodic protection system in a prototype ship. Applied Mathematics and Computation, 26(2):119– 134, 1988. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

Zang:1987:NTL

[Zan87] Zhu Quan Zang. A note on the theorem of Lakshmikantham and Leela. Applied Mathematics and Computation, 23(3):269– 275, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zaslavsky:1993:AAM

[Zas93]  Leonid Yu. Zaslavsky.  An

adaptive algebraic multigrid for multigroup neutron diffusion reactor core calculations. Applied Mathematics and Computation, 53(1):13–26, January 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:1997:CSA

[ZB97] Jun Zhang and John A. Belward. Chebyshev series approximations for the Bessel function Y_{n}(z) of complex argument. Applied Mathematics and Computation, 88(2–3):275– 286, December 30, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
88&issue=2-3&aid=9600335.

Zaghrout:1991:GGB

[ZED91] A. A. S. Zaghrout and N. El-Dawaliby. Generalizations of Gronwall-Bellman-Reid inequalities. Applied Mathematics and Computation, 41(2 (part II)): 129–143, ???? 1991. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zeldin:1987:TAL

[Zel87] Yevgeny Zeldin. Topological analysis of large networks. Applied Mathematics and Computation, 23(3):193–210, ???? 1987. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zemanian:1986:MMN

[Zem86] A. H. Zemanian. Multilevel marketing networks in lessdeveloped countries. Applied Mathematics and Computation, 18(3):185–210, ???? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zeng:1999:HAF

[Zen99] Pan Zeng. High-accuracy formula for discrete calculation of Fourier transforms. Applied Mathematics and Computation, 106(2–3):117–140, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/18/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/18/
article.pdf.

Zufiria:1990:CMF

[ZG90] Pedro J. Zufiria and Ramesh S.

Guttalu. A computational method for finding all the roots of a vector function. Applied Mathematics and Computation, 35(1 (part I)):13–59, January 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zang:1986:SMM

[ZH86] Thomas A. Zang and M. Yousuff Hussaini. On spectral multigrid methods for the timedependent Navier–Stokes equations. Applied Mathematics and Computation, 19(1–4):359– 372, ???? 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Second Copper Mountain conference on multigrid methods (Copper Mountain, Colo., 1985).

Zhang:1993:SVH

[ZH93] Li Qing Zhang and Guo Qiang Han. Structure variable homotopy methods for solving nonlinear systems. Applied Mathematics and Computation, 59(2–3): 177–191, December 1993. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:1989:SRI

[Zha89]  Wei-Bin Zhang.  Short-run

inventory oscillations in the Eckalbar disequilibrium macro model. Applied Mathematics and Computation, 33(1 (part I)): 53–67, ???? 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:1991:AEG

[Zha91] Li Qing Zhang. Asymptotic expansion and geometric properties of the spline collocation periodic solution of an ODE system. Applied Mathematics and Computation, 42(3):209– 221, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:1995:SCT

[Zha95] B. G. Zhang. Sturm comparison theorem of difference equations. Applied Mathematics and Computation, 72(2–3): 277–284, October 1, 1995. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
72&issue=2-3&aid=9400201.

Zhang:1996:NWW

[Zha96a] Hong Zhang. A note on windowing for the waveform relaxation method. Applied Mathematics and Computation, 76 (1):49–63, April 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
76&issue=1&aid=9500149.

Zhang:1996:AFP

[Zha96b] Jun Zhang. Acceleration of fivepoint red-black Gauss–Seidel in multigrid for Poisson equation. Applied Mathematics and Computation, 80(1):73– 93, November 1, 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1996&volume=
80&issue=1&aid=9500276.

Zhang:1997:MRS

[Zha97a] Jun Zhang. Minimal residual smoothing in multi-level iterative method. Applied Mathematics and Computation, 84 (1):1–25, June 1, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
84&issue=1&aid=9600050.

Zhang:1997:RST

[Zha97b] Jun Zhang. Residual scaling techniques in multigrid. I. Equivalence proof. Applied Mathematics and Computation, 86(2–3):283–303, October 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
86&issue=2-3&aid=9600194.

Zhao:1997:EFF

[Zha97c] Yun-Bin Zhao. Exceptional families and finite-dimensional variational inequalities over polyhedral convex sets. Applied Mathematics and Computation, 87(2– 3):111–126, December 1, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1997&volume=
87&issue=2-3&aid=9600224.

Zhang:1998:RST

[Zha98a] Jun Zhang. Residual scaling techniques in multigrid. II. Practical applications. Applied Mathematics and Computation, 90(2–3):229–252, March 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/90/2-3/
9700400.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
90&issue=2-3&aid=9700400.

Zhang:1998:TGA

[Zha98b] Jun Zhang. Two-grid analysis of minimal residual smoothing as a multigrid acceleration technique. Applied Mathematics and Computation, 96 (1):27–45, October 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/1/6118.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.

                                        cgi?year=1998&volume=96&issue=
                                        1&aid=6118.

Zhang:1998:VCV

[Zha98c] Jun Zhang. VML: a class of virtual multi-level iterative methods for solving partial differential equations. Applied Mathematics and Computation, 92(1):29–48, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/92/1/6029.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=92&issue=
                                        1&aid=6029.

Zhao:1998:GCF

[Zha98d] Jinxi Zhao. The generalized Cholesky factorization method for saddle point problems. Applied Mathematics and Computation, 92(1):49–58, June 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1998/92/1/6030.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1998&volume=92&issue=
                                        1&aid=6030.

Zhang:1999:ASP

[Zha99a] Jun Zhang. Acceleration and stabilization properties of minimal residual smoothing technique in multigrid. Applied Mathematics and Computation, 100(2–3):151–168, May 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/100/2-
3/6213.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1999&volume=
100&issue=2-3&aid=6213.

Zhao:1999:DOS

[Zha99b] Yun-Bin Zhao. D-orientation sequences for continuous functions and nonlinear complementarity problems. Applied Mathematics and Computation, 106 (2–3):221–235, December ??, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.elsevier.nl/gej-ng/
10/9/12/71/18/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/9/12/71/18/27/
article.pdf.

Zheng:1993:MMC

[Zhe93] Chen Zheng. A mathematical model and computer program of economic evaluation and prediction. Applied Mathematics and Computation, 59(1):63– 72, November 1993. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Han:1999:FFF

[zHhT99] Zheng zhong Han and Yu hua Tang. Fuzzy forecast of flood disaster caused by solar proton flares. Applied Mathematics and Computation, 98 (1):83–89, January 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/98/1/6156.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=98&issue=
                                        1&aid=6156.

Zhou:1989:SSB

[Zho89] Bin Zhou. A-stable and Lstable block implicit one-step methods with modified algorithm. Applied Mathematics and Computation, 31(Special issue): 419–436, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Zhu:1994:SMF

[Zhu94] Jianping Zhu. Simulating multiphase flow in porous media on KSR-1 parallel computers. Applied Mathematics and Computation, 65(1–3):19–31, September/October 15, 1994. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhu:1998:AAN

[Zhu98] Wen-Xing Zhu. An approximate algorithm for nonlinear integer programming. Applied Mathematics and Computation, 93(2–3):183–193, July 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/93/2-3/
6085.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=93&issue=
2-3&aid=6085.

Zia:1989:EPT

[Zia89] Lee L. Zia. Estimation of parameters in two-dimensional transport equations. Applied Mathematics and Computation, 30 (1 (part I)):19–47, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zielinski:1984:AFP

[Zie84] Przemyslaw Zielinski. An application of the Fokker–Planck equation in stochastic reservoir theory. Applied Mathematics and Computation, 15(2): 123–136, ???? 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zaghrout:1998:CBT

[ZK98] A. A. S. Zaghrout and F. M.

Kandil. Competition between three microbial populations for a single limiting resource in continuous culture. Applied Mathematics and Computation, 92 (2–3):271–281, June 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/92/2-3/
9710027.pdf;   http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
92&issue=2-3&aid=9710027.

Zouyousefain:1990:SRD

[ZL90] M. Zouyousefain and S. Leela. Stability results for difference equations of Volterra type. Applied Mathematics and Computation, 36(1 (part I)):51– 61, March 1990. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zhang:1998:ASP

[ZL98a] Wenfei Zhang and Xiaojiang Li. Analytic solution of poromechanical problems in a hollow axisymmetric domain. Applied Mathematics and Computation, 97(2–3):209–221, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6077.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6077.

Zhu:1998:AMB

[ZL98b] Song-Ping Zhu and Huan-Wen Liu. On the application of the multiquadric bases in conjunction with the LTDRM method to solve nonlinear diffusion equations. Applied Mathematics and Computation, 96(2–3):

161–175, November 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/2-3/
6126.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
2-3&aid=6126.

Zhu:1998:FWA

[ZL98c] Tonglin Zhu and Wei Lin. Fast wavelet algorithm of the Poisson integral. Applied Mathematics and Computation, 96(2– 3):127–144, November 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/96/2-3/
6122.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=96&issue=
2-3&aid=6122.

Zlatev:1989:ATV

[Zla89] Zahari Zlatev. Advances in the theory of variable stepsize variable formula methods for ordinary differential equations. Applied Mathematics and Computation, 31(Special issue): 209–250, ???? 1989. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). Proceedings of the 1986 ODE Conference: Numerical ordinary differential equations (Albuquerque, NM, 1986).

Zhang:1991:GMV

[ZM91] Hong Zhang and Xiao-Yi Min. On the Galerkin method with vector basis functions. Applied Mathematics and Computation, 42(1 (part I)):85–103, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Zorzano:1999:NST

[ZMV99] M. P. Zorzano, H. Mais, and L. Vazquez. Numerical solution of two dimensional Fokker– Planck equations. Applied Mathematics and Computation, 98(2– 3):109–117, February 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/98/2-3/
                                        6163.pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.
                                        cgi?year=1999&volume=98&issue=
                                        2-3&aid=6163.

Zheng:1999:MMU

[ZPY99] Chen Zheng, He Ping, and Chen Yan. A mathematical model using artificial neural networks to forecast shares tendency. Applied Mathematics and Computation, 99(1):71–76, March 1, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/amc/sub/1999/99/1/6174.
                                        pdf;   http://www.elsevier.
                                        com/cgi-bin/cas/tree/store/
                                        amc/cas_sub/browse/browse.

cgi?year=1999&volume=99&issue=
1&aid=6174.

Zaslavsky:1998:AMT

[ZS98] Leonid Y. Zaslavsky and Tamar Schlick. An adaptive multigrid technique for evaluating long-range forces in biomolecular simulations. Applied Mathematics and Computation, 97 (2–3):237–250, December 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/2-3/
6140.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
2-3&aid=6140.

Zheng:1999:SSI

[ZS99] Shiming Zheng and Fangyu Sun. Some simultaneous iterations for finding all zeros of a polynomial with high order convergence. Applied Mathematics and Computation, 99 (2–3):233–240, March 15, 1999. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1999/99/2-3/
6193.pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1999&volume=99&issue=
2-3&aid=6193.

Zhang:1998:LAT

[ZSL98] Y. G. Zhang, M. Sugisaka, and X. J. Li. Lifelike artificial trees based on Growth Iterated Function System. Applied Mathematics and Computation, 91(1):3–8, April 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/91/1/6010.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=91&issue=
1&aid=6010.

Zhou:1995:NAO

[ZW95a] Jing Zhou and Luther W. White. Numerical approximation of optimal control of fluid equations with buoyancy. Applied Mathematics and Computation, 71(1):47–68, August 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=1&aid=9400139.

Zhou:1995:OCF

[ZW95b] Jing Zhou and Luther W. White. Optimal control of fluid equations with buoyancy. Applied Mathematics and Computation, 71(1):15–45, August 1, 1995. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1995&volume=
71&issue=1&aid=9400138.

Zeng:1998:SAS

[ZZ98] Jinping Zeng and Shuzi Zhou.

Schwarz algorithm for the solution of variational inequalities with nonlinear source terms. Applied Mathematics and Computation, 97(1):23–35, December 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.elsevier.com/cas/tree/
store/amc/sub/1998/97/1/6136.
pdf;   http://www.elsevier.
com/cgi-bin/cas/tree/store/
amc/cas_sub/browse/browse.
cgi?year=1998&volume=97&issue=
1&aid=6136.

Zhong:1998:SAT

[ZZZ98] Wanxie Zhong, Xinglai Zhuang, and Jianping Zhu. A selfadaptive time integration algorithm for solving partial differential equations. Applied Mathematics and Computation, 89 (1–3):295–312, January 1, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/amc/cas_sub/browse/
browse.cgi?year=1998&volume=
89&issue=1-3&aid=9781664.

Differential equations and computational simulations, II (Mississippi State, MS, 1995).